A Complete Bibliography of Publications in Mathematics of Computation, 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

02 January 2024 Version 1.42

Title word cross-reference

(1; e) [Sij12]. (2, 1) [SSV14]. (2^{n}k + 1)_{n≥1} [CLPM16]. (ℓ, ℓ) [CR15]. 1 [ABBM18, BFZ10, CZ11, DKMW13, DHMG11, EHR18, IPZ15, KMPW10, Ort11, Rau16, Sør16, Tao14, XZ10]. 1, 2, 3 [LWCI13]. 1, 2, 4 [LWCI13]. 10^{1500} [OR12]. 10^{24} [CS10]. 11 [FG17, HKÖ11]. 12 [CLO14]. 16 [Zie10a]. 17 [GY19]. 1^{k} + 2^{k} + ⋯ + (m − 1)^{k} = m^{k} [GMZ11]. 2 [AN15, AK12, BKN15, Cer11, CLTZ12, CFLTL14, Cos10, CR15, Fis17, GM13b, KP12, LZ16, LWZ18, MP17, MRSC19, OS13]. 2k [CH13, HZ17]. 2m [WX19c, WX13]. 2 × 2 [JLM13]. 3 [ALS17b, BHLL19, BZ18, DD19, HFC18, Kop19b]. 3 × 3 × 3 [BH13]. 3z^{p} [DF14]. 4 [BZ18]. 4 · 10^{18} [OHP14]. 4 × 4 [BvH11]. 72 [NP14]. 8 [KÖ18]. > dim R [Yen11]. _{2}H_{2} [Chu11]. _{3}F_{2} [Chu12]. _{5}F_{4} [CZ14a]. b [Ser17]. b_{1}^{2} + b_{2}^{2} + 2^{n^{2}} [PZ11]. C^{0} [BGNyS11]. C^{1} [DKS15]. C^{s} [ADL11]. D [BJM17, FN11, Gra19a]. D_{ℓ} [Coh15]. E_{8} [GH15]. E_{p−3} [Meš14]. ℓ [CVZ19]. ℓ^{1} [YMO13]. ℓ_{1} [CD18, BCKM18]. ϵ [JK17, BHH^{+}12]. exp(1/e) [TK12]. f [BZ18]. F_{2} [HMS11]. F_{4}(2) [BHLL19]. F_{q}[t] [Kir12]. G^{2} [JKK^{+}10]. γ_{k}(a) [KC11a]. H(curl) [Cal16]. H(div) [EFP10]. H(curl) [CW17]. H(div) [CW17]. H^{1} [HFC18, DT11]. H^{k} [HZ15]. hp [HW18, LCQ17, SSW16, qWlGjY17, ZqCt16, FW11, MR10, SW10]. ∞ [Was14]. k [BM12, CH13, CS13a, RR16]. K3 [Cor10]. k ≥ 6 [Zha11a]. Kp^{n} + 1 [GOMS15]. L [BF18, Büt15, Lan11, Ram16, Rob15, RTwaabRS16, ST18a, Tru15]. L(1/3) [Bia14]. L^{1} [KRS14, Xu14a, BFZ10, GLL12]. L^{2} [CCLX17, Che12, DE13]. L^{∞}(L^{2}) [BM11]. L^{p} [LS17, MNPW10]. L_{∞}(0, T; L_{2}(Ω)) [LMP15]. L_{m}(s) [CFJ12]. Λ [BÇS15]. ≤ 2 [AK12]. M [Azi19, CFS17]. m = n + 1 [WX19c]. 𝔽_{p^{m}} [Zie10b]. ℚ [BGR19, DF14, BF18]. ℚ(ζ_{5}) [ZX16]. ℝ^{3} [GL15]. ℝ^{d} [GKS13, GKS17b, HNRW18]. ℝ^{n} [WX19c]. ℤ_{2k} [HM14]. d [CZ11]. GSp_{4} [Dem14]. H^{1} [BS15b]. P^{1} [Tui16]. Q [Mor19]. C [CM13]. H [FMP16]. L_{0} [ZDL13]. N [DI15, ABE^{+}16, CFO^{+}15, Har14, ZC13]. O(1/n^{2}) [TY19]. O(log^{2}(N)) [GOM11]. p [BÇS15, BMS16, Cre14, DMPP19, Dum17, Gra19b, Meš14, MT19, PV15, Rob15, Sut11b, ZY15, Zra10b]. p − 1 [Zra10b, Zra10a]. p = 13 [Sza11]. p^{2} + b^{2} + 2^{n} [PZ11]. p^{k} [Zie10b]. P_{1} [HFC18]. p_{g} = 0 [BP12a]. p_{g} = q = 1 [Rit10]. π [CZ14a, Chu11]. π(x) [Büt16, FKBJ17, Pla15, SD11]. π(x) − li(x) [STD15, SD10]. ψ(x) [Büt18, FK15, FK18a]. ψ, θ [Dus16]. q [LL15, WO10]. q > 1 [WO10]. Q_{k} [HZ17]. QR [BEG16, DOZ13, CSV12]. R [CSV12, HP12]. R[X] [Yen11]. R^{3} [ZHX11]. R^{n} [HZ15, WX13]. S [BB17]. s > 1 [ADL11]. S_{γ}(n) [KC11a]. SL(3, Z) [Mez11]. SL_{3}(ℂ) × SL_{3}(ℂ) × SL_{3}(ℂ) [BH13]. SL_{4}(ℤ) [AGM10]. \sqrt{2} [KT10]. T [DGKS16]. t_{ϵ} [ZC18]. τ [KZ18]. θ [dGO14]. θ(x) − x [PT16]. u [BHM^{+}11, BHM^{+}12a]. V [SZ18]. φ [FLM14]. w [KZ18]. W_{∞}^{1} [DLSW12]. W_{p}^{1} [Li17]. x [Dus16]. x < 10^{10^{13}} [SD11]. x^{2} − 1 [LN11, LN14]. x^{5} + y^{5} = 2z^{p} [DF14]. x^{5} + y^{5} = dz^{p} [BD10a]. X_{0}^{∗}(N) [BR19]. X_{1}(5) [KK10]. X_{1}(6) [KK10]. X_{1}(N) [Baa10]. x_{1} + x_{2} + ⋯ + x_{k} + c = x_{k+1} [ABE^{+}16]. Z[u] [BHM^{+}11, BHM^{+}12a]. ζ(s) [SD11]. |Δ| < 2^{40} [MJ16].

-abelian [RR16]. -adaptive [HW18]. -adaptivity [MR10]. -adic [BÇS15, BMS16, CVZ19, Dum17, Gra19b, KZ18, MT19, Rob15, Ser17, BÇS15]. -based [DOZ13]. -Bernstein [WO10]. -codes [HM14]. -color [ABE^{+}16]. -conforming [HFC18]. -conjecture [CH13]. -cube [KÖ18]. -curves [Sij12, BF18, DF14]. -D [EHR18, IPZ15, XZ10]. -decompositions [CFS17]. -descent [CFO^{+}15]. -descents [Cre14]. -designs [ZC18]. -dGFEM [SSW16]. -dimensional [NP14]. -Discontinuous [FW11]. -error [BFZ10]. -estimates [DE13]. -finite [BJM17, DKS15]. -formulae [Chu11]. -fractional [ABBM18]. -function [FLM14]. -functions [CM13, Rob15, BF18, Büt15, Lan11, RTwaabRS16, ST18a, Tru15]. -groups [DMPP19, Sut11b, dGO14]. -HDG [LCQ17]. -integrability [SZ18]. -inverses [LWCI13].

-isogenies [CR15]. -linear [HMS11, HP12]. -matrix [FMP16, Azi19]. -modular [BHLL19]. -Optimal [SW10]. -polytopes [BZ18]. -power [AK12]. -projection [Che12]. -radius [BM12, CLTZ12]. -ramified [PV15]. -reflection [MP17]. -regularized [BCKM18]. -series [Chu11, Chu12, Ram16]. -singular [LL15]. -spectral [ZqCt16]. -spheres [BZ18]. -stage [Cer11, GM13b]. -sum [CZ14a]. -Superconvergence [HFC18]. -Sylvester [DGKS16]. -term [DI15]. -th [CS13a, Har14]. -th-order [WX13]. -torsion [CFLTL14, DD19, Fis17, Zie10b]. -tuples [ZC13]. -units [BB17]. -variate [Was14]. -vectors [BZ18]. -version [qWlGjY17]. -wave [FN11].

12-split [CLR13].

Abel [MN19a]. abelian [AK12, BMS16, Bis15, HL16, JdRV14, RR16, Sut11b]. absolute [FR15a]. Absolutely [FX13, ÁB17, BHS15, LCQ17, Sch17]. absorbing [BLW10]. absorption [BDG^{+}19, GSV17]. Accelerated [GK11, TY17, LW16]. Accelerating [CZ14a, Fan14]. Acceleration [HC11, HC15, Adc11]. Accuracy [JXR12, CL10, GSS16, JvSRV14, XY13]. accuracy-conserving [JvSRV14]. Accuracy-enhancement [JXR12]. Accurate [Ye18, AH15a, BFS18, CCL^{+}19, CPBG19, CPSV18b, ET10, JLM13, Kop14]. achieve [CQS14]. achieved [LPSH11]. Acoustic [LMMR19, CQB14]. acting [JKdR16]. action [BH13]. active [CD18]. adapted [MC12]. Adaptive [BD15, Bör18, CGS18, CNV14, CGS15, CDHM12, DVY15, GV15, HMS14, Sfa13, BLW10, BBSV19, BGGG17, BMBO13, CR11a, CD15, CNZ16, DFR12, GM14, GM17, Git13, HW18, KSU16, Kre12, KG18, OWZD18, XZ15, ZCS^{+}12]. adaptivity [Bar15b, MR10]. Addenda [Alk11, BHM^{+}12a]. Additive [Vab14, GLW19, HK18a, Vab12]. adelic [CFLTL14]. adic [BÇS15, BMS16, CVZ19, Dum17, Gra19b, KZ18, MT19, Rob15, Ser17, BÇS15]. adjacent [KZ18]. adjoint [PMH18]. admissible [BCL^{+}11]. Advances [BBC10]. advection [BNMP18, HLL16, YDk12]. AEDG [LW18b]. AFEM [MMN11]. affine [Alf10, DJ12a, GLM19b, Kir10, SY16]. aggregation [CB16]. AGM [Dup11]. algebra [IKRS12, Kem10]. Algebraic [BY12, GPOS14, AZ18, Bat15, BHM12b, Dem14, DK15, HZ13, HMT19, Lez14, McK11, NS13, RVX19, RSTV10, Sid12]. algebras [BJ11, DJ12b, PS14]. algebroid [Shi13]. Algorithm [Han15, ABF^{+}10, ABBR17, BO15, Bia14, BST11, BJ15, BHH^{+}12, Cal16, Chè13, FCW19, FJK^{+}15, FK18b, GGG11, GOP16, GCL19, Har10, Har14, Hia16b, HY15, JLM13, JLH13, Joh15, KCL14, LM17, LW16, LGY15, Mor13, NJZ17, PPTZ13, ZY14, Zra10a]. Algorithmic [BD13, PS19]. Algorithms [DDLR15, HZ13, MN14, SW13, AGHS14, AZ11, BGH14, BOP17, BP12b, BCCW16, CPSV18a, CKS19, GJ19, GMS12, Har11, Kus18, LZ18, Yan17, CFO^{+}15, IKRS12, OWZD18]. aliquot [Boo18]. Allen [BM11]. Almost [KB16, Kön17]. along [He18]. alpha [DEJ14]. alternate [MS16]. alternating [TY19, YY13]. alternative [Hia16a]. Alvero [CLO14]. amicable [NP19]. amoebas [FMMdW19, TdW15]. amortized [Hia11]. amortized-complexity [Hia11]. Ampère [BCM16, BGNyS11, NNZ19]. Anadromic [LK12]. analog [RGNS18]. analogue [BMS16]. Analogues [MS16]. analyses [Mus18]. Analysis [AV14, BPT10, COT17, CCQ17, CSV12, CC14, CL16, CGN^{+}11, GH18, GZ13, GOS11, HS15, JLZ19, KCL14, KSX17, Kre12, OS13, SS16, Wan10a, Arg11, Bar10, BCFG17, BJV18, BGM15, BP12b, BGP14, Cao15, CCS12, CT10, CLWW16, CS13a, CGS10, CC12b, CQS14, CDHM12, DE10, DGM^{+}16, DPV19, DJTZ13, DRS11, FJS16, GMSC^{+}17, GM10, GQ14, Gud10, GMS12, HK17, HMP13, JLM13, KO12, Kop19a, KPY18, Li15a, LS17, LZ17, Li18, LY17, MRSC19, MS10, OMS15, Ols12, Ort11, RS16, Sid10, TV10, Voh10, XZ10, Gav10a]. Analytic [Gau17, LP17c, AH15a, Büt18, FKBJ17, HLJ16, KSW17, YF15]. Analytical [Chu12]. analytically [Pla15, SW11]. Analyticity [Li19]. Ando [BMP10]. angular [BC13a]. anisotropic [AADL11, CG16, Cao15, CDHM12, DI15, DT11, EE18, HP10, HZ17, KHX14, Kop14, Li18, RS16, Sty14]. annihilators [ST18a]. anomalous [CLAT12]. ANOVA [GKS17b, GKS13, GKS17a]. any [AADL11]. AOR [MYS12]. AP [BDFL17]. Apostol [Bay11, BC13b, NRV12]. appearing [KC11a]. Application [CMZ17, GÖS18, APR15, BJV18, BL12, CLTZ12, DE10, DL15, FA14a, KSX17, Źra19, Alf10]. applications [AH15b, BB16, BNS15, BH16a, BMM18, CSW16, Fre18, HYZ14, JR18, JZ11, MRSC19, Seg17, Ste11, Gav10b]. applied [CMTZ16, DHMG11, HS13]. approach [AHS17, CJ15, DS17, Gav10b, GJ19, HLZ15, HW18, HLJ16, Kim17, KM16, KT18, LM13, RR13, Sog15, Sog16, ZY15, ZZWZ15]. approaches [HK17]. approximants [FMP16]. Approximate [GI16, MPT18, BO15, BCL^{+}11, Pan17]. Approximating [TdW15, HHT19, PPTZ13, Was14]. Approximation [BG11, BMR16, GM14, GM17, HP10, Kve10, Rah14, Wan10b, AN15, AL16, BGN17, Bar10, BM11, Bar16, BCFG17, BKMW11, BJM17, BP15, BPT10, BTDG13, CCDE19, CDFS13, CDTW18, Chr18, CMS11, DPR14, Deb11, DI15, DLSW12, FMMdW19, FHN19, Gal19a, Gal19b, GGG11, GO16, HLZ15, HZ17, Hir13, HJW18, Jag12, KS19, KP14, LW18a, LZ15, LPSH11, LR19b, LZ18, MNZ13, MNPW10, Mon10b, NT16, OS13, RS17a, RV14, SS15, SS16, UP14, WX12, WMxY17, Wen13, YF15]. approximations [AA11, AD13, BC17, ByTC13, BLT13, BFZ10, BCL^{+}11, CH19, CPBG19, CW10, GV15, Gon13, GLNP15, GSRM12, GL12, GSS16, JK17, KRS14, LS10a, LLX11, LWL19, LA14, NJZ17, ST18b, TZD15, Wan10a, ZC18, ZD14, dDHZZ14]. Arbitrary [FSX19, KSX17, Fon11, KHX14, Sid12, YF15]. Archimedean [KN12]. argument [Tru12]. arising [CGO10, MC13]. Arithmetic [ABBR17, Sij12, DFH18, FR15b, Mül14, NV17, Pag15, SW13]. arrays [BH13]. Artificial [BNDHV15]. Artin [AS15, PM11]. aspects [CS13b, LPD13]. assimilation [BH18]. associated [BS15b, GV11]. assumptions [Büt16]. asymmetric [BGN17, LZ15]. Asymptotic [IL11, NRV12, ND19, Sid10, Sid11, SZW11, BFS18, BLT13, KC11a, KC11b, LWZ18, PR15, SZ18, Wei19]. asymptotic-preserving [BLT13]. Asymptotics [GS10, IPZ15]. atomistic [OS13]. atomistic/continuum [OS13]. attached [RSS12, RTwaabRS16]. augmented [COT17, HYZ14, MYS12, YY13]. Automatic [BGU19, GLLZ14]. automorphism [LRS16]. automorphisms [GPOS14, HKW17]. auxiliary [CHH18, KCL14]. Average [KtR10, Was14, BB12, Shp11]. averaged [COS14, Spa17]. Avoidability [RR16]. avoiders [Kus18]. avoiding [ADL11]. axisymmetric [CGO10, Li15a]. axon [SS16].

B [CLR13, DPR14, Sdi19]. B-spline [Sdi19]. B-spline-like [CLR13]. B-splines [DPR14]. babystep [Hit18]. back [BI19, HLL16]. Backward [AK16, AMR^{+}19, DPV19, GT16, JX14]. bad [BDFP15]. balance [CPSV18b]. balanced [AG16, BC16, BCL15, CPPT10, DZBK16, Han10]. band [OS16a]. banks [HZ13, Han15]. barotropic [BGM15, MO13]. barrier [AZ11]. barycentric [Kle13, RGB14, RS17b, RS19, WHV14]. base [Ane19, KT10, TK12]. based [ABBR17, BO15, BCFG17, CZ12, CHZ17, CCS12, CD18, CGN17, CM14, CGS10, CDHM12, CLR13, DOZ13, FA14b, FR15b, HH13, HYZ14, HFC18, KPY18, LZ17, LM13, MP12, MSW16, NKK17, OS13, RW18, RVX19, TWO18, Voh10, Yan17, ZDL13, ZGFD14]. bases [BD11, CM19, CDFS13, GVW16, HH13, HNRW18, JZ11, JD14, KP11, La 15, LNRW18, Mel18, NRW17, SW17, Ste11]. basis [AJ14, CSV12, CLR13, DZ16, MNPW10, NH11, UP14]. Bass [JdRV14]. Baxter [GV17]. Bayesian [Mat19, ST18b]. BDDC [Cal16, OWZD18]. be [Kop14, Kop19b]. beam [BMR16, LRT13]. beams [CCS12]. Belı [Sij12]. Bell [MNW10]. Bellman [BFZ10]. Bellman’s [Kry13]. Beltrami [BP12b, HP12, MMN11]. BEM [FMP16, FHK17, HS15, MSS11]. bending [FHN19]. Bernoulli [Bay11, BC13b, Har10, Har14, NRV12]. Bernstein [MNPW10, Not15, Spa13, WO10]. Best [DLSW12]. Bethe [GCL19]. Better [BBKZ16]. between [ALS17a, LOX14, Ten10]. bi [CH13, FN11]. bi- [CH13]. bi-wave [FN11]. biased [MSW16]. bidegree [SSV14]. Bielliptic [BR19]. bifurcation [SY16]. biharmonic [EGHL12, LL17]. bilateral [Chu11]. billion [HHO17]. binary [BER17, CFvdG19, LOX14, Pro17, RJS12]. biomembranes [BGN17]. biorthogonal [Han10, LS10c]. bipartite [HÖ14]. biperiodic [BLW10]. Birch [BMS16, Ble12]. Biregular [Qur19]. bisection [MC12]. Bittner [LZ13]. bivariate [BL12, FA14a, KS15]. blob [CB16, HL17, LY17]. block [LSR19, ZN19]. block-centered [LSR19]. BMR [MP17]. body [FGMS12]. Boltzmann [CGP12, HY15]. Bombieri [AH15b]. Book [Gav10a, Gav10b, Gav11, Gra11, Hai10, Kem10, Naz10]. Borcherds [BEF16]. border [KP11]. Borwein [DJŠ18]. boson [HY15]. bound [AH14, BJ15, CW15, FJK11, SGD11, Sfa13, Ten10, Tho10, Tru12, Tru15, UP14]. boundaries [CPBG19]. boundary [ALS17b, APR15, AFF^{+}17, BR18, BHL18, CQT11, CGN17, CQS14, CG11, DKMW13, GHS10, Gav11, He18, HW18, Li15b, LLS17b, Lic19a, LR19b, LS10b, MS10, NL19, QS16, Ste11, WKN13, ZZ14]. bounded [Bar15b, BCPR14, DK15, FW14, Kru16, LWCI13, WX19a]. Boundedness [HS11, Spi17]. Bounding [CMSC10, Büt18]. bounds [BS18, BBR12, CG14, CGN17, DL13, FK15, FK18a, FS12, Fre18, GP17, GM18, Hur18, JKLM17, JR18, LLS15, LLS17a, Lep16, LP18b, LXX14, Lou11, MFRV18, Mel18, Mor11, Nie15, PR15, PM11, Spa13]. Boussinesq [AD13]. box [BBC10]. braces [GV17]. braid [GMS19]. branch [vdBLM10]. Brauer [Gra19b]. Bregman [YMO13]. Brent [BJ15]. Brownian [JLH13]. Bruijn [SGD11]. Brumer [GRT15, GRT04]. Buffa [NKK17]. Burgers [ALS17a, HLR13]. Burmeister [PKLC14].

cable [LLX11]. Cahn

[BM11, CS18, CLWW16, WZZZ19]. Calculating [AM11, FKBJ17]. calculations [CL16]. calculus [AL17b]. can [GKS17a]. canceling [YDk12]. cannot [ELSW18, Kop19b]. canonical [CLRR12, HJSZ18, Mül14, Tho10, Wel17]. capture [DZBK16]. capturing [CJLW18]. Carlo [AH14, BSSW14, HPS17, He18, KSS^{+}17, MSM14, MS12]. Carmichael [AGHS14, CLPM16, Zha11b, Zha15].

Cartan [FG17]. Cartesian [NM17]. Casas [CLO14]. case [Deb11, EGHL10, GH18, LMY12, TY19, WO10, Was14]. Castelnuovo [Has12]. Cauchy [Bur17, Pan17]. cell [BDFL17, CJLW18]. cell-centered [BDFL17]. cells [DiP18, GH15]. cellular [ZGFD14]. centered [BDFL17, CPV14, LSR19]. central [KMF17, MP12, RTwaabRS16]. central-difference [KMF17]. certain [Bia14, Büt15, DMPP19, JST14, RSS12, SZW11]. Certification [Mas18]. Chabauty [BBM17]. change [PT16]. changes [ABF^{+}10, FK18b]. Character [BKS15, BB16, BHLL19, CFvdG19]. characteristic [GY19, KZ18, Tui19, VJS14]. characterizations [SX16]. Chebyshev [AHHR15, BJM17, LWL19, NP17, SH11]. Chebyshev-type [AHHR15]. check [Azi19]. chemo [GGRBRG19]. chemo-repulsion [GGRBRG19]. Chevalley [BHLL19]. Chinese [Sut11a]. Choice [CGH^{+}16]. Choosing [RS10]. Chow [DDLR15]. Christiansen [NKK17]. Christoffel [NJZ17]. chromatic [KÖ18]. Chudnovsky [ABBR17]. CIR [Alf10]. circle [Peh11]. circles [RSTV10]. circuit [ITT12]. circular [GX17]. Circumscribed [BST11]. clamped [BySZ12]. Class [KM19, Kuc11, AK12, Akr16, AZ11, BJV18, Bia14, BMPR16, DM14, Die11, EFH19, GL15, GM18, GMS12, GSS16, Jin10, Kon14, KPY18, LOX14, Lou11, Mel16, Mil15, MJ16, ST18a, Spa13, Str14b, Sut11a, TZD15, Vab12, Źra19]. classes [Ang16, BGU19, BCCM13, GM14, GM17, PS14, SZW11, ULS12]. Classical [CKS19, AD13, ADGP13, BFS18, Kve10, LX19, LZ18, RS17b, RS19, WHV14]. classification [BP12a]. Classifying [dGO14, KMSwaAbMDS12, RR13]. cloaks [LHY15]. Cluster [ZN19]. clusters [BGGG17]. CM [RS10]. coalescing [DP15]. coamoebas [TdW15]. Coarse [KPRBT14, HSW10]. Coarse-graining [KPRBT14]. cochain [FW14]. coderivatives [LW18a]. codes [MC13, HM14]. coding [AHS17]. coding-theoretic [AHS17]. coefficient [JL12, JLZ19]. coefficients [AH15a, BHW13, BH16b, CDS10, Kve10, Lan11, LS17, MS14, NT16, Sán19, Sid11, ZY15, dDHZZ14]. cofinite [FJK11]. Cohen [COT15]. Cohomology [AGM10]. colleague [NP17]. collecting [Ang16]. collision [HY15]. Collocation [KMPW10, CT10, DKMW13, DGM^{+}16, HMT19, NJZ17, qWlGjY17, ZqCt16]. color [ABE^{+}16]. Combinatorics [SV12]. combined [ES16]. combing [GMS19]. Combining [ALL17]. comments [HP13]. Commuting [CQ14, Mat19]. compact [Dem14, HNRW18, LPRY10, LPSH11]. companion [BEG16]. companion-like [BEG16]. comparison [LMP15, SZ18]. compatible [PS16]. compensation [HLL16]. complementarity [PKLC14, ZY14]. complements [MP19, Cof14]. Complete [Bat15, KL13, BGH14, Fuk11]. completely [KL10, LA14]. completeness [Büt15]. completion [BHLL19]. complex [ASSW16, BAS16, BLS13, HZ14, HP16, JKLM17, JB19, NKK17, SD11]. complexes [Nei15]. complexity [BGH14, Hia11]. component [BCL15]. components [Boo18, LW12, LW15]. Composite [Mor19, Pet14, BK12, Gui19, IL11, WLZ17]. composites [BS15a]. compositum [DLRNS18]. compressed [BO15, BMBO13, CDTW18, BNMP18]. compressible [EGHL10, GHLM18, KK11, Wil18]. compression [Bör18]. Compressive [RS17a]. Computable [Sch17, BHS15, MSM14, TU18]. Computation [BBC14, BB16, BAFG18, Cin15, HL16, KK14, Kön17, LPRY10, Lez14, TK12, AL17a, AMR^{+}19, Bia14, BCPR14, CMSV19, DY12, Dem14, Fen18, FR15b, Fuk11, GM15, Has12, JLM13, JX14, LMS11, LRS12, LW12, LW15, Nar14, OHP14, RSS12, Sut11b, ZY15, dR11]. Computational [CS13b, Kem10, LPD13]. Computations [Hur18, KS17, NM12, Pan17, Pla16, RS14]. compute [CJ15, DRS11, Hia11, IKMF17, KPSY18, Mor13]. computer [PPTZ13]. computers [ZGFD14]. Computing [AJ14, BBM17, BF15, BGR19, Bis15, BJ11, BDD^{+}19, Bru13, Che16, CDDM18, CR15, DD10, DK15, GPR13, GV11, HKL16, HKW17, JB19, KC15, Kru16, Kuz15, Lab18, Lai16, Ler12, Lom19, MX18, McK11, MN19a, Mül14, NP12, Olv11, PS14, Pag15, PV15, Pla15, Rau16, Rob15, RY15, ST18a, Ser19, Str14b, Sut11a, Tho10, Tui19, VJS14, Wel17, XY13, PDSV15, BCCW16, DFH18, GVW16, Gra11, Har10, Har14, NN16, NP17, Ye18, Źra19, BCKM18]. concavity [CC13b]. concentration [KS19]. concentration-dependent [KS19]. concerning [Pla16]. condition [BR18, Ngu16, SS15, TV10, ZD14]. Conditional [LLS15, MMN11, LLS17a]. conditionally [BzCS12]. conditioned [Ye18]. Conditioning [KHX14, WHL19]. Conditions [CQS12, CS13c, BNDHV15, GT16, HP10, LLS17b, Lic19a, MS10, NL19, PMH18, Ste11]. conductivity [ABF^{+}10, AKLZ12]. conductor [DD19, GM15, Mil15, PM11, Sad14]. cone [PKLC14, ZY14]. cones [ZZWZ15]. confederate [NN16]. confluent [BJLP19]. Conformal [LPD13, LPRY10]. Conforming [GN14, CW17, CS14, HFC18, HZ15, Nei15]. conformity [CQS14]. congruence [AGM10]. conjecture [AKR18, AT16, BMS16, Ble12, BD13, CLO14, CH13, COT15, CR11b, GRT04, GRT15, HS12, Hur18, MP17, OHP14, StR14a, Sza11]. conjectures [Gra19b, MSV19]. conjugates [DHJ17]. conjunction [KZ18]. connected [Boo18]. conquer [CL16]. conservation [BCG16, BGP14, CZ11, CJ13, CMR10, CJLW18, Dro10, FSX19, GI16, HMS14, IPZ15, KSX17, LZ16, MS12, Sfa13, WX19a, Xu14b, YDk12]. Conservative [BCKX13, CY18]. conserving [JvSRV14]. consistency [ST18b]. consistent [BCM16, GMP14]. constant [BS16, BD13, CIL15, DMPP19, DEJ14, FJK11, Gal19b, MNZ13, SGD11, Wei19]. constants [AL17a, AH15b, BTDG13, FLM14, JB19, KC11a, KC11b, Mor11]. constrained [Bar16, Ish17, MPT18]. Constraints [CLO14, LRS12, OWZD18]. construct [FK18b]. constructed [COQ10]. Constructing [AGHS14, BM12, Kon14, LOX12, Sut12, Han15]. construction [ÁB17, CXZ15, GH13, LA14]. Constructively [BTDG13]. contact [CHR15]. contain [McN15]. context [CSV12]. continuation [BO15, LP17b]. Continued [GM13b, CVZ19, GMZ11, Jag12, LP18a, MRW10, MT19]. continuity [BTDG13]. continuous [BG11, LL17, Yao13]. continuous/discontinuous [LL17]. continuum [AH17, OS13]. contour [LFS15]. Contraction [CNZ16, MMN11]. contrast [CGH10]. Control [LZ16, APR15, Bar15b, BER17, CGN17, DVY15, GP14, CG16]. controls [BMR16, LR19b]. convection [CC14, CMZ17, FKS12, GLL12, GO16, HH13, IW13, JXR12, KRS14, Liu15, LW18b, IW13]. convection-diffusion [CMZ17, FKS12, GLL12, GO16, JXR12, KRS14, Liu15]. convection-diffusion-reaction [HH13, IW13]. Convection-dominated [IW13]. Convergence [Adc11, ALS17a, ALS17b, BCG16, BCS16, BP12b, CG16, CD15, CT10, CLWW16, CXZ16, CMR10, DKMW13, FJK^{+}15, GLL12, GHLM18, GMS12, HS13, KK11, KG18, LGY15, MS10, MO13, Mon10b, NNZ19, WNK18, XZ15, ZCS^{+}12, AL16, Arg11, BFS18, BI19, BAS16, BGGG17, BDFL17, CH19, Che12, CD10, DZ16, DLPW11, Fan12, GL15, GLW19, Jin13, KO12, KP14, Li10, Li15b, LSR19, LY17, MP12, NT16, QzSZ15, Str13, TY19, WX12, WX19b, XWZ13, Xu14a, Yao13, ZZ14, dW11]. Convergent [BP10, BJLP19, EGHL10, GG19, LS18, Mon10a]. converging [ADL11]. convex [BL12, DHYZ17, GL12, HYZ14, HLZ15, LS17, TY17]. convex-dense [BL12]. Convolution [BR18, LX19]. coordinates [NM17, RGB14]. corner [FKS12]. correct [RS10]. correction [BHM12b, BHL18, COQ10, HLL16, LX15, Pet17]. Corrigenda [BHM^{+}12a, Alk11]. Corrigendum [FK18a, GRT15, HC15, LLS17a, Ram19, Shi19]. Costabel [QS16]. counterexample [AT16]. counterexamples [CLO14]. Counting [BS15a, HÖ14, Shi86, Shi19, Tui16, Zha11b, Coh15, KtR10, Kus18, Tru15]. counts [Zha15]. coupled [GOS11]. Coupling [BCL15, ALS17a, CGH^{+}16, FHK17, GHS10, HS15, MSS11, NKK17]. Covariants [CFvdG19]. Covering [SSV17, BHLL19, MC13]. coverings [CLTZ12, FS12]. covers [RS14]. Cox [HKL16]. Crandall [Cof14]. Crank [AKPZ15, Ing13, RS16, Xu14a]. Cremona [SV12]. criteria [MRW10]. criterion [CVZ19, KC11a, Shi13]. critical [Che16, HO12, Tru12, Yao13]. cross [GLLZ14]. cross-derivatives [GLLZ14]. Crouzeix [CGS15, DL15, Wan19]. crystallographic [BK12]. crystals [GGG11]. cubature [SX14, vDE19]. cube [He18, KÖ18, Lan11]. cubes [CQ14]. cubic [BCST19a, BHM^{+}11, BHM^{+}12a, DLRNS18, DN19, Fan12, HP16, JKK^{+}10, JKL11a, Jeo16, Kim17, Lou11, Mor13, RJS12, ZQY12]. cubical [AA14]. Cui [CS13b]. Cullen [GOM11]. Cuntz [DJ12b]. curl [MRSC19, RV14, ZHX11]. curse [HNUW14]. curve [Baa10, DGP10, FG17, Fis17, FK18b, KSW19, RS10]. curves [ASSW16, Ane19, Baa10, BBG19, BBM17, BR19, BB17, BGR19, BBLP13, Bru13, BF18, Che16, CFLTL14, Cos10, CR15, CFO^{+}15, Cre14, DS18, DLRNS18, DN19, Die11, DF14, DD19, FFS^{+}13, FS12, GJLR18, JKK^{+}10, JKL11a, JKL11b, Jeo16, KK10, Kir10, KZ18, LRS16, LP17b, MS13, MN19a, MS16, Mor19, Nel15, OS19, RS14, Sad14, SS14, Shi13, Sij12, SW13, Sut12, Tho10, Tui16, ULS12, VJS14, Wut18, Zie10b, vdBLM10]. cusp [FK18b, PY15]. cut [BHL18]. CWENO [CPSV18b]. cycles [CLTZ12, HÖ14]. cyclic [DN19, Jeo16, LRS16, LM17, NP12]. Cyclotomic [Xia18, AM11, FLM14, Kar13, Mil15]. cylindrical [LHY15].

D [BKN15, HFC18, LWZ18, ABBM18, AN15, ALS17b, DHMG11, EHR18, IPZ15, KP12, Kop19b, LZ16, MRSC19, Ort11, OS13, XZ10]. DAEs [DKMW13, KMPW10]. damped [AG16]. Darboux [BCCW16]. Darcy [ALS17b, CGHW14, GOS11, GRBT16, MZ10, SW11]. Darmon [GM15]. data [Bat15, BH18, CQT11, CS13a, CM14, DS17, EFP10, FCW19, GLL12, Gon13, MS12]. Davidson [WX19b]. decay [CM12]. decimal [MNZ13]. decomposition [AH17, AZ11, BDG^{+}19, BCNS12, CGHW14, Chè13, GX17, GSV17, GKS13, GKS17a, GKS17b, GM15, HK17, HSW10, IW13, JgLW13, KPY18, LZ17, Yan17]. decompositions [CFS17, KSWW10, Rah14]. Decoupled [MZ10]. decoupling [GGG11]. Dedekind [BF15, GM16b, GM16c, Lou11, Tru15]. deferred [BHM12b, COQ10]. deficiencies [SY16]. defined [Bar15b, Ler12]. definite [BzCS12, Kir12]. deflection [dVMR19]. degenerate [BHW13, CG16, GK11, KRS14]. degree [BCPR14, CLO14, CH13, CC14, CW17, CFvdG19, CZ14b, FR15b, GJ19, GRT04, GRT15, Gui19, KM19, Lan11, OS10, RSS12, Sør16, Zie10a]. degree-3 [Lan11]. degrees [AK12, LOX14]. Dejean [CR11b]. Del [Qur19]. Delaunay [DI15]. delays [ZqCt16]. Delsarte [Hei12]. deluxe [Cal16, OWZD18]. d’enfants [KS17]. dense [BL12]. density [DFH18, FA14a, Ish17, LS18, Ram16]. dependent [CGHW14, CMQ13, DJL19, FM18, GRBT16, JvSRV14, JLZ19, KS19, LZ17, Ols12, PS16, WLZS19, ZD14]. depth [IKMF17]. derivation [NH11]. derivative [Gal19a, Kop19a, Naz10]. derivative-free [Naz10]. derivatives [ELSW18, GO16, GLLZ14, ST18a]. derived [NRV15]. descent [CFO^{+}15, LRS16, MS13, MPT18]. descents [Cre14, Fis17]. description [Bur10]. Design [FGMS12, MMN11]. designs [BSSW14, ZC18]. desingularized [NM12]. dessins [KS17]. determinant [ELSW18]. Determinantal [LP17b]. determinants [Bor10, JLM13]. Determination [Hut15, KPSY18]. Deterministic [Hit17, TCH12, ASSW16, AG14, BKS15, CH14, FFS^{+}13, Hia16b, Hit18, Zra10a]. DG [HS15]. dGFEM [SSW16]. diagonal [Els12]. diagonally [Azi19]. diagrams [CQ14, JK17]. diaphonies [Ser17]. Diffeomorphic [Sdi19]. Difference [RS19, AK16, BC13a, BS18, BJKM11, CY18, CLWW16, CG11, Fen18, GÖS18, GK11, HFC18, KRS14, KMF17, Kry13, La 15, LSR19, LLX11, LR19b, SV14, Ten10, TZD15, Vab12, WNK18, WX19a, Xia18, YDk12]. difference/spectral [LLX11]. differences [CWX16, Hit17, Kry13, LRS12, LFS15, MF11]. differencing [JLQZ18]. differentiable [KPW14, Yao13]. Differential [Chr18, AZ18, AA14, Bar16, BHM12b, BTDG13, CSW16, CR16, Deb11, FHN17, GG17, GK19, GT16, Hai10, HMT19, HLR13, HLJ16, HJW18, JLPR15, KSU16, KHOLT14, LK12, MDK13, MN19b, Mus13, NT16, RW18, RS16, Sán19, WX13, qWlGjY17, WKN13, WX19c, Xu14a, Yan17, Ye18]. differential-algebraic [AZ18, HMT19]. differentials [FG17]. Differentiation [dHAL11, HWZ14, SV14, XWZ13]. differents [Nar14]. diffusion [BNMP18, CC14, CNOS16, CMZ17, CFS17, DJ13, DJTZ13, DJL19, EFP10, EHR18, FKS12, GLL12, GO16, GSRM12, HH13, HPS17, IW13, JXR12, KRS14, KP12, KSS^{+}17, LNRW18, Liu15, LW18b, Mus18, SS15, TZD15, WHL19]. diffusion-reaction [BNMP18]. Diffusive [LMY12]. digit [KZ18]. digital [DK10]. digits [Kir10, MNZ13]. dilation [BK12]. Dilational [HA11]. dilogarithm [O’S16b]. Dimension [DiP18, Bis15, BFZ10, CQT11, CCD16, KSX17, Mou14, NV17, RPR12, Sán19, Sør16]. dimensional [AKPZ15, AJ14, Bar19, BCL15, BPT10, BCNS12, CZ16, CLAT12, CCLX17, CDTW18, CM14, CMR10, DKS15, DJL19, GM10, Gne12, Gra11, HKK13, LMS11, LMY12, LZ15, LR19b, LW18b, LS10b, NP14, Rah14, RS17a, WX19a, Xu14b]. dimensionality [HNUW14]. dimensions [Adc11, Cal16, CC12b, Gra19a, GY19, JXR12, Kop19a, Nei15, WNK18]. dimer [GOP16]. dimer-type [GOP16]. diminishing [BNS15, CG16]. Diophantine [BH16a, Nie15, Ula14]. Dirac [Chr18]. direct [GHS10, JKdR16, Liu15]. direction [Gra19b, GMS12, TY19, YY13]. Directional [AHHR15, Arg11]. Dirichlet [CFJ12, CGN17, LA14, MS10, PS19, Ram16, RSS12, Tru15]. Dirichlet-to-Neumann [MS10]. discontinuities [Kve10]. Discontinuous [Abd12, AW11, CZ16, CW10, ES16, FN11, FW11, BCS16, BCKX13, BD15, CGS18, CCQ17, CGP12, CMZ17, CCLX17, CNZ16, CD16, DJL19, DT11, FX13, FM18, FSX19, GH18, GHS10, GV15, GMP14, Gud10, GP14, He18, HMP13, HS19a, HW18, HMS14, JXR12, JvSRV14, KM15, KSX17, KG18, LL17, Liu15, LM18, MSW16, Mus13, NM17, SW10, WLZS19, Wil18, XZ10, YPP13, dDHZZ14, DE10]. discontinuously [JKdR16]. discovery [BGU19]. discrepancies [Sch17, WW10]. discrepancy [AH14, ÁB17, Jin13]. Discrete [AGS16, DE10, HK17, Nei15, ADGP13, BCL^{+}11, BKN15, CLWW16, Cin15, Die11, GPR13, GT16, GJ19, GGRBRG19, Gui19, Hak13, HJHM15, JX14, KCL14, KMF17, KP14, KW19, LP17a, Li19, RS19, Sut11b, VJS14]. discretisation [RW18]. discretization [Alf10, AdVV13, BP10, BCM16, BDFL17, DVY15, GH18, GLW19, HS19a, KO12, KZ11, LXX14, LM15, SW11, Wan10b]. discretizations [ALL17, AKPZ15, CPV14, ETX11, MS10]. discretized [CS18, GCL19]. discriminants [JW12, Nar14]. Disk [Kir10]. Disk-like [Kir10]. dislocations [GM10]. dispersive [BJV18]. displacement [BySZ12]. Disproof [HS12]. dissipative [HS13, LMMR19, Wan10b]. distance [FM12, LPD13]. Distinct [HZ14]. distributed [XZ15]. Distribution [LSW14, FLM14, JST14, Shp11]. distributions [Hoa16, SZW11, ST18b]. Divergence [CS14, Ste11, Zha11a, BCNS12, FHN17, Gal19b, GN14, SW11, WW18]. Divergence-conforming [CS14]. Divergence-free [Ste11, Zha11a, BCNS12, GN14, SW11]. divide [CL16]. divide-and-conquer [CL16]. Divided [MF11, LFS15]. Divisibility [Pap15, Mah14]. divisible [FLM14, Meš14]. divisor [BBR12]. divisors [KC15, PY14, Źra19]. Dixon [Chu12]. does [BKMW11]. Domain [BBSV19, BDG^{+}19, GSV17, AA11, AKPZ15, CGHW14, GX17, HK17, HSW10, JgLW13, LHY15, QS16, Yan17]. domains [CPBG19, GMS12, Li18, Lic19b, ZZJZ18]. dominant [Azi19]. dominated [IW13]. double [Bur10, CD15, CMS11, Rit10]. double-well [CD15, CMS11]. Doubling [GCL19]. Dougall [Chu11, CZ14a]. Douglas [DZ16]. DPG [BDS18, FHK17, FHN19, GQ14]. dream [HKW17]. driven [KHOLT14]. Dual [FJS16, Han10, Han15, WW18, YMO13]. Duality [DK10, BKN15, DJ12a, LMP15, FJS16]. dwindling [CS15a]. dyadic [Alk10, Alk11]. Dyer [BMS16, Ble12, Che16]. dynamic [BOP17]. Dynamical [HR10, GPOS14, OS10, Wan10b]. dynamics [BGN17, CGH^{+}16, EE18, GM10, JLH13, LRS12, LZ18].

easily [LLP16]. ECM [BBLP13]. edge

[BLW10, ZCS^{+}12]. Edwards [BBLP13]. effect [GKS13, GKS17b, Pet17]. Effective [FR15b, BMP10, CVZ19, FJK11, KC11b]. Efficient [BCCW16, Has12, KZ11, RVX19, AG14, KSU16, ZY14]. Eigenanalysis [ZGFD14]. Eigenanalysis-based [ZGFD14]. eigenforms [KPSY18]. eigenmode [Chr18]. eigenproblems [SX16]. eigensolvers [SX16, ZN19]. Eigenvalue [CM12, Mel18, BGGG17, BG11, CGS15, DPV19, GCL19, HC11, HC15, LX15, Nak11, Sha19, WX19b]. eigenvalues [AMR^{+}19, CG14, DP15, GV11, JST14, NP17, SL17, SX16, Ye18]. eigenvectors [AMR^{+}19]. eight [JD14]. Eisenstein [BI10, KPSY18, Lan11]. elastic [BPT10, CXZ16]. elasticity [BNDHV15, CHH18, CHR15, CGG10, DL15, QSS18]. electromagnetic [AKKL12]. electromagnetics [CGO10]. electronic [CL16]. element [Abd12, AV14, AA11, AGS16, ABHV11, AD16, AKPZ15, APR15, AA14, AL17b, BP10, BLW10, BGN17, Bar10, BCFG17, BHW13, BD15, BySZ12, Bur17, BHL18, BH18, Cao15, CR11a, CD15, CT12, CWD14, CS13a, CPBG19, CR16, CGH10, CGG10, CD16, CL10, DKS15, DGS11, DLSW12, DJTZ13, DT11, Dua14, DRS11, DE13, ETX11, EFP10, ES16, EGHL10, FN11, FHN17, FKS12, GM14, GM17, GHS10, GOS11, GRBT16, Gav11, GK19, Gon13, Gud10, GZZ17, GL12, GSS16, HFC18, Hir13, HW18, HZ15, HX19, JLTZ17, KHX14, KS19, Kre12, KW19, LHY15, LS17, Li18, Li19, LMNN18, MS10, NM17, OS16a, SL17, Voh10, WX13, WY14, WW18, WHL19, Wan19, WZZZ19, WX19c, XZ15, ZHX11, ZCS^{+}12, ZD14, CG16]. element-finite [EGHL10]. Elementary [GM15, BJLP19]. elements [BCST19a, BG11, CHZ17, CH13, CW17, CQ14, GN14, GO18, GS19, HZ17, Kop14, Kop19b, LP17a, LPSH11, Pet14, RGB14, Sty14, Zha11a, dVMR19]. Eliminating [Pet17]. elimination [GJ19]. Elkies [KSW19]. ellipsoid [BST11]. Elliptic [Ane19, CFLTL14, ASSW16, Abd12, AV14, AN15, AdVV13, Baa10, BC17, BBB16, BGR19, BHW13, BP15, Bur17, CCDE19, CGS18, CZ10, Che16, CNOS16, CGH10, CQS12, CD16, CFO^{+}15, Cre14, DS18, DLRNS18, DN19, DGP10, FFS^{+}13, FHN17, FS12, Fis17, Fuk11, Fuk12, FK18b, Git13, GJLR18, Gud10, GP14, GK11, HPS17, Hei12, HW18, HSW10, JKL11a, JKL11b, Jeo16, JZ11, KZ11, Kry13, Li15b, LFS15, MP14, Mah14, MS13, MS16, Mor19, Mül10, OS16a, Pet14, DD17, RS10, RS14, Sad14, SSW16, SW10, SW13, Sut12, Tho10, ULS12, WY14, WW18, WKN13, Wut18, YPP13, dDHZZ14]. elliptical [LHY15]. embedding [LOX14, RPR12]. embeddings [BDD^{+}19, SX14]. Empirical [OHP14, FCW19, Gau17]. employment [BKM18]. endomorphism [Bis15, CMSV19, FK18b, Lom19]. endpoint [Sid11, Sid12]. energies [EE18]. Energy [GMP14, JLQZ18, LSR19, NL19, ALL17, COS14, CGN17, DGS11, GS10, GGRBRG19, HY15, Ish17, LMP15, LRS12, LGY15, OS13, PPTZ13]. energy-based [OS13]. enforcing [CGG10]. Engquist [CMR10]. enhancement [JXR12]. enriched [ABHV11]. entanglement [Mor19]. Entropic [CJLW18]. Entropy [CNPT10, ABBSM16, BGM15, BC16, BGP14, CG16, GPOS14, Wil18, vzG19]. Entropy-satisfying [CNPT10, BC16]. entrywise [XY13]. Enumeration [BvH11, EW16, KÖP15, VY19, Bur10, HS19b]. epitaxial [JLQZ18, QzSZ15]. epsilon [BD13]. equal [BD10b, RPR12]. equation [AH17, ABE^{+}16, ALS17a, AG16, AL16, AKPZ15, BR18, BC13a, BCLZ14, BJKM11, BCKX13, BC14, BGNyS11, BDFL17, BMR16, CCOV14, CC12a, CLAT12, CLWW16, CH19, CQB14, CB16, FN11, FW11, FX13, GMZ11, GLL12, GZ13, GM10, GV17, Hak13, HH13, HFC18, HKRT11, HL17, HP12, KM15, KP14, LLX11, LR19b, LY17, MRW10, MS10, NNZ19, PZ17, WNK18, WZZZ19, Xu14a, YDk12, ZZJZ18]. Equations [Baa10, AN15, Akr13, AK16, Akr16, ALL17, AZ18, AD16, AHHR15, BP10, BM11, Bar16, BFS18, BLT13, BC16, BH16a, BD10a, BCCM13, BFZ10, BDG^{+}19, BGP14, BNDHV15, BCL15, BCNS12, BDS18, BHM12b, BTDG13, CZ12, CG16, CZ16, CCQ17, CMTZ16, CCL^{+}19, CT10, CQT11, CC14, CSW16, CCD16, CGP12, CMZ17, CCLX17, CS13a, Chr18, CW10, CDDM18, CMQ13, CMS11, DM14, DJ13, Deb11, DEMS19, DL13, DE10, DF14, DHMG11, DGKS16, EFH19, EHR18, Fan12, Fan14, FHN17, Fen18, GHLM18, GLS11, GRBT16, GLM^{+}19a, GG17, GT16, Gon13, GL15, GG19, GMS12, HP10, HMT19, HS13, HMP13, HR10, HS19a, HLR13, HKK13, HLL16, HLJ16, HJW18, IW13, Ing13, ITT12, JXR12, JZ11, JLTZ17, KHX14, KRS14, KSU16, Kim17, KHOLT14, KW19, Kry13, LS18]. equations [LMY12, LZ15, LK12, Li15a, LLS17b, LS17, LZ17, Li19, Liu15, LW18b, LWZ18, LCQ17, LM15, MSW16, Mon10b, MR10, Mus13, Mus18, NT16, Nie15, NL19, OMS15, DD17, PS16, QS16, RW18, RS17a, RVX19, RS16, Rit10, SS15, SS16, Seg17, SY16, Ste11, TU18, Ten10, TZD15, Vab14, WX13, qWlGjY17, WW18, WKN13, Wen13, Wil18, Wu10b, WX19c, YPP13, Yan17, ZHX11, ZqCt16, ZCS^{+}12, Zie10a, Hai10]. equiangular [GY19]. equidistribution [CS13b, HK17]. equivalence [BBB16]. Equivariant [BD11, Ble12]. Erdos [GMZ11, PY14]. Ergodic [Jor12]. Errata [Coh12, GM17, LN14, BR05, BR11]. Erratum [LW15, Szm13]. Error [AADL11, AG16, AD13, AD16, Bar15b, BI19, CS18, CGN17, Gon13, HMP13, HL17, JKLM17, Kop19a, LRT13, LW18b, Spa13, Ten10, YDk12, AW11, AC18, APR15, BC13a, BS18, BM11, BER17, BFZ10, BJ15, CHZ17, CY18, CPV14, CCS12, CMTZ16, CT12, CGS10, CMQ13, CQS14, CZ14b, DM10, DL13, DVY15, DPV19, DJTZ13, FA14a, Gud10, GP14, GSRM12, GL12, HLL16, JvSRV14, JLQZ18, KM15, KRS14, Kop19b, KZ11, LL17, LS10a, Li18, LP18b, LXX14, Liu15, LWL19, MSW16, MR10, Mus18, Not15, RS16, Sty14, TU18, Tru15, UP14, Voh10]. errors [BKM18, GG17, JR18, LR19a, Rah14]. Escott [Cal13]. essential [GMRL18]. essentially [XY13]. estimate [CHZ17, CPV14, CZ14b, HL17, Ram16]. Estimates [Dus16, Sán19, AADL11, ALL17, AG16, AC18, AD13, AD16, APR15, AFF^{+}17, BC13a, BM11, BI19, BFZ10, CS18, CY18, CMTZ16, CT12, Che12, CMQ13, DM10, DGS11, DE13, Gon13, GSRM12, GL12, HJHM15, JvSRV14, JLQZ18, KM15, KRS14, Kop19b, KZ11, Kry13, LL15, LMP15, LR19a, LL17, LS10a, LRT13, Liu15, LW18b, LWL19, MSW16, MNPW10, NRV12, Ram15, Ram19, Sty14, TU18, WO10, WKN13, Wei19, ZN19, ZD14]. Estimating [Büt16, HMT19, Zha15]. estimation [AW11, MR10]. Euclidean [Cer11, GM13b, Lez14, LM17, Nak11]. Euler [Bay11, BGM15, CNPT10, FLM14, HS13, Meš14, NRV12, NT16, Pro17, Sid12, Źra19]. Eulerian [FSX19, KSX17]. Evaluating [BL14, Mez11, Nel15]. evaluation [BS16, BKM18, Bor10, Doo15, Dup11, Esp10, FA14a, KM16, LP17c, NS18]. evaluations [GLLZ14]. even [NP14, OHP14, VJS14]. event [LM18]. Eventual [DZ16]. Every [CLB18, Tao14, GKS17a]. evidence [Ble12, HO12, MSV19]. evolution [CMS11, ET10, HS13, KP14, KW19]. evolutionary [Hai10, Vab14]. evolving [DEMS19, DE13, LM15]. Exact [Coh15, vDE19]. exactly [TU18]. exactness [dVG16]. examples [AKR18]. exceed [CS10]. exist [SR11]. Existence [BCST19a, FMP16, HM14, Mül10]. exotic [XW10]. expansion [FA14b, IL11, NRV15]. expansions [Bay11, BJM17, BJLP19, Fuk12, KZ18, LS10b, ND19, RGNS18, Sid11, Sid12, WLZ17]. experimental [NS13, RR13]. experiments [CHR15]. Explicit [BBR12, BGG^{+}17, CFO^{+}15, GM16a, GM16b, GM18, Lep16, LRS16, MS13, NM12, Ram15, RS17b, WHV14, AH15b, Akr13, Akr16, Alk10, Alk11, AZ18, BGP14, DZBK16, GGH15, HS13, Kön17, Ram16, SZ18, WLZS19, Ram19]. Exploring [FH16]. exponent [AK12, AJ14, WW10]. Exponential [DLPW11, HJW18, BFS18, BH16a, HH13, JLQZ18, XWZ13, ZZ14]. exponential-type [BFS18]. exponentials [KT10, TK12, XY13]. extended [BBC14, Chu12, DPR14]. Extension [HJHM15, AGS16, CM14, DGS12, DL15, FJS16, Gui19, HZ13, Kle13, LZ13, NP12, Źra19]. extensions [Ane19, BCST19a, CDFS13, CQS14, DS14, GRT04, GRT15, LPD13, MS14, PR15, RSTV10]. exterior [AL17b]. external [BJKM11]. extra [Chu12]. extrapolated [Ing13]. extrapolation [Sid10]. extrapolative [KT18]. extrema [WHV14]. extremal [HM14, NP14]. Extreme [MFRV18, SX16]. extremely [Ye18].

F [Coh12, Szm13]. factor

[BD11, CSV12, DEJ14, LN11, LN14, Wei19].

factored [Cha18, LLP16]. factoring [IKRS12, Zra10b]. Factorization [Cos10, BOP17, CH14, Hia16b, Hit17, Hit18, Wel17, Źra19, dHAL11]. factorizations [BL12]. Factors [BR98, BR05, BR11, BKK19, OR14]. Faltings [GMRL18]. Families [JKL11a, JKL11b, Jeo16, GRT04, GRT15]. family [Kle13, LK12, SVVR14, Zie10a]. Fast [AL17a, AMR^{+}19, BFZ16, BCPR14, CHH18, CT19, DY12, Dup11, Gav11, HMS11, JX14, Kus18, LFS15, MZ16, Pan17, SWX19, TWO18, XW10, Azi19, BO15, CPPT10, DOZ13, Fuk11, HY15, JLH13, Joh15, LMS11, PPTZ13, YMO13]. Faster [CH14, Gui19, Har11, HvdH19, Hit18]. FD [MDK13]. FD-method [MDK13]. Fekete [BCL^{+}11]. FEM [BGGG17, COT17, CGS15, MSS11, Mus18]. Feng [DFGSL13]. Fermat [BR05, BD10a, BR98, BR11, CD17, CT19, DF14, LP18b]. Fermat-type [BD10a, DF14]. fewest [LZ18]. FFT [HvdH19, SH11]. Fictitious [BBSV19]. Field [Bar15a, BBT15, BCST19a, BCST19b, CS18, COS14, CCL^{+}19, DGP10, Fon11, Kru16, PV15, Sch10, Tre12]. Fields [Sha19, AK12, Ane19, ABBR17, Bia14, BCCW16, BKS15, BC15, BCNS12, BS15b, BDD^{+}19, Bru13, Cer11, CIL15, CM19, Cha18, Coh15, DLRNS18, DN19, DS14, FLM14, GP17, GJLR18, Gui19, GM13b, HZ14, HZ13, JKL11a, JKL11b, Jeo16, JW12, KKP10, KK10, KM19, Kuc11, Lez14, LM17, LOX12, Lou11, Mil15, Mor13, Mor19, MJ16, OWZD18, PR15, RGNS18, Rob15, RJS12, Sdi19, Sut12, SwaAbLCW11, Tho10, VJS14, Xia18]. fifteen [KKP10]. figure [MSM14]. filter [CS15a, HZ13, Han15]. filtered [BI19]. filtering [GÖS18]. filters [BK12]. filtration [LR19b]. find [TCH12]. Finding [HHT19, BLS13, BKS15, CZ10, Kus18, Yao13]. Finite [AA11, APR15, AA14, AL17b, BGN17, BJKM11, Boo18, BySZ12, CG16, CWX16, EGHL12, FHN17, Hir13, KS19, LLX11, SV14, Abd12, AV14, ALS17a, AGS16, ABHV11, Ane19, AD16, AKPZ15, ABBR17, BCST19a, BP10, BC13a, BS18, Bar10, BCFG17, BCG16, BJM17, BHW13, BG11, BD15, BCST19b, BKS15, BKN15, BPT10, BDD^{+}19, BCPR14, Bru13, Bur17, BHL18, BH18, CY18, CPV14, Cao15, CR11a, CD15, CT12, CWD14, CXZ15, CLWW16, CW17, CSO13, CS13a, CPBG19, CR16, CGH10, CD16, CL10, CG11, DKS15, DGS11, DLSW12, DS14, DJTZ13, DT11, Dua14, DRS11, DE13, ETX11, EFP10, ES16, EGHL10, FN11, FKS12, GLL12, GM14, GM17, GOS11, GRBT16, GK19, Gon13, Gud10, Gui19, GZZ17, GL12, GSS16, GO18, GS19, GK11, HHT19, HH13, HZ17, HFC18, HW18, HZ15]. finite [HX19, HL16, JdRV14, JJK^{+}15, JLTZ17, KHX14, Kop14, Kop19b, Kre12, KW19, Kry13, LP17a, LSXZ12, LHY15, LS17, Li18, Li19, LSR19, LOX12, LMNN18, LR19b, MS10, MS12, NM17, NPPY12, OS16a, Pet14, PR15, RGB14, RGNS18, Sut11b, Sut12, SwaAbLCW11, VJS14, Voh10, WX13, WY14, WNK18, WW18, WHL19, Wan19, WX19a, WZZZ19, WX19c, Xia18, XZ15, Zha11a, ZHX11, ZCS^{+}12, ZD14]. finite-difference [Kry13, LR19b]. finite-precision [BCPR14]. First [CS13a, BCCW16, CPPT10, ES16, JD14, KMPW10, Kop14, PT16, RW18, Sha19]. first-order [ES16, Kop14]. Fischer [PKLC14]. Fisher’s [Hak13]. five [Tao14]. fixed [BKK19, BST11]. fixed-point [BST11]. Floater [Kle13]. floating [JKLM17, JR18]. floating-point [JKLM17, JR18]. flow [ALS17b, CPV14, CGN^{+}11, CS13c, DKS15, DL15, MO13, WLZS19]. flows [BH18, CS18, CS14, KS19, LPSH11, LSR19, MZ16]. fluid [BP10, HKK13, MO13, WLZS19]. fluidic [BGN17]. fluids [KS19, LMMR19]. Flux [CZ11, BCG16, WX19a, Xu14b, XZ15]. flux-splitting [BCG16]. fluxes [MSW16]. FMA [JKLM17]. folded [NM17]. foliations [CJ15]. following [GKL^{+}17, vdBLM10]. form [BGGG17, CHH18, FHN17, Gal19a, Kir12, KC11a, KZ18, PZ11, WW18]. forming [CMS11]. forms [AA14, CR16, Chr18, CFvdG19, Dem14, Dum17, FK18b, GP17, GK19, GV11, JST14, Kir12, Lep16, LOX14, Mez11, Nel15, PY15, Rau16, RJS12, RSS12, RTwaabRS16, ZY15]. formula [BSO12, BMPR16, KC11b, KtR10, dR11]. formulae [AK16, BC13b, Chu11, Chu12, Not15, RS17b, Spa13, Tru15]. Formulas [Coh12, MOS66, RTwaabRS16, Hia16a, LP17c, MS14, MS16, Peh11, Spa17, Szm13]. formulation [CZ12, CC12b, EFP10, FHN19, GGG11, JLPR15, Voh10, dVMR19]. formulation-based [Voh10]. formulations [BDS18]. forth [HLL16]. four [KT10, MNZ13]. Fourier [Adc11, AH15a, BY12, Bat15, Bay11, BH16b, CNV14, DY12, DJ12a, FA14b, JX14, Kve10, Lan11, MS14]. fourth [EFH19, GV15, MR10, ZHX11]. fourth-order [MR10]. Fowler [BC14]. fractal [Dro10, DJ12a, Jor12]. fraction [LP18a, MRW10]. fractional [ABBM18, BP15, CSW16, CNOS16, Chr18, CJ13, EHR18, GLM19b, GLW19, JLPR15, Kop19a, LLX11, LWL19, MN19b, Mus18, Sha19, TZD15, ZZJZ18]. fractional-derivative [Kop19a]. fractions [CVZ19, GMZ11, GM13b, Jag12, MT19]. frame [ZDL13]. framelet [Han15]. framelets [HJSZ18]. framework [ASSW16, FM18, GVW16, KT18]. Fredholm [Bor10, DM14]. free [BR19, Bar16, BHSP11, BJ11, BCNS12, GN14, LRS12, Naz10, SW11, Ste11, Yen11, Zha11a, Ste11]. Free-slip [Ste11]. Freeden [CS13b]. freeness [MP17]. frequency [BDG^{+}19, GSV17]. Freud [Gau17]. Frey [BB17]. Friedrichs [DT11, Sog15, Sog16]. friendly [ULS12]. Frobenius [RPR12]. front [BCS16, CJLW18, Mon10b]. Frozen [LZ18, Jin10]. Fuchsian [FJK11]. Full [GLM19b, CFLTL14]. fully [BC16, BGNyS11, CS18, CLWW16, DJ13, GOS11, GGRBRG19, KP14, KW19]. fully-mixed [GOS11]. Function [BzCS12, LSW14, Alk10, Alk11, BSO12, BF15, BT15, BB12, CFJ12, CWX16, CM14, CC13b, Cof14, CW15, Doo15, FA14a, FLM14, GKS17a, Hia11, HO12, Hur18, Ish17, KK10, KM16, Kuz15, Kve10, LP18a, LM13, NJZ17, NRV15, ND19, PKLC14, PY14, Pro17, Ram15, RJS12, RY15, Tru12, dR11, Bar15a, Ram19]. functional [DE10, LM13, NRV15, Yao13]. functionals [dHAL11]. Functions [FG17, Szm13, AH15a, ADL11, Bar15b, BY12, Bat15, BJM17, BKM18, BMPR16, BS15b, BL14, BF18, BJLP19, Büt15, Büt16, CSW16, Chè13, CDTW18, Coh12, DT11, Dup11, Fre18, GM16b, GM16c, HNUW14, KL10, KL13, KPW14, KSWW10, KSW17, Lan11, LP17a, LWL19, LFS15, Lou11, LA14, MOS66, Mez11, MNPW10, MN19b, MF11, NH11, Not15, Ram15, Ram19, RTwaabRS16, ST18a, Sid11, SZW11, Tru15, Tui19, Wan10a, Was14, YF15, vDE19, CM13, Rob15, Gra11]. Fundamental [BH13]. Fundamentality [BHM^{+}11, BHM^{+}12a]. Further [JLM13, Jin13].

Galerkin [Abd12, AW11, AD13, AD16, BCS16, BCKX13, CGS18, CNV14, CZ16, CCQ17, CGP12, CMZ17, CCLX17, CW10, CNZ16, CD16, DE10, DJL19, DT11, ES16, FW11, FX13, FKS12, FM18, FSX19, GH18, GMSC^{+}17, GHS10, GV15, GMP14, Git13, GP14, GSRM12, HMP13, HS19a, HW18, HMS14, JXR12, JvSRV14, KM15, KSU16, KSX17, KG18, KW19, LL17, LNRW18, Liu15, MSW16, Mus13, NRW17, PS16, RS17a, SW10, TU18, WY14, WW18, WLZS19, WZZZ19, Wil18, XZ10, YPP13, dDHZZ14]. Galerkin-finite [AD16]. Galois [CFLTL14, Fen18, JW12, Kön17, LRS16, Mas18, Mor19]. gamma [ND19]. gaps [OHP14]. Garsia [HP13]. gas [CGH^{+}16]. Gauss [BS16, CMZ17, Doo15, JR13, Jin10, Jin13, Not15, PS16, WLZ17]. Gauss-compatible [PS16]. Gaussian [BI10, BD11, Cal13, KSW17, LLP16, LRT13, LZ18, Mon10a, Spa13, Spa17, ST18b]. Gegenbauer [XWZ13]. genera [KN12]. General [CFS17, BDFL17, DL15, EGHL12, Fuk11, GT16, GN14, Jin13, LLS17b, Ngu16, DD17, WHL19]. Generalised [GM13a]. Generalization [YF15, NH11]. Generalized [AH15a, CSW16, AKLZ12, AKKL12, BR98, BR05, BR11, BCKX13, Cer11, CIL15, CMZ17, CD17, CT19, DFGSL13, DK10, Gal17, GI16, GOM11, KM15, KCL14, KS19, Mel18, Nak11, PKLC14, RGB14, Sid10, Spa17]. Generalizing [RSTV10]. generated [BS15b, DFGSL13, HZ14, LS10c, Shp11]. generates [MC12]. Generating [Cha18, LLP16, CFJ12]. generation [GLM19b, HK17]. Generators [CM13, KK10, BJ11, GPOS14, GM18, HMS11, JJK^{+}15, OS10]. generic [Tui19]. Genocchi [Bay11]. genus [Cos10, CR15, DD19, Lom19, Rau16, VJS14, Zie10b]. genus-2 [Lom19]. Geometric [BCL^{+}11, BMP10, Lom19, McN15]. geometrically [Bar16, GO18]. Gerschgorin [Nak11]. giantstep [Hit18]. Gibbs [ADL11, BKMW11]. Ginzburg [LZ17]. given [BMPR16, Coh15, DGP10, FLM14, FK18b]. Glimm [CJLW18]. Global [vdBLM10, DS17, Fon11]. GNFS [BBKZ16]. Godunov [BC16]. Godunov-type [BC16]. Goldbach [OHP14]. Good [Sør16, AKR18]. Gordon [BS18, BFS18, BJKM11]. governed [ET10]. Graded [AN15, DLSW12, Kop19a, Li17]. Gradient [CD18, Cao15, DL13, LPSH11, LSR19, ZN19].

Gradient-based [CD18]. gradient-type [ZN19]. gradings [La 15]. graining [KPRBT14]. Gram [BB12]. Gramain [MNZ13]. Gramian [FJS16]. graph [Boo18, Cin15]. graphs [DKS15, HÖ14, MC13, MMN11]. Grassmannian [LMT10]. greater [OR12, Tao14]. Greedy [MC12, MZ16, GS10]. green [Gra19a]. GRH [GM16a, GM16b, GM16c, IKRS12, KSW19, Pla16]. grid [HC11, HC15]. grids [BGM15, DGS11, HZ15, JX14, WX19c, Zha11a]. Gröbner [BD11, CM19, GVW16, La 15]. Gross [BC13a]. ground [PPTZ13]. group [Bia14, BJ11, BJK11, BHLL19, Fen18, GM18, JdRV14, JJK^{+}15, LRS16, MJ16]. groups [AK12, BJK11, Bru13, Cor10, DS18, DMPP19, DFH18, Die11, FJK11, HL16, JdRV14, JJK^{+}15, JKdR16, JW12, JST14, Kön17, LPSH11, MP17, MNU15, NV17, Pag15, Pap15, RR13, ST18a, Sut11b, dGO14]. growth [CMS11, HP10, JLQZ18, OS10, QzSZ15]. Guaranteed [CG14].

Hadamard [AHS17]. Half [LLS17b]. Half-space [LLS17b]. Hall [AKR18]. halting [FK18b]. Halton [HPS17]. Hamilton [BFZ10, CMQ13, DEMS19]. Hamiltonian [BFZ16, COS14, HÖ14, MO13]. Hankel [Fre18, TWO18]. harmonic [Bar10, BDG^{+}19, BG11, CS10, HMP13, LP17a, LCQ17, LGY15, ZCS^{+}12]. Harnack [LP17a]. hashing [FFS^{+}13]. Hasse [BH16a]. Hasse-type [BH16a]. having [FR15a]. HDG [CCS12, CC12a, CC14, CCD16, CGS10, CGN^{+}11, CC12b, CQS12, CS13c, CS14, CQS14, CQB14, CFS17, GLNP15, LCQ17, QSS18]. heat [AL16, BDFL17, CC12a]. Hecke [GP17, GV11]. Heegner [BÇS15, DDLR15]. height [DK15, FS12, GMRL18, Kru16, Tho10]. heights [BÇS15, Mül14, Wel17]. Hele [CLWW16]. Hellegouarch [BB17]. Helmholtz [BCLZ14, BCFG17, FW11, FX13, GSV17, MS10, Pet17]. Hermite [BKMW11, CIL15, FGMS12]. Hermitian [DP15, Gre15, KKP10, SX16]. Hessian [GZZ17]. Heston [Alf10]. heterogeneous [Abd12, AV14, BCFG17]. Heuristic [LM13]. Heuristics [Gra19b]. hierarchical [BBB16, KZ11].

High [Alf10, dR11, ABBM18, BCCM13, BDG^{+}19, CDTW18, CR16, COQ10, CGH10, CM14, DS14, GGH15, GSV17, HJSZ18, LMS11, LFS15, LCQ17, RS17a, SW11, Sør16, VJS14, WX19a, Wen13, Xu14b, XY13, CCDE19, DD17]. high-contrast [CGH10]. high-dimensional [CDTW18, RS17a]. high-frequency [BDG^{+}19, GSV17]. high-genus [VJS14]. high-order [BCCM13, SW11, Wen13, CCDE19, DD17]. Higher [Fis17, CQT11, CXZ15, CPBG19, FL18, GM14, GM17, Gra11, GM15, HMT19, KN12, MSV19]. higher-dimensional [Gra11]. higher-index [HMT19]. higher-order [CXZ15, CPBG19, FL18]. highest [CH13]. highly [CMTZ16, IL11, MX18, NRW17, XW10]. highly-oscillatory [CMTZ16]. Hilbert [CC13a, GP17, Gne12, GV11, Olv11, Sut11a, Tab17, Yao13]. Hill [CD10]. Hilliard [CLWW16, CS18, WZZZ19]. Hirota’s [BHH^{+}12]. Hodge [BCNS12, LW18a, LZ17]. hodograph [JKK^{+}10]. homogeneous [CCL^{+}19]. homogenization [AV14]. homotopies [HSW11]. Homotopy [RR13, BCKM18, DYY14, LP17b]. hopping [LZ18]. Hormann [Kle13]. Huber [FJK11]. Hurwitz [Kön17, Tab17]. hybrid [DYY14, ITT12, PVV17, CCDE19, DD17]. hybrid-mixed [PVV17]. Hybridizable [CD16, CCQ17, CNZ16, GH18, Wil18]. hydrodynamic [JLH13]. hydrostatic [ABBSM16]. hyperbolic [AW11, BCG16, BLT13, BCL15, BDFL17, CG11, DHMG11, GI16, JKdR16, JgLW13, LP17b, MSW16, MS12, Sfa13, Xu14b]. hypercircles [Tab17]. hypercube [Ste11]. hyperelliptic [BBM17, FFS^{+}13, LRS16, VJS14]. hypergeometric [BJLP19, Doo15, SZW11]. hypersurfaces [KT18, Ser19, Tui19]. hypocoercivity [PZ17]. hypothesis [Dus16].

IBVPs [CNPT10]. ideal

[AK12, AO14, Bia14, GM16b, PS14]. ideals [Cha18, GM16a, La 15, MN14]. identification [CZ11, JL12]. identities [RS19]. if [Azi19]. Igusa [BL14, Str14b]. II [Ble12, HM14, AA11, AKKL12, BJ11, CC14, EGHL10, FM12, LPD13, Spa17]. III [AGM10, CFO^{+}15, DGS12, FR15a]. ill [CSS15, Ye18]. ill-conditioned [Ye18]. ill-posed [CSS15]. image [HYZ14, ZDL13]. images [Mor19]. imaginary [AK12, BC15, CIL15, DGP10, DHJ17, KKP10, Kuc11, MJ16]. imaging [AKLZ12, AKKL12]. IMEX [DHMG11]. immersed [Li15b]. immiscible [CPV14]. impact [SD11]. impedance [BR18]. Implicit [Akr13, BEG16, Akr16, BHM12b, CCOV14, HS13, HS19a, MF11, WLZS19]. implicit-explicit [HS13, WLZS19]. Improved [CMTZ16, LP18b, StR14a, AGHS14, CSO13, Hur18, SGD11, Tru12, Tru15, UP14]. Improvements [Tru11]. inclusion [ABF^{+}10, AKLZ12, AKKL12]. incompatible [CQT11]. incomplete [ND19]. incompressible [BP10, BH18, CS18, CPV14, CCQ17, DE10, DL15, HKK13, LS18, NL19, RVX19, WLZS19]. increasing [JvSRV14]. independent [BS15b]. index [BJV18, DKMW13, HMT19, ITT12, KS19, KMPW10]. Indifferentiable [FFS^{+}13]. individual [Gui19]. Inequalities [ADGR12, HA11, KMF17, RS19, Sdi19]. inequality [ABBSM16, DT11, LP17a]. inert [Tre12]. inertial [CJ15]. inexact [BKM18, KM16, KM16]. Inf [GO18, BTDG13, Gal19b, BKN15]. Inf-sup [GO18, BKN15]. Infinite [Gne12, BMM18, CZ14a, NH11, SVVR14]. Infinite-dimensional [Gne12]. infinitely [GKS17a, KPW14]. infrastructure [Fon11]. Inhomogeneous [HLZ15]. initial [CQT11, CG11, MS12, Spi17]. initial-boundary [CG11]. inner [Dem14]. inputs [KCL14, LP18b]. instability [HMT19]. instationary [Ste11]. integer [CDS10, Chu12, CLB18, CH14, CT19, Gre15, HvdH19, Hia16b, Hit18, Riv16, Źra19]. integers [Cal13, LLP16, McK11, McN15, PZ11, Sch10, ZC13]. integrability [HJW18, SZ18]. Integral [COQ10, ALS17b, AFF^{+}17, BBM17, BSO12, BJK11, CM12, CT10, DM14, DY12, FA14a, Fuk11, GL15, JdRV14, JJK^{+}15, Kuz15, LP17c, OMS15, QS16, ZqCt16]. integrals [BBC10, BCCW16, DDLR15, FA14b, Fre18, Fuk12, IL11, Lai16, MX18, Sid12, WLZ17, XW10]. integrands [He18]. integrate [BBD^{+}11]. integration [BSSW14, DLPW11, Gne12, GKS13, GKS17b, HNUW14, HS19a, HWZ14, JB19, KPW14, KT18, KSW17, LFS15, XW10]. integrator [MMV17]. integrators [BFS18, BFZ16, COQ10, GLM^{+}19a]. integro [Mus13, RS16, qWlGjY17, Xu14a]. integro-differential [Mus13, RS16, qWlGjY17, Xu14a]. interactions [JLH13, KPRBT14]. interface [ABF^{+}10, CHZ17, CGS18, CGH10, CGH^{+}16, GSS16, Li15b, Pet14]. interfaces [MSS11]. interfacial [ALS17b]. Interior [Kry13, Sid11, HS15]. interpolants [Kle13]. interpolation [AADL11, BNS15, CSO13, FGMS12, Hoa16, JKK^{+}10, Li18, RS17b, Sty14, SH11, WHV14, XWZ13, dVG16]. interpolations [ByTC13]. interpolatory [HA11, LS10c]. intersection [BGH14, KS15, Mül14]. interval [FR15b, GPR13, Gav10a, KC15]. intervals [DFGSL13, NH11]. intrinsic [BGH14]. Introduction [Gav10a, Naz10, Kem10]. invariant [AJ14]. invariants [BOP17, BH13, Cin15, HL16, Kon14]. Inverse [HJHM15, AFF^{+}17, BCLZ14, BMBO13, CSS15, DFR12, HNRW18, Jin10, Mat19, MZ16, Olv11, WKN13, ZZWZ15]. inverses [FMP16, LWCI13, Ye18]. inversion [BC13b]. investigation [NS13]. inviscid [ALS17a]. involving [CZ14a, JLPR15, Ram15, Ram19, Ula14]. irrational [NS13]. Irreducibility [Shi13]. irreducible [FM12, LOX12]. Irregular [BH11, HHO17, Cal16, NT16]. isentropic [EGHL10]. isogenies [CR15, MS16]. isogeny [BLS12, MS13, ULS12]. isogeometric [DGM^{+}16, GMSC^{+}17]. Isolating [Pla17]. isometric [SX14]. isomorphism [VY19]. isomorphisms [BDD^{+}19]. issues [DHMG11]. iterated [DDLR15, RS19]. Iteration [vzG19, DZ16, RVX19]. iterations [Dup11]. iterative [CGHW14, LGY15, MRSC19, Yan17]. iteratively [Jin13]. IV [BCL15]. Iwasawa [SW13].

Jacobi [BFZ10, ADGR12, ByTC13, CT10, CSW16, CMQ13, CC13b, DEMS19, LS10a, MN19a, Rau16, SWX19, qWlGjY17, WX19b]. Jacobi-weighted [LS10a]. Jacobian [CMSV19, Lom19, VJS14].

Jacobians [CR15]. Jacobi’s [Lab18]. Jacobsthal [CW15, HS12]. Jaeger [Fre18]. Janko [BJK11]. Jin [CJLW18]. jump [dDHZZ14]. jumps [GLM19b]. justification [LZ18].

Kaczmarz [LW16]. Kahan [JLM13]. KAM [BFZ16]. Katsurada [KPSY18]. Kazhdan [GH15]. KdV [HKRT11, ZZJZ18]. Keller [LWZ18, HL17, LY17]. kernel [CT10, HNRW18, LNRW18, NRW17, RW18, ZX16]. kernel-based [RW18]. kernels [BzCS12, KSW17, Mus13, qWlGjY17]. kind [Fuk11, KMPW10]. Kinetic [ABBSM16, BGM15, CH19, LLS17b]. Kirchhoff [FHN19]. Klein [BS18, BFS18, BJKM11]. Kleinian [Pag15]. Kloosterman [AG14]. Knapsack [MNU15]. Koblitz [KZ18]. Kolmogorov [BC17, PZ17]. Korobov [DLPW11, KPW14]. Korteweg [BCKX13, GMP14, KM15]. KPP [Hak13]. KPP-theory [Hak13]. Kronecker [FLM14]. Kronrod [MS14]. Krylov [BHM12b, MN19b]. Kummer [NM12]. Kurepa’s [AT16]. Kutta [AZ18, BGG^{+}17, COS14, COQ10, ET10, GGH15, HS11].

L. [Sza11]. L1 [Kop19a]. lacunary [Ler12, Mel16]. Lagrange [Gav10b, BG11, DHYZ17, MN10, TU18]. Lagrange-type [MN10]. Lagrangian [DJ13, FSX19, HYZ14, KSX17, YY13]. Laguerre [ByTC13]. laminated [DRS11]. Lanczos [Li10]. Landau [FJK^{+}15, LZ17, Mor11]. Langevin [LRS12]. Laplace [BP12b, LW18a, Li15a, MMN11]. Laplace-type [Li15a]. Laplacian [ABBM18, Cin15, Sha19]. Large [IPZ15, Bör18, CNPT10, DGKS16, Dus16, FW11, FX13, KM19, SX16]. large-scale [DGKS16, SX16]. Large-time [IPZ15]. larger [Tui19]. largest [LL15, LN11, LN14]. Late [BLT13]. Late-time [BLT13]. Late-time/stiff-relaxation [BLT13]. Latin [HKÖ11]. Lattice [SS15, CSV12, CMSC10, HMS11, KPRBT14, OS13, Sør16]. Lattices [BEF16, BC15, HHT19, KKP10, NP14]. Lavrentiev [PMH18]. law [Hir13, KS19, LWCI13]. laws [BCG16, BCL15, CZ11, CJ13, CMR10, CJLW18, CPSV18b, Dro10, FSX19, GI16, HMS14, IPZ15, KSX17, LZ16, MS12, Sfa13, WX19a, Xu14b, YDk12]. Lax [Sog15, Sog16]. layer [FMP16, ZZ14]. layers [BLW10, BJV18]. LDG [ZZ14]. Least [HMT19, LPSH11, BHW13, BCL^{+}11, LLS15, LLS17a, NJZ17, TV10, Tre12]. Least-squares [HMT19, LPSH11]. Legendre [Sid11, WX12, qWlGjY17]. Lehmer [Shi19, DMPP19, Shi86]. Leray [DD17]. Lerch [BSO12, NRV15]. level [ABHV11, FG17, KT18, LX15, MS12, WX19b]. levels [BR19]. Levenberg [Fan12, Fan14]. Levin’s [ÁB17]. Lévy [KHOLT14]. Li [BMP10, KC11a]. liars [BS15a]. Lie [LPSH11]. lifting [BDM10]. liftings [MP19]. like [BEG16, Chu11, CLR13, Kir10]. limit [BS18, CY18]. limited [LWL19]. limiters [MP12, WX19a, Xu14b]. limiting [HLL16]. Lind [DMPP19]. Lindelöf [NRV15]. line [CLRR12, CS15a, DM14, HO12, MN10, Tru12, Wel17]. Linear [BT15, CWD14, GT16, Kop14, AW11, AZ18, AKPZ15, Bar19, BGP14, BNDHV15, CHZ17, Cao15, CZ16, CCD16, CHH18, CGP12, CMZ17, CDDM18, DJ13, DZ16, DI15, DL15, DJTZ13, ES16, FHN17, Fen18, GGH15, Gud10, HK18a, HMT19, HMS11, HS19a, HP12, Kop19b, KP14, Lab18, Lep16, LP17a, LW18b, MSW16, QSS18, Sid10, SW10, SZ18, UP14, WKN13, Xu14a, YDk12, ZY14]. linearization [DVY15]. linearizations [NN16]. Linearized [YY13, DPV19, ZZJZ18]. linearly [Ing13]. lines [BB12, CCOV14, GY19]. linesearch [GOP16]. linkages [GKL^{+}17]. Lions [DD17]. Liouville [BHS15, MDK13]. Lipschitz [AA11, HNRW18, Lic19b, SS15, Yao13]. liquid [GGG11]. list [Büt15]. Littlewood [DJŠ18]. LLL [CSV12]. Local [AFF^{+}17, CMQ13, DGS11, DL13, FW14, FS12, LW18a, Nar14, WLZS19, dVG16, CZ10, CMZ17, DM10, FX13, GHS10, GM10, LNRW18, LW12, LW15, Pet17, CZ16]. Localization [GG17, MP14]. Localized [AH17, Lin17, HNRW18, LNRW18, NRW17]. locally [AGS16, AG16, HS19a, Yao13]. location [LM18]. log [CC13b, HPS17, Qur19]. log-concavity [CC13b]. log-normal [HPS17]. Logarithm [Kop19b, CWX16, Die11, GJ19, KCL14]. logarithms [GPR13, Gui19, Sut11b, VJS14]. lognormal [KSS^{+}17]. long [KPRBT14, RR16]. long-range [KPRBT14]. Lopsided [FMMdW19]. Lotka [BHH^{+}12]. Love [FHN19]. low [ÁB17, BC17, ETX11, Sch10, Sør16]. low-order [ETX11]. low-rank [BC17]. Lower [LXX14, PM11, AL17b, CG14, CDTW18, Lep16, SGD11, Tho10]. lower-order [AL17b]. lumped [CT12]. Lusztig [GH15]. Lyapunov [AJ14].

M. [ÁB17]. Maass [JST14, Mez11]. MAC [GHLM18]. Maclaurin [Sid12]. made [CGG10]. magic [BvH11]. magnetic [CCL^{+}19]. magnetohydrodynamics [BP10]. Magnus [Coh12, Szm13]. Mahler [CDDM18, FR15a, Kim17]. maintaining [YDk12]. manifold [DS17]. manifolds [CJ15]. many [BEF16, GH18, GKS17a, KPW14]. many-subdomain [GH18]. map [Ang16, MN19a, Tui16]. mapping [ZGFD14]. maps [AJ14, Bar10, KN12, LGY15, Sij12, SV12]. marching [WLZS19]. mark [Alk10, Alk11]. marks [NP12]. Marquardt [Fan12, Fan14]. Maruyama [NT16]. mass [CT12, GSRM12]. Masser [MNZ13]. massively [GMS12]. matched [BLW10, BJV18]. matchings [MC13]. Mathematical [LZ17, Szm13, BL17, Coh12, LZ18, MOS66]. Mathias [BMP10]. matrices [AHS17, BMM18, DP15, DGM^{+}16, DOZ13, FMP16, GMSC^{+}17, Gre15, JR13, LL15, MN19a, NP17, Pan17, TWO18, XY13, Ye18]. Matricial [BK12]. matrix [AMR^{+}19, Azi19, BBB16, BMP10, Cin15, DD10, GY19, GM15, HZ13, HO12, LPSH11, LK12, Mel18, Riv16, dHAL11, FMP16]. Matsubara [Esp10]. max [CZ10]. Maximal [LS17, ALL17, Ane19, CFLTL14, Hei12, Li19]. Maximum [MP12, Kop19b, Li19, Xu14b]. maximum-norm [Li19]. Maxwell [BDG^{+}19, BG11, BCNS12, HMP13, HS19a, LCQ17, PS16, ZCS^{+}12]. may [Kop14]. Mazur [Che16]. McMillan [BJ15]. MDP [GT16]. mean [BMP10]. means [BAFG18]. measure [AJ14, CGH^{+}16, FR15a, Gon13, MFRV18]. measure-preserving [AJ14]. measurements [ABF^{+}10, AKKL12]. measures [DJ12a, DJ12b, Jor12]. media [BCFG17, BJV18]. medium [GG19]. meets [CKS19]. Meissel [Shi19, Shi86]. merit [MSM14, PKLC14]. Mersenne [BI10, GM13a]. Mertens [Hur18, StR14a]. Mesh [ZD14, AN15, FKS12, HK17, HZ17, HK18b, MO13, NKK17, Sfa13]. meshes [AGS16, AA14, BCL^{+}11, BS15b, BDFL17, Cao15, CC14, CWD14, CW17, CSO13, DLSW12, DL15, EGHL12, FSX19, Gra19a, GN14, HFC18, KHX14, Kop19a, Li17, Mou14, DD17, WHL19, Wan19]. meshless [LNRW18]. metamaterials [BJV18]. Method [CCOV14, Abd12, AV14, AH17, ABHV11, ALS17b, AD16, BLW10, BS18, ByHl10, BC16, BHW13, BAS16, BLS13, BKN15, BHH^{+}12, BCKM18, BMBO13, Bur17, BHL18, Bur10, Büt15, Büt18, CY18, CCDE19, Cao15, CR11a, COS14, CCQ17, CT12, CZ10, CS15a, CL16, CXZ16, CMZ17, CD18, CHR15, CGH10, CJ13, CZ14b, CL10, CNPT10, CB16, CD10, DKS15, DGS11, DP15, DYY14, Dro10, DJTZ13, DJL19, Dua14, DRS11, DE13, ELSW18, ES16, Fan12, Fan14, FKBJ17, FSX19, GLS11, GRBT16, Git13, GÖS18, GQ14, GLNP15, Gra11, HH13, HYZ14, HPS17, HFC18, Hia11, Hit18, HR10, HC11, HC15, HLL16, HL17, HK18b, JgLW13, KK11, KSU16, KSX17, KP12, Kop19a, Kre12, KW19, LL17, LNRW18, Li10, LSXZ12, Li15b, Li18, LSR19, LMNN18, LR19b, Liu15, LY17, LW18b, LWZ18, LCQ17, MDK13]. method [MYS12, MO13, Mus13, NRW17, NM17, NNZ19, OS16a, Ort11, DD17, QSS18, RW18, RS10, SW11, Seg17, SL17, Shi86, Shi19, TY19, Tru11, WY14, qWlGjY17, WW18, WHL19, WX19b, Wen13, Wil18, XZ10, Xu14a, XZ15, YMO13, Yao13, ZHX11, ZqCt16, ZZ14, BBSV19, Gav11]. methods [ABBM18, Akr13, Akr16, AZ18, AC18, Arg11, AHHR15, BC13a, BCCM13, BGP14, BD15, BC14, BGNyS11, BySZ12, BGG^{+}17, BHM12b, BTDG13, BH18, CZ12, CGS18, CNV14, CGHW14, CZ16, CD15, CCS12, CC12a, CMTZ16, CCL^{+}19, CT10, CLAT12, CC14, CXZ15, CNOS16, CCD16, CCLX17, CPBG19, COQ10, CGS10, CGN^{+}11, CC12b, CQS12, CS13c, CS14, CQS14, CQB14, CNZ16, CD16, CFS17, DPR14, DLSW12, DE10, DGM^{+}16, DGKS16, DT11, ET10, FN11, FW11, FX13, FHN17, FNO12, FKS12, FM18, GX17, GH18, GMSC^{+}17, GHS10, GOS11, GMP14, GL15, GGH15, Gud10, GP14, GZZ17, HK18a, Hai10, HMP13, Hoa16, HSW10, HX19, HS11, Jin10, Jin13, JLTZ17, KM15, KMPW10, KPY18, KG18, KSS^{+}17, LK12, MSW16, MS12, MN19b, MPT18, OMS15, PVV17, SW10, TCH12, TY17, Voh10, WMxY17]. methods [WNK18, WLZS19, YPP13, YY13, ZCS^{+}12, BCKX13]. metric [Nak11]. metrical [Jag12]. metrized [Cin15]. MHD [HX19]. microfiche [BR98]. Midpoint [MRW10]. million [BH11]. mimetic [AdVV13]. min [CZ10]. min-max-orthogonal [CZ10]. Mindlin [Dua14, DRS11, dVMR19]. Minimal [CW17, Err14, WX13, BP12a, COS14, HZ15, MMV17]. minimal-variable [MMV17]. minimality [KZ18]. minimax [Yao13]. minimization [BKM18, CD15, HYZ14, Ish17, KM16, LM13, YMO13, YY13, ZDL13]. minimizing [FGMS12]. minimum [CLTZ12, GMRL18, Lez14]. Minkowski [Alk11, Alk10]. Minkowskian [KMSwaAbMDS12]. Mixed [CZ12, GGG11, AGS16, ABHV11, BGGG17, COT17, CR11a, CHH18, CD16, DiP18, DYY14, EFP10, GOS11, GRBT16, KK11, Lic19a, LZ18, MSS11, MZ10, NM17, PVV17, SSW16, Voh10, WY14]. mixed-FEM [COT17]. MNT [ULS12]. Möbius [BC13b]. modal [ABF^{+}10, GÖS18]. model [Bar15b, BER17, BD11, CS18, CH19, CGH^{+}16, DZBK16, FHN19, GGG11, GGRBRG19, GSRM12, JLQZ18, LM13, MZ10, QzSZ15]. modeling [FN11]. modelling [GM10]. models [Alf10, Baa10, BGM15, DJTZ13, HO12, HX19, MS16, Qur19]. modes [DRS11]. modified [Adc11, Fan12, Fan14, Ten10, WLZ17]. Modular [BLS12, Baa10, BMS16, BR19, BHLL19, CFvdG19, Dem14, Dup11, FG17, GP17, GV11, KK10, Kim17, LP18a, Mas18, Nel15, OS19, RSS12, Wut18, ZY15]. modularity [DGP10]. module [IKMF17, PV15]. modules [BJ11, Yen11]. moduli [Err14]. modulo [GG14, MNW10, ZY15, Zra10b]. modulus [Fre18]. Moebius [Ram19, Ram15]. MOLS [EW16]. moment [CS13a, RY15]. moments [Alk10, Alk11, HJSZ18, HO12]. momentum [BC13a]. Monge [BCM16, BGNyS11, NNZ19]. monoids [BOP17]. monomial [KTA18, SV12]. Monotone [BCM16, ET10, HR10, LA14, Ten10]. monotonic [KL10]. Monotonicity [Gau17, KL13]. Monte [BSSW14, AH14, HPS17, He18, KSS^{+}17, MSM14, MS12]. moonshine [JST14]. Mordell [BBC14, BB16, Kuz15]. Mori [HKW17]. morphisms [Hut15]. Moser [GMZ11]. most [Tao14]. motion [GKL^{+}17]. motions [FGMS12]. moving [FSX19, HK18b]. Mullen [COT15]. multi [BP10, Bar19, BCL15, CGHW14, CMR10, LZ15, LX15, MS12]. multi-component [BCL15]. multi-dimensional [Bar19, BCL15, CMR10, LZ15]. multi-fluid [BP10]. multi-level [LX15, MS12]. multi-physics [CGHW14]. Multidimensional [LSW14, CG11, JX14, MT19]. multiframelets [Han10]. Multigrid [CGO10, Ols12, BP12b, Li15a]. Multilevel [CNOS16, DFR12, KSS^{+}17, LSXZ12, dDHZZ14]. multimodular [Har10]. Multiple [DLZ11, Abd12, CZ10, JXR12, MS14, Zha15]. multiple-dimensions [JXR12]. multiples [DJŠ18]. multiplication [ASSW16, ABBR17, CT19, HvdH19, JKLM17, PR15]. multiplicative [BCG16]. multiplier [DHYZ17, Gav10b]. multipliers [AK16, TY19]. multipole [Gav11]. multiresolution [CDHM12, HMS14]. Multiscale [CH19, Abd12, AV14, CGH10, IW13, KPY18, MP14, PVV17]. multistatic [AKKL12]. Multistep [BHH^{+}12, Akr13, Akr16, BGG^{+}17, HK18a]. Multivariate [KPW14, KSW17, Chè13, DLPW11, FA14b, Han10, KSWW10]. multiwavelets [Han10]. Mumford [BER17, Has12].

Naghdi [NM17]. Naor [LSW14]. narrow [OS16a]. narrow-band [OS16a]. Navier [CZ12, COT17, CCQ17, CW10, DE10, GHLM18, GLS11, GMP14, GMS12, Ing13, LS18, NL19, RVX19, TU18, Wil18]. nearest [MRW10]. negative [BJV18, XY13]. nematic [GGG11]. nerve [SS16]. nets [DK10]. networks [ZGFD14]. Neumann [DS18, MS10]. newform [LW12, LW15]. Newman [DJŠ18, SGD11]. Newton [Arg11, BAS16, BLS13, Dup11, HW18, Jin10, Jin13, Seg17]. Newton-discontinuous-Galerkin [HW18]. Newtonian [KS19]. NFS [BL17]. Nicolson [AKPZ15, Ing13, RS16, Xu14a]. ninth [BD10b].

Nitsche [CHR15]. NLS [BFS18]. No [DHJ17, McN15, Wel17]. node [Hoa16]. nodes [CH13, MS14, SH11]. noise [GLW19]. Non [KZ18, Yan17, AA11, AFF^{+}17, BP12b, BGP14, CGHW14, CCL^{+}19, CHR15, DJ13, DZBK16, DHJ17, FHN17, FG17, GM10, GKS17a, HFC18, HS15, JT19, LLS15, LLS17a, LNRW18, Mat19, MZ10, Mül10, Mus13, Pla17, SZ18, WW18, XY13]. non-adjacent [KZ18]. non-commuting [Mat19]. non-divergence [FHN17, WW18]. non-existence [Mül10]. non-explicit [DZBK16]. non-homogeneous [CCL^{+}19]. Non-iterative [Yan17, CGHW14]. non-linear [BGP14, DJ13, SZ18]. non-Lipschitz [AA11]. non-local [AFF^{+}17, GM10, LNRW18]. Non-minimality [KZ18]. non-negative [XY13]. non-real [DHJ17]. non-residue [LLS15, LLS17a]. non-residues [JT19]. non-smooth [GKS17a, Mus13]. non-split [FG17]. non-stationary [MZ10]. non-symmetric [CHR15, HS15]. non-trivial [Pla17]. non-uniform [HFC18]. non-variational [BP12b]. nonabelian [KM19]. noncollinear [Kir10]. Nonconforming [CHZ17, WX19c, Bur17, BH18, CGS15, CC14, LMNN18, Wan19, ZHX11]. nonconservative [CPPT10]. nonconvex [CD15, LZ17, Li19]. noncooperative [CZ10]. noncrystallographic [CS15b]. noncylindrical [AKPZ15]. nondivergence [Gal19a]. nonlinear [Akr13, Akr16, ADL11, AL16, BR18, BJKM11, BCL15, BGNyS11, CG16, CMTZ16, CS13a, CMS11, DLZ11, Deb11, DKMW13, ET10, Fan12, Fan14, GI16, GGH15, HP10, HR10, HJW18, Jin10, Kre12, MPT18, NL19, OS10, QzSZ15, RVX19, Seg17, Shp11, SX16, WMxY17, YPP13, ZqCt16]. nonlinearity [HLR13]. nonlocal [BC14, DJTZ13, DJL19, Mon10b]. Nonnegativity [GG19, JLTZ17]. Nonoverlapping [HSW10, GX17]. nonresidues [SwaAbLCW11]. nonsingular [Azi19]. nonsingularity [HK18b]. nonsmooth [LS17]. nonstandard [HP10]. nonstationary [BP10]. Nonsymmetric [MSS11]. nontrivial [CM13]. nonuniformly [CNOS16]. norm [Cer11, DLSW12, FL18, Kop19b, Li19, MPT18, Not15, WO10, YY13]. norm-Euclidean [Cer11]. normal [ÁB17, BHS15, Bur10, FA14a, FA14b, HPS17, Kir12, Sch17]. normality [NS13]. Note [GKS17b]. novel [FA14b, Mus18, NRW17]. Nuclear [FL18, YY13]. Number [JW12, ABE^{+}16, AK12, BBT15, BKK19, BHS15, Bia14, BC15, Cer11, Cha18, Coh15, DHJ17, FW11, FX13, GPOS14, GLM19b, HZ13, HMS11, Har14, HKÖ11, JKL11a, JKL11b, Jeo16, Kem10, KM19, KÖ18, Kru16, Lez14, Lou11, LCQ17, Meš14, Ngu16, OR14, OS10, PV15, RGNS18, Rob15, Sch10, TV10, Tao14, Tho10, ULS12, ZD14]. numbers [AGHS14, ÁB17, BC13b, BR98, BR05, BR11, CLPM16, CS10, CD17, DFGSL13, Dit14, DK15, FNO12, GM13a, GOM11, GOMS15, HP13, Har10, KM19, Kuc11, LOX14, Lou11, Mil15, MNW10, NP19, NS13, Nie15, OR12, Pap15, Sch17, Shp11, Wu10a, Zha11b, Zha15]. Numerical [Bar10, Ble12, BP15, CLAT12, DM14, Doo15, DRS11, Gal19a, GO16, Hai10, HWZ14, HP12, JL12, LZ15, MSV19, MNZ13, Pla16, PZ17, RV14, RSS12, Wut18, ZZJZ18, ABBM18, BOP17, BFZ10, Bor10, BAFG18, CHR15, DFGSL13, DP15, Dro10, FH16, GZ13, GM10, HNUW14, HJW18, JZ11, JLZ19, LK12, LZ17, RPR12, Spi17, SZ18, HO12]. numerics [HLJ16, IPZ15]. Nyström [GL15, OMS15].

Oberhettinger [Coh12, Szm13]. objective [KM16]. oblique [Gal19a]. Obrechkoff [ADGR12]. obstacle [AdVV13, BySZ12, GP14, KZ11]. obstacles [BCS16]. obstruction [LRS16]. occurrences [Kus18]. Odd [CS10, Nie15, OR12, Dit14, FNO12, FLM14, OR14, Tao14]. ODEs [LM18]. one [Adc11, AJ14, CZ16, CCD16, CCLX17, DOZ13, DJL19, GM10, KSX17, KZ18, LMY12, LR19b, LW18b, LS10b, SS15, Spa13, WX19a, Xu14b]. one-dimensional [AJ14, CZ16, CCLX17, DJL19, GM10, LMY12, LR19b, LW18b, LS10b, WX19a, Xu14b]. one-sided [SS15]. one-space [CCD16]. only [KSW19, Kop14]. operations [JR18]. Operator [HKRT11, HLR13, HKK13, AZ18, BNS15, BCM16, BzCS12, BP12b, Che12, CS13a, ET10, FMP16, HY15, LMY12, MDK13, MRSC19, MMN11, RS17a, RV14, Vab12, Ver14, WO10]. operator-difference [Vab12]. operator-valued [BzCS12]. operators [AGS16, AFF^{+}17, BP15, CCDE19, CM12, CNOS16, DY12, DGS12, Gal17, Git13, HR10, JLPR15, KMF17, LWCI13, Mat19, MN19b, WKN13, Ye18]. Optimal [BC13a, BGGG17, BSSW14, Che12, DI15, GGH15, KHOLT14, LS10a, LR19b, Liu15, LWL19, MSW16, PMH18, SW10, APR15, BM11, BBSV19, CR11a, CPSV18a, Mus18, Str13, dVG16, JR18]. Optimality [LM13, LMNN18, ZCS^{+}12]. Optimally [CPBG19, MC12]. Optimization [ABF^{+}10, BBT15, BO15, CS15a, CD18, HLZ15, LGY15, TY17, ZQY12, Naz10]. Optimized [BCCM13, GX17]. Optimizing [Bur10]. orbit [GLM19b]. orbits [LPSH11, dGO14]. order [AADL11, ABBM18, Alf10, AC18, AL17b, BCCM13, BSSW14, CSS15, CPPT10, CLAT12, CH13, CXZ15, CCD16, CS13a, CPBG19, CR16, COQ10, CQS12, EW16, ETX11, EFH19, ES16, EHR18, FHN17, FL18, GM14, GM17, GV15, GGH15, GK11, HKÖ11, JLPR15, KP12, Kop14, KL10, KP14, Kry13, LP18a, LWCI13, LFS15, MSV19, MR10, PKLC14, QzSZ15, RW18, SSW16, SW11, Str13, TZD15, WX13, WY14, WW18, WX19a, Wen13, WX19c, Xu14b, ZY14, ZZWZ15, ZHX11, CCDE19, DD17]. Orders [DS18, BJ11]. orthogonal [AH17, ADGP13, ByTC13, BMPR16, CLRR12, CZ10, HZ13, IW13, KS15, Kve10, MFRV18, RS17b, WHV14]. oscillations [ADL11]. oscillators [XW10]. oscillatory [CMTZ16, IL11, MX18, XW10]. osculatory [RS17b]. Osher [CMR10]. other [Cof14, Riv16]. overlapping [KP12, Yan17].

P [Coh12, Szm13]. Padé [BKMW11, YF15]. pair [KtR10]. Pairing [IJ13, ULS12]. pairing-friendly [ULS12]. pairs [CLTZ12]. palindromes [CLB18]. paper [Cof14]. para [CLRR12, MFRV18]. para-orthogonal [CLRR12, MFRV18]. parabolic [Akr13, AK16, Akr16, ALL17, BCCM13, CG16, CZ16, CT12, CQT11, DM10, GV15, Gon13, GO16, GK11, JL12, KSU16, LMP15, LS17, Li19, RS16, UP14, Wen13, Yan17, ZD14]. Parallel [CGHW14, GMS12, HYZ14, Yan17, ZGFD14]. parallelepipeds [SR11, SVVR14]. parameter [LM13]. parameterized [SY16]. parameters [AKKL12, Chu12, DP15]. parametric [BC17, RS17a, Tab17]. Parametrized [Xu14b]. Paramodular [PY15, RTwaabRS16]. Part [Ble12, BB12, DHJ17, AA11, AKLZ12, AKKL12, CC14, CFS17, DGS12, EGHL10, NH11]. partial [Bar16, BTDG13, Büt16, Deb11, ELSW18, FHN17, GG17, HLR13, KSU16, La 15, RW18, WX13, WKN13, WX19c, Yan17, Zra10b]. particle [ALS17a, HL17, LY17, MO13]. particle-mesh [MO13]. partition [CWX16, Pro17]. past [BM11]. patches [KS15]. paths [BCKM18]. pattern [CMS11, Kus18]. pattern-forming [CMS11]. PDE [HK18b]. PDES [NRW17, BC17, ES16, GLW19, HPS17, OS16a]. Pearcey [LP17c]. Pell [MRW10]. Pellet [Mel16]. penalized [GGG11]. Penalty [HS15, WMxY17, AC18, BGNyS11]. penalty-projection [AC18]. pencils [BEG16, Ver14]. Perfect [BC15, BSL13, SR11, Dit14, FNO12, Nie15, OR12, OR14, SVVR14]. perfectly [BLW10, BJV18]. peridynamic [DJTZ13]. Period [RS14, Alk10, Alk11, MN19a, MNW10]. periodic [CJ13, KSU16]. periodicity [CVZ19]. periods [Lai16, Ser19]. permanent [ELSW18]. permutation [BGU19]. permutations [Kus18, NPPY12, Shp11]. Perron [Wu10a]. persistence [JK17]. perspective [PS19]. Perturbation [CSV12, BCG16, LR19a, ZZWZ15]. perturbed [FM18, GO16, HS13, KP12, XZ10, ZZ14]. perturbing [DP15]. Petrov [RS17a]. Peyre [DEJ14]. Pezzo [Qur19]. phase [CS18, CPV14, DVY15, Fre18]. phase-field [CS18]. phenomenon [BKMW11]. physics [CGHW14, Coh12, MOS66, Szm13]. Picard [Bru13]. pick [Riv16]. Piecewise [BS15b, CDFS13, AH15a, BY12, Bat15, DT11, LP17a, LXX14]. piecewise-smooth [Bat15]. Pisot [DHJ17, HP16]. Pitaevskii [BC13a]. plain [HvdH19]. Planar [GKL^{+}17, BCCW16, Gal19a, JKK^{+}10, Mou14]. plane [BBG19]. planes [Rit10]. plate [DRS11, FHN19]. plates [BySZ12, Dua14, Xu14a, dVMR19]. plus [CL10]. PML [BPT10, CXZ16]. PN [Hut15]. Poincaré [DT11, JJK^{+}15]. Point [BGH14, AH14, BST11, DKMW13, Fis17, HH13, JKLM17, JR18, LS10b, MSM14]. points [Baa10, BÇS15, BBM17, BLS13, BCL^{+}11, DDLR15, Els12, GS10, GM15, HHT19, HPS17, Hut15, Kru16, Ler12, OS19, Tui16, Yao13]. Pointwise [GL12, DM10, DL13, Kop14]. Poisson [HFC18]. polarization [AKLZ12, AKKL12]. polarized [Cin15]. Pollard [Zra10a]. pollution [Pet17]. polygonal [Wan19]. polygons [LZ17, RGB14]. polyharmonic [Gal17]. polyhedra [GL12, LS17, Li19, SSW16]. polyhedral [Li18, MSS11]. polylogarithm [RGNS18]. polylogarithms [BB16]. Polynomial [CDTW18, DGS12, GMS19, AZ11, AJ14, BBD^{+}11, BL17, BDM10, BCCW16, BKS15, BCPR14, COS14, Che12, CR15, DYY14, DPV19, DJŠ18, FM12, GY19, HLZ15, HLJ16, Kve10, LP17b, LS10a, NN16, Ngu16, OS10, RJS12, SH11, TWO18, WHV14, dVG16]. polynomials [AS15, ADGR12, ADGP13, AM11, AMR^{+}19, BBT15, BBKZ16, Bar15a, Bay11, BC13b, BL12, BAS16, BLS13, BCST19b, BCCW16, BLS12, CDS10, CLRR12, DJŠ18, Err14, FR15a, Gau17, GG14, HSW11, IKRS12, Kön17, KS15, Ler12, LXX14, LOX12, LX19, LS10c, MFRV18, Mel16, Mel18, MC15, NRV12, NS18, Pan17, RY15, RS17b, Sán19, SZW11, Str14b, Sut11a, Ula14, WHV14, Zra10b]. polytopal [CW17, DiP18]. polytopes [BZ18]. polytopial [CPBG19]. porosity [GRBT16]. porous [GG19]. porous-medium [GG19]. Portrait [KT10]. posed [AKPZ15, BTDG13, CSS15, OS16a]. posedness [GM10, LHY15]. posets [CS15b]. position [JvSRV14]. position-dependent [JvSRV14]. Positive [LS10c, Peh11, BzCS12, BC16, CM12, CLB18, KL10, McK11, SD10, STD15, SX14, ZC13]. Positivity [CGP12, LWZ18, LP17a]. Positivity-preserving [CGP12, LWZ18]. post [JvSRV14]. post-processing [JvSRV14]. Posterior [ST18b]. posteriori [BM11, BER17, CHZ17, CPV14, CMQ13, CZ14b, DM10, DL13, DVY15, DJTZ13, GP14, KM15, LMP15, MR10, Ort11, RS16, Voh10, WKN13]. postprocessing [Cao15]. potentials [CMS11, LMS11]. Powell [CLR13]. power [AK12, Ang16, Doo15, Har11, Hir13, Joh15, KS19, KB16, Mah14, Wu10b]. power-law [Hir13, KS19]. powers [BP15, BD10b, Hit17, Sad14]. practical [FKBJ17, GQ14]. Precise [Fuk11]. precision [BCPR14, dR11]. Preconditioned [SX16, WX19b]. preconditioners [BBB16, CHH18, dDHZZ14]. preconditioning [BDG^{+}19, DFR12, EFP10, GOP16, GSV17, LSXZ12]. preimage [KO12]. preorders [KTA18]. preperiodic [Hut15]. prescribed [GKL^{+}17, JKL11a, JKL11b, Jeo16, Sut12]. presence [CZ11, LZ16]. Presentations [JKdR16]. preservation [JLTZ17]. preserving [AJ14, Bar19, BLT13, BGG^{+}17, COS14, CGP12, GGH15, GG19, HK18a, HX19, LWZ18, NKK17, Xu14b]. pressure [CZ12, CC12b, GRBT16, LMNN18]. pressure-robust [LMNN18]. primal [OWZD18, Voh10, WW18]. primal-dual [WW18]. primality [ASSW16, GOM11, GOMS15, LP18b]. Prime [Mah14, AS15, BKK19, FLM14, GG14, GM16a, GM16b, JD14, KC15, KtR10, LN11, LN14, Mil15, MNW10, OR14, OHP14, RGNS18, Sad14, SW17, Tre12]. prime-pair [KtR10]. primes [BI10, BDFP15, BH11, CGH14, CT19, HHO17, HvdH19, Meš14, TCH12, Tao14]. Primitive [BCST19b, BCST19a, BKK19, COT15, JT19]. principle [BH16a, FJS16, HK17, Jin13, MP12, Xu14b]. priori [CQS14, LL17, Ort11, Voh10]. priors [CSS15]. Probabilistic [AH14, BLS13, LL15]. probability [FA14a, FA14b]. problem [ABHV11, BCLZ14, Bar15b, BBR12, BG11, BD15, BKN15, BPT10, BySZ12, Bur17, CC13a, Cal13, COT17, CD15, CT12, CGN17, CC12b, CGH^{+}16, DVY15, Die11, EFP10, EGHL10, EGHL12, FKS12, GOS11, GCL19, GL12, Ish17, KCL14, KP12, Kop19a, Kre12, LL17, LMMR19, Ols12, PKLC14, RPR12, SW11, XZ10, Xu14b, ZZWZ15, ZZ14, ETX11, LMNN18]. problems [Abd12, AV14, AdVV13, BCFG17, BBB16, BHW13, BST11, BCKM18, BNMP18, CHZ17, CSS15, CGS18, CGS15, CXZ16, CPSV18a, CHR15, CGH10, CQS12, CD16, CG11, DFR12, DM10, DKMW13, DPV19, DJL19, ET10, FM18, Gal19a, Gav10b, GV15, GI16, GO16, GSV17, Gud10, GP14, GSS16, GK11, HW18, HSW10, HC11, HC15, Jin10, JL12, JLPR15, KZ11, KSS^{+}17, LMP15, LLS15, LLS17a, LW18a, Li15b, LX15, LS10b, MP14, MDK13, Mat19, MNU15, Nak11, Pet14, Pet17, SSW16, Spi17, SW10, UP14, WY14, WHL19, WX19b, WKN13, ZQY12, ZY14, ZD14, dDHZZ14]. procedures [Gra11]. process [Alf10, HS19b, Sid10, ST18b, SH11]. processes [BMM18, HJW18]. processing [HYZ14, JvSRV14]. producing [AS15]. product [AGHS14, BP12a, CDFS13, FA14a, GH13, JdRV14, Sad14, SSV14, ZC13]. product-quotient [BP12a]. products [BEF16, JKdR16]. programming [AZ11, DHYZ17, WMxY17, ZZWZ15]. progression [McN15]. Projection [Bar16, DGKS16, AC18, BI19, CCS12, CGS10, Li17, Che12]. projection-based [CCS12, CGS10]. Projection-free [Bar16]. projections [CMZ17, FW14, Lic19a, Lic19b]. projective [Bru13, Kru16, Tui19, Wel17]. projector [MN10]. prolate [Wan10a]. Proof [BDFL17, BD13, CH13, COT15, CR11b, PS19]. propagation [BCFG17, BCS16, DK10, Mon10b]. proper [IW13, PY14]. properties [AK16, Akr16, BMP10, CH19, Gau17, HP16, HJW18, KL13, NH11, Ser17, Wan10b]. property [CNZ16, DLSW12, MMN11, SZ18]. Prouhet [Cal13]. Proving [DGP10, ASSW16, Büt15]. Pseudo [LSW14]. Pseudo-Random [LSW14]. pseudoprimes [JD14, Mül10, SW17]. pseudorandom [GPOS14, GLM19b, HMS11, OS10, Shp11]. pseudospectral [GLS11]. pseudostress [CZ12]. pseudostress-pressure-velocity [CZ12]. PSLQ [FCW19]. Puiseux [WLZ17]. pursuit [DZ16]. Pythagorean [JKK^{+}10]. Pythagorean-hodograph [JKK^{+}10].

QBD [BMM18]. QMC [BSSW14]. Quadratic [JT19, Kle16, OS19, RGB14, BBM17, BCST19b, BC15, BF18, CIL15, Coh15, CMS11, DGP10, GP17, Gre15, GM13b, KKP10, Kir12, Kuc11, LLS15, LLS17a, LOX14, MJ16, SwaAbLCW11, Tre12, ZZWZ15]. quadrature [BR18, GLS11, IL11, JR13, LS10c, MS14, Mon10a, Not15, Peh11, Spa13]. quadrilaterals [ByHl10]. quandles [VY19]. quantum [LZ18, PPTZ13]. quantum-classical [LZ18]. quartic [BCST19a, Els12, GJLR18, HZ14, JKL11b, Lou11]. Quasi [AHS17, BM11, He18, KSS^{+}17, LMNN18, ByTC13, BMM18, DL15, HPS17, Lab18, MSM14, SH11, BSSW14]. quasi-Chebyshev [SH11]. quasi-incompressible [DL15]. quasi-linear [Lab18]. Quasi-Monte [He18, KSS^{+}17, HPS17, MSM14]. Quasi-optimal [BM11]. Quasi-optimality [LMNN18]. quasi-orthogonal [ByTC13]. quasi-Toeplitz [BMM18]. Quasi-unbiased [AHS17]. quasicontinuum [Ort11]. quasilinear [AV14, ALL17, GLM^{+}19a]. quasinonlocal [Ort11]. quasiseparable [DOZ13]. Quaternion [MC15, PS14]. question [Alk10, Alk11]. questions [Riv16]. quincunx [HJSZ18]. quintic [Dum17, LM17]. quotient [BP12a].

R [Coh12, Szm13]. Rachford [DZ16]. racks [HS19b, VY19]. Radau [CMZ17, Not15]. radiative [GZ13]. radii [HLJ16]. radius [BM12, CLTZ12]. Rado [ABE^{+}16]. Ramanujan [Chu11]. Ramanujan-like [Chu11]. ramified [PV15]. Random [LSW14, BCLZ14, Cha18, EFP10, Git13, HO12, HL17, LL15, LLP16, LP18b, LY17, MS12, Ngu16, Riv16]. randomized [KW19, LW16]. range [Gau17, KPRBT14]. Rank [SY16, BC17, DOZ13, Fon11, Hei12, KSW19, PR15, Sør16, Yen11]. rank- [Sør16]. Rao [DFGSL13]. rapidly [Mon10a]. rate [NT16, TY19]. rates [BI19, CDFS13, GLW19, Li10, PMH18, WX12]. Rational [Els12, BDFP15, BCCW16, CM13, Chè13, DLRNS18, FGMS12, Fis17, GJLR18, Hut15, JR13, Kle13, Lai16, NH11, Pap15, SSV17, YF15, PDSV15]. Raviart [CGS15, AADL11, CZ14b, DL15, OWZD18, Sty14, Wan19]. ray [Kuc11]. Rayleigh [Gal19b]. reaction [BNMP18, DHMG11, HH13, IW13, KP12, SS15]. reaction-diffusion [KP12]. Real [Mil15, BGH14, BB12, CLRR12, DM14, DHJ17, GP17, KK14, MN10, Rob15, Tre12]. realizations [LMY12]. reciprocal [Har11, NP19]. recombination [Chè13]. reconstructing [ABF^{+}10]. reconstruction [AH15a, AKLZ12, AKKL12, ABBSM16, BY12, Bat15, BDFP15, MZ16, Sfa13, XZ15]. reconstructions [CPSV18b]. Recovering [GG14]. recovery [Cao15, GZZ17]. rectangular [CH13, HZ15, Sty14]. recurrence [BMPR16]. Red [Gra19a]. Rédei [Sza11]. reduced [LRS16, UP14]. reducing [Źra19]. Reduction [BL12, CSV12, CMSC10, HMS11, Pap15, ZY15]. Reductions [Ver14]. refined [AGS16]. refinement [DS17, Gra19a, dVG16]. refinements [BS15b]. reflection [MP17]. Regeneration [HSW11]. regime [BS18, CY18]. region [SD10, STD15]. regions [HNRW18, ND19]. regression [GT16]. regular [KT10, TK12]. Regularity [ABBM18, EHR18, Li15a, ALL17, Has12, Li19, LWL19]. regularization [DVY15, KO12, PMH18]. regularized [BCKM18, Jin10, Jin13]. regulator [Bia14]. regulators [BÇS15]. Reingold [LSW14]. Reissner [Dua14, DRS11, dVMR19]. related [BO15, ByTC13, Büt16, IKRS12, KL13, LLS15, LLS17a, MRSC19, Wei19]. relation [LOX14]. relations [BT15, NRV15]. relative [JR18, Mor13, XY13]. relaxation [BLT13, CL10, CNPT10, CJLW18]. relaxations [JgLW13]. reliable [KZ11]. remainder [BBR12, Sut11a]. remarks [BL17]. removed [Kop19b]. repdigits [BSL13]. repetitions [RR16]. representation [Cof14]. representations [CFLTL14, DD10, LPRY10, LP17b, Mas18]. representatives [PS14]. repulsion [GGRBRG19]. Residual [CHZ17]. Residual-based [CHZ17]. residue [Ang16, BF15, LLS15, LLS17a, Źra19]. residues [JT19, Lou11]. resolved [AKLZ12, AKKL12]. resolvent [Coh15]. restoration [BKM18, ZDL13, KM16]. result [Jag12, Źra19]. results [BCST19a, Cof14, Gau17, Jin13, LWCI13, StR14a]. reverse [LWCI13]. reversion [Joh15]. review [Gav10a, Gav10b, Gav11, Gra11, Hai10, Kem10, Naz10]. revisited [BDM10, BCKM18, GMZ11, GM13a, GS19]. RH [Büt16]. Rham [Nei15]. Riccati [LK12]. Richardson [Sid10]. Riemann [BSO12, BB12, Cof14, Dus16, GI16, Hia11, Hia16a, KK14, LPRY10, Tru12, dR11]. Riemann-Siegel [dR11]. Riesz [JZ11]. rigid [FGMS12, Qur19]. Rigorous [BJM17, CMSV19, HLJ16, GJ19, vdBLM10]. ring [CMSV19, FK18b, Lom19]. rings [Bis15, BJK11, Gre15, HKL16, JdRV14, JJK^{+}15]. Ripa [DZBK16]. Ritz [Gal19b, Li17]. Robin [JL12]. robust [BM11, LMNN18, ZN19]. robustness [PVV17]. rods [Xu14a]. role [CD17]. Root [BBT15, BKS15, CS15b, Har11, JW12]. rootfinding [NP17]. roots [BLS13, CMSC10, COT15, DS14, JT19, NN16, RS17b, SwaAbLCW11, WHV14]. rotating [GLS11]. rotation [BC13a, FGMS12]. ruled [SSV17]. rules [BGU19, DK10, LS10c, SX14, Sør16, WLZ17, vDE19]. Runge [AZ18, BGG^{+}17, COS14, COQ10, ET10, GGH15, HS11].

Sabin [CLR13]. saddle [GOP16, Yao13]. Saint [ABBSM16]. Saint-Venant [ABBSM16]. Salpeter [GCL19]. same [DHJ17, ZC13]. sampling [AH15a, CJLW18]. satisfying [BC16, BMP10, CNPT10]. SAV [LSR19]. scalar [BCG16, BGP14, BCL15, CZ11, CJLW18, DHMG11, IPZ15, LZ16, WX19a, Xu14b]. scale [BMBO13, DGKS16, JgLW13, MZ16, NNZ19, SX16]. scales [Abd12, AH17, DJ12a, Jor12]. Scaling [CPSV18a, Cal16, OWZD18]. scattering [BLW10, BPT10, CXZ16]. scheme [ADL11, AC18, ABBSM16, Bar10, CS18, CG16, CLWW16, CSO13, CMR10, DHYZ17, EGHL10, GLL12, GHLM18, GT16, HS13, Ing13, LS18, LX15, Mon10a, Sog15, Sog16, TU18, WZZZ19, YDk12]. schemes [ALS17a, Alf10, AG16, Bar19, BCG16, BGM15, BCS16, BCL15, BSSW14, CGP12, CJLW18, CG11, DJ13, DZBK16, DKMW13, DHMG11, EGHL12, GG19, GGRBRG19, GK11, HMS14, JLQZ18, KPRBT14, KHOLT14, LZ15, MP12, Mon10b, MZ10, PS16, Pro17, QzSZ15, RVX19, SZ18, Ten10, Vab12, Vab14, WX19a, Xu14b]. Schottky [CKS19]. Schrödinger [AKPZ15, CMTZ16]. Schrödinger-type [AKPZ15]. Schwarz [GX17, GH18, KP12, Yan17]. Schwarzian [Seg17]. Scott [GS19]. SDEs [SZ18]. Search [AKR18, CS15a, CGH14, GOP16, Meš14]. Searching [AT16, BGH14]. Second [AC18, Cre14, EFH19, CQS12, FHN17, Fuk11, GK11, KP12, Kry13, PKLC14, QzSZ15, SSW16, Str13, TZD15, WY14, WW18, ZY14, ZZWZ15]. Second-order [AC18, CQS12, KP12, Kry13, PKLC14, QzSZ15, SSW16, ZY14, ZZWZ15]. sectional [MP19]. sector [FR15a]. seeds [BAFG18, Zha11b, Zha15]. Segal [LWZ18]. Segel [HL17, LY17]. Seidel [GY19]. Selecting [Bar15a]. selection [BL17, JLQZ18, LM13, OWZD18, QzSZ15]. self [Ang16, Kir10, YDk12]. self-affine [Kir10]. self-canceling [YDk12]. self-power [Ang16]. Semi [BMM18, BHM12b, DJ13, AZ18, KK11, Li19, Yao13]. semi-differentiable [Yao13]. semi-discrete [Li19]. semi-explicit [AZ18]. Semi-implicit [BHM12b]. Semi-infinite [BMM18]. Semi-Lagrangian [DJ13]. semi-stationary [KK11]. semidefinite [AZ11]. Semigroup [CG11]. semigroups [BAFG18, DFGSL13, FH16, RPR12]. semilinear [HW18, JgLW13, KP12, KW19, Wen13]. semilocal [Arg11]. Semimatching [CC14]. semiregular [NPPY12]. semisimple [dGO14]. sensing [BO15, BMBO13, CDTW18, RS17a]. separable [BFZ16, DHYZ17, HYZ14, SY16]. separate [ELSW18]. septic [LM17]. sequence [CLPM16, MFRV18]. sequences [BM12, CLTZ12, GLM19b, Mah14, Sid10]. sequential [DHYZ17]. serendipity [GK19, RGB14]. Series [Cof14, Fuk12, Adc11, Alk10, Alk11, BH16b, CFJ12, CZ14a, FA14b, Har11, Joh15, KPSY18, Lan11, LA14, RSS12, Sid11, Chu11, Chu12, Ram16]. set [BHSP11, BLS13, CD18, HHT19, HLZ15, KT18, NPPY12, Sdi19]. sets [AH14, BKK19, BHSP11, CDTW18, MSM14, McN15, Xia18]. Setzer [DS18]. several [Ram15, Ram19, Wei19]. sextic [BH16b]. sextics [CFvdG19]. Shafarevich [Cor10, DS18]. Shah [BER17]. shallow [AD16, BC16]. shallow-water [BC16]. Shanks’ [BHH^{+}12]. Shape [AKLZ12, AKKL12]. Sharp [Fre18, GLW19, LR19a, Mor11, Sty14, SD10]. sharper [STD15]. Sharply [DM10, DGS11]. Sharpness [Li10]. Shaw [CLWW16]. shear [dVMR19]. shear-deflection [dVMR19]. shell [NM17]. shifted [ELSW18]. Shimura [Nel15]. Shishkin [FKS12]. shock [CJLW18]. shocks [CZ11, LZ16]. shooting [LMT10]. Short [MC13, KPRBT14]. SIAC [JvSRV14]. sided [SS15]. Siegel [Gra19b, BSO12, Hia16a, KPSY18, Lan11, RSS12, dR11]. Sieve [FNO12, BBT15, Gra11, Sch10, Bar15a]. sieving [Kle16]. sign [FK18b, PT16]. simple [BJK11, FMP16, GMS12, HSW10, Kön17]. simple-layer [FMP16]. simplex [BBD^{+}11, Che12, CLR13, FSX19, SX14]. simplicial [CWD14, Gra19a, WHL19, WX19c]. simplified [GJ19]. simulation [ITT12, JLH13, KHOLT14, LHY15]. simulations [CSO13]. Sinc [OMS15]. singular [BGH14, BHW13, CT10, DKMW13, Err14, GL15, Hir13, LL15, MDK13, TV10, WLZ17, qWlGjY17]. singularities [BM11, FKS12, He18, Qur19, SSV14, Sid11, Sid12, PDSV15]. singularity [KMPW10]. singularly [FM18, GO16, KP12, XZ10, ZZ14]. Skew [GV17]. slicing [Lin17]. slip [Ste11]. slope [JLQZ18, MP12, QzSZ15]. Small [BZ18, Gre15, BHSP11, BLS13, EW16, JW12, McK11, Zha11b, Zha15]. Small-span [Gre15]. smallest [LL15, Wu10a]. Smooth [CM14, TV10, BBG19, BY12, Bat15, BKMW11, CPBG19, GKS17a, HNUW14, Mus13, Nei15, qWlGjY17, vdBLM10]. Smoothed [Lic19a, Lic19b, GM16a]. smoothing [GKS13, GKS17b]. smoothness [CSS15, HJSZ18, JvSRV14, Zra10b]. smoothness-increasing [JvSRV14]. Sobolev [AN15, Chr18, HZ17, LS10a, LWL19]. Sobolev-type [LWL19]. solubility [FS12]. solution [BCFG17, BJKM11, BCKM18, CCOV14, DFR12, EHR18, GO16, HP12, Ish17, JLZ19, NNZ19, Spi17, ZZJZ18, vdBLM10]. solutions [CZ10, CJ15, CDDM18, GI16, HLJ16, JXR12, JvSRV14, JZ11, Kry13, LMY12, LS17, Li19, MRSC19]. solvable [JW12]. solved [DF14]. solver [GZ13]. solvers [CPPT10, DOZ13]. Solving [BD10a, AHHR15, CG16, CLAT12, DYY14, GG17, HSW11, HR10, HLL16, MDK13, MRW10, NRW17, Seg17, Ste11, TZD15, WX19a, Wu10b, Xu14b, BNMP18]. Some [BL17, Cer11, CT12, HP13, KMF17, KL10, NRV15, BCCM13, CPPT10, GRT04, GRT15, KL13, LMY12, LS10c, NS18, OS10, Pla17, Ula14, Vab14]. Soni [Coh12, Szm13]. source [BCLZ14, BJKM11, CGH^{+}16, PMH18]. space [BMBO13, CCD16, CHH18, CGO10, DL15, ES16, GG17, GLW19, GK11, HS19a, HSW10, HY15, KO12, KSU16, Kru16, LLS17b, MZ16, TZD15, WNK18, Yao13]. space-time [GLW19, KSU16]. spaces [AN15, AGS16, BzCS12, CGN17, CR16, Chr18, DLPW11, Gne12, GH13, GLM19b, HKW17, HZ17, HJHM15, HZ15, JKdR16, Kön17, KPW14, LZ15, LS10a, LWL19, Mou14, RR13, WX13, WX19c, dVG16]. span [CDS10, Gre15]. Sparse [MS12, DFR12, GH13, JX14, MZ16, NS18]. sparsity [GPOS14]. spatial [AL16]. Special [Szm13, Coh12, KL13, MOS66, MDK13]. specified [GRT04, GRT15]. Spectral [ByHl10, CCDE19, CMS11, DGM^{+}16, DJ12b, GMSC^{+}17, GLNP15, HP16, MRSC19, BBB16, CLRR12, CM13, CT10, CH19, CJ13, GÖS18, GSRM12, Hoa16, LMT10, LLX11, SV14, SL17, Wan10a, qWlGjY17, XWZ13, ZqCt16]. spectral-collocation [CT10]. spectrum [Lin17, RV14, Str13]. speed [BAS16]. sphere [BSSW14, CM12, CS13b, NRW17, ZC18]. spheres [BZ18, GLS11, MMV17, Wen13, Sha19]. Spherical [ZC18, MNPW10, MC15, ZQY12]. spheroidal [Wan10a]. spin [RTwaabRS16]. spline [CLR13, JKK^{+}10, Mou14, Sdi19]. splines [CLR13, DPR14, DI15, DiP18]. split [AG16, CLR13, FG17, MP19, YMO13]. Splitting [BC14, BFS18, BCG16, BCCM13, CMTZ16, EFH19, Gal17, GMS12, HYZ14, HKRT11, HLR13, HKK13, RVX19, Vab12, Vab14]. Spoof [Dit14]. square [BR19, CMSC10, Har11, KÖ18, MRW10, PS19, SwaAbLCW11]. square-free [BR19]. squares [BvH11, BHW13, BCL^{+}11, HMT19, HKÖ11, LPSH11, NJZ17, Shi86, Shi19, TdW15]. Squaring [Kar13]. Stability [Akr16, BCFG17, BGP14, DOZ13, QzSZ15, Spi17, AK16, BGM15, BKN15, BGG^{+}17, BDS18, CG11, DHMG11, GGH15, GO18, HK18a, HS11, JLQZ18, Li17, Li19, LSR19, NN16, SZ18, ZD14]. Stabilized [BH18, Bur17, ETX11, ES16, MRSC19]. Stable [Gal17, AH15a, AMR^{+}19, Azi19, CCLX17, FX13, GGRBRG19, LFS15, LCQ17, MN14, NP17, NL19, Wil18]. Stably [Yen11]. stage [COS14, Cer11, DPR14, GM13b]. staggered [BGM15, LS18]. standard [AD16]. Stanley [IKMF17]. Stark [GRT15, GRT04, MSV19]. Stark-type [MSV19]. starting [BLS13]. state [CCQ17, PPTZ13]. states [DZBK16]. Stationarity [Bar19]. stationary [CZ12, CCD16, GHLM18, HX19, KK11, MZ10, Wan10b]. Statistical [Ser17, Wan10b]. statistics [CS13b, ULS12]. steady [CCQ17, DZBK16, KS19]. steady-state [CCQ17]. Stechkin [BS16]. Stefan [DVY15]. Steiner [KÖP15]. Stephan [QS16]. stepping [Ing13]. steps [CNPT10]. Stieltjes [AL17a, JB19, KC11a, KC11b]. stiff [DHMG11]. stiff-relaxation [BLT13]. Stiffly [ET10]. still [STD15]. Stochastic [Sog15, AL16, AZ11, BCG16, BMM18, CS13a, Deb11, GG17, Git13, GT16, GG19, GLW19, HJW18, KPRBT14, KHOLT14, KP14, KW19, NT16, SS15, SS16, Sog16, WMxY17]. Stokes [GMP14, BD15, BKN15, CZ12, COT17, CGHW14, CCQ17, CW10, CGN^{+}11, CC12b, CS13c, CS14, DL13, DE10, DL15, ETX11, EGHL10, Gal17, GHLM18, GLS11, GOS11, GMS12, GL12, GN14, GO18, Ing13, KK11, Kre12, LS18, LMNN18, MZ10, NL19, Ols12, RVX19, SL17, Ste11, TU18, Wil18]. strategy [CD18, HLL16]. streamline [FKS12]. stress [CGG10]. stresses [QSS18]. Strong [HK18a, JD14, KP14, NT16, SW17, BS15a, BGG^{+}17, CCL^{+}19, GGH15, HS11, KCL14, QSS18]. Strong-stability-preserving [HK18a]. strongly [MN14]. structural [BGU19]. Structure [HX19, NKK17, Sut11b, Alf10, BB16, CL16, Han10, JR13]. Structure-preserving [HX19, NKK17]. Structured [DPV19, SWX19]. structures [BLW10, IKRS12]. study [BNMP18, LHY15]. Sturm [GP17, MDK13]. sub [CJLW18, Gau17]. sub-cell [CJLW18]. sub-range [Gau17]. subclass [DOZ13]. Subdiffusion [JLZ19, CLAT12, JLTZ17]. subdivision [ADL11, Pro17, dW11]. subdomain [GH18]. subdomains [Cal16, CQS14]. subgroup [Che16]. subgroups [AGM10, DLRNS18, JKL11a, JKL11b, Kar13]. Subideal [KP11]. subject [BKM18, GLW19, KSW19]. sublattices [KMSwaAbMDS12]. subquadratic [Har14]. subscale [Pet17]. Subsequence [dW11]. subset [AGHS14]. subset-product [AGHS14]. subsonic [BS18, CY18]. subspace [KPY18, MN19b]. subspaces [SY16]. subsystems [KÖP15]. such [Meš14]. sum [AG14, BHSP11, CMSC10, CZ14a, CLB18, KC11a, KB16, NP19, PY14, Tao14, ZC13]. sum-of-proper-divisors [PY14]. sum-set [BHSP11]. summation [Mon10a]. summations [Doo15, LR19a]. summatory [Ram15, Ram19]. Sums [BKK19, BB17, BBC14, BB16, BS16, BKS15, BD10b, COT15, Esp10, Hit17, Mon10a, Shi86, Shi19, TdW15]. sup [BTDG13, Gal19b, GO18, BKN15]. super [KT10, TK12]. super-exponentials [KT10, TK12]. Supercloseness [ETX11]. superconductors [FN11]. Superconvergence [Cao15, CZ16, CFS17, HFC18, CC12a, CH13, CWD14, CQS12, CS13c, CQB14, ETX11, HZ17, SV14, XZ10]. Superconvergent [CCD16, JvSRV14, YPP13, Mus13]. superelliptic [MN19a]. superpositions [LRT13]. supersingular [FK18b]. supplement [BR98]. support [MFRV18]. surface [ALS17b, Bur10, DEMS19, DE13, EE18, KK14, LZ18, NM12]. surfaces [Bar10, BP12a, CD16, Cor10, Els12, GL15, Hei12, LPRY10, LL17, LM15, OS16a, Qur19, SSV14, SSV17, PDSV15]. suspended [RR13]. Swinnerton [BMS16, Ble12, Che16]. Swinnerton-Dyer [BMS16, Ble12, Che16]. switches [LZ18]. Sylvester [DGKS16]. symbol [DGM^{+}16, GMSC^{+}17, MRSC19]. symbols [Wut18]. Symmetric [CHR15, HJSZ18, DYY14, Fuk12, Han15, HS15, Lan11, Li10, QSS18, Ula14, vDE19]. symmetrizable [AW11]. symmetry [CGG10]. symplectic [MMV17]. system [AD13, ABBSM16, BS18, BG11, BHH^{+}12, CY18, CMR10, DYY14, DOZ13, GMP14, KK11, QS16]. systems [AW11, Bar19, BFZ16, BCPR14, CGHW14, CPPT10, COS14, CZ10, CS15b, DLZ11, DM14, DYY14, Gal17, GI16, GPOS14, GV11, HSW11, Hir13, HR10, JgLW13, KÖP15, KPRBT14, KB16, MYS12, MPT18, Ngu16, OS10, Ula14, Vab14, Wan10b, Wu10b]. Syzygies [SSV14]. Szego [CLRR12, Not15, Spa13].

T [Mou14]. T-meshes [Mou14]. Table [BR05, BR11, Coh12, BHLL19, NP12]. Tabulation [RJS12, MJ16]. Tailored [HH13]. Taking [DS14, SwaAbLCW11]. tame [ZX16]. tamely [LRS16]. Tarry [Cal13]. task [ZGFD14]. Tate [Cor10, DS18]. Taylor [LS10b]. techniques [BCL15]. Teichmüller [LGY15]. tension [ALS17b]. tensor [CDFS13, GH13, MS12, PR15, SSV14]. tensors [AKLZ12, AKKL12, FL18]. term [Alf10, BBR12, BMPR16, BJ15, DI15, GKS17a]. terms [AL17b, BJLP19, CGH^{+}16, DHMG11, Mah14, McN15]. test [AG14, Azi19, GOM11, GOMS15, LP18b]. tetrahedra [AADL11]. tetrahedral [Zha11a]. th [CS13a, Har14, WX13, WX19c]. their [AH15a, CSW16, Mil15, Ten10]. theorem [AH15b, ADGR12, GL15, GM16b, HNRW18, KKP10, LZ13, Nak11, PS19, PY14, RGNS18, Sut11a, Tab17, Gra19b, JJK^{+}15, Mel16]. Theorems [Szm13, Coh12, GM16a, MOS66]. theoretic [AHS17]. Theoretical [OMS15, BNMP18]. Theory [YF15, ABBM18, BBC10, BBC14, CHR15, CFS17, DK10, Hak13, HO12, Kem10, KTA18, Mül14, SW13]. theta [BH16b, CC13b, Kuz15, Lab18]. third [CCD16]. third-order [CCD16]. Thomas [AADL11, CZ14b, OWZD18, Sty14]. three [BMPR16, BPT10, CLB18, CC12b, DP15, Kop19a, McN15, Nei15, PS19, RPR12]. three-square [PS19]. Threshold [EE18, HMT19]. Thue [Kim17, Zie10a]. tight [HJSZ18]. tiles [Kir10]. Time [LM18, ALL17, AG16, BO15, BDG^{+}19, BG11, CGHW14, CC12a, CMQ13, CQB14, CL10, CNPT10, CR15, DJL19, ES16, FM18, GOM11, GLW19, GSRM12, HMP13, HS19a, IPZ15, Ing13, JLZ19, JLQZ18, KSU16, Lab18, LHY15, LZ17, LCQ17, Mus18, Ols12, PS16, QS16, WLZS19, ZZJZ18, ZCS^{+}12, Wan10b]. time-dependent [CGHW14, CMQ13, DJL19, FM18, JLZ19, LZ17, PS16, WLZS19]. time-domain [LHY15]. time-fractional [GLW19, Mus18]. time-harmonic [BDG^{+}19, HMP13, LCQ17, ZCS^{+}12]. time-marching [WLZS19]. time-periodic [KSU16]. time-split [AG16]. time-stepping [Ing13]. time-steps [CNPT10]. Time-transformations [LM18]. time/stiff [BLT13]. times [KPW14]. Timoshenko [CCS12]. TNT [CQ14]. Toeplitz [BMM18, TWO18]. tools [DE10]. topological [FR15b, KK14]. Tornheim [BBC14, BB16]. Torsion [BJK11, DLRNS18, DN19, Baa10, CFLTL14, DD19, Fis17, GJLR18, JKL11a, JKL11b, Jeo16, Ler12, PV15, Sut12, Zie10b]. Total [WX19a, BNS15, Sfa13]. totally [HZ14, McK11, Rob15]. totient [Źra19]. trace [Che12, HJHM15, KZ18, McK11]. traces [BDM10]. tracking [CJ15]. Tractability [ITT12, DLPW11, Was14]. tractable [Sdi19]. Trading [IKRS12]. transcendent [NRV15]. transform [JX14, Olv11]. transformation [BHH^{+}12]. transformations [CLRR12, CM13, LM18]. transforms [SWX19, TWO18]. transition [ND19]. transport [BDS18, CGP12, CPSV18a, GZ13, GM10]. transpose [CCOV14]. Treatment [CQT11, DM14]. tree [FH16]. Trefftz [HMP13]. Trefftz-discontinuous [HMP13]. triangle [KS15, LS10a, NV17]. triangular [CMR10, GN14, OS13, SL17, dHAL11]. triangulations [CDHM12, DI15, Kop14, MC12]. trifocal [AO14]. Trigonometric [GLM^{+}19a, DYY14, Peh11]. Trimmed [GK19]. triple [KÖP15]. triples [CLTZ12]. trivial [Pla17, ZX16]. tropical [CKS19]. truncated [Kuz15, Rah14, RGNS18]. Tschirnhaus [SS14]. tuples [ZC13]. Turing [Tru11]. twelve [LP18a, SW17]. twisted [RTwaabRS16]. twists [BBG19]. Two [DPR14, LS10b, NNZ19, ABHV11, AKPZ15, BS15a, BB17, Bis15, BCNS12, BD11, CS18, Cal16, CPV14, CLAT12, CCLX17, Chu12, DKS15, DVY15, DHJ17, HHO17, HKK13, HC11, HC15, JgLW13, Kop19a, KZ18, RSS12, Sad14, Shi86, Shi19, TK12, WNK18, WX19b, YDk12, Zie10b]. two-dimensional [AKPZ15, BCNS12, CLAT12, DKS15, HKK13]. two-factor [BD11]. two-grid [HC11, HC15]. two-level [ABHV11, WX19b]. two-phase [CS18, CPV14, DVY15]. Two-point [LS10b]. Two-scale [NNZ19, JgLW13]. Two-stage [DPR14]. two-way [CCLX17]. type [AKPZ15, AHHR15, BFS18, BC13b, BC16, BH16a, BD10a, DS18, DF14, GH15, GOP16, GO16, Hia16a, KK14, KMF17, KP14, Li15a, LWL19, LX19, MN10, MSV19, WLZ17, Wan19, ZN19, HM14].

ultraweak [FHN19]. unbalanced [CPSV18a]. unbiased [AHS17]. unbounded [ZZJZ18]. Unconditional [MJ16]. Unconditionally [GGRBRG19]. unconstrained [CS15a]. unfitted [GO18, OS16a]. Unified [Voh10, CJ15, FM18]. Uniform [BS18, CY18, CC12a, CQB14, GSRM12, XZ10, Xu14a, ZZ14, BJM17, BDFL17, HFC18, Kop19a, PR15]. Uniform-in-time [CC12a, CQB14, GSRM12]. Uniformly [BFS18, CCL^{+}19, CPSV18b]. unimodular [NP14]. unit [BHM^{+}11, BHM^{+}12a, Fon11, HZ14, He18, JJK^{+}15, Peh11]. unitary [FJS16]. units [BB17, BJK11, HP16, JdRV14]. unity [BTDG13]. Universal [CLTZ12, BLS13, KKP10]. unstructured [CSO13, CM14]. updating [DHYZ17]. Upper [Lou11, BBR12, CW15, Nie15, PR15, Tru12, Tru15, Yao13]. Upwind [HS19a, MSW16]. upwind-biased [MSW16]. Urban’s [LZ13]. use [BDFP15]. Using [KPSY18, Zra10b, ASSW16, AJ14, BBM17, BJM17, BBLP13, Cin15, CQS14, CT19, Dup11, GMZ11, HNRW18, HvdH19, JB19, LNRW18, LR19b, MRW10, Mül14, NRW17, RGB14, SW13, Tab17, Tui16, Wan10a, WLZ17, dHAL11, vdBLM10]. Uzawa [Kre12, TY17].

vacancy [OS13]. valuations [CM19]. value [BHL18, CG11, DKMW13, HW18, LS10b, Spi17, TV10, WKN13]. valued [BzCS12]. Values [FLM14, Lan11, BCST19b, Dus16, LL15, RTwaabRS16]. vanilla [HvdH19]. vanishing [BMR16, CJ13, HJSZ18, IPZ15, ZqCt16]. Variable [CSS15, ADGP13, CLAT12, CC14, CZ14b, LS18, MMV17, ZqCt16]. variable-degree [CC14, CZ14b]. variable-order [CLAT12]. variables [GKS17a]. Variant [PY14, AH15b, CC13a, MYS12]. variants [CHR15]. variate [Was14]. variation [Bar15b, BNS15, Sfa13, WX19a]. Variational [IW13, JLPR15, LM15, Bar15b, BP12b, Gav10b, KPY18, Sog15, Sog16, SX16]. varieties [BMS16, BGH14, Bis15]. variety [AO14, Ler12]. varying [DGS11]. vector [AC18, BCCW16, BCNS12, BS15b, COS14, OWZD18, Sdi19]. vectors [Bör18, BZ18]. velocity [CZ12, CC12b]. Vélu [MS16]. Venant [ABBSM16]. verification [CLO14, OHP14]. Verifying [Sza11]. version [Tab17, qWlGjY17, Zra10a]. versions [GM16b]. vertex [CPV14]. vertex-centered [CPV14]. very [ÁB17]. via [ADGR12, BB17, BLS12, CMSC10, CDTW18, CJLW18, DDLR15, DY12, GLM19b, HK17, Kuz15, LP17b, LFS15, LGY15, NN16, NP17, RGNS18, RJS12]. vibration [DRS11, LMMR19]. Vinogradov [AH15b]. Virtual [dVMR19]. viscoelastic [Xu14a]. viscosity [BGP14, BD15, BMR16, CJ13, IPZ15, NNZ19]. Vlasov [CCL^{+}19, CGP12]. Vogelius [GS19]. volcano [IJ13]. volcanoes [BLS12]. Volterra [BHH^{+}12, CT10, KP14, LX19, Mus13, OMS15, qWlGjY17, ZqCt16]. Volterra-type [LX19]. Volume [CG16, ALS17a, BCG16, BKN15, CPV14, CXZ15, CSO13, EGHL10, EGHL12, GLL12, LMS11, LSXZ12, MS12, WHL19]. volumetric [AZ11]. vorticity [CC12b]. vorticity-velocity-pressure [CC12b]. Vries [BCKX13, KM15].

W [Coh12, Szm13]. Walk [Sha19]. Wang [FJK^{+}15]. Waring [CC13a]. Wasserstein [LPD13]. water [AD16, BC16]. Watson [Chu12]. wave [AH17, AG16, BR18, BLW10, BPT10, CCOV14, CXZ16, CCLX17, CQB14, DY12, FN11, FW11, FX13, GLM^{+}19a, HP10, LZ15, LR19b, LCQ17, LM15, Wan10a, WNK18]. wavelet [CDFS13, HZ13, KSU16, Ste11, ZDL13]. wavelets [BK12, JZ11]. way [CCLX17]. Weak [AL16, Deb11, BFZ16, CGG10, WY14, WW18, WZZZ19]. weak-KAM [BFZ16]. weakly [AHS17, Azi19, BCL^{+}11, CT10, GL15, Lic19b, qWlGjY17]. Weierstrass [SS14]. weight [Not15, Sch10]. weighted [AN15, BHW13, CGO10, Gne12, HZ17, KPW14, LS10a, MC13, NJZ17]. weights [RS19, SV14, Spa13, WHV14]. Well [DZBK16, GM10, LHY15, AG16, BKMW11, BC16, BTDG13, CD15, CPPT10, CMS11, HHT19, BCL15]. Well-balanced [DZBK16, AG16, BC16, CPPT10, BCL15]. well-posed [BTDG13]. Well-posedness [GM10, LHY15]. where [SD10, STD15]. which [McN15]. Whipple [Chu12]. white [GLW19]. Whittaker [Mez11]. whole [GG17, GK11]. whose [BKM18, Sad14]. width [KZ18]. width- [KZ18]. widths [BC17]. Wiener [NH11]. Williams [Zha15]. Willmore [DKS15]. Wilson [CGH14]. Wirtinger [NRV15]. without [Dus16, JLQZ18, QzSZ15, SwaAbLCW11]. Witten [BBC14, BB16]. worst [TY19]. worst-case [TY19]. Writing [JdRV14].

Xin [CJLW18].

Yang [GV17].

Zakharov [BS18, CY18]. Zariski [DFH18]. Zernike [MC15]. Zero [BHSP11, HJHM15, Tru15]. zero-counting [Tru15]. Zero-sum [BHSP11]. zeroes [BT15]. Zeros [ADGP13, GM16c, O’S16b, AG14, ADGR12, BCPR14, Büt15, DLZ11, Dum17, FR15a, Gau17, GG14, MFRV18, Pla17, SZW11, SD11, CLRR12]. zeta [BSO12, BF15, BT15, BB12, Cof14, GM16b, GM16c, Hia11, HO12, Lou11, Pla17, RY15, Tru12, Tru15, Tui19, dR11]. zeta-function [BT15, Tru12]. zeta-functions [Tru15].

References

Acosta:2011:FEA

[AA11] Gabriel Acosta and María G. Armentano. Finite element approximations in a non-Lipschitz domain: Part II. Mathematics of Computation, 80(276): 1949–1978, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                                               tronic). URL http://www.ams.
                                                               org/journals/mcom/2011-80-
                                                               276/S0025-5718-2011-02481-
                                                               6/;   http://www.ams.org/
                                                               journals/mcom/2011-80-276/
                                                               S0025-5718-2011-02481-6/S0025-
                                                               5718-2011-02481-6.pdf;   http:
                                                               //www.ams.org/mathscinet-
                                                               getitem?mr=2813345.

Arnold:2014:FED

[AA14] Douglas N. Arnold and Gerard Awanou. Finite element differential forms on cubical meshes. Mathematics of Computation, 83(288): 1551–1570, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-288/
S0025-5718-2013-02783-4;
http://www.ams.org/journals/
mcom/2014-83-288/S0025-5718-
2013-02783-4/S0025-5718-2013-
02783-4.pdf.

Acosta:2011:EER

[AADL11] Gabriel Acosta, Thomas Apel, Ricardo G. Durán, and Ariel L. Lombardi. Error estimates for Raviart–Thomas interpolation of any order on anisotropic tetrahedra. Mathematics of Computation, 80(273):141– 163, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02406-
8/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02406-8/S0025-
5718-2010-02406-8.pdf.

Alvarez:2017:MLC

[ÁB17] Nicolás Álvarez and Verónica Becher. M. Levin’s construction of absolutely normal numbers with very low discrepancy. Mathematics of Computation, 86(308):2927–2946, 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-
308/S0025-5718-2017-03188-

                                              4;   http://www.ams.org/journals/
                                              mcom/2017-86-308/S0025-5718-
                                              2017-03188-4/S0025-5718-2017-
                                              03188-4.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Alvarez%
                                              2C%20Nicolas;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=368040.

Acosta:2018:RTH

[ABBM18] Gabriel Acosta, Juan Pablo Borthagaray, Oscar Bruno, and Martín Maas. Regularity theory and high order numerical methods for the (1D)-fractional Laplacian. Mathematics of Computation, 87(312):1821– 1857, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-312/
                                              S0025-5718-2017-03276-2;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-312/S0025-5718-
                                              2017-03276-2/S0025-5718-2017-
                                              03276-2.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Acosta%
                                              2C%20Gabriel;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Bruno%
                                              2C%20Oscar;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1171838;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              1193956.

Atighehchi:2017:AFF

[ABBR17] Kevin Atighehchi, Stéphane Ballet, Alexis Bonnecaze, and Robert Rolland. Arithmetic in finite fields based on the Chudnovsky–Chudnovsky multiplication algorithm. Mathematics of Computation, 86 (308):2975–3000, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-308/
S0025-5718-2017-03230-0;
http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03230-0/S0025-5718-2017-
03230-0.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1147020;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
352581;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=361503;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
654144.

Audusse:2016:KEI

[ABBSM16] Emmanuel Audusse, François Bouchut, Marie-Odile Bristeau, and Jacques Sainte-Marie. Kinetic entropy inequality and hydrostatic reconstruction scheme for the Saint-Venant system. Mathematics of Computation, 85(302):2815– 2837, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-302/
S0025-5718-2016-03099-9;
http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-
2016-03099-9/S0025-5718-2016-
03099-9.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Sainte-
Marie%2C%20Jacques;   http:
//www.ams.org/mathscinet/

                                              search/author.html?mrauthid=
                                              314037;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=41730;   http:
                                              //www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              717952.

Abdulle:2012:DGF

[Abd12] Assyr Abdulle. Discontinuous Galerkin finite element heterogeneous multiscale method for elliptic problems with multiple scales. Mathematics of Computation, 81(278):687– 713, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-278/
                                              S0025-5718-2011-02527-5;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-278/S0025-5718-
                                              2011-02527-5/S0025-5718-2011-
                                              02527-5.pdf.

Adhikari:2016:CRN

[ABE^{+}16] S. D. Adhikari, L. Boza, S. Eliahou, J. M. Marín, M. P. Revuelta, and M. I. Sanz. On the n-color Rado number for the equation x_{1} + x_{2} + ⋯ + x_{k} + c = x_{k+1}. Mathematics of Computation, 85(300): 2047–2064, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-300/
                                              S0025-5718-2015-03034-8;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-300/S0025-5718-
                                              2015-03034-8/S0025-5718-2015-
                                              03034-8.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=216209;

http://www.ams.org/mathscinet/
search/author.html?mrauthid=
23140;   http://www.ams.org/
mathscinet/search/author.
html?mrauthid=357631;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
718653;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=772279;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
962033.

Ammari:2010:OAR

[ABF^{+}10] Habib Ammari, Elena Beretta, Elisa Francini, Hyeonbae Kang, and Mikyoung Lim. Optimization algorithm for reconstructing interface changes of a conductivity inclusion from modal measurements. Mathematics of Computation, 79(271): 1757–1777, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-10-02344-6/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
10-02344-6/S0025-5718-10-
02344-6.pdf.

Allendes:2011:TLE

[ABHV11] Alejandro Allendes, Gabriel R.

Barrenechea, Erwin Hernández, and Frédéric Valentin. A twolevel enriched finite element method for a mixed problem. Mathematics of Computation, 80(273):11–41, January 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-

273/S0025-5718-2010-02364-
6/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02364-6/S0025-
5718-2010-02364-6.pdf.

Angot:2018:EEV

[AC18] Philippe Angot and Rima Cheaytou. On the error estimates of the vector penaltyprojection methods: Secondorder scheme. Mathematics of Computation, 87(313):2159– 2187, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-313/
S0025-5718-2017-03309-3;
http://www.ams.org/journals/
mcom/2018-87-313/S0025-5718-
2017-03309-3/S0025-5718-2017-
03309-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Cheaytou%
2C%20Rima;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=328099.

Antonopoulos:2013:EEG

[AD13] D. C. Antonopoulos and V. A. Dougalis. Error estimates for Galerkin approximations of the “classical” Boussinesq system. Mathematics of Computation, 82(282):689– 717, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02663-9;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02663-9/S0025-5718-2012-
02663-9.pdf.

Antonopoulos:2016:EES

[AD16] D. C. Antonopoulos and V. A.

Dougalis. Error estimates for the standard Galerkin-finite element method for the shallow water equations. Mathematics of Computation, 85(299): 1143–1182, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2015-03040-3;
http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2015-03040-3/S0025-5718-2015-
03040-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=59415;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
670004.

Adcock:2011:CAM

[Adc11] Ben Adcock. Convergence acceleration of modified Fourier series in one or more dimensions. Mathematics of Computation, 80(273):225–261, January 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02393-
2/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02393-2/S0025-
5718-2010-02393-2.pdf.

Area:2013:ZCO

[ADGP13] Iván Area, Dimitar K. Dimitrov, Eduardo Godoy, and Vanessa G. Paschoa. Zeros of classical orthogonal polynomials of a discrete variable. Mathematics of Computation, 82(282): 1069–1095, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02646-9;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02646-9/S0025-5718-2012-
02646-9.pdf.

Area:2012:IZJ

[ADGR12] Iván Area, Dimitar K. Dimitrov, Eduardo Godoy, and Fernando R. Rafaeli. Inequalities for zeros of Jacobi polynomials via Obrechkoff’s theorem. Mathematics of Computation, 81(278):991–1004, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2012-81-
278/S0025-5718-2011-02553-
6;   http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02553-6/S0025-5718-2011-
02553-6.pdf.

Amat:2011:NSS

[ADL11] S. Amat, K. Dadourian, and J. Liandrat. On a nonlinear subdivision scheme avoiding Gibbs oscillations and converging towards C^{s} functions with s > 1. Mathematics of Computation, 80(274):959–971, April 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02434-
2/;   http://www.ams.org/
journals/mcom/2011-80-274/

S0025-5718-2010-02434-2/S0025-
5718-2010-02434-2.pdf.

Antonietti:2013:MDE

[AdVV13] Paola F. Antonietti, Lourenco Beirão da Veiga, and Marco Verani. A mimetic discretization of elliptic obstacle problems. Mathematics of Computation, 82(283): 1379–1400, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2013-02670-1;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-
2013-02670-1/S0025-5718-2013-
02670-1.pdf.

Aurada:2017:LIE

[AFF^{+}17] M. Aurada, M. Feischl, T. Führer, M. Karkulik, J. M. Melenk, and D. Praetorius. Local inverse estimates for non-local boundary integral operators. Mathematics of Computation, 86 (308):2651–2686, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-308/
S0025-5718-2017-03175-6;
http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03175-6/S0025-5718-2017-
03175-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1017746;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
613978;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=702616;   http:
//www.ams.org/mathscinet/

                                              search/author.html?mrauthid=
                                              965785;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=965821;   http:
                                              //www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              965857.

Ahmadi:2014:EDT

[AG14] Omran Ahmadi and Robert Granger. An efficient deterministic test for Kloosterman sum zeros. Mathematics of Computation, 83 (285):347–363, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-285/
                                              S0025-5718-2013-02705-6;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-285/S0025-5718-
                                              2013-02705-6/S0025-5718-2013-
                                              02705-6.pdf.

Amadori:2016:EEW

[AG16] Debora Amadori and Laurent Gosse. Error estimates for well-balanced and timesplit schemes on a locally damped wave equation. Mathematics of Computation, 85 (298):601–633, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-298/
                                              S0025-5718-2015-03043-9;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-298/S0025-5718-
                                              2015-03043-9/S0025-5718-2015-
                                              03043-9.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=352024;
                                               http://www.ams.org/mathscinet/

search/author.html?mrauthid=
611045.

Alford:2014:CCN

[AGHS14] W. R. Alford, Jon Grantham, Steven Hayman, and Andrew Shallue. Constructing Carmichael numbers through improved subset-product algorithms. Mathematics of Computation, 83(286):899–915, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2014-83-
286/S0025-5718-2013-02737-
8;   http://www.ams.org/journals/
mcom/2014-83-286/S0025-5718-
2013-02737-8/S0025-5718-2013-
02737-8.pdf.

Ash:2010:CCS

[AGM10] Avner Ash, Paul E. Gunnells, and Mark McConnell. Cohomology of congruence subgroups of SL_{4}(ℤ). III. Mathematics of Computation, 79(271): 1811–1831, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-10-02331-8/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
10-02331-8/S0025-5718-10-
02331-8.pdf.

Ainsworth:2016:DEO

[AGS16] Mark Ainsworth, Johnny Guzmán, and Francisco-Javier Sayas. Discrete extension operators for mixed finite element spaces on locally refined meshes. Mathematics of Computation, 85(302):

2639–2650, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2016-85-
302/S0025-5718-2016-03074-
4;   http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-
2016-03074-4/S0025-5718-2016-
03074-4.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=261514;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
621885;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=775211.

Aistleitner:2014:PDB

[AH14] Christoph Aistleitner and Markus Hofer. Probabilistic discrepancy bound for Monte Carlo point sets. Mathematics of Computation, 83(287): 1373–1381, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-287/
S0025-5718-2013-02773-1;
http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02773-1/S0025-5718-2013-
02773-1.pdf.

Adcock:2015:GSS

[AH15a] Ben Adcock and Anders C. Hansen. Generalized sampling and the stable and accurate reconstruction of piecewise analytic functions from their Fourier coefficients. Mathematics of Computation, 84 (291):237–270, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-291/
S0025-5718-2014-02860-3;
http://www.ams.org/journals/
mcom/2015-84-291/S0025-5718-
2014-02860-3/S0025-5718-2014-
02860-3.pdf.

Akbary:2015:VBV

[AH15b] Amir Akbary and Kyle Hambrook. A variant of the Bombieri–Vinogradov theorem with explicit constants and applications. Mathematics of Computation, 84(294):1901– 1932, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/
2015-84-294/S0025-5718-2014-
02919-0;   http://www.ams.
org/journals/mcom/2015-84-
294/S0025-5718-2014-02919-
0/S0025-5718-2014-02919-0.
pdf.

Abdulle:2017:LOD

[AH17] Assyr Abdulle and Patrick Henning. Localized orthogonal decomposition method for the wave equation with a continuum of scales. Mathematics of Computation, 86 (304):549–587, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03114-2;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03114-2/S0025-5718-2016-
03114-2.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Henning%

2C%20Patrick;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=662986.

Argyros:2015:DCT

[AHHR15] I. K. Argyros, M. A. Hernández, S. Hilout, and N. Romero. Directional Chebyshev-type methods for solving equations. Mathematics of Computation, 84 (292):815–830, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-292/
S0025-5718-2014-02906-2;
http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02906-2/S0025-5718-2014-
02906-2.pdf.

Araya:2017:QUH

[AHS17] Makoto Araya, Masaaki Harada, and Sho Suda. Quasi-unbiased Hadamard matrices and weakly unbiased Hadamard matrices: a coding-theoretic approach. Mathematics of Computation, 86(304):951–984, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03122-1;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03122-1/S0025-5718-2016-
03122-1.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Harada%
2C%20Masaaki;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Suda%
2C%20Sho;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=609178.

Aston:2014:CIM

[AJ14] Philip J. Aston and Oliver Junge. Computing the invariant measure and the Lyapunov exponent for one-dimensional maps using a measure-preserving polynomial basis. Mathematics of Computation, 83(288): 1869–1902, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-288/
S0025-5718-2013-02811-6;
http://www.ams.org/journals/
mcom/2014-83-288/S0025-5718-
2013-02811-6/S0025-5718-2013-
02811-6.pdf.

Ahn:2012:IAN

[AK12] Jeoung-Hwan Ahn and Soun-Hi Kwon. The imaginary abelian number fields of 2-power degrees with ideal class groups of exponent ≤ 2. Mathematics of Computation, 81(277):533– 554, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02509-3;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02509-3/S0025-5718-2011-
02509-3.pdf.

Akrivis:2016:BDF

[AK16] Georgios Akrivis and Emmanuil Katsoprinakis. Backward difference formulae: New multipliers and stability properties for parabolic equations. Mathematics of Computation, 85(301):2195–2216, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-301/
S0025-5718-2015-03055-5;
http://www.ams.org/journals/
mcom/2016-85-301/S0025-5718-
2015-03055-5/S0025-5718-2015-
03055-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Katsoprinakis%
2C%20Emmanuil;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=24080.

Ammari:2012:GPTb

[AKKL12] Habib Ammari, Hyeonbae Kang, Eunjoo Kim, and June-Yub Lee. The generalized polarization tensors for resolved imaging. Part II: Shape and electromagnetic parameters reconstruction of an electromagnetic inclusion from multistatic measurements. Mathematics of Computation, 81(278):839– 860, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02534-2;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02534-2/S0025-5718-2011-
02534-2.pdf.

Ammari:2012:GPTa

[AKLZ12] Habib Ammari, Hyeonbae Kang, Mikyoung Lim, and Habib Zribi. The generalized polarization tensors for resolved imaging. Part I: Shape reconstruction of a conductivity inclusion. Mathematics of Computation, 81(277):367– 386, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02533-0;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02533-0/S0025-5718-2011-
02533-0.pdf.

Antonopoulou:2015:CNF

[AKPZ15] D. C. Antonopoulou, G. D.

Karali, M. Plexousakis, and G. E. Zouraris. Crank–Nicolson finite element discretizations for a two-dimensional linear Schrödinger-type equation posed in a noncylindrical domain. Mathematics of Computation, 84(294): 1571–1598, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-294/
S0025-5718-2014-02900-1;
http://www.ams.org/journals/
mcom/2015-84-294/S0025-5718-
2014-02900-1/S0025-5718-2014-
02900-1.pdf.

Akrivis:2013:IEM

[Akr13] Georgios Akrivis. Implicit– explicit multistep methods for nonlinear parabolic equations. Mathematics of Computation, 82(281):45–68, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2013-82-
281/S0025-5718-2012-02628-
7;   http://www.ams.org/journals/

mcom/2013-82-281/S0025-5718-
2012-02628-7/S0025-5718-2012-
02628-7.pdf.

Akrivis:2016:SPI

[Akr16]  Georgios Akrivis.  Stability

properties of implicit–explicit multistep methods for a class of nonlinear parabolic equations. Mathematics of Computation, 85(301):2217–2229, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-301/
S0025-5718-2015-03070-1;
http://www.ams.org/journals/
mcom/2016-85-301/S0025-5718-
2015-03070-1/S0025-5718-2015-
03070-1.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=24080.

Aanderaa:2018:SGE

[AKR18] Stål Aanderaa, Lars Kristiansen, and Hans Kristian Ruud. Search for good examples of Hall’s conjecture. Mathematics of Computation, 87(314): 2903–2914, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-314/
S0025-5718-2018-03298-7;
http://www.ams.org/journals/
mcom/2018-87-314/S0025-5718-
2018-03298-7/S0025-5718-2018-
03298-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Aanderaa%
2C%20Stal;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Ruud%
2C%20Hans%20Kristian;   https:

//www.ams.org/mathscinet/
search/authors.html?mrauthid=
617768.

Andersson:2016:WCS

[AL16] Adam Andersson and Stig Larsson. Weak convergence for a spatial approximation of the nonlinear stochastic heat equation. Mathematics of Computation, 85(299): 1335–1358, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2015-03016-6;
http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2015-03016-6/S0025-5718-2015-
03016-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Andersson%
2C%20Adam;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=245008.

Adell:2017:FCS

[AL17a] José A. Adell and Alberto Lekuona. Fast computation of the Stieltjes constants. Mathematics of Computation, 86 (307):2479–2492, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-307/
S0025-5718-2017-03176-8;
http://www.ams.org/journals/
mcom/2017-86-307/S0025-5718-
2017-03176-8/S0025-5718-2017-
03176-8.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=340766;
 http://www.ams.org/mathscinet/

                                              search/author.html?mrauthid=
                                              663604.

Arnold:2017:FEE

[AL17b] Douglas N. Arnold and Lizao Li. Finite element exterior calculus with lower-order terms. Mathematics of Computation, 86 (307):2193–2212, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-307/
                                              S0025-5718-2016-03158-0;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-307/S0025-5718-
                                              2016-03158-0/S0025-5718-2016-
                                              03158-0.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=1051998;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              27240.

Alfonsi:2010:HOD

[Alf10] Aurélien Alfonsi. High order discretization schemes for the CIR process: Application to affine term structure and Heston models. Mathematics of Computation, 79(269):209– 237, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-269/
                                              S0025-5718-09-02252-2/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-269/S0025-5718-
                                              09-02252-2/S0025-5718-09-
                                              02252-2.pdf.

Alkauskas:2010:MQM

[Alk10]  Giedrius Alkauskas.  The

Minkowski question mark function: explicit series for the dyadic period function and moments. Mathematics of Computation, 79(269):383–418, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-269/
S0025-5718-09-02263-7/;
http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02263-7/S0025-5718-09-

02263-7.pdf. See addenda and corrigenda [Alk11].

Alkauskas:2011:ACM

[Alk11]  Giedrius Alkauskas.  Ad-

denda and corrigenda to “The Minkowski question mark function: explicit series for the dyadic period function and moments”. Mathematics of Computation, 80(276):2445– 2454, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02479-
8/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02479-8/S0025-
5718-2011-02479-8.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813370.  See

[Alk10].

Akrivis:2017:CMR

[ALL17] Georgios Akrivis, Buyang Li, and Christian Lubich. Combining maximal regularity and energy estimates for time discretizations of quasilinear parabolic equations. Mathematics of Computation, 86(306):

1527–1552, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-306/
                                              S0025-5718-2017-03228-2;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-306/S0025-5718-
                                              2017-03228-2/S0025-5718-2017-
                                              03228-2.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Lubich%
                                              2C%20Christian;   http://
                                              www.ams.org/mathscinet/search/
                                              author.html?mrauthid=24080;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              910552.

Aguillon:2017:CFV

[ALS17a] Nina Aguillon, Frédéric Lagoutière, and Nicolas Seguin. Convergence of finite volume schemes for the coupling between the inviscid Burgers equation and a particle. Mathematics of Computation, 86(303):157– 196, January 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-303/
                                              S0025-5718-2016-03082-3;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-303/S0025-5718-
                                              2016-03082-3/S0025-5718-2016-
                                              03082-3.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=1025886;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              648607;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=696531.

Ambrose:2017:CBI

[ALS17b] David M. Ambrose, Yang Liu, and Michael Siegel. Convergence of a boundary integral method for 3D interfacial Darcy flow with surface tension. Mathematics of Computation, 86 (308):2745–2775, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-308/
S0025-5718-2017-03196-3;
http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03196-3/S0025-5718-2017-
03196-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Siegel%
2C%20Michael;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1009087;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
720777.

Arnold:2011:CCP

[AM11] Andrew Arnold and Michael Monagan. Calculating cyclotomic polynomials. Mathematics of Computation, 80(276): 2359–2379, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02467-
1/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02467-1/S0025-
5718-2011-02467-1.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813365.

Aurentz:2019:FBS

[AMR^{+}19] Jared Aurentz, Thomas Mach, Leonardo Robol, Raf Vandebril, and David S. Watkins. Fast and backward stable computation of eigenvalues and eigenvectors of matrix polynomials. Mathematics of Computation, 88(315): 313–347, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03338-5;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03338-5/S0025-5718-2018-
03338-5.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1012086;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1069123;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=180870;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
720650;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=901224.

Adler:2015:GMA

[AN15] James H. Adler and Victor Nistor. Graded mesh approximation in weighted Sobolev spaces and elliptic equations in 2D. Mathematics of Computation, 84(295): 2191–2220, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-295/
S0025-5718-2015-02934-2;
http://www.ams.org/journals/

mcom/2015-84-295/S0025-5718-
2015-02934-2/S0025-5718-2015-
02934-2.pdf.

Anema:2019:ECM

[Ane19] A. S. I. Anema. Elliptic curves maximal over extensions of finite base fields. Mathematics of Computation, 88(315): 453–465, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03342-7;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03342-7/S0025-5718-2018-
03342-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1051027.

Anghel:2016:SPM

[Ang16]  Catalina V. Anghel.  The

self-power map and collecting all residue classes. Mathematics of Computation, 85 (297):379–399, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-297/
S0025-5718-2015-02978-0;
http://www.ams.org/journals/
mcom/2016-85-297/S0025-5718-
2015-02978-0/S0025-5718-2015-
02978-0.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Anghel%
2C%20Catalina%20V..

Aholt:2014:ITV

[AO14] Chris Aholt and Luke Oeding. The ideal of the trifocal variety. Mathematics of Computation, 83(289):

2553–2574, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-289/
S0025-5718-2014-02842-1;
http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-
2014-02842-1/S0025-5718-2014-
02842-1.pdf.

Apel:2015:FEE

[APR15] Thomas Apel, Johannes Pfefferer, and Arnd Rösch. Finite element error estimates on the boundary with application to optimal control. Mathematics of Computation, 84 (291):33–70, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-291/
S0025-5718-2014-02862-7;
http://www.ams.org/journals/
mcom/2015-84-291/S0025-5718-
2014-02862-7/S0025-5718-2014-
02862-7.pdf.

Argyros:2011:SCA

[Arg11] Ioannis K. Argyros. A semilocal convergence analysis for directional Newton methods. Mathematics of Computation, 80(273): 327–343, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02398-
1/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02398-1/S0025-
5718-2010-02398-1.pdf.

Akbary:2015:APP

[AS15] Amir Akbary and Keilan Scholten. Artin prime producing polynomials. Mathematics of Computation, 84(294): 1861–1882, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-294/
S0025-5718-2014-02902-5;
http://www.ams.org/journals/
mcom/2015-84-294/S0025-5718-
2014-02902-5/S0025-5718-2014-
02902-5.pdf.

Abatzoglou:2016:FDP

[ASSW16] Alexander Abatzoglou, Alice Silverberg, Andrew V. Sutherland, and Angela Wong. A framework for deterministic primality proving using elliptic curves with complex multiplication. Mathematics of Computation, 85(299): 1461–1483, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2015-03001-4;
http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2015-03001-4/S0025-5718-2015-
03001-4.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Abatzoglou%
2C%20Alexander;   http://
www.ams.org/mathscinet/search/
author.html?authorName=Silverberg%
2C%20Alice;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Wong%
2C%20Angela;   http://www.

                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=852273.

Andrejic:2016:SCK

[AT16] Vladica Andrejić and Milos Tatarevic. Searching for a counterexample to Kurepa’s conjecture. Mathematics of Computation, 85(302):3061–3068, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-302/
                                              S0025-5718-2016-03098-7;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-302/S0025-5718-
                                              2016-03098-7/S0025-5718-2016-
                                              03098-7.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Tatarevic%
                                              2C%20Milos;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=789950.

Abdulle:2014:AFE

[AV14] Assyr Abdulle and Gilles Vilmart. Analysis of the finite element heterogeneous multiscale method for quasilinear elliptic homogenization problems. Mathematics of Computation, 83(286):513–536, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-286/
                                              S0025-5718-2013-02758-5;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-286/S0025-5718-
                                              2013-02758-5/S0025-5718-2013-
                                              02758-5.pdf.

Adjerid:2011:DGE

[AW11] Slimane Adjerid and Thomas Weinhart. Discontinuous Galerkin error estimation for linear symmetrizable hyperbolic systems. Mathematics of Computation, 80(275): 1335–1367, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2011-02460-
9/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2011-02460-9/S0025-
5718-2011-02460-9.pdf.

Ariyawansa:2011:CPV

[AZ11] K. A. Ariyawansa and Yuntao Zhu. A class of polynomial volumetric barrier decomposition algorithms for stochastic semidefinite programming. Mathematics of Computation, 80(275): 1639–1661, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2010-02449-
4/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2010-02449-4/S0025-
5718-2010-02449-4.pdf.

Altmann:2018:RKM

[AZ18] R. Altmann and C. Zimmer. Runge–Kutta methods for linear semi-explicit operator differential-algebraic equations. Mathematics of Computation, 87(310):149–174, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03270-1;

http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03270-1/S0025-5718-2017-
03270-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Zimmer%
2C%20C.;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=977251.

Azimzadeh:2019:FST

[Azi19] Parsiad Azimzadeh. A fast and stable test to check if a weakly diagonally dominant matrix is a nonsingular M-matrix. Mathematics of Computation, 88(316):783– 800, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03347-6;
https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03347-6/S0025-5718-2018-
03347-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1097701.

Baaziz:2010:EMC

[Baa10] Houria Baaziz. Equations for the modular curve X_{1}(N) and models of elliptic curves with torsion points. Mathematics of Computation, 79(272):2371– 2386, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02332-X/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02332-X/S0025-5718-10-
02332-X.pdf.

Bras-Amoros:2018:CNS

[BAFG18] Maria Bras-Amorós and Julio Fernández-González. Computation of numerical semigroups by means of seeds. Mathematics of Computation, 87(313):2539– 2550, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-313/
S0025-5718-2018-03292-6;
http://www.ams.org/journals/
mcom/2018-87-313/S0025-5718-
2018-03292-6/S0025-5718-2018-
03292-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Bras-
Amoros%2C%20Maria;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
704910.

Bartels:2010:NAF

[Bar10] Sören Bartels. Numerical analysis of a finite element scheme for the approximation of harmonic maps into surfaces. Mathematics of Computation, 79(271): 1263–1301, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-09-02300-X/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
09-02300-X/S0025-5718-09-
02300-X.pdf.

Barbulescu:2015:SPF

[Bar15a] Razvan Barbulescu. Selecting polynomials for the Function Field Sieve. Mathematics of Computation, 84(296):2987– 3012, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-296/
S0025-5718-2015-02940-8;
http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02940-8/S0025-5718-2015-
02940-8.pdf.

Bartels:2015:ECA

[Bar15b] Sören Bartels. Error control and adaptivity for a variational model problem defined on functions of bounded variation. Mathematics of Computation, 84(293):1217–1240, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2015-84-
293/S0025-5718-2014-02893-
7;   http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02893-7/S0025-5718-2014-
02893-7.pdf.

Bartels:2016:PFA

[Bar16]  Sören Bartels.  Projection-

free approximation of geometrically constrained partial differential equations. Mathematics of Computation, 85(299): 1033–1049, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2015-03008-7;
http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2015-03008-7/S0025-5718-2015-
03008-7.pdf;   http://www.
ams.org/mathscinet/search/

author.html?authorName=Bartels%
2C%20Soren.

Barsukow:2019:SPS

[Bar19]  Wasilij Barsukow.  Station-

arity preserving schemes for multi-dimensional linear systems. Mathematics of Computation, 88(318):1621–1645, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

www.ams.org/journals/mcom/
2019-88-318/S0025-5718-2018-
03394-4;   https://www.ams.
org/journals/mcom/2019-88-
318/S0025-5718-2018-03394-
4/S0025-5718-2018-03394-4.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Barsukow%
2C%20Wasilij.

Bilarev:2016:SCN

[BAS16] Todor Bilarev, Magnus Aspenberg, and Dierk Schleicher. On the speed of convergence of Newton’s method for complex polynomials. Mathematics of Computation, 85 (298):693–705, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-02985-8;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-02985-8/S0025-5718-2015-
02985-8.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Aspenberg%
2C%20Magnus;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Bilarev%
2C%20Todor;   http://www.

                                              ams.org/mathscinet/search/
                                              author.html?authorName=Schleicher%
                                              2C%20Dierk.

Batenkov:2015:CAR

[Bat15]  Dmitry Batenkov.  Com-

plete algebraic reconstruction of piecewise-smooth functions from Fourier data. Mathematics of Computation, 84(295): 2329–2350, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-295/
                                              S0025-5718-2015-02948-2;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-295/S0025-5718-
                                              2015-02948-2/S0025-5718-2015-
                                              02948-2.pdf.

Bayad:2011:FEA

[Bay11]  A. Bayad.  Fourier expan-

sions for Apostol–Bernoulli, Apostol–Euler and Apostol– Genocchi polynomials. Mathematics of Computation, 80(276): 2219–2221, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              276/S0025-5718-2011-02476-
                                              2/;   http://www.ams.org/
                                              journals/mcom/2011-80-276/
                                              S0025-5718-2011-02476-2/S0025-
                                              5718-2011-02476-2.pdf;   http:
                                              //www.ams.org/mathscinet-
                                              getitem?mr=2813356.

Broughan:2012:GLA

[BB12] Kevin A. Broughan and A. Ross Barnett. Gram lines and the average of the real part of the Riemann zeta function. Mathematics of Computation, 81(279): 1669–1679, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2011-02565-2;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2011-02565-2/S0025-5718-2011-
02565-2.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Broughan%
2C%20Kevin%20A;   http://
www.ams.org/mathscinet/search/
authors.html?authorName=Barnett%
2C%20A.%20Ross.

Bailey:2016:CSC

[BB16] D. H. Bailey and J. M. Borwein. Computation and structure of character polylogarithms with applications to character Mordell–Tornheim– Witten sums. Mathematics of Computation, 85(297): 295–324, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-297/
S0025-5718-2015-02974-3;
http://www.ams.org/journals/
mcom/2016-85-297/S0025-5718-
2015-02974-3/S0025-5718-2015-
02974-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Borwein%
2C%20J.%20M;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=29355.

Bennett:2017:STU

[BB17] Michael A. Bennett and Nicolas Billerey. Sums of two S-units via Frey–Hellegouarch curves.

Mathematics of Computation, 86(305):1375–1401, 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2017-86-
                                              305/S0025-5718-2016-03129-
                                              4;   http://www.ams.org/journals/
                                              mcom/2017-86-305/S0025-5718-
                                              2016-03129-4/S0025-5718-2016-
                                              03129-4.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Bennett%
                                              2C%20Michael%20A;   http:
                                              //www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              823614.

Bebendorf:2016:SEH

[BBB16] M. Bebendorf, M. Bollhöfer, and M. Bratsch. On the spectral equivalence of hierarchical matrix preconditioners for elliptic problems. Mathematics of Computation, 85(302):2839– 2861, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-302/
                                              S0025-5718-2016-03086-0;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-302/S0025-5718-
                                              2016-03086-0/S0025-5718-2016-
                                              03086-0.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Bollhofer%
                                              2C%20M;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?authorName=Bratsch%2C%
                                              20M;   http://www.ams.org/
                                              mathscinet/search/author.
                                              html?mrauthid=656638.

Bailey:2010:ATB

[BBC10] D. H. Bailey, J. M. Borwein, and R. E. Crandall. Advances in the theory of box integrals. Mathematics of Computation, 79(271): 1839–1866, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-10-02338-0/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
10-02338-0/S0025-5718-10-
02338-0.pdf.

Bailey:2014:CTE

[BBC14] David H. Bailey, Jonathan M.

Borwein, and Richard E. Crandall. Computation and theory of extended Mordell–Tornheim– Witten sums. Mathematics of Computation, 83(288): 1795–1821, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-288/
S0025-5718-2014-02768-3;
http://www.ams.org/journals/
mcom/2014-83-288/S0025-5718-
2014-02768-3/S0025-5718-2014-
02768-3.pdf.

Baldoni:2011:HIP

[BBD^{+}11] Velleda Baldoni, Nicole Berline, Jesus A. De Loera, Matthias Köppe, and Michèle Vergne. How to integrate a polynomial over a simplex. Mathematics of Computation, 80(273): 297–325, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02378-

6/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02378-6/S0025-
5718-2010-02378-6.pdf.

Badr:2019:TSP

[BBG19] Eslam Badr, Francesc Bars, and Elisa Lorenzo García. On twists of smooth plane curves. Mathematics of Computation, 88(315): 421–438, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03317-8;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03317-8/S0025-5718-2018-
03317-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Bars%
2C%20Francesc;   https://
www.ams.org/mathscinet/search/
authors.html?authorName=Lorenzo%
20Garcia%2C%20Elisa;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
1078860.

Bai:2016:BPG

[BBKZ16] Shi Bai, Cyril Bouvier, Alexander Kruppa, and Paul Zimmermann. Better polynomials for GNFS. Mathematics of Computation, 85 (298):861–873, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-03048-8;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-03048-8/S0025-5718-2015-
03048-8.pdf;   http://www.

ams.org/mathscinet/search/
author.html?authorName=Kruppa%
2C%20Alexander;   http://
www.ams.org/mathscinet/search/
author.html?mrauthid=1073627;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
1075898;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=273776.

Bernstein:2013:EUE

[BBLP13] Daniel J. Bernstein, Peter Birkner, Tanja Lange, and Christiane Peters. ECM using Edwards curves. Mathematics of Computation, 82(282): 1139–1179, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02633-0;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02633-0/S0025-5718-2012-
02633-0.pdf.

Balakrishnan:2017:CIP

[BBM17] Jennifer S. Balakrishnan, Amnon Besser, and J. Steffen Müller. Computing integral points on hyperelliptic curves using quadratic Chabauty. Mathematics of Computation, 86(305):1403–1434, 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-
305/S0025-5718-2016-03130-
0;   http://www.ams.org/journals/
mcom/2017-86-305/S0025-5718-
2016-03130-0/S0025-5718-2016-
03130-0.pdf;   http://www.
ams.org/mathscinet/search/

                                              author.html?mrauthid=364540;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              895560;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=910890.

Berkane:2012:EUB

[BBR12] D. Berkane, O. Bordellès, and O. Ramaré. Explicit upper bounds for the remainder term in the divisor problem. Mathematics of Computation, 81(278): 1025–1051, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-278/
                                              S0025-5718-2011-02535-4;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-278/S0025-5718-
                                              2011-02535-4/S0025-5718-2011-
                                              02535-4.pdf.

Berrone:2019:OAF

[BBSV19] Stefano Berrone, Andrea Bonito, Rob Stevenson, and Marco Verani. An optimal adaptive Fictitious Domain Method. Mathematics of Computation, 88(319): 2101–2134, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-319/
                                              S0025-5718-2019-03414-2;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-319/S0025-5718-
                                              2019-03414-2/S0025-5718-2019-
                                              03414-2.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=310898;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              679290;   https://www.ams.
                                              org/mathscinet/search/authors.

html?mrauthid=704488;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
783728.

Bai:2015:ROP

[BBT15] Shi Bai, Richard P. Brent, and Emmanuel Thomé. Root optimization of polynomials in the number field sieve. Mathematics of Computation, 84 (295):2447–2457, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-295/
S0025-5718-2015-02926-3;
http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-
2015-02926-3/S0025-5718-2015-
02926-3.pdf.

Bao:2013:OEE

[BC13a] Weizhu Bao and Yongyong Cai. Optimal error estimates of finite difference methods for the Gross–Pitaevskii equation with angular momentum rotation. Mathematics of Computation, 82(281):99– 128, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-281/
S0025-5718-2012-02617-2;
http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-
2012-02617-2/S0025-5718-2012-
02617-2.pdf.

Bayad:2013:MIF

[BC13b] A. Bayad and J. Chikhi.

Möbius inversion formulae for Apostol–Bernoulli type polynomials and numbers. Mathematics of Computation, 82 (284):2327–2332, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-284/
S0025-5718-2013-02699-3;
http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02699-3/S0025-5718-2013-
02699-3.pdf.

Bouharguane:2014:SMN

[BC14] Afaf Bouharguane and Rémi Carles. Splitting methods for the nonlocal Fowler equation. Mathematics of Computation, 83(287):1121–1141, 2014. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2014-83-
287/S0025-5718-2013-02757-
3;   http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02757-3/S0025-5718-2013-
02757-3.pdf.

Braun:2015:PLI

[BC15] Oliver Braun and Renaud Coulangeon. Perfect lattices over imaginary quadratic number fields. Mathematics of Computation, 84(293): 1451–1467, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02891-3;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02891-3/S0025-5718-2014-
02891-3.pdf.

Berthon:2016:FWB

[BC16] Christophe Berthon and Christophe Chalons. A fully well-balanced, positive and entropy-satisfying Godunov-type method for the shallow-water equations. Mathematics of Computation, 85 (299):1281–1307, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2015-03045-2;
http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2015-03045-2/S0025-5718-2015-
03045-2.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Chalons%
2C%20Christophe;   http://
www.ams.org/mathscinet/search/
author.html?mrauthid=654277.

Bachmayr:2017:KWL

[BC17] Markus Bachmayr and Albert Cohen. Kolmogorov widths and low-rank approximations of parametric elliptic PDEs. Mathematics of Computation, 86(304):701–724, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03132-4;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03132-4/S0025-5718-2016-
03132-4.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=308419;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
881952.

Blanes:2013:OHO

[BCCM13] S. Blanes, F. Casas, P. Chartier, and A. Murua. Optimized high-order splitting methods for some classes of parabolic equations. Mathematics of Computation, 82(283):1559– 1576, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2012-02657-3;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-
2012-02657-3/S0025-5718-2012-
02657-3.pdf.

Bostan:2016:EAC

[BCCW16] Alin Bostan, Guillaume Chèze, Thomas Cluzeau, and Jacques-Arthur Weil. Efficient algorithms for computing rational first integrals and Darboux polynomials of planar polynomial vector fields. Mathematics of Computation, 85(299): 1393–1425, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2015-03007-5;
http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2015-03007-5/S0025-5718-2015-
03007-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Cheze%
2C%20Guillaume;   http://
www.ams.org/mathscinet/search/
author.html?authorName=Cluzeau%
2C%20Thomas;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Weil%

2C%20Jacques-Arthur;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
725685.

Barucq:2017:SAH

[BCFG17] Hélène Barucq, Théophile Chaumont-Frelet, and Christian Gout. Stability analysis of heterogeneous Helmholtz problems and finite element solution based on propagation media approximation. Mathematics of Computation, 86(307): 2129–2157, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-307/
S0025-5718-2016-03165-8;
http://www.ams.org/journals/
mcom/2017-86-307/S0025-5718-
2016-03165-8/S0025-5718-2016-
03165-8.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Barucq%
2C%20Helene;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Chaumont-
Frelet%2C%20Theophile;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
623513.

Bauzet:2016:CFS

[BCG16] C. Bauzet, J. Charrier, and T. Gallouët. Convergence of flux-splitting finite volume schemes for hyperbolic scalar conservation laws with a multiplicative stochastic perturbation. Mathematics of Computation, 85(302):2777–2813, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-302/
                                              S0025-5718-2016-03084-7;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-302/S0025-5718-
                                              2016-03084-7/S0025-5718-2016-
                                              03084-7.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Bauzet%
                                              2C%20C;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?authorName=Gallouet%
                                              2C%20T;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=968621.

Bringmann:2018:HMR

[BCKM18] Bjoern Bringmann, Daniel Cremers, Felix Krahmer, and Michael Moeller. The homotopy method revisited: Computing solution paths of ℓ_{1}-regularized problems. Mathematics of Computation, 87(313):2343– 2364, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-313/
                                              S0025-5718-2017-03287-7;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-313/S0025-5718-
                                              2017-03287-7/S0025-5718-2017-
                                              03287-7.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Bringmann%
                                              2C%20Bjoern;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=644986;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              845768;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=974311.

Bona:2013:CDG

[BCKX13] J. L. Bona, H. Chen, O. Karakashian, and Y. Xing. Conservative, discontinuous Galerkin–methods for the generalized Korteweg– de Vries equation. Mathematics of Computation, 82(283):1401– 1432, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2013-02661-0;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-
2013-02661-0/S0025-5718-2013-
02661-0.pdf.

Bos:2011:GWA

[BCL^{+}11] L. Bos, J.-P. Calvi, N. Levenberg, A. Sommariva, and M. Vianello. Geometric weakly admissible meshes, discrete least squares approximations and approximate Fekete points. Mathematics of Computation, 80 (275):1623–1638, July 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2011-02442-
7/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2011-02442-7/S0025-
5718-2011-02442-7.pdf.

Boutin:2015:CTN

[BCL15] Benjamin Boutin, Frédéric Coquel, and Philippe G. LeFloch. Coupling techniques for nonlinear hyperbolic equations. IV. Well-balanced schemes for scalar multi-dimensional and multi-component laws. Mathematics of Computation, 84 (294):1663–1702, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-294/
S0025-5718-2015-02933-0;
http://www.ams.org/journals/
mcom/2015-84-294/S0025-5718-
2015-02933-0/S0025-5718-2015-
02933-0.pdf.

Bao:2014:IRS

[BCLZ14] Gang Bao, Shui-Nee Chow, Peijun Li, and Haomin Zhou. An inverse random source problem for the Helmholtz equation. Mathematics of Computation, 83(285):215–233, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-285/
S0025-5718-2013-02730-5;
http://www.ams.org/journals/
mcom/2014-83-285/S0025-5718-
2013-02730-5/S0025-5718-2013-
02730-5.pdf.

Benamou:2016:MCD

[BCM16] Jean-David Benamou, Francis Collino, and Jean-Marie Mirebeau. Monotone and consistent discretization of the Monge– Ampère operator. Mathematics of Computation, 85(302):2743– 2775, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-302/
S0025-5718-2016-03080-X;
http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-

2016-03080-X/S0025-5718-2016-
03080-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Mirebeau%
2C%20Jean-Marie;   http://
www.ams.org/mathscinet/search/
author.html?mrauthid=292227;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
326754.

Brenner:2012:HDD

[BCNS12] S. C. Brenner, J. Cui, Z. Nan, and L.-Y. Sung. Hodge decomposition for divergence-free vector fields and two-dimensional Maxwell’s equations. Mathematics of Computation, 81(278): 643–659, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02540-8;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02540-8/S0025-5718-2011-
02540-8.pdf.

Briquel:2014:FCZ

[BCPR14] Irénée Briquel, Felipe Cucker, Javier Peña, and Vera Roshchina. Fast computation of zeros of polynomial systems with bounded degree under finiteprecision. Mathematics of Computation, 83(287):1279–1317, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/
2014-83-287/S0025-5718-2013-
02765-2;   http://www.ams.
org/journals/mcom/2014-83-
287/S0025-5718-2013-02765-

                                              2/S0025-5718-2013-02765-2.
                                              pdf.

Balakrishnan:2015:AHH

[BÇS15] Jennifer S. Balakrishnan, Mirela Çiperiani, and William Stein. padic heights of Heegner points and Λ-adic regulators. Mathematics of Computation, 84 (292):923–954, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-292/
                                              S0025-5718-2014-02876-7;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-292/S0025-5718-
                                              2014-02876-7/S0025-5718-2014-
                                              02876-7.pdf.

Bokanowski:2016:CDG

[BCS16] Olivier Bokanowski, Yingda Cheng, and Chi-Wang Shu. Convergence of discontinuous Galerkin schemes for front propagation with obstacles. Mathematics of Computation, 85(301):2131–2159, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-301/
                                              S0025-5718-2015-03072-5;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-301/S0025-5718-
                                              2015-03072-5/S0025-5718-2015-
                                              03072-5.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=242268;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              605144;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=811395.

Bailey:2019:ERP

[BCST19a] Geoff Bailey, Stephen D. Cohen, Nicole Sutherland, and Tim Trudgian. Existence results for primitive elements in cubic and quartic extensions of a finite field. Mathematics of Computation, 88(316):931– 947, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03357-9;
https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03357-9/S0025-5718-2018-
03357-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=50360;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
618788;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=909247;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
975175.

Booker:2019:PVQ

[BCST19b] Andrew R. Booker, Stephen D.

Cohen, Nicole Sutherland, and Tim Trudgian. Primitive values of quadratic polynomials in a finite field. Mathematics of Computation, 88(318): 1903–1912, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-318/
S0025-5718-2018-03390-7;
https://www.ams.org/journals/
mcom/2019-88-318/S0025-5718-
2018-03390-7/S0025-5718-2018-

                                              03390-7.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=50360;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              672596;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=909247;   https:
                                              //www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              975175.

Billerey:2010:SFT

[BD10a] Nicolas Billerey and Luis V. Dieulefait. Solving Fermat-type equations x^{5} +y^{5} = dz^{p}. Mathematics of Computation, 79(269): 535–544, January 2010. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2010-79-
                                              269/S0025-5718-09-02294-7/
                                              ;   http://www.ams.org/journals/
                                              mcom/2010-79-269/S0025-5718-
                                              09-02294-7/S0025-5718-09-
                                              02294-7.pdf.

Bremner:2010:ESN

[BD10b] A. Bremner and Jean-Joël Delorme. On equal sums of ninth powers. Mathematics of Computation, 79(269):603– 612, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-269/
                                              S0025-5718-09-02288-1/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-269/S0025-5718-
                                              09-02288-1/S0025-5718-09-
                                              02288-1.pdf.

Brouwer:2011:EGB

[BD11] Andries E. Brouwer and Jan Draisma. Equivariant Gröbner bases and the Gaussian twofactor model. Mathematics of Computation, 80(274): 1123–1133, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02415-
9/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02415-9/S0025-
5718-2010-02415-9.pdf.

Bley:2013:APE

[BD13] Werner Bley and Ruben Debeerst. Algorithmic proof of the epsilon constant conjecture. Mathematics of Computation, 82(284):2363–2387, 2013. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2013-82-
284/S0025-5718-2013-02691-
9;   http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02691-9/S0025-5718-2013-
02691-9.pdf.

Bonito:2015:AFE

[BD15] Andrea Bonito and Denis Devaud. Adaptive finite element methods for the Stokes problem with discontinuous viscosity. Mathematics of Computation, 84(295):2137–2162, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2015-84-
295/S0025-5718-2015-02935-
4;   http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-

                                              2015-02935-4/S0025-5718-2015-
                                              02935-4.pdf.

Brieulle:2019:CIE

[BDD^{+}19] Ludovic Brieulle, Luca De Feo, Javad Doliskani, Jean-Pierre Flori, and Éric Schost. Computing isomorphisms and embeddings of finite fields. Mathematics of Computation, 88 (317):1391–1426, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

                                              www.ams.org/journals/mcom/
                                              2019-88-317/S0025-5718-2018-
                                              03363-4;   https://www.ams.
                                              org/journals/mcom/2019-88-
                                              317/S0025-5718-2018-03363-
                                              4/S0025-5718-2018-03363-4.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Brieulle%
                                              2C%20Ludovic;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1041035;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              672551;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=923705;   https:
                                              //www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              962856.

Buet:2017:PUC

[BDFL17] Christophe Buet, Bruno Després, Emmanuel Franck, and Thomas Leroy. Proof of uniform convergence for a cellcentered AP discretization of the hyperbolic heat equation on general meshes. Mathematics of Computation, 86(305): 1147–1202, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-305/
S0025-5718-2016-03131-2;
http://www.ams.org/journals/
mcom/2017-86-305/S0025-5718-
2016-03131-2/S0025-5718-2016-
03131-2.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1088146;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
293142;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=604962;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
961618.

Bohm:2015:UBP

[BDFP15] Janko Böhm, Wolfram Decker, Claus Fieker, and Gerhard Pfister. The use of bad primes in rational reconstruction. Mathematics of Computation, 84(296): 3013–3027, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-296/
S0025-5718-2015-02951-2;
http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02951-2/S0025-5718-2015-
02951-2.pdf.

Bonazzoli:2019:DDP

[BDG^{+}19] M. Bonazzoli, V. Dolean, I. G.

Graham, E. A. Spence, and P.-H. Tournier. Domain decomposition preconditioning for the high-frequency timeharmonic Maxwell equations with absorption. Mathematics of Computation, 88(320):2559– 2604, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-320/
                                              S0025-5718-2019-03447-6;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-320/S0025-5718-
                                              2019-03447-6/S0025-5718-2019-
                                              03447-6.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Tournier%
                                              2C%20P.-H.;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1192481;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              691157;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=76020;   https:
                                              //www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              858295.

Bernardi:2010:LPT

[BDM10] Christine Bernardi, Monique Dauge, and Yvon Maday. The lifting of polynomial traces revisited. Mathematics of Computation, 79(269): 47–69, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-269/
                                              S0025-5718-09-02259-5/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-269/S0025-5718-
                                              09-02259-5/S0025-5718-09-
                                              02259-5.pdf.

Broersen:2018:SDF

[BDS18] D. Broersen, W. Dahmen, and R. P. Stevenson. On the stability of DPG formulations of transport equations. Mathematics of Computation, 87(311):

1051–1082, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-311/
S0025-5718-2017-03242-7;
http://www.ams.org/journals/
mcom/2018-87-311/S0025-5718-
2017-03242-7/S0025-5718-2017-
03242-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Broersen%
2C%20D.;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=310898;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
54100.

Bruinier:2016:LMB

[BEF16] Jan Hendrik Bruinier, Stephan Ehlen, and Eberhard Freitag. Lattices with many Borcherds products. Mathematics of Computation, 85(300): 1953–1981, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-300/
S0025-5718-2015-03059-2;
http://www.ams.org/journals/
mcom/2016-85-300/S0025-5718-
2015-03059-2/S0025-5718-2015-
03059-2.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Bruinier%
2C%20Jan%20Hendrik;   http:
//www.ams.org/mathscinet/
search/author.html?authorName=
Freitag%2C%20Eberhard;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
918800.

Boito:2016:ICL

[BEG16] P. Boito, Y. Eidelman, and L. Gemignani. Implicit QR for companion-like pencils. Mathematics of Computation, 85(300): 1753–1774, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-300/
                                              S0025-5718-2015-03020-8;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-300/S0025-5718-
                                              2015-03020-8/S0025-5718-2015-
                                              03020-8.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=234370;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              315051;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=840065.

Berkels:2017:PEC

[BER17] Benjamin Berkels, Alexander Effland, and Martin Rumpf. A posteriori error control for the binary Mumford– Shah model. Mathematics of Computation, 86(306): 1769–1791, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-306/
                                              S0025-5718-2016-03138-5;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-306/S0025-5718-
                                              2016-03138-5/S0025-5718-2016-
                                              03138-5.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Berkels%
                                              2C%20Benjamin;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Effland%

2C%20Alexander;   http://
www.ams.org/mathscinet/search/
author.html?mrauthid=604100.

Belabas:2015:CRD

[BF15] Karim Belabas and Eduardo Friedman. Computing the residue of the Dedekind zeta function. Mathematics of Computation, 84(291):357–369, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/
2015-84-291/S0025-5718-2014-
02843-3;   http://www.ams.
org/journals/mcom/2015-84-
291/S0025-5718-2014-02843-
3/S0025-5718-2014-02843-3.
pdf.

Bruin:2018:FQQ

[BF18] Peter Bruin and Andrea Ferraguti. On L-functions of quadratic ℚ-curves. Mathematics of Computation, 87 (310):459–499, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03217-8;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03217-8/S0025-5718-2017-
03217-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Bruin%
2C%20Peter;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1156160.

Baumstark:2018:UAE

[BFS18] Simon Baumstark, Erwan Faou, and Katharina Schratz. Uniformly accurate exponentialtype integrators for Klein– Gordon equations with asymptotic convergence to the classical NLS splitting. Mathematics of Computation, 87(311): 1227–1254, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-311/
S0025-5718-2017-03263-4;
http://www.ams.org/journals/
mcom/2018-87-311/S0025-5718-
2017-03263-4/S0025-5718-2017-
03263-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Baumstark%
2C%20Simon;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Schratz%
2C%20Katharina;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=656335.

Bokanowski:2010:EEN

[BFZ10] Olivier Bokanowski, Nicolas Forcadel, and Hasnaa Zidani. L^{1}-error estimates for numerical approximations of Hamilton– Jacobi–Bellman equations in dimension 1. Mathematics of Computation, 79(271):1395– 1426, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-10-02311-2/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
10-02311-2/S0025-5718-10-
02311-2.pdf.

Bouillard:2016:FWK

[BFZ16] Anne Bouillard, Erwan Faou, and Maxime Zavidovique. Fast weak-KAM integrators for separable Hamiltonian systems. Mathematics of Computation, 85(297):85–117, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-297/
S0025-5718-2015-02976-7;
http://www.ams.org/journals/
mcom/2016-85-297/S0025-5718-
2015-02976-7/S0025-5718-2015-
02976-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Bouillard%
2C%20Anne;   http://www.ams.
org/mathscinet/search/author.
html?authorName=Faou%2C%20Erwan;
http://www.ams.org/mathscinet/
search/author.html?authorName=
Zavidovique%2C%20Maxime.

Bonito:2011:AEP

[BG11] Andrea Bonito and Jean-Luc Guermond. Approximation of the eigenvalue problem for the time harmonic Maxwell system by continuous Lagrange finite elements. Mathematics of Computation, 80(276): 1887–1910, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02464-
6/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02464-6/S0025-
5718-2011-02464-6.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813343.

Bresten:2017:ESS

[BGG^{+}17] Christopher Bresten, Sigal Gottlieb, Zachary Grant, Daniel Higgs, David I. Ketcheson, and Adrian Németh. Explicit strong stability preserving multistep Runge–Kutta methods. Mathematics of Computation, 86 (304):747–769, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03115-4;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03115-4/S0025-5718-2016-
03115-4.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Bresten%
2C%20Christopher;   http:
//www.ams.org/mathscinet/
search/author.html?authorName=
Grant%2C%20Zachary;   http:
//www.ams.org/mathscinet/
search/author.html?authorName=
Ketcheson%2C%20David%20I;
 http://www.ams.org/mathscinet/
search/author.html?authorName=
Nemeth%2C%20Adrian;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
358958;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=906696.

Boffi:2017:OCA

[BGGG17] Daniele Boffi, Dietmar Gallistl, Francesca Gardini, and Lucia Gastaldi. Optimal convergence of adaptive FEM for eigenvalue clusters in mixed form. Mathematics of Computation, 86 (307):2213–2237, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-307/

S0025-5718-2017-03212-9;
http://www.ams.org/journals/
mcom/2017-86-307/S0025-5718-
2017-03212-9/S0025-5718-2017-
03212-9.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1020312;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
348743;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=71735;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
747613.

Bank:2014:PSR

[BGH14] Bernd Bank, Marc Giusti, and Joos Heintz. Point searching in real singular complete intersection varieties: algorithms of intrinsic complexity. Mathematics of Computation, 83 (286):873–897, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-286/
S0025-5718-2013-02766-4;
http://www.ams.org/journals/
mcom/2014-83-286/S0025-5718-
2013-02766-4/S0025-5718-2013-
02766-4.pdf.

Berthelin:2015:KSS

[BGM15] Florent Berthelin, Thierry Goudon, and Sebastian Minjeaud. Kinetic schemes on staggered grids for barotropic Euler models: entropy–stability analysis. Mathematics of Computation, 84(295):2221–2262, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2015-84-

                                              295/S0025-5718-2015-02957-
                                              3;   http://www.ams.org/journals/
                                              mcom/2015-84-295/S0025-5718-
                                              2015-02957-3/S0025-5718-2015-
                                              02957-3.pdf.

Barrett:2017:FEA

[BGN17] John W. Barrett, Harald Garcke, and Robert Nürnberg. Finite element approximation for the dynamics of asymmetric fluidic biomembranes. Mathematics of Computation, 86 (305):1037–1069, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-305/
                                              S0025-5718-2016-03162-2;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-305/S0025-5718-
                                              2016-03162-2/S0025-5718-2016-
                                              03162-2.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=31635;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              352477;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=698349.

Brenner:2011:PMF

[BGNyS11] Susanne C. Brenner, Thirupathi Gudi, Michael Neilan, and Li yeng Sung. C^{0} penalty methods for the fully nonlinear Monge–Ampère equation. Mathematics of Computation, 80(276):1979–1995, October 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              276/S0025-5718-2011-02487-
                                              7/;   http://www.ams.org/

journals/mcom/2011-80-276/
S0025-5718-2011-02487-7/S0025-
5718-2011-02487-7.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813346.

Bonito:2014:SAE

[BGP14] Andrea Bonito, Jean-Luc Guermond, and Bojan Popov. Stability analysis of explicit entropy viscosity methods for non-linear scalar conservation equations. Mathematics of Computation, 83(287): 1039–1062, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-287/
S0025-5718-2013-02771-8;
http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02771-8/S0025-5718-2013-
02771-8.pdf.

Bennett:2019:CEC

[BGR19] Michael A. Bennett, Adela Gherga, and Andrew Rechnitzer. Computing elliptic curves over ℚ. Mathematics of Computation, 88(317):1341– 1390, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-317/
S0025-5718-2018-03370-1;
https://www.ams.org/journals/
mcom/2019-88-317/S0025-5718-
2018-03370-1/S0025-5718-2018-
03370-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Gherga%
2C%20Adela;   https://www.
ams.org/mathscinet/search/

                                              authors.html?mrauthid=339361;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              626723.

Bean:2019:ADS

[BGU19] Christian Bean, Bjarki Gudmundsson, and Henning Ulfarsson. Automatic discovery of structural rules of permutation classes. Mathematics of Computation, 88(318):1967– 1990, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-318/
                                              S0025-5718-2018-03386-5;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-318/S0025-5718-
                                              2018-03386-5/S0025-5718-2018-
                                              03386-5.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1146492;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              1168490;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=848375.

Buhler:2011:IPM

[BH11] J. P. Buhler and D. Harvey. Irregular primes to 163 million. Mathematics of Computation, 80(276):2435– 2444, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              276/S0025-5718-2011-02461-
                                              0/;   http://www.ams.org/
                                              journals/mcom/2011-80-276/
                                              S0025-5718-2011-02461-0/S0025-
                                              5718-2011-02461-0.pdf;   http:

//www.ams.org/mathscinet-
getitem?mr=2813369.

Bremner:2013:FIA

[BH13] Murray R. Bremner and Jiaxiong Hu. Fundamental invariants for the action of SL_{3}(ℂ) × SL_{3}(ℂ) × SL_{3}(ℂ) on 3 × 3 × 3 arrays. Mathematics of Computation, 82(284): 2421–2438, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-284/
S0025-5718-2013-02706-8;
http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02706-8/S0025-5718-2013-
02706-8.pdf.

Bertok:2016:HTP

[BH16a] Csanád Bertók and Lajos Hajdu. A Hasse-type principle for exponential Diophantine equations and its applications. Mathematics of Computation, 85 (298):849–860, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-03002-6;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-03002-6/S0025-5718-2015-
03002-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Bertok%
2C%20Csanad;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Hajdu%
2C%20Lajos.

Broker:2016:FCS

[BH16b] Reinier Bröker and Jeff Hoffstein. Fourier coefficients of sextic theta series. Mathematics of Computation, 85(300): 1901–1927, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-300/
                                              S0025-5718-2015-03044-0;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-300/S0025-5718-
                                              2015-03044-0/S0025-5718-2015-
                                              03044-0.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=759393;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              87085.

Burman:2018:SNF

[BH18] Erik Burman and Peter Hansbo. Stabilized nonconforming finite element methods for data assimilation in incompressible flows. Mathematics of Computation, 87(311):1029– 1050, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-311/
                                              S0025-5718-2017-03255-5;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-311/S0025-5718-
                                              2017-03255-5/S0025-5718-2017-
                                              03255-5.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=269716;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              602430.

Brezinski:2012:MAS

[BHH^{+}12] Claude Brezinski, Yi He, Xing-Biao Hu, Michela Redivo-Zaglia, and Jian-Qing Sun. Multistep ϵ–algorithm, Shanks’ transformation, and the Lotka– Volterra system by Hirota’s method. Mathematics of Computation, 81(279):1527– 1549, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2011-02554-8;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2011-02554-8/S0025-5718-2011-
02554-8.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Brezinski%
2C%20Claude;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=He%
2C%20Yi;   http://www.ams.
org/mathscinet/search/authors.
html?authorName=Hu%2C%20Xing-
Biao;   http://www.ams.org/
mathscinet/search/authors.
html?authorName=Redivo-Zaglia%
2C%20Michela;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Sun%
2C%20Jian-Qing.

Burman:2018:CFE

[BHL18] Erik Burman, Peter Hansbo, and Mats G. Larson. A cut finite element method with boundary value correction. Mathematics of Computation, 87 (310):633–657, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/

                                              journals/mcom/2018-87-310/
                                              S0025-5718-2017-03240-3;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-310/S0025-5718-
                                              2017-03240-3/S0025-5718-2017-
                                              03240-3.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=269716;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              602430;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=648688.

Breuer:2019:CMC

[BHLL19] Thomas Breuer, Gerhard Hiss, Frank Lübeck, and Klaus Lux. The completion of the 3modular character table of the Chevalley group F_{4}(2) and its covering group. Mathematics of Computation, 88(320):3023– 3040, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-320/
                                              S0025-5718-2019-03435-X;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-320/S0025-5718-
                                              2019-03435-X/S0025-5718-2019-
                                              03435-X.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=117185;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              350539;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=362381;   https:
                                              //www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              86475.

Beers:2011:FCU

[BHM^{+}11] J. Beers, D. Henshaw, C. K. McCall, S. B. Mulay, and M. Spindler. Fundamentality of a cubic unit u for ℤ[u]. Mathematics of Computation, 80(273): 563–578, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02383-
X/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02383-X/S0025-
5718-2010-02383-X.pdf. See

corrigenda and addenda [BHM^{+}12a].

Beers:2012:CAF

[BHM^{+}12a] J. Beers, D. Henshaw, C. K.

McCall, S. B. Mulay, and M. Spindler. Corrigenda and addenda to “Fundamentality of a cubic unit u for ℤ[u]”. Mathematics of Computation, 81(280):2383– 2387, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-280/
S0025-5718-2012-02501-4;
http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-
2012-02501-4/S0025-5718-2012-

02501-4.pdf. See [BHM^{+}11].

Bu:2012:SIK

[BHM12b] Sunyoung Bu, Jingfang Huang, and Michael L. Minion. Semiimplicit Krylov deferred correction methods for differential algebraic equations. Mathematics of Computation, 81(280):2127– 2157, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/

                                              journals/mcom/2012-81-280/
                                              S0025-5718-2012-02564-6;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-280/S0025-5718-
                                              2012-02564-6/S0025-5718-2012-
                                              02564-6.pdf.

Becher:2015:CAN

[BHS15] Verónica Becher, Pablo Ariel Heiber, and Theodore A. Slaman. A computable absolutely normal Liouville number. Mathematics of Computation, 84 (296):2939–2952, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2015-84-
                                              296/S0025-5718-2015-02964-
                                              0;   http://www.ams.org/journals/
                                              mcom/2015-84-296/S0025-5718-
                                              2015-02964-0/S0025-5718-2015-
                                              02964-0.pdf.

Bhowmik:2011:ZSF

[BHSP11] Gautami Bhowmik, Immanuel Halupczok, and Jan-Christoph Schlage-Puchta. Zero-sum free sets with small sum-set. Mathematics of Computation, 80(276): 2253–2258, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              276/S0025-5718-2011-02385-
                                              9/;   http://www.ams.org/
                                              journals/mcom/2011-80-276/
                                              S0025-5718-2011-02385-9/S0025-
                                              5718-2011-02385-9.pdf;   http:
                                              //www.ams.org/mathscinet-
                                              getitem?mr=2813358.

Bidwell:2013:WLS

[BHW13] S. Bidwell, M. E. Hassell, and C. R. Westphal. A weighted least squares finite element method for elliptic problems with degenerate and singular coefficients. Mathematics of Computation, 82(282):673– 688, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02659-7;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02659-7/S0025-5718-2012-
02659-7.pdf.

Berrizbeitia:2010:GME

[BI10] Pedro Berrizbeitia and Boris Iskra. Gaussian Mersenne and Eisenstein Mersenne primes. Mathematics of Computation, 79(271):1779–1791, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2010-79-
271/S0025-5718-10-02324-0/
;   http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
10-02324-0/S0025-5718-10-
02324-0.pdf.

Beckmann:2019:EEC

[BI19] Matthias Beckmann and Armin Iske. Error estimates and convergence rates for filtered back projection. Mathematics of Computation, 88(316):801– 835, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03343-9;

                                              https://www.ams.org/journals/
                                              mcom/2019-88-316/S0025-5718-
                                              2018-03343-9/S0025-5718-2018-
                                              03343-9.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Beckmann%
                                              2C%20Matthias;   https://
                                              www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=600018.

Biasse:2014:AIC

[Bia14]  Jean-François Biasse.  An

L(1/3) algorithm for ideal class group and regulator computation in certain number fields. Mathematics of Computation, 83(288):2005–2031, 2014. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2014-83-
                                              288/S0025-5718-2014-02651-
                                              3;   http://www.ams.org/journals/
                                              mcom/2014-83-288/S0025-5718-
                                              2014-02651-3/S0025-5718-2014-
                                              02651-3.pdf.

Bisson:2015:CER

[Bis15] Gaetan Bisson. Computing endomorphism rings of abelian varieties of dimension two. Mathematics of Computation, 84 (294):1977–1989, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-294/
                                              S0025-5718-2015-02938-X;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-294/S0025-5718-
                                              2015-02938-X/S0025-5718-2015-
                                              02938-X.pdf.

Bley:2011:CGF

[BJ11] Werner Bley and Henri Johnston. Computing generators of free modules over orders in group algebras II. Mathematics of Computation, 80(276): 2411–2434, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02488-
9/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02488-9/S0025-
5718-2011-02488-9.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813368.

Brent:2015:BET

[BJ15] Richard P. Brent and Fredrik Johansson. A bound for the error term in the Brent– McMillan algorithm. Mathematics of Computation, 84 (295):2351–2359, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-295/
S0025-5718-2015-02931-7;
http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-
2015-02931-7/S0025-5718-2015-
02931-7.pdf.

Bovdi:2011:TUI

[BJK11] V. A. Bovdi, E. Jespers, and A. B. Konovalov. Torsion units in integral group rings of Janko simple groups. Mathematics of Computation, 80(273): 593–615, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02376-

                                              2/;   http://www.ams.org/
                                              journals/mcom/2011-80-273/
                                              S0025-5718-2010-02376-2/S0025-
                                              5718-2010-02376-2.pdf.

Berikelashvili:2011:FDS

[BJKM11] G. Berikelashvili, O. Jokhadze, S. Kharibegashvili, and B. Midodashvili. Finite difference solution of a nonlinear Klein– Gordon equation with an external source. Mathematics of Computation, 80(274): 847–862, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              274/S0025-5718-2010-02416-
                                              0/;   http://www.ams.org/
                                              journals/mcom/2011-80-274/
                                              S0025-5718-2010-02416-0/S0025-
                                              5718-2010-02416-0.pdf.

Bujanda:2019:CEC

[BJLP19] Blanca Bujanda, José, L. López, and Pedro J. Pagola. Convergent expansions of the confluent hypergeometric functions in terms of elementary functions. Mathematics of Computation, 88(318):1773– 1789, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-318/
                                              S0025-5718-2018-03389-0;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-318/S0025-5718-
                                              2018-03389-0/S0025-5718-2018-
                                              03389-0.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Lopez%
                                              2C%20Jose%20L.;   https://

www.ams.org/mathscinet/search/
authors.html?mrauthid=636519;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
806866.

Benoit:2017:RUA

[BJM17] Alexandre Benoit, Mioara Joldeş, and Marc Mezzarobba. Rigorous uniform approximation of D-finite functions using Chebyshev expansions. Mathematics of Computation, 86 (305):1303–1341, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-305/
S0025-5718-2016-03135-X;
http://www.ams.org/journals/
mcom/2017-86-305/S0025-5718-
2016-03135-X/S0025-5718-2016-
03135-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Benoit%
2C%20Alexandre;   http://
www.ams.org/mathscinet/search/
author.html?authorName=Mezzarobba%
2C%20Marc;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=931466.

Becache:2018:APM

[BJV18] Éliane Bécache, Patrick Joly, and Valentin Vinoles. On the analysis of perfectly matched layers for a class of dispersive media and application to negative index metamaterials. Mathematics of Computation, 87(314):2775– 2810, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-314/

                                              S0025-5718-2018-03307-5;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-314/S0025-5718-
                                              2018-03307-5/S0025-5718-2018-
                                              03307-5.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Becache%
                                              2C%20Eliane;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1135885;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              234723.

Blanchard:2012:MFC

[BK12] Jeffrey D. Blanchard and Ilya A. Krishtal. Matricial filters and crystallographic composite dilation wavelets. Mathematics of Computation, 81(278): 905–922, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-278/
                                              S0025-5718-2011-02518-4;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-278/S0025-5718-
                                              2011-02518-4/S0025-5718-2011-
                                              02518-4.pdf.

Bayless:2019:SPS

[BKK19] Jonathan Bayless, Paul Kinlaw, and Dominic Klyve. Sums over primitive sets with a fixed number of prime factors. Mathematics of Computation, 88(320):3063–3077, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-320/
                                              S0025-5718-2019-03416-6;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-320/S0025-5718-

2019-03416-6/S0025-5718-2019-
03416-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=769072;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
776121;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=902693.

Birgin:2018:EIR

[BKM18] E. G. Birgin, N. Krejić, and J. M. Martínez. On the employment of inexact restoration for the minimization of functions whose evaluation is subject to errors. Mathematics of Computation, 87(311): 1307–1326, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-311/
S0025-5718-2017-03246-4;
http://www.ams.org/journals/
mcom/2018-87-311/S0025-5718-
2017-03246-4/S0025-5718-2017-
03246-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Birgin%
2C%20E.%20G.;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Krejic%
2C%20N.;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=120570.

Beckermann:2011:HWD

[BKMW11] Bernhard Beckermann, Valeriy Kalyagin, Ana C. Matos, and Franck Wielonsky. How well does the Hermite–Padé approximation smooth the Gibbs phenomenon? Mathematics of Computation, 80(274): 931–958, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              274/S0025-5718-2010-02411-
                                              1/;   http://www.ams.org/
                                              journals/mcom/2011-80-274/
                                              S0025-5718-2010-02411-1/S0025-
                                              5718-2010-02411-1.pdf.

Boyer:2015:ISS

[BKN15] Franck Boyer, Stella Krell, and Flore Nabet. Inf-Sup stability of the discrete duality finite volume method for the 2D Stokes problem. Mathematics of Computation, 84(296):2705– 2742, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-296/
                                              S0025-5718-2015-02956-1;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-296/S0025-5718-
                                              2015-02956-1/S0025-5718-2015-
                                              02956-1.pdf.

Bourgain:2015:CSD

[BKS15] Jean Bourgain, Sergei V. Konyagin, and Igor E. Shparlinski. Character sums and deterministic polynomial root finding in finite fields. Mathematics of Computation, 84(296):2969– 2977, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-296/
                                              S0025-5718-2015-02946-9;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-296/S0025-5718-
                                              2015-02946-9/S0025-5718-2015-
                                              02946-9.pdf.

Berthomieu:2012:RBP

[BL12] Jérémy Berthomieu and Grégoire Lecerf. Reduction of bivariate polynomials from convexdense to dense, with application to factorizations. Mathematics of Computation, 81(279): 1799–1821, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2011-02562-7;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2011-02562-7/S0025-5718-2011-
02562-7.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Berthomieu%
2C%20Jeremy;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Lecerf%
2C%20Gregoire.

Broker:2014:EIF

[BL14] Reinier Bröker and Kristin Lauter. Evaluating Igusa functions. Mathematics of Computation, 83(290):2977–2999, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2014-83-
290/S0025-5718-2014-02816-
0;   http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02816-0/S0025-5718-2014-
02816-0.pdf.

Barbulescu:2017:SMR

[BL17] Razvan Barbulescu and Armand Lachand. Some mathematical remarks on the polynomial selection in NFS. Mathematics of Computation, 86(303):397– 418, January 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-303/
S0025-5718-2016-03112-9;
http://www.ams.org/journals/
mcom/2017-86-303/S0025-5718-
2016-03112-9/S0025-5718-2016-
03112-9.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Barbulescu%
2C%20Razvan;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1110511.

Bley:2012:NEE

[Ble12] Werner Bley. Numerical evidence for the equivariant Birch and Swinnerton-Dyer conjecture (Part II). Mathematics of Computation, 81(279):1681– 1705, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2012-02572-5;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2012-02572-5/S0025-5718-2012-
02572-5.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Bley%
2C%20Werner.

Broker:2012:MPI

[BLS12] Reinier Bröker, Kristin Lauter, and Andrew V. Sutherland. Modular polynomials via isogeny volcanoes. Mathematics of Computation, 81(278):1201– 1231, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02508-1;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02508-1/S0025-5718-2011-
02508-1.pdf.

Bollobas:2013:SPU

[BLS13] Béla Bollobás, Malte Lackmann, and Dierk Schleicher. A small probabilistic universal set of starting points for finding roots of complex polynomials by Newton’s method. Mathematics of Computation, 82(281):443– 457, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-281/
S0025-5718-2012-02640-8;
http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-
2012-02640-8/S0025-5718-2012-
02640-8.pdf.

Berthon:2013:LTS

[BLT13] Christophe Berthon, Philippe G.

LeFloch, and Rodolphe Turpault. Late-time/stiff-relaxation asymptotic-preserving approximations of hyperbolic equations. Mathematics of Computation, 82(282):831–860, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2013-82-
282/S0025-5718-2012-02666-
4;   http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02666-4/S0025-5718-2012-
02666-4.pdf.

Bao:2010:AEE

[BLW10] Gang Bao, Peijun Li, and Haijun Wu. An adaptive edge element method with perfectly matched absorbing layers for wave scattering by biperiodic structures. Mathematics of Computation, 79(269): 1–34, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-269/
                                              S0025-5718-09-02257-1/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-269/S0025-5718-
                                              09-02257-1/S0025-5718-09-
                                              02257-1.pdf.

Bartels:2011:QOR

[BM11] Sören Bartels and Rüdiger Müller. Quasi-optimal and robust a posteriori error estimates in L^{∞}(L^{2}) for the approximation of Allen–Cahn equations past singularities. Mathematics of Computation, 80(274): 761–780, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              274/S0025-5718-2010-02444-
                                              5/;   http://www.ams.org/
                                              journals/mcom/2011-80-274/
                                              S0025-5718-2010-02444-5/S0025-
                                              5718-2010-02444-5.pdf.

Blackburn:2012:CRS

[BM12] Simon R. Blackburn and James F. McKee. Constructing k-radius sequences. Mathematics of Computation, 81(280): 2439–2459, October 2012. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2012-81-
280/S0025-5718-2011-02510-
X;   http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-
2011-02510-X/S0025-5718-2011-
02510-X.pdf.

Burger:2013:AIS

[BMBO13] Martin Burger, Michael Möller, Martin Benning, and Stanley Osher. An adaptive inverse scale space method for compressed sensing. Mathematics of Computation, 82(281):269– 299, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-281/
S0025-5718-2012-02599-3;
http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-
2012-02599-3/S0025-5718-2012-
02599-3.pdf.

Bini:2018:SIQ

[BMM18] Dario A. Bini, Stefano Massei, and Beatrice Meini. Semiinfinite quasi-Toeplitz matrices with applications to QBD stochastic processes. Mathematics of Computation, 87(314): 2811–2830, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-314/
S0025-5718-2018-03301-4;
http://www.ams.org/journals/
mcom/2018-87-314/S0025-5718-
2018-03301-4/S0025-5718-2018-
03301-4.pdf;   https://www.
ams.org/mathscinet/search/

                                              authors.html?authorName=Massei%
                                              2C%20Stefano;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=367501;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              37060.

Bini:2010:EMG

[BMP10] Dario A. Bini, Beatrice Meini, and Federico Poloni. An effective matrix geometric mean satisfying the Ando–Li–Mathias properties. Mathematics of Computation, 79(269):437–452, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-269/
                                              S0025-5718-09-02261-3/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-269/S0025-5718-
                                              09-02261-3/S0025-5718-09-
                                              02261-3.pdf.

Bracciali:2016:COF

[BMPR16] C. F. Bracciali, J. H. McCabe, T. E. Pérez, and A. Sri Ranga. A class of orthogonal functions given by a three term recurrence formula. Mathematics of Computation, 85(300): 1837–1859, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-300/
                                              S0025-5718-2015-03041-5;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-300/S0025-5718-
                                              2015-03041-5/S0025-5718-2015-
                                              03041-5.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Bracciali%
                                              2C%20C%20F;   http://www.

ams.org/mathscinet/search/
author.html?mrauthid=122065;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
238837;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=321333.

Bugariu:2016:ACB

[BMR16] Ioan Florin Bugariu, Sorin Micu, and Ionel Rovenţa. Approximation of the controls for the beam equation with vanishing viscosity. Mathematics of Computation, 85(301):2259– 2303, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-301/
S0025-5718-2016-03064-1;
http://www.ams.org/journals/
mcom/2016-85-301/S0025-5718-
2016-03064-1/S0025-5718-2016-
03064-1.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Micu%
2C%20Sorin;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Roventa%
2C%20Ionel;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1056928.

Balakrishnan:2016:AAC

[BMS16] Jennifer S. Balakrishnan, J. Steffen Müller, and William A. Stein. A p-adic analogue of the conjecture of Birch and Swinnerton-Dyer for modular abelian varieties. Mathematics of Computation, 85(298): 983–1016, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/

                                              journals/mcom/2016-85-298/
                                              S0025-5718-2015-03029-4;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-298/S0025-5718-
                                              2015-03029-4/S0025-5718-2015-
                                              03029-4.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=679996;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              895560;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=910890.

Bonnaillie-NoAogl:2015:ACL

[BNDHV15] Virginie Bonnaillie-Noël, Marc Dambrine, Frédéric Hérau, and Grégory Vial. Artificial conditions for the linear elasticity equations. Mathematics of Computation, 84(294): 1599–1632, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-294/
                                              S0025-5718-2014-02901-3;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-294/S0025-5718-
                                              2014-02901-3/S0025-5718-2014-
                                              02901-3.pdf.

Brugiapaglia:2018:TSC

[BNMP18] Simone Brugiapaglia, Fabio Nobile, Stefano Micheletti, and Simona Perotto. A theoretical study of COmpRessed Solv-ING for advection–diffusionreaction problems. Mathematics of Computation, 87 (309):1–38, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-309/
                                              S0025-5718-2017-03209-9;

http://www.ams.org/journals/
mcom/2018-87-309/S0025-5718-
2017-03209-9/S0025-5718-2017-
03209-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Brugiapaglia%
2C%20Simone;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Nobile%
2C%20Fabio;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Perotto%
2C%20Simona;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=637026.

Bartels:2015:TVD

[BNS15] Sören Bartels, Ricardo H. Nochetto, and Abner J. Salgado. A total variation diminishing interpolation operator and applications. Mathematics of Computation, 84(296): 2569–2587, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-296/
S0025-5718-2015-02942-1;
http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02942-1/S0025-5718-2015-
02942-1.pdf.

Barekat:2015:TCB

[BO15] Farzin Barekat and Stanley Osher. A time continuation based fast approximate algorithm for compressed sensing related optimization. Mathematics of Computation, 84(296): 2791–2822, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-296/

                                              S0025-5718-2015-02972-X;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-296/S0025-5718-
                                              2015-02972-X/S0025-5718-2015-
                                              02972-X.pdf.

Booker:2018:FCC

[Boo18] Andrew R. Booker. Finite connected components of the aliquot graph. Mathematics of Computation, 87(314):2891– 2902, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-314/
                                              S0025-5718-2018-03299-9;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-314/S0025-5718-
                                              2018-03299-9/S0025-5718-2018-
                                              03299-9.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=672596.

Barron:2017:DAF

[BOP17] Thomas Barron, Christopher O’Neill, and Roberto Pelayo. On dynamic algorithms for factorization invariants in numerical monoids. Mathematics of Computation, 86(307): 2429–2447, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-307/
                                              S0025-5718-2016-03160-9;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-307/S0025-5718-
                                              2016-03160-9/S0025-5718-2016-
                                              03160-9.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Barron%
                                              2C%20Thomas;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=1019989;

http://www.ams.org/mathscinet/
search/author.html?mrauthid=
1058872.

Bornemann:2010:NEF

[Bor10] Folkmar Bornemann. On the numerical evaluation of Fredholm determinants. Mathematics of Computation, 79(270): 871–915, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02280-7/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02280-7/S0025-5718-09-
02280-7.pdf.

Borm:2018:ACL

[Bör18] Steffen Börm. Adaptive compression of large vectors. Mathematics of Computation, 87 (310):209–235, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03203-8;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03203-8/S0025-5718-2017-
03203-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=678579.

Banas:2010:CFE

[BP10] Lubomír Banas and Andreas Prohl. Convergent finite element discretization of the multi-fluid nonstationary incompressible magnetohydrodynamics equations. Mathematics of Computation, 79(272):1957– 1999, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02341-0/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02341-0/S0025-5718-10-
02341-0.pdf.

Bauer:2012:CMP

[BP12a] I. Bauer and R. Pignatelli. The classification of minimal product-quotient surfaces with p_{g} = 0. Mathematics of Computation, 81(280):2389– 2418, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-280/
S0025-5718-2012-02604-4;
http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-
2012-02604-4/S0025-5718-2012-
02604-4.pdf.

Bonito:2012:CAV

[BP12b] Andrea Bonito and Joseph E. Pasciak. Convergence analysis of variational and nonvariational multigrid algorithms for the Laplace–Beltrami operator. Mathematics of Computation, 81(279):1263– 1288, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2011-02551-2;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2011-02551-2/S0025-5718-2011-
02551-2.pdf;   http://www.

ams.org/mathscinet/search/
authors.html?authorName=Bonito%
2C%20Andrea;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Pasciak%
2C%20Joseph%20E..

Bonito:2015:NAF

[BP15] Andrea Bonito and Joseph E.

Pasciak. Numerical approximation of fractional powers of elliptic operators. Mathematics of Computation, 84(295): 2083–2110, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-295/
S0025-5718-2015-02937-8;
http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-
2015-02937-8/S0025-5718-2015-
02937-8.pdf.

Bramble:2010:AFP

[BPT10] James H. Bramble, Joseph E.

Pasciak, and Dimitar Trenev. Analysis of a finite PML approximation to the three dimensional elastic wave scattering problem. Mathematics of Computation, 79(272):2079– 2101, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02355-0/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02355-0/S0025-5718-10-
02355-0.pdf.

Bjorn:1998:FGF

[BR98] Anders Björn and Hans Riesel.

Factors of generalized Fermat numbers. With microfiche supplement. Mathematics of Computation, 67(221):441– 446, January 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
jourcgi/jour-pbprocess?fn=
110&arg1=S0025-5718-98-00891-
6&u=/mcom/1998-67-221/. See

errata [BR05, BR11].

Bjorn:2005:TEF

[BR05] Anders Björn and Hans Riesel. Table errata to “Factors of generalized Fermat numbers”. Mathematics of Computation, 74(252):2099, October 2005. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.
ams.org/mcom/2005-74-252/
S0025-5718-05-01816-8/;
http://www.ams.org/mcom/2005-
74-252/S0025-5718-05-01816-
8/S0025-5718-05-01816-8.dvi;
 http://www.ams.org/mcom/
2005-74-252/S0025-5718-05-
01816-8/S0025-5718-05-01816-
8.pdf;   http://www.ams.org/
mcom/2005-74-252/S0025-5718-
05-01816-8/S0025-5718-05-
01816-8.ps;   http://www.
ams.org/mcom/2005-74-252/
S0025-5718-05-01816-8/S0025-
5718-05-01816-8.tex.  See

[BR98, BR11].

Bjorn:2011:TEF

[BR11] Anders Björn and Hans Riesel. Table errata 2 to “Factors of generalized Fermat numbers”. Mathematics of Computation, 80(275):1865–1866, July 2011.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-10-02371-9/
;   http://www.ams.org/journals/
mcom/2011-80-275/S0025-5718-
10-02371-9/S0025-5718-10-
02371-9.pdf.  See [BR98,

BR05].

Banjai:2018:CQW

[BR18] Lehel Banjai and Alexander Rieder. Convolution quadrature for the wave equation with a nonlinear impedance boundary condition. Mathematics of Computation, 87(312):1783– 1819, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2017-03279-8;
http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03279-8/S0025-5718-2017-
03279-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Banjai%
2C%20Lehel;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1126143.

Bars:2019:BMC

[BR19] Francesc Bars and Josep González Rovira. Bielliptic modular curves X_{0}^{∗}(N) with squarefree levels. Mathematics of Computation, 88(320):2939– 2957, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-320/
S0025-5718-2019-03424-5;

                                              https://www.ams.org/journals/
                                              mcom/2019-88-320/S0025-5718-
                                              2019-03424-5/S0025-5718-2019-
                                              03424-5.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=319937;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              647724.

Bruin:2013:CPG

[Bru13] Peter Bruin. Computing in Picard groups of projective curves over finite fields. Mathematics of Computation, 82(283):1711– 1756, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-283/
                                              S0025-5718-2012-02650-0;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-283/S0025-5718-
                                              2012-02650-0/S0025-5718-2012-
                                              02650-0.pdf.

Bach:2015:CCT

[BS15a] Eric Bach and Andrew Shallue. Counting composites with two strong liars. Mathematics of Computation, 84(296):3069– 3089, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-296/
                                              S0025-5718-2015-02949-4;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-296/S0025-5718-
                                              2015-02949-4/S0025-5718-2015-
                                              02949-4.pdf.

Brenner:2015:PHM

[BS15b] Susanne C. Brenner and Li-Yeng Sung. Piecewise H^{1} functions and vector fields associated with meshes generated by independent refinements. Mathematics of Computation, 84 (293):1017–1036, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02866-4;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02866-4/S0025-5718-2014-
02866-4.pdf.

Banks:2016:GSE

[BS16] William D. Banks and Igor E.

Shparlinski. On Gauss sums and the evaluation of Stechkin’s constant. Mathematics of Computation, 85(301):2569–2581, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-301/
S0025-5718-2015-03056-7;
http://www.ams.org/journals/
mcom/2016-85-301/S0025-5718-
2015-03056-7/S0025-5718-2015-
03056-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=192194;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
336964.

Bao:2018:UEB

[BS18] Weizhu Bao and Chunmei Su. Uniform error bounds of a finite difference method for the Klein–Gordon–Zakharov system in the subsonic limit regime. Mathematics of Computation, 87(313):2133–2158, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-313/
                                              S0025-5718-2017-03278-6;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-313/S0025-5718-
                                              2017-03278-6/S0025-5718-2017-
                                              03278-6.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Bao%
                                              2C%20Weizhu;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=955028.

Broughan:2013:PR

[BSL13] Kevin Broughan, Sergio Guzman Sanchez, and Florian Luca. Perfect repdigits. Mathematics of Computation, 82(284): 2439–2459, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-284/
                                              S0025-5718-2013-02682-8;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-284/S0025-5718-
                                              2013-02682-8/S0025-5718-2013-
                                              02682-8.pdf.

Balanzario:2012:RSI

[BSO12] Eugenio P. Balanzario and Jorge Sánchez-Ortiz. Riemann–Siegel integral formula for the Lerch zeta function. Mathematics of Computation, 81(280):2319– 2333, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-280/
                                              S0025-5718-2011-02566-4;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-280/S0025-5718-

2011-02566-4/S0025-5718-2011-
02566-4.pdf.

Brauchart:2014:QDO

[BSSW14] J. S. Brauchart, E. B. Saff, I. H.

Sloan, and R. S. Womersley. QMC designs: Optimal order Quasi Monte Carlo integration schemes on the sphere. Mathematics of Computation, 83 (290):2821–2851, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02839-1;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02839-1/S0025-5718-2014-
02839-1.pdf.

Boonyasiriwat:2011:CEA

[BST11] C. Boonyasiriwat, K. Sikorski, and C. Tsay. Circumscribed ellipsoid algorithm for fixedpoint problems. Mathematics of Computation, 80(275): 1703–1723, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2010-02443-
3/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2010-02443-3/S0025-
5718-2010-02443-3.pdf.

Best:2015:LRZ

[BT15] D. G. Best and T. S. Trudgian. Linear relations of zeroes of the zeta-function. Mathematics of Computation, 84 (294):2047–2058, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-294/
S0025-5718-2014-02916-5;
http://www.ams.org/journals/
mcom/2015-84-294/S0025-5718-
2014-02916-5/S0025-5718-2014-
02916-5.pdf.

Bui-Thanh:2013:CWP

[BTDG13] Tan Bui-Thanh, Leszek Demkowicz, and Omar Ghattas. Constructively well-posed approximation methods with unity inf–sup and continuity constants for partial differential equations. Mathematics of Computation, 82(284): 1923–1952, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-284/
S0025-5718-2013-02697-X;
http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02697-X/S0025-5718-2013-
02697-X.pdf.

Burton:2010:ODD

[Bur10] Benjamin A. Burton. Optimizing the double description method for normal surface enumeration. Mathematics of Computation, 79(269):453– 484, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-269/
S0025-5718-09-02282-0/;
http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02282-0/S0025-5718-09-
02282-0.pdf.

Burman:2017:SNF

[Bur17]  Erik Burman.  A stabi-

lized nonconforming finite element method for the elliptic Cauchy problem. Mathematics of Computation, 86(303): 75–96, January 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-303/
S0025-5718-2016-03092-6;
http://www.ams.org/journals/
mcom/2017-86-303/S0025-5718-
2016-03092-6/S0025-5718-2016-
03092-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Burman%
2C%20Erik.

Buthe:2015:MPC

[Büt15]  Jan Büthe.  A method

for proving the completeness of a list of zeros of certain L-functions. Mathematics of Computation, 84(295): 2413–2431, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-295/
S0025-5718-2015-02922-6;
http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-
2015-02922-6/S0025-5718-2015-
02922-6.pdf.

Buthe:2016:ERF

[Büt16] Jan Büthe. Estimating π(x) and related functions under partial RH assumptions. Mathematics of Computation, 85(301):2483– 2498, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-301/
                                              S0025-5718-2015-03060-9;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-301/S0025-5718-
                                              2015-03060-9/S0025-5718-2015-
                                              03060-9.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=1017601.

Buthe:2018:AMB

[Büt18] Jan Büthe. An analytic method for bounding ψ(x). Mathematics of Computation, 87(312):1991– 2009, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-312/
                                              S0025-5718-2017-03264-6;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-312/S0025-5718-
                                              2017-03264-6/S0025-5718-2017-
                                              03264-6.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1017601.

Beck:2011:EMS

[BvH11] Matthias Beck and Andrew van Herick. Enumeration of 4 × 4 magic squares. Mathematics of Computation, 80(273):617– 621, January 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2011-80-273/
                                              S0025-5718-10-02347-1/;
                                              http://www.ams.org/journals/
                                              mcom/2011-80-273/S0025-5718-
                                              10-02347-1/S0025-5718-10-
                                              02347-1.pdf.

Batenkov:2012:AFR

[BY12] Dmitry Batenkov and Yosef Yomdin. Algebraic Fourier reconstruction of piecewise smooth functions. Mathematics of Computation, 81(277):277– 318, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02539-1;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02539-1/S0025-5718-2011-
02539-1.pdf.

Ben-yu:2010:SMQ

[ByHl10] Guo Ben-yu and Jia Hongli. Spectral method on quadrilaterals. Mathematics of Computation, 79(272):2237– 2264, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02329-X/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02329-X/S0025-5718-10-
02329-X.pdf.

Brenner:2012:FEM

[BySZ12] Susanne C. Brenner, Li yeng Sung, and Yi Zhang. Finite element methods for the displacement obstacle problem of clamped plates. Mathematics of Computation, 81(279): 1247–1262, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2012-02602-0;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-

                                              2012-02602-0/S0025-5718-2012-
                                              02602-0.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Brenner%
                                              2C%20Susanne%20C;   http:
                                              //www.ams.org/mathscinet/
                                              search/authors.html?authorName=
                                              Sung%2C%20Li-yeng;   http:
                                              //www.ams.org/mathscinet/
                                              search/authors.html?authorName=
                                              Zhang%2C%20Yi.

Ben-yu:2013:JLQ

[ByTC13] Guo Ben-yu, Sun Tao, and Zhang Chao. Jacobi and Laguerre quasi-orthogonal approximations and related interpolations. Mathematics of Computation, 82(281):413– 441, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-281/
                                              S0025-5718-2012-02614-7;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-281/S0025-5718-
                                              2012-02614-7/S0025-5718-2012-
                                              02614-7.pdf.

Brinkmann:2018:SVS

[BZ18] Philip Brinkmann and Günter M. Ziegler. Small f-vectors of 3-spheres and of 4-polytopes. Mathematics of Computation, 87(314):2955–2975, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2018-87-
                                              314/S0025-5718-2018-03300-
                                              2;   http://www.ams.org/journals/
                                              mcom/2018-87-314/S0025-5718-
                                              2018-03300-2/S0025-5718-2018-
                                              03300-2.pdf;   https://www.

ams.org/mathscinet/search/
authors.html?authorName=Ziegler%
2C%20Gunter%20M.;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
1199626.

Berschneider:2012:FSC

[BzCS12] Georg Berschneider, Wolfgang zu Castell, and Stefan J. Schrödl. Function spaces for conditionally positive definite operator-valued kernels. Mathematics of Computation, 81(279): 1551–1569, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2011-02552-4;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2011-02552-4/S0025-5718-2011-
02552-4.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Berschneider%
2C%20Georg;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=zu%
20Castell%2C%20Wolfgang;
http://www.ams.org/mathscinet/
search/authors.html?authorName=
Schrodl%2C%20Stefan%20J..

Caley:2013:PTE

[Cal13] Timothy Caley. The Prouhet– Tarry–Escott problem for Gaussian integers. Mathematics of Computation, 82(282):1121– 1137, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02532-4;
http://www.ams.org/journals/

                                              mcom/2013-82-282/S0025-5718-
                                              2012-02532-4/S0025-5718-2012-
                                              02532-4.pdf.

Calvo:2016:BAD

[Cal16] Juan G. Calvo. A BDDC algorithm with deluxe scaling for H(curl) in two dimensions with irregular subdomains. Mathematics of Computation, 85 (299):1085–1111, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-299/
                                              S0025-5718-2015-03028-2;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-299/S0025-5718-
                                              2015-03028-2/S0025-5718-2015-
                                              03028-2.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Calvo%
                                              2C%20Juan%20G..

Cao:2015:SAL

[Cao15] Weiming Cao. Superconvergence analysis of the linear finite element method and a gradient recovery postprocessing on anisotropic meshes. Mathematics of Computation, 84 (291):89–117, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-291/
                                              S0025-5718-2014-02846-9;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-291/S0025-5718-
                                              2014-02846-9/S0025-5718-2014-
                                              02846-9.pdf.

Craig:2016:BMA

[CB16] Katy Craig and Andrea L. Bertozzi. A blob method for the aggregation equation. Mathematics of Computation, 85(300): 1681–1717, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-300/
S0025-5718-2015-03033-6;
http://www.ams.org/journals/
mcom/2016-85-300/S0025-5718-
2015-03033-6/S0025-5718-2015-
03033-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Craig%
2C%20Katy;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=265966.

Chabaud:2012:UTS

[CC12a] Brandon Chabaud and Bernardo Cockburn. Uniform-in-time superconvergence of HDG methods for the heat equation. Mathematics of Computation, 81 (277):107–129, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2012-81-
277/S0025-5718-2011-02525-
1;   http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02525-1/S0025-5718-2011-
02525-1.pdf.

Cockburn:2012:AHM

[CC12b] Bernardo Cockburn and Jintao Cui. An analysis of HDG methods for the vorticityvelocity-pressure formulation of the Stokes problem in three dimensions. Mathematics of Computation, 81(279):1355– 1368, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-279/
                                              S0025-5718-2011-02575-5;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-279/S0025-5718-
                                              2011-02575-5/S0025-5718-2011-
                                              02575-5.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Cockburn%
                                              2C%20Bernardo;   http://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Cui%
                                              2C%20Jintao.

Cai:2013:NVH

[CC13a] Tianxin Cai and Deyi Chen. A new variant of the Hilbert– Waring problem. Mathematics of Computation, 82(284): 2333–2341, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-284/
                                              S0025-5718-2013-02685-3;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-284/S0025-5718-
                                              2013-02685-3/S0025-5718-2013-
                                              02685-3.pdf.

Coffey:2013:LCJ

[CC13b] Mark W. Coffey and George Csordas. On the log-concavity of a Jacobi theta function. Mathematics of Computation, 82(284):2265–2272, 2013. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2013-82-
                                              284/S0025-5718-2013-02681-
                                              6;   http://www.ams.org/journals/
                                              mcom/2013-82-284/S0025-5718-
                                              2013-02681-6/S0025-5718-2013-
                                              02681-6.pdf.

Chen:2014:AVD

[CC14] Yanlai Chen and Bernardo Cockburn. Analysis of variabledegree HDG methods for convection–diffusion equations. Part II: Semimatching nonconforming meshes. Mathematics of Computation, 83 (285):87–111, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-285/
S0025-5718-2013-02711-1;
http://www.ams.org/journals/
mcom/2014-83-285/S0025-5718-
2013-02711-1/S0025-5718-2013-
02711-1.pdf.

Chen:2016:SHM

[CCD16] Yanlai Chen, Bernardo Cockburn, and Bo Dong. Superconvergent HDG methods for linear, stationary, third-order equations in one-space dimension. Mathematics of Computation, 85(302):2715–2742, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-302/
S0025-5718-2016-03091-4;
http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-
2016-03091-4/S0025-5718-2016-
03091-4.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Chen%
2C%20Yanlai;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Cockburn%
2C%20Bernardo;   http://www.
ams.org/mathscinet/search/

                                              author.html?authorName=Dong%
                                              2C%20Bo.

Calo:2019:SAE

[CCDE19] Victor Calo, Matteo Cicuttin, Quanling Deng, and Alexandre Ern. Spectral approximation of elliptic operators by the Hybrid High-Order method. Mathematics of Computation, 88(318): 1559–1586, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-318/
                                              S0025-5718-2018-03405-6;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-318/S0025-5718-
                                              2018-03405-6/S0025-5718-2018-
                                              03405-6.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Cicuttin%
                                              2C%20Matteo;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Deng%
                                              2C%20Quanling;   https://
                                              www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=349433;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              816582.

Chartier:2019:UAM

[CCL^{+}19] Philippe Chartier, Nicolas Crouseilles, Mohammed Lemou, Florian Méhats, and Xiaofei Zhao. Uniformly accurate methods for Vlasov equations with non-homogeneous strong magnetic field. Mathematics of Computation, 88(320):2697– 2736, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-320/

S0025-5718-2019-03436-1;
https://www.ams.org/journals/
mcom/2019-88-320/S0025-5718-
2019-03436-1/S0025-5718-2019-
03436-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1045425;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
335517;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=355223;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
601414;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=695280.

Cheng:2017:SDG

[CCLX17] Yingda Cheng, Ching-Shan Chou, Fengyan Li, and Yulong Xing. L^{2} stable discontinuous Galerkin methods for one-dimensional two-way wave equations. Mathematics of Computation, 86(303):121– 155, January 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-303/
S0025-5718-2016-03090-2;
http://www.ams.org/journals/
mcom/2017-86-303/S0025-5718-
2016-03090-2/S0025-5718-2016-
03090-2.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=718718;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
761305;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=782690;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
811395.

Causley:2014:MLT

[CCOV14] Matthew Causley, Andrew Christlieb, Benjamin Ong, and Lee Van Groningen. Method of lines transpose: an implicit solution to the wave equation. Mathematics of Computation, 83 (290):2763–2786, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-290/
                                              S0025-5718-2014-02834-2;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-290/S0025-5718-
                                              2014-02834-2/S0025-5718-2014-
                                              02834-2.pdf.

Cesmelioglu:2017:AHD

[CCQ17] Aycil Cesmelioglu, Bernardo Cockburn, and Weifeng Qiu. Analysis of a hybridizable discontinuous Galerkin method for the steady-state incompressible Navier–Stokes equations. Mathematics of Computation, 86 (306):1643–1670, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-306/
                                              S0025-5718-2016-03195-6;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-306/S0025-5718-
                                              2016-03195-6/S0025-5718-2016-
                                              03195-6.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Cockburn%
                                              2C%20Bernardo;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=845089;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              864513.

Celiker:2012:PBE

[CCS12] Fatih Celiker, Bernardo Cockburn, and Ke Shi. A projection-based error analysis of HDG methods for Timoshenko beams. Mathematics of Computation, 81(277):131– 151, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02522-6;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02522-6/S0025-5718-2011-
02522-6.pdf.

Curtis:2010:CHM

[CD10] Christopher W. Curtis and Bernard Deconinck. On the convergence of Hill’s method. Mathematics of Computation, 79(269):169–187, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2010-79-
269/S0025-5718-09-02277-7/
;   http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02277-7/S0025-5718-09-
02277-7.pdf.

Carstensen:2015:CAF

[CD15] Carsten Carstensen and Georg Dolzmann. Convergence of adaptive finite element methods for a nonconvex double-well minimization problem. Mathematics of Computation, 84 (295):2111–2135, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-295/
S0025-5718-2015-02947-0;
http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-
2015-02947-0/S0025-5718-2015-
02947-0.pdf.

Cockburn:2016:HDG

[CD16] Bernardo Cockburn and Alan Demlow. Hybridizable discontinuous Galerkin and mixed finite element methods for elliptic problems on surfaces. Mathematics of Computation, 85(302): 2609–2638, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2016-85-
302/S0025-5718-2016-03093-
8;   http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-
2016-03093-8/S0025-5718-2016-
03093-8.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Cockburn%
2C%20Bernardo;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=693541.

Cosgrave:2017:RGF

[CD17] John B. Cosgrave and Karl Dilcher. A role for generalized Fermat numbers. Mathematics of Computation, 86 (304):899–933, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03111-7;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03111-7/S0025-5718-2016-
03111-7.pdf;   http://www.

ams.org/mathscinet/search/
author.html?authorName=Cosgrave%
2C%20John%20B;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Dilcher%
2C%20Karl.

Cheng:2018:GBM

[CD18] Wanyou Cheng and Yu-Hong Dai. Gradient-based method with active set strategy for ℓ_{1} optimization. Mathematics of Computation, 87(311): 1283–1305, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-311/
S0025-5718-2017-03238-5;
http://www.ams.org/journals/
mcom/2018-87-311/S0025-5718-
2017-03238-5/S0025-5718-2017-
03238-5.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=620453;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
828685.

Chyzak:2018:CSL

[CDDM18] Frédéric Chyzak, Thomas Dreyfus, Philippe Dumas, and Marc Mezzarobba. Computing solutions of linear Mahler equations. Mathematics of Computation, 87(314):2977– 3021, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-314/
S0025-5718-2018-03359-2;
http://www.ams.org/journals/
mcom/2018-87-314/S0025-5718-
2018-03359-2/S0025-5718-2018-
03359-2.pdf;   https://www.

                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Chyzak%
                                              2C%20Frederic;   https://
                                              www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=1051219;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              342673;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=904789.

Chegini:2013:PTP

[CDFS13] Nabi Chegini, Stephan Dahlke, Ulrich Friedrich, and Rob Stevenson. Piecewise tensor product wavelet bases by extensions and approximation rates. Mathematics of Computation, 82(284): 2157–2190, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-284/
                                              S0025-5718-2013-02694-4;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-284/S0025-5718-
                                              2013-02694-4/S0025-5718-2013-
                                              02694-4.pdf.

Cohen:2012:AMA

[CDHM12] Albert Cohen, Nira Dyn, Frédéric Hecht, and Jean-Marie Mirebeau. Adaptive multiresolution analysis based on anisotropic triangulations. Mathematics of Computation, 81(278):789– 810, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-278/
                                              S0025-5718-2011-02495-6;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-278/S0025-5718-

2011-02495-6/S0025-5718-2011-
02495-6.pdf.

Capparelli:2010:SPI

[CDS10] Stefano Capparelli, Alberto Del Fra, and Carlo Sciò. On the span of polynomials with integer coefficients. Mathematics of Computation, 79(270): 967–981, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02292-3/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02292-3/S0025-5718-09-
02292-3.pdf.

Chkifa:2018:PAC

[CDTW18] Abdellah Chkifa, Nick Dexter, Hoang Tran, and Clayton G. Webster. Polynomial approximation via compressed sensing of high-dimensional functions on lower sets. Mathematics of Computation, 87(311): 1415–1450, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-311/
S0025-5718-2017-03272-5;
http://www.ams.org/journals/
mcom/2018-87-311/S0025-5718-
2017-03272-5/S0025-5718-2017-
03272-5.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1003967;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1160932;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=820027;   https:

                                              //www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              956091.

Cerri:2011:SGE

[Cer11] Jean-Paul Cerri. Some generalized Euclidean and 2-stage Euclidean number fields that are not norm-Euclidean. Mathematics of Computation, 80(276): 2289–2298, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              276/S0025-5718-2011-02468-
                                              3/;   http://www.ams.org/
                                              journals/mcom/2011-80-276/
                                              S0025-5718-2011-02468-3/S0025-
                                              5718-2011-02468-3.pdf;   http:
                                              //www.ams.org/mathscinet-
                                              getitem?mr=2813361.

Chen:2012:GFD

[CFJ12] William Y. C. Chen, Neil J. Y. Fan, and Jeffrey Y. T. Jia. The generating function for the Dirichlet series L_{m}(s). Mathematics of Computation, 81(278): 1005–1023, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-278/
                                              S0025-5718-2011-02520-2;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-278/S0025-5718-
                                              2011-02520-2/S0025-5718-2011-
                                              02520-2.pdf.

Corwin:2014:ECF

[CFLTL14] David Corwin, Tony Feng, Zane Kun Li, and Sarah Trebat-Leder. Elliptic curves with full 2-torsion and maximal adelic Galois representations. Mathematics of Computation, 83 (290):2925–2951, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02804-4;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02804-4/S0025-5718-2014-
02804-4.pdf.

Cremona:2015:EDE

[CFO^{+}15] J. E. Cremona, T. A. Fisher, C. O’Neil, D. Simon, and M. Stoll. Explicit n-descent on elliptic curves III. Algorithms. Mathematics of Computation, 84(292):895–922, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-292/
S0025-5718-2014-02858-5;
http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02858-5/S0025-5718-2014-
02858-5.pdf.

Cockburn:2017:SDP

[CFS17] Bernardo Cockburn, Guosheng Fu, and Francisco Javier Sayas. Superconvergence by Mdecompositions. Part I: General theory for HDG methods for diffusion. Mathematics of Computation, 86(306): 1609–1641, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-306/
S0025-5718-2016-03140-3;
http://www.ams.org/journals/

                                              mcom/2017-86-306/S0025-5718-
                                              2016-03140-3/S0025-5718-2016-
                                              03140-3.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Cockburn%
                                              2C%20Bernardo;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Fu%
                                              2C%20Guosheng;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=621885.

Clery:2019:CBS

[CFvdG19] Fabien Cléry, Carel Faber, and Gerard van der Geer. Covariants of binary sextics and modular forms of degree 2 with character. Mathematics of Computation, 88(319):2423– 2441, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-319/
                                              S0025-5718-2019-03412-9;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-319/S0025-5718-
                                              2019-03412-9/S0025-5718-2019-
                                              03412-9.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Clery%
                                              2C%20Fabien;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=194375;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              64735.

Coulombel:2011:SSF

[CG11] Jean-François Coulombel and Antoine Gloria. Semigroup stability of finite difference schemes for multidimensional hyperbolic initial-boundary value problems. Mathematics of Computation, 80(273):165–203, January 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-10-02368-9/;
http://www.ams.org/journals/
mcom/2011-80-273/S0025-5718-
10-02368-9/S0025-5718-10-
02368-9.pdf.

Carstensen:2014:GLB

[CG14] Carsten Carstensen and Joscha Gedicke. Guaranteed lower bounds for eigenvalues. Mathematics of Computation, 83 (290):2605–2629, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02833-0;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02833-0/S0025-5718-2014-
02833-0.pdf.

Cances:2016:CNE

[CG16] Clément Cancès and Cindy Guichard. Convergence of a nonlinear entropy diminishing Control Volume Finite Element scheme for solving anisotropic degenerate parabolic equations. Mathematics of Computation, 85(298):549–580, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-02997-4;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-02997-4/S0025-5718-2015-
02997-4.pdf;   http://www.

                                              ams.org/mathscinet/search/
                                              author.html?authorName=Cances%
                                              2C%20Clement;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Guichard%
                                              2C%20Cindy.

Cockburn:2010:NEE

[CGG10] Bernardo Cockburn, Jayadeep Gopalakrishnan, and Johnny Guzmán. A new elasticity element made for enforcing weak stress symmetry. Mathematics of Computation, 79(271): 1331–1349, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-271/
                                              S0025-5718-10-02343-4/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-271/S0025-5718-
                                              10-02343-4/S0025-5718-10-
                                              02343-4.pdf.

Chu:2010:NMF

[CGH10] C.-C. Chu, I. G. Graham, and T.-Y. Hou. A new multiscale finite element method for high-contrast elliptic interface problems. Mathematics of Computation, 79(272): 1915–1955, October 2010. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2010-79-
                                              272/S0025-5718-2010-02372-
                                              5/;   http://www.ams.org/
                                              journals/mcom/2010-79-272/
                                              S0025-5718-2010-02372-5/S0025-
                                              5718-2010-02372-5.pdf.

Costa:2014:SWP

[CGH14] Edgar Costa, Robert Gerbicz, and David Harvey. A search for Wilson primes. Mathematics of Computation, 83(290): 3071–3091, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02800-7;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02800-7/S0025-5718-2014-
02800-7.pdf.

Coquel:2016:CMS

[CGH^{+}16] Frédéric Coquel, Edwige Godlewski, Khalil Haddaoui, Claude Marmignon, and Florent Renac. Choice of measure source terms in interface coupling for a model problem in gas dynamics. Mathematics of Computation, 85(301): 2305–2339, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2016-85-
301/S0025-5718-2016-03063-
X;   http://www.ams.org/journals/
mcom/2016-85-301/S0025-5718-
2016-03063-X/S0025-5718-2016-
03063-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Coquel%
2C%20Frederic;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Haddaoui%
2C%20Khalil;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=266191;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
652914;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=938391.

Cao:2014:PNI

[CGHW14] Yanzhao Cao, Max Gunzburger, Xiaoming He, and Xiaoming Wang. Parallel, noniterative, multi-physics domain decomposition methods for time-dependent Stokes– Darcy systems. Mathematics of Computation, 83(288): 1617–1644, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-288/
                                              S0025-5718-2014-02779-8;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-288/S0025-5718-
                                              2014-02779-8/S0025-5718-2014-
                                              02779-8.pdf.

Cockburn:2011:AHM

[CGN^{+}11] Bernardo Cockburn, Jayadeep Gopalakrishnan, Ngoc Cuong Nguyen, Jaume Peraire, and Francisco-Javier Sayas. Analysis of HDG methods for Stokes flow. Mathematics of Computation, 80(274): 723–760, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              274/S0025-5718-2010-02410-
                                              X/;   http://www.ams.org/
                                              journals/mcom/2011-80-274/
                                              S0025-5718-2010-02410-X/S0025-
                                              5718-2010-02410-X.pdf.

Chowdhury:2017:EBD

[CGN17] Sudipto Chowdhury, Thirupathi Gudi, and A. K. Nandakumaran. Error bounds for a Dirichlet boundary control problem based on energy spaces. Mathematics of Computation, 86(305): 1103–1126, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-305/
S0025-5718-2016-03125-7;
http://www.ams.org/journals/
mcom/2017-86-305/S0025-5718-
2016-03125-7/S0025-5718-2016-
03125-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Chowdhury%
2C%20Sudipto;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Gudi%
2C%20Thirupathi;   http://
www.ams.org/mathscinet/search/
author.html?authorName=Nandakumaran%
2C%20A.%20K..

Copeland:2010:MWS

[CGO10] Dylan M. Copeland, Jayadeep Gopalakrishnan, and Minah Oh. Multigrid in a weighted space arising from axisymmetric electromagnetics. Mathematics of Computation, 79(272): 2033–2058, October 2010. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2010-79-
272/S0025-5718-2010-02384-
1/;   http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-2010-02384-1/S0025-
5718-2010-02384-1.pdf.

Cheng:2012:PPD

[CGP12] Yingda Cheng, Irene M. Gamba, and Jennifer Proft. Positivitypreserving discontinuous Galerkin schemes for linear Vlasov– Boltzmann transport equations.

Mathematics of Computation, 81(277):153–190, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2012-81-
                                              277/S0025-5718-2011-02504-
                                              4;   http://www.ams.org/journals/
                                              mcom/2012-81-277/S0025-5718-
                                              2011-02504-4/S0025-5718-2011-
                                              02504-4.pdf.

Cockburn:2010:PBE

[CGS10] Bernardo Cockburn, Jayadeep Gopalakrishnan, and Francisco-Javier Sayas. A projectionbased error analysis of HDG methods. Mathematics of Computation, 79(271):1351– 1367, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-271/
                                              S0025-5718-10-02334-3/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-271/S0025-5718-
                                              10-02334-3/S0025-5718-10-
                                              02334-3.pdf.

Carstensen:2015:ANC

[CGS15] Carsten Carstensen, Dietmar Gallistl, and Mira Schedensack. Adaptive nonconforming Crouzeix–Raviart FEM for eigenvalue problems. Mathematics of Computation, 84 (293):1061–1087, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-293/
                                              S0025-5718-2014-02894-9;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-293/S0025-5718-

2014-02894-9/S0025-5718-2014-
02894-9.pdf.

Cangiani:2018:ADG

[CGS18] Andrea Cangiani, Emmanuil H.

Georgoulis, and Younis A. Sabawi. Adaptive discontinuous Galerkin methods for elliptic interface problems. Mathematics of Computation, 87(314): 2675–2707, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-314/
S0025-5718-2018-03322-1;
http://www.ams.org/journals/
mcom/2018-87-314/S0025-5718-
2018-03322-1/S0025-5718-2018-
03322-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Georgoulis%
2C%20Emmanuil%20H.;   https:
//www.ams.org/mathscinet/
search/authors.html?authorName=
Sabawi%2C%20Younis%20A.;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
757643.

Chen:2013:HOS

[CH13] Chuanmiao Chen and Shufang Hu. The highest order superconvergence for bi-k degree rectangular elements at nodes: a proof of 2k-conjecture. Mathematics of Computation, 82(283): 1337–1355, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2012-02653-6;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-

                                              2012-02653-6/S0025-5718-2012-
                                              02653-6.pdf.

Costa:2014:FDI

[CH14] Edgar Costa and David Harvey. Faster deterministic integer factorization. Mathematics of Computation, 83 (285):339–345, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-285/
                                              S0025-5718-2013-02707-X;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-285/S0025-5718-
                                              2013-02707-X/S0025-5718-2013-
                                              02707-X.pdf.

Chen:2019:MCP

[CH19] Zheng Chen and Cory D. Hauck. Multiscale convergence properties for spectral approximations of a model kinetic equation. Mathematics of Computation, 88(319):2257– 2293, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-319/
                                              S0025-5718-2018-03399-3;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-319/S0025-5718-
                                              2018-03399-3/S0025-5718-2018-
                                              03399-3.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1055259;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              748066.

Charles:2018:GRF

[Cha18]  Zachary Charles.  Generat-

ing random factored ideals in number fields. Mathematics of Computation, 87(312):2047– 2056, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2017-03283-X;
http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03283-X/S0025-5718-2017-
03283-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1002462.

Chernov:2012:OCE

[Che12] Alexey Chernov. Optimal convergence estimates for the trace of the polynomial L^{2}-projection operator on a simplex. Mathematics of Computation, 81(278): 765–787, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02513-5;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02513-5/S0025-5718-2011-
02513-5.pdf.

Cheze:2013:RAD

[Chè13] Guillaume Chèze. A recombination algorithm for the decomposition of multivariate rational functions. Mathematics of Computation, 82(283):1793– 1812, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2012-02658-5;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-

                                              2012-02658-5/S0025-5718-2012-
                                              02658-5.pdf.

Chen:2016:CMS

[Che16] Hao Chen. Computing the Mazur and Swinnerton-Dyer critical subgroup of elliptic curves. Mathematics of Computation, 85(301):2499–2514, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-301/
                                              S0025-5718-2015-03057-9;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-301/S0025-5718-
                                              2015-03057-9/S0025-5718-2015-
                                              03057-9.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Chen%
                                              2C%20Hao.

Chen:2018:FAS

[CHH18] Long Chen, Jun Hu, and Xuehai Huang. Fast auxiliary space preconditioners for linear elasticity in mixed form. Mathematics of Computation, 87(312):1601– 1633, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-312/
                                              S0025-5718-2017-03285-3;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-312/S0025-5718-
                                              2017-03285-3/S0025-5718-2017-
                                              03285-3.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=714525;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              735779;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=854280.

Chouly:2015:SNS

[CHR15] Franz Chouly, Patrick Hild, and Yves Renard. Symmetric and non-symmetric variants of Nitsche’s method for contact problems in elasticity: theory and numerical experiments. Mathematics of Computation, 84(293): 1089–1112, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02913-X;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02913-X/S0025-5718-2014-
02913-X.pdf.

Christiansen:2018:EAD

[Chr18] Snorre H. Christiansen. On eigenmode approximation for Dirac equations: Differential forms and fractional Sobolev spaces. Mathematics of Computation, 87(310):547–580, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2018-87-
310/S0025-5718-2017-03233-
6;   http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03233-6/S0025-5718-2017-
03233-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=663397.

Chu:2011:DBS

[Chu11] Wenchang Chu. Dougall’s bilateral _{2}H_{2}-series and Ramanujanlike π-formulae. Mathematics of Computation, 80(276): 2223–2251, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02474-
9/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02474-9/S0025-
5718-2011-02474-9.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813357.

Chu:2012:AFE

[Chu12] Wenchang Chu. Analytical formulae for extended _{3}F_{2}- series of Watson–Whipple– Dixon with two extra integer parameters. Mathematics of Computation, 81(277):467– 479, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02512-3;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02512-3/S0025-5718-2011-
02512-3.pdf.

Cai:2017:RBP

[CHZ17] Zhiqiang Cai, Cuiyu He, and Shun Zhang. Residual-based a posteriori error estimate for interface problems: Nonconforming linear elements. Mathematics of Computation, 86 (304):617–636, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03151-8;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-

2016-03151-8/S0025-5718-2016-
03151-8.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=He%
2C%20Cuiyu;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=235961;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
704861.

Chan:2015:GHC

[CIL15] Wai Kiu Chan, María Inés Icaza, and Emilio A. Lauret. A generalized Hermite constant for imaginary quadratic fields. Mathematics of Computation, 84(294): 1883–1900, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-294/
S0025-5718-2015-02903-2;
http://www.ams.org/journals/
mcom/2015-84-294/S0025-5718-
2015-02903-2/S0025-5718-2015-
02903-2.pdf.

Cinkir:2015:CPM

[Cin15] Zubeyir Cinkir. Computation of polarized metrized graph invariants by using discrete Laplacian matrix. Mathematics of Computation, 84(296):2953– 2967, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-296/
S0025-5718-2015-02981-0;
http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02981-0/S0025-5718-2015-
02981-0.pdf.

Cifani:2013:SVV

[CJ13] Simone Cifani and Espen R. Jakobsen. On the spectral vanishing viscosity method for periodic fractional conservation laws. Mathematics of Computation, 82(283):1489– 1514, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-283/
                                              S0025-5718-2013-02690-7;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-283/S0025-5718-
                                              2013-02690-7/S0025-5718-2013-
                                              02690-7.pdf.

Chung:2015:UAC

[CJ15] Y.-M. Chung and M. S. Jolly. A unified approach to compute foliations, inertial manifolds, and tracking solutions. Mathematics of Computation, 84(294): 1729–1751, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-294/
                                              S0025-5718-2014-02904-9;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-294/S0025-5718-
                                              2014-02904-9/S0025-5718-2014-
                                              02904-9.pdf.

Coquel:2018:ESC

[CJLW18] Frédéric Coquel, Shi Jin, Jian-Guo Liu, and Li Wang. Entropic sub-cell shock capturing schemes via Jin–Xin relaxation and Glimm front sampling for scalar conservation laws. Mathematics of Computation, 87(311): 1083–1126, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-311/
S0025-5718-2017-03253-1;
http://www.ams.org/journals/
mcom/2018-87-311/S0025-5718-
2017-03253-1/S0025-5718-2017-
03253-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Coquel%
2C%20Frederic;   https://
www.ams.org/mathscinet/search/
authors.html?authorName=Jin%
2C%20Shi;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=233036;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
239750.

Chua:2019:SAC

[CKS19] Lynn Chua, Mario Kummer, and Bernd Sturmfels. Schottky algorithms: Classical meets tropical. Mathematics of Computation, 88(319):2541– 2558, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-319/
S0025-5718-2018-03406-8;
https://www.ams.org/journals/
mcom/2019-88-319/S0025-5718-
2018-03406-8/S0025-5718-2018-
03406-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1037521;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1125499;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=238151.

Connors:2010:AFE

[CL10] J. Connors and W. Layton. On the accuracy of the finite element method plus time relaxation. Mathematics of Computation, 79(270): 619–648, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-270/
                                              S0025-5718-09-02316-3/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-270/S0025-5718-
                                              09-02316-3/S0025-5718-09-
                                              02316-3.pdf.

Chen:2016:ADC

[CL16] Jingrun Chen and Jianfeng Lu. Analysis of the divide-and-conquer method for electronic structure calculations. Mathematics of Computation, 85(302): 2919–2938, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2016-85-
                                              302/S0025-5718-2016-03066-
                                              5;   http://www.ams.org/journals/
                                              mcom/2016-85-302/S0025-5718-
                                              2016-03066-5/S0025-5718-2016-
                                              03066-5.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Chen%
                                              2C%20Jingrun;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Lu%
                                              2C%20Jianfeng.

Chen:2012:NMS

[CLAT12] Chang-Ming Chen, F. Liu, V. Anh, and I. Turner. Numerical methods for solving a two-dimensional variableorder anomalous subdiffusion equation. Mathematics of Computation, 81(277):345–366, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02447-6;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02447-6/S0025-5718-2011-
02447-6.pdf.

Cilleruelo:2018:EPI

[CLB18] Javier Cilleruelo, Florian Luca, and Lewis Baxter. Every positive integer is a sum of three palindromes. Mathematics of Computation, 87(314): 3023–3055, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-314/
S0025-5718-2017-03221-X;
http://www.ams.org/journals/
mcom/2018-87-314/S0025-5718-
2017-03221-X/S0025-5718-2017-
03221-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=292544;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
410370;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=630217.

Castryck:2014:CCC

[CLO14] Wouter Castryck, Robert Laterveer, and Myriam Ounaïes. Constraints on counterexamples to the Casas–Alvero conjecture and a verification in degree 12. Mathematics of Computation, 83(290): 3017–3037, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02809-3;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02809-3/S0025-5718-2014-
02809-3.pdf.

Cilleruelo:2016:CNS

[CLPM16] Javier Cilleruelo, Florian Luca, and Amalia Pizarro-Madariaga. Carmichael numbers in the sequence (2^{n}k + 1)_{n≥1}. Mathematics of Computation, 85 (297):357–377, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-297/
S0025-5718-2015-02982-2;
http://www.ams.org/journals/
mcom/2016-85-297/S0025-5718-
2015-02982-2/S0025-5718-2015-
02982-2.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Luca%
2C%20Florian;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Pizarro-
Madariaga%2C%20Amalia;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
292544.

Cohen:2013:BSL

[CLR13] Elaine Cohen, Tom Lyche, and Richard F. Riesenfeld. A Bspline-like basis for the Powell– Sabin 12-split based on simplex splines. Mathematics of Computation, 82(283):1667– 1707, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/

journals/mcom/2013-82-283/
S0025-5718-2013-02664-6;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-
2013-02664-6/S0025-5718-2013-
02664-6.pdf.

Castillo:2012:SPO

[CLRR12] Kenier Castillo, Regina Litz Lamblém, Fernando Rodrigo Rafaeli, and Alagacone Sri Ranga. Szegő and paraorthogonal polynomials on the real line: Zeros and canonical spectral transformations. Mathematics of Computation, 81(280):2229– 2249, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-280/
S0025-5718-2012-02593-2;
http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-
2012-02593-2/S0025-5718-2012-
02593-2.pdf.

Chee:2012:UCM

[CLTZ12] Yeow Meng Chee, San Ling, Yin Tan, and Xiande Zhang. Universal cycles for minimum coverings of pairs by triples, with application to 2radius sequences. Mathematics of Computation, 81(277):585– 603, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02473-7;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-

2011-02473-7/S0025-5718-2011-
02473-7.pdf.

Chen:2016:CAF

[CLWW16] Wenbin Chen, Yuan Liu, Cheng Wang, and Steven M. Wise. Convergence analysis of a fully discrete finite difference scheme for the Cahn–Hilliard–Hele– Shaw equation. Mathematics of Computation, 85(301):2231– 2257, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-301/
S0025-5718-2015-03052-X;
http://www.ams.org/journals/
mcom/2016-85-301/S0025-5718-
2015-03052-X/S0025-5718-2015-
03052-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Chen%
2C%20Wenbin;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Liu%
2C%20Yuan;   http://www.ams.
org/mathscinet/search/author.
html?authorName=Wise%2C%20Steven%
20M;   http://www.ams.org/
mathscinet/search/author.
html?mrauthid=652762.

Castro:2012:EDP

[CM12] M. H. Castro and V. A. Menegatto. Eigenvalue decay of positive integral operators on the sphere. Mathematics of Computation, 81(280):2303– 2317, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-280/
S0025-5718-2012-02595-6;
http://www.ams.org/journals/

mcom/2012-81-280/S0025-5718-
2012-02595-6/S0025-5718-2012-
02595-6.pdf.

Castillo:2013:GRS

[CM13] Kenier Castillo and Francisco Marcellán. Generators of rational spectral transformations for nontrivial C-functions. Mathematics of Computation, 82(282): 1057–1068, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02655-X;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02655-X/S0025-5718-2012-
02655-X.pdf.

Chui:2014:SFE

[CM14] Charles K. Chui and H. N.

Mhaskar. Smooth function extension based on high dimensional unstructured data. Mathematics of Computation, 83 (290):2865–2891, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02819-6;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02819-6/S0025-5718-2014-
02819-6.pdf.

Chan:2019:GBF

[CM19] Andrew J. Chan and Diane Maclagan. Gröbner bases over fields with valuations. Mathematics of Computation, 88(315): 467–483, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03321-X;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03321-X/S0025-5718-2018-
03321-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Chan%
2C%20Andrew%20J.;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
607134.

Cockburn:2013:LPE

[CMQ13] Bernardo Cockburn, Ivan Merev, and Jianliang Qian. Local a posteriori error estimates for time-dependent Hamilton– Jacobi equations. Mathematics of Computation, 82(281):187– 212, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-281/
S0025-5718-2012-02610-X;
http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-
2012-02610-X/S0025-5718-2012-
02610-X.pdf.

Coclite:2010:CEO

[CMR10] G. M. Coclite, S. Mishra, and N. H. Risebro. Convergence of an Engquist–Osher scheme for a multi-dimensional triangular system of conservation laws. Mathematics of Computation, 79(269):71– 94, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-269/

S0025-5718-09-02251-0/;
http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02251-0/S0025-5718-09-
02251-0.pdf.

Condette:2011:SAP

[CMS11] Nicolas Condette, Christof Melcher, and Endre Süli. Spectral approximation of patternforming nonlinear evolution equations with double-well potentials of quadratic growth. Mathematics of Computation, 80(273):205–223, January 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-10-02365-3/
;   http://www.ams.org/journals/
mcom/2011-80-273/S0025-5718-
10-02365-3/S0025-5718-10-
02365-3.pdf.

Cheng:2010:BSS

[CMSC10] Qi Cheng, Xianmeng Meng, Celi Sun, and Jiazhe Chen. Bounding the sum of square roots via lattice reduction. Mathematics of Computation, 79(270): 1109–1122, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02304-7/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02304-7/S0025-5718-09-
02304-7.pdf.

Costa:2019:RCE

[CMSV19] Edgar Costa, Nicolas Mascot, Jeroen Sijsling, and John Voight. Rigorous computation of the endomorphism ring of a Jacobian. Mathematics of Computation, 88(317):1303– 1339, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-317/
S0025-5718-2018-03373-7;
https://www.ams.org/journals/
mcom/2019-88-317/S0025-5718-
2018-03373-7/S0025-5718-2018-
03373-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1040021;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1041071;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=727424;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
974789.

Chartier:2016:IEE

[CMTZ16] Philippe Chartier, Florian Méhats, Mechthild Thalhammer, and Yong Zhang. Improved error estimates for splitting methods applied to highlyoscillatory nonlinear Schrödinger equations. Mathematics of Computation, 85(302):2863– 2885, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-302/
S0025-5718-2016-03088-4;
http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-
2016-03088-4/S0025-5718-2016-
03088-4.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Zhang%

2C%20Yong;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=335517;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
601414;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=661917.

Cheng:2017:AGG

[CMZ17] Yao Cheng, Xiong Meng, and Qiang Zhang. Application of generalized Gauss– Radau projections for the local discontinuous Galerkin method for linear convectiondiffusion equations. Mathematics of Computation, 86(305): 1233–1267, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-305/
S0025-5718-2016-03141-5;
http://www.ams.org/journals/
mcom/2017-86-305/S0025-5718-
2016-03141-5/S0025-5718-2016-
03141-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1102194;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
637183;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=998988.

Chen:2016:MMN

[CNOS16] Long Chen, Ricardo H. Nochetto, Enrique Otárola, and Abner J. Salgado. Multilevel methods for nonuniformly elliptic operators and fractional diffusion. Mathematics of Computation, 85(302):2583– 2607, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-302/
                                              S0025-5718-2016-03089-6;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-302/S0025-5718-
                                              2016-03089-6/S0025-5718-2016-
                                              03089-6.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Otarola%
                                              2C%20Enrique;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=131850;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              735779;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=847180.

Coquel:2010:ESR

[CNPT10] Frédéric Coquel, Quang Long Nguyen, Marie Postel, and Quang Huy Tran. Entropysatisfying relaxation method with large time-steps for Euler IBVPs. Mathematics of Computation, 79(271):1493– 1533, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-271/
                                              S0025-5718-10-02339-2/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-271/S0025-5718-
                                              10-02339-2/S0025-5718-10-
                                              02339-2.pdf.

Canuto:2014:AFG

[CNV14] C. Canuto, R. H. Nochetto, and M. Verani. Adaptive Fourier– Galerkin methods. Mathematics of Computation, 83(288): 1645–1687, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-288/
S0025-5718-2013-02781-0;
http://www.ams.org/journals/
mcom/2014-83-288/S0025-5718-
2013-02781-0/S0025-5718-2013-
02781-0.pdf.

Cockburn:2016:CPA

[CNZ16] Bernardo Cockburn, Ricardo H.

Nochetto, and Wujun Zhang. Contraction property of adaptive hybridizable discontinuous Galerkin methods. Mathematics of Computation, 85(299): 1113–1141, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2015-03014-2;
http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2015-03014-2/S0025-5718-2015-
03014-2.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Cockburn%
2C%20Bernardo;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=131850;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
977369.

Coffey:2014:SRR

[Cof14] Mark W. Coffey. Series representation of the Riemann zeta function and other results: Complements to a paper of Crandall. Mathematics of Computation, 83(287): 1383–1395, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-287/

                                              S0025-5718-2013-02755-X;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-287/S0025-5718-
                                              2013-02755-X/S0025-5718-2013-
                                              02755-X.pdf.

Cohl:2012:TEF

[Coh12] Howard S. Cohl. Table errata to “Formulas and theorems for the special functions of mathematical physics” by W. Magnus, F. Oberhettinger & R. P. Soni (1966). Mathematics of Computation, 81(280): 2251, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-280/
                                              S0025-5718-2012-02612-3;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-280/S0025-5718-
                                              2012-02612-3/S0025-5718-2012-

02612-3.pdf. See [MOS66].

Cohen:2015:ECN

[Coh15] Henri Cohen. Exact counting of D_{ℓ} number fields with given quadratic resolvent. Mathematics of Computation, 84 (294):1933–1951, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-294/
                                              S0025-5718-2015-02920-2;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-294/S0025-5718-
                                              2015-02920-2/S0025-5718-2015-
                                              02920-2.pdf.

Christlieb:2010:IDC

[COQ10] Andrew Christlieb, Benjamin Ong, and Jing-Mei Qiu. Integral deferred correction methods constructed with high order Runge–Kutta integrators. Mathematics of Computation, 79(270):761–783, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2010-79-
270/S0025-5718-09-02276-5/
;   http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02276-5/S0025-5718-09-
02276-5.pdf.

Corn:2010:TSG

[Cor10] Patrick Corn. Tate–Shafarevich groups and K3 surfaces. Mathematics of Computation, 79(269): 563–581, January 2010. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2010-79-
269/S0025-5718-09-02264-9/
;   http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02264-9/S0025-5718-09-
02264-9.pdf.

Cosset:2010:FGC

[Cos10] Romain Cosset. Factorization with genus 2 curves. Mathematics of Computation, 79(270): 1191–1208, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02295-9/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02295-9/S0025-5718-09-
02295-9.pdf.

Celledoni:2014:MSE

[COS14] Elena Celledoni, Brynjulf Owren, and Yajuan Sun. The minimal stage, energy preserving Runge– Kutta method for polynomial Hamiltonian systems is the averaged vector field method. Mathematics of Computation, 83 (288):1689–1700, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-288/
                                              S0025-5718-2014-02805-6;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-288/S0025-5718-
                                              2014-02805-6/S0025-5718-2014-
                                              02805-6.pdf.

Cohen:2015:PCC

[COT15] Stephen D. Cohen, Tomás Oliveira e Silva, and Tim Trudgian. A proof of the conjecture of Cohen and Mullen on sums of primitive roots. Mathematics of Computation, 84(296): 2979–2986, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-296/
                                              S0025-5718-2015-02950-0;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-296/S0025-5718-
                                              2015-02950-0/S0025-5718-2015-
                                              02950-0.pdf.

Camano:2017:AAM

[COT17] Jessika Camaño, Ricardo Oyarzúa, and Giordano Tierra. Analysis of an augmented mixed-FEM for the Navier–Stokes problem. Mathematics of Computation, 86(304):589–615, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03124-5;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03124-5/S0025-5718-2016-
03124-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Oyarzua%
2C%20Ricardo;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Tierra%
2C%20Giordano;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=962509.

Cheung:2019:OAH

[CPBG19] James Cheung, Mauro Perego, Pavel Bochev, and Max Gunzburger. Optimally accurate higher-order finite element methods for polytopial approximations of domains with smooth boundaries. Mathematics of Computation, 88(319): 2187–2219, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-319/
S0025-5718-2019-03415-4;
https://www.ams.org/journals/
mcom/2019-88-319/S0025-5718-
2019-03415-4/S0025-5718-2019-
03415-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1230885;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
38390;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=78360;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
869976.

Castro:2010:SFW

[CPPT10] Manuel J. Castro, Alberto Pardo, Carlos Parés, and E. F. Toro. On some fast well-balanced first order solvers for nonconservative systems. Mathematics of Computation, 79(271):1427– 1472, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-271/
                                              S0025-5718-09-02317-5/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-271/S0025-5718-
                                              09-02317-5/S0025-5718-09-
                                              02317-5.pdf.

Chizat:2018:SAU

[CPSV18a] Lénaïc Chizat, Gabriel Peyré, Bernhard Schmitzer, and François-Xavier Vialard. Scaling algorithms for unbalanced optimal transport problems. Mathematics of Computation, 87(314): 2563–2609, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-314/
                                              S0025-5718-2018-03303-8;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-314/S0025-5718-
                                              2018-03303-8/S0025-5718-2018-
                                              03303-8.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Chizat%
                                              2C%20Lenaic;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Peyre%
                                              2C%20Gabriel;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Vialard%
                                              2C%20Francois-Xavier;   https:

//www.ams.org/mathscinet/
search/authors.html?mrauthid=
940527.

Cravero:2018:CUA

[CPSV18b] I. Cravero, G. Puppo, M. Semplice, and G. Visconti. CWENO: Uniformly accurate reconstructions for balance laws. Mathematics of Computation, 87(312): 1689–1719, October 2018. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2018-87-
312/S0025-5718-2017-03273-
7;   http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03273-7/S0025-5718-2017-
03273-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1152946;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
324467;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=355386;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
638685.

Cances:2014:PEE

[CPV14] Clément Cancès, Iuliu Sorin Pop, and Martin Vohralík. An a posteriori error estimate for vertex-centered finite volume discretizations of immiscible incompressible two-phase flow. Mathematics of Computation, 83(285):153–188, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-285/
S0025-5718-2013-02723-8;
http://www.ams.org/journals/

                                              mcom/2014-83-285/S0025-5718-
                                              2013-02723-8/S0025-5718-2013-
                                              02723-8.pdf.

Cockburn:2014:CDT

[CQ14] Bernardo Cockburn and Weifeng Qiu. Commuting diagrams for the TNT elements on cubes. Mathematics of Computation, 83(286):603–633, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-286/
                                              S0025-5718-2013-02729-9;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-286/S0025-5718-
                                              2013-02729-9/S0025-5718-2013-
                                              02729-9.pdf.

Cockburn:2014:UTS

[CQB14] Bernardo Cockburn and Vincent Quenneville-Bélair. Uniformin-time superconvergence of the HDG methods for the acoustic wave equation. Mathematics of Computation, 83 (285):65–85, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-285/
                                              S0025-5718-2013-02743-3;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-285/S0025-5718-
                                              2013-02743-3/S0025-5718-2013-
                                              02743-3.pdf.

Cockburn:2012:CSH

[CQS12] Bernardo Cockburn, Weifeng Qiu, and Ke Shi. Conditions for superconvergence of HDG methods for second-order elliptic problems. Mathematics of Computation, 81(279):

1327–1353, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2011-02550-0;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2011-02550-0/S0025-5718-2011-
02550-0.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Cockburn%
2C%20Bernardo;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Qiu%
2C%20Weifeng;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Shi%
2C%20Ke.

Cockburn:2014:PEA

[CQS14] Bernardo Cockburn, Weifeng Qiu, and Manuel Solano. A priori error analysis for HDG methods using extensions from subdomains to achieve boundary conformity. Mathematics of Computation, 83(286): 665–699, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-286/
S0025-5718-2013-02747-0;
http://www.ams.org/journals/
mcom/2014-83-286/S0025-5718-
2013-02747-0/S0025-5718-2013-
02747-0.pdf.

Chen:2011:TII

[CQT11] Qingshan Chen, Zhen Qin, and Roger Temam. Treatment of incompatible initial and boundary data for parabolic equations in higher dimension. Mathematics of Computation, 80(276):

2071–2096, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              276/S0025-5718-2011-02469-
                                              5/;   http://www.ams.org/
                                              journals/mcom/2011-80-276/
                                              S0025-5718-2011-02469-5/S0025-
                                              5718-2011-02469-5.pdf;   http:
                                              //www.ams.org/mathscinet-
                                              getitem?mr=2813349.

Carstensen:2011:OAM

[CR11a] Carsten Carstensen and Hella Rabus. An optimal adaptive mixed finite element method. Mathematics of Computation, 80(274):649–667, April 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              274/S0025-5718-2010-02397-
                                              X/;   http://www.ams.org/
                                              journals/mcom/2011-80-274/
                                              S0025-5718-2010-02397-X/S0025-
                                              5718-2010-02397-X.pdf.

Currie:2011:PDC

[CR11b] James Currie and Narad Rampersad. A proof of Dejean’s conjecture. Mathematics of Computation, 80(274): 1063–1070, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              274/S0025-5718-2010-02407-
                                              X/;   http://www.ams.org/
                                              journals/mcom/2011-80-274/
                                              S0025-5718-2010-02407-X/S0025-
                                              5718-2010-02407-X.pdf.

Cosset:2015:CIP

[CR15] Romain Cosset and Damien Robert. Computing (ℓ, ℓ)- isogenies in polynomial time on Jacobians of genus 2 curves. Mathematics of Computation, 84(294):1953–1975, 2015. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2015-84-
294/S0025-5718-2014-02899-
8;   http://www.ams.org/journals/
mcom/2015-84-294/S0025-5718-
2014-02899-8/S0025-5718-2014-
02899-8.pdf.

Christiansen:2016:HOF

[CR16] Snorre H. Christiansen and Francesca Rapetti. On high order finite element spaces of differential forms. Mathematics of Computation, 85 (298):517–548, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-02995-0;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-02995-0/S0025-5718-2015-
02995-0.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=602457;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
663397.

Creutz:2014:SDE

[Cre14] Brendan Creutz. Second pdescents on elliptic curves. Mathematics of Computation, 83(285):365–409, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-285/
                                              S0025-5718-2013-02713-5;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-285/S0025-5718-
                                              2013-02713-5/S0025-5718-2013-
                                              02713-5.pdf.

Cohen:2010:OHN

[CS10] Graeme L. Cohen and Ronald M. Sorli. Odd harmonic numbers exceed 10^{24}. Mathematics of Computation, 79(272):2451– 2460, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-272/
                                              S0025-5718-10-02337-9/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-272/S0025-5718-
                                              10-02337-9/S0025-5718-10-
                                              02337-9.pdf.

Chernov:2013:FOM

[CS13a] Alexey Chernov and Christoph Schwab. First order k-th moment finite element analysis of nonlinear operator equations with stochastic data. Mathematics of Computation, 82 (284):1859–1888, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-284/
                                              S0025-5718-2013-02692-0;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-284/S0025-5718-
                                              2013-02692-0/S0025-5718-2013-
                                              02692-0.pdf.

Choirat:2013:CAC

[CS13b] Christine Choirat and Raffaello Seri. Computational aspects of Cui–Freeden statistics for equidistribution on the sphere. Mathematics of Computation, 82(284):2137–2156, 2013. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2013-82-
284/S0025-5718-2013-02698-
1;   http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02698-1/S0025-5718-2013-
02698-1.pdf.

Cockburn:2013:CSH

[CS13c] Bernardo Cockburn and Ke Shi.

Conditions for superconvergence of HDG methods for Stokes flow. Mathematics of Computation, 82(282):651– 671, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02644-5;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02644-5/S0025-5718-2012-
02644-5.pdf.

Cockburn:2014:DCH

[CS14] Bernardo Cockburn and Francisco-Javier Sayas. Divergenceconforming HDG methods for Stokes flows. Mathematics of Computation, 83(288): 1571–1598, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-288/
S0025-5718-2014-02802-0;
http://www.ams.org/journals/

                                              mcom/2014-83-288/S0025-5718-
                                              2014-02802-0/S0025-5718-2014-
                                              02802-0.pdf.

Chen:2015:DFL

[CS15a] Yannan Chen and Wenyu Sun. A dwindling filter line search method for unconstrained optimization. Mathematics of Computation, 84 (291):187–208, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-291/
                                              S0025-5718-2014-02847-0;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-291/S0025-5718-
                                              2014-02847-0/S0025-5718-2014-
                                              02847-0.pdf.

Cuntz:2015:RPN

[CS15b] Michael Cuntz and Christian Stump. On root posets for noncrystallographic root systems. Mathematics of Computation, 84(291):485–503, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-291/
                                              S0025-5718-2014-02841-X;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-291/S0025-5718-
                                              2014-02841-X/S0025-5718-2014-
                                              02841-X.pdf.

Cai:2018:EEF

[CS18] Yongyong Cai and Jie Shen. Error estimates for a fully discretized scheme to a Cahn– Hilliard phase-field model for two-phase incompressible flows. Mathematics of Computation, 87(313):2057–2090, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/
2018-87-313/S0025-5718-2017-
03280-4;   http://www.ams.
org/journals/mcom/2018-87-
313/S0025-5718-2017-03280-
4/S0025-5718-2017-03280-4.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=257933;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
819002.

Chenoweth:2013:IIS

[CSO13] Samuel K. M. Chenoweth, Julio Soria, and Andrew Ooi. An improved interpolation scheme for finite volume simulations on unstructured meshes. Mathematics of Computation, 82(282): 803–830, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02639-1;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02639-1/S0025-5718-2012-
02639-1.pdf.

Calvetti:2015:VOS

[CSS15] Daniela Calvetti, Erkki Somersalo, and Ruben Spies. Variable order smoothness priors for illposed inverse problems. Mathematics of Computation, 84 (294):1753–1773, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-294/
S0025-5718-2014-02909-8;

                                              http://www.ams.org/journals/
                                              mcom/2015-84-294/S0025-5718-
                                              2014-02909-8/S0025-5718-2014-
                                              02909-8.pdf.

Chang:2012:PAF

[CSV12] Xiao-Wen Chang, Damien Stehlé, and Gilles Villard. Perturbation analysis of the QR factor R in the context of LLL lattice basis reduction. Mathematics of Computation, 81(279): 1487–1511, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-279/
                                              S0025-5718-2012-02545-2;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-279/S0025-5718-
                                              2012-02545-2/S0025-5718-2012-
                                              02545-2.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Chang%
                                              2C%20Xiao-Wen;   http://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Stehle%
                                              2C%20Damien;   http://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Villard%
                                              2C%20Gilles.

Chen:2016:GJF

[CSW16] Sheng Chen, Jie Shen, and Li-Lian Wang. Generalized Jacobi functions and their applications to fractional differential equations. Mathematics of Computation, 85(300): 1603–1638, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-300/
                                              S0025-5718-2015-03035-X;

http://www.ams.org/journals/
mcom/2016-85-300/S0025-5718-
2015-03035-X/S0025-5718-2015-
03035-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Chen%
2C%20Sheng;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=257933;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
681795.

Chen:2010:CAJ

[CT10] Yanping Chen and Tao Tang.

Convergence analysis of the Jacobi spectral-collocation methods for Volterra integral equations with a weakly singular kernel. Mathematics of Computation, 79(269):147– 167, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-269/
S0025-5718-09-02269-8/;
http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02269-8/S0025-5718-09-
02269-8.pdf.

Chatzipantelidis:2012:SEE

[CT12] R. D. Lazarov P. Chatzipantelidis and V. Thomée. Some error estimates for the lumped mass finite element method for a parabolic problem. Mathematics of Computation, 81(277): 1–20, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02503-2;
http://www.ams.org/journals/

                                              mcom/2012-81-277/S0025-5718-
                                              2011-02503-2/S0025-5718-2011-
                                              02503-2.pdf.

Covanov:2019:FIM

[CT19] Svyatoslav Covanov and Emmanuel Thomé. Fast integer multiplication using generalized Fermat primes. Mathematics of Computation, 88(317):1449– 1477, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-317/
                                              S0025-5718-2018-03367-1;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-317/S0025-5718-
                                              2018-03367-1/S0025-5718-2018-
                                              03367-1.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Thome%
                                              2C%20Emmanuel;   https://
                                              www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=1105937.

Capuano:2019:ECP

[CVZ19] Laura Capuano, Francesco Veneziano, and Umberto Zannier. An effective criterion for periodicity of ℓ-adic continued fractions. Mathematics of Computation, 88(318): 1851–1882, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-318/
                                              S0025-5718-2018-03385-3;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-318/S0025-5718-
                                              2018-03385-3/S0025-5718-2018-
                                              03385-3.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1146921;
                                               https://www.ams.org/mathscinet/

search/authors.html?mrauthid=
186540;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=966417.

Chrysafinos:2010:DGA

[CW10] Konstantinos Chrysafinos and Noel J. Walkington. Discontinuous Galerkin approximations of the Stokes and Navier– Stokes equations. Mathematics of Computation, 79(272):2135– 2167, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02348-3/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02348-3/S0025-5718-10-
02348-3.pdf.

Costello:2015:UBJ

[CW15] Fintan Costello and Paul Watts.

An upper bound on Jacobsthal’s function. Mathematics of Computation, 84(293): 1389–1399, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02896-2;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02896-2/S0025-5718-2014-
02896-2.pdf.

Chen:2017:MDC

[CW17] Wenbin Chen and Yanqiu Wang. Minimal degree H(curl) and H(div) conforming finite elements on polytopal meshes. Mathematics of Computation, 86(307):2053–2087, 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-
307/S0025-5718-2016-03152-
X;   http://www.ams.org/journals/
mcom/2017-86-307/S0025-5718-
2016-03152-X/S0025-5718-2016-
03152-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Chen%
2C%20Wenbin;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=670715.

Chen:2014:LFE

[CWD14] Jie Chen, Desheng Wang, and Qiang Du. Linear finite element superconvergence on simplicial meshes. Mathematics of Computation, 83(289): 2161–2185, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-289/
S0025-5718-2014-02810-X;
http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-
2014-02810-X/S0025-5718-2014-
02810-X.pdf.

Chen:2016:FDL

[CWX16] William Y. C. Chen, Larry X. W. Wang, and Gary Y. B. Xie. Finite differences of the logarithm of the partition function. Mathematics of Computation, 85(298):825–847, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-02999-8;
http://www.ams.org/journals/

mcom/2016-85-298/S0025-5718-
2015-02999-8/S0025-5718-2015-
02999-8.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Chen%
2C%20William%20Y.%20C;   http:
//www.ams.org/mathscinet/
search/author.html?authorName=
Xie%2C%20Gary%20Y.%20B;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
845775.

Chen:2015:CHO

[CXZ15] Zhongying Chen, Yuesheng Xu, and Yuanyuan Zhang. A construction of higher-order finite volume methods. Mathematics of Computation, 84 (292):599–628, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-292/
S0025-5718-2014-02881-0;
http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02881-0/S0025-5718-2014-
02881-0.pdf.

Chen:2016:CPM

[CXZ16] Zhiming Chen, Xueshuang Xiang, and Xiaohui Zhang. Convergence of the PML method for elastic wave scattering problems. Mathematics of Computation, 85(302):2687–2714, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-302/
S0025-5718-2016-03100-2;
http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-
2016-03100-2/S0025-5718-2016-

                                              03100-2.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Chen%
                                              2C%20Zhiming;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Zhang%
                                              2C%20Xiaohui;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=999656.

Cai:2018:UEE

[CY18] Yongyong Cai and Yongjun Yuan. Uniform error estimates of the conservative finite difference method for the Zakharov system in the subsonic limit regime. Mathematics of Computation, 87(311): 1191–1225, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-311/
                                              S0025-5718-2017-03269-5;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-311/S0025-5718-
                                              2017-03269-5/S0025-5718-2017-
                                              03269-5.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=819002;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              970837.

Chen:2010:LMM

[CZ10] Xianjin Chen and Jianxin Zhou. A local min-max-orthogonal method for finding multiple solutions to noncooperative elliptic systems. Mathematics of Computation, 79(272):2213– 2236, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-272/

S0025-5718-10-02336-7/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02336-7/S0025-5718-10-
02336-7.pdf.

Castro:2011:FIS

[CZ11] Carlos Castro and Enrique Zuazua. Flux identification for 1-d scalar conservation laws in the presence of shocks. Mathematics of Computation, 80(276): 2025–2070, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02465-
8/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02465-8/S0025-
5718-2011-02465-8.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813348.

Cai:2012:MMS

[CZ12] Zhiqiang Cai and Shun Zhang.

Mixed methods for stationary Navier–Stokes equations based on pseudostress-pressure-velocity formulation. Mathematics of Computation, 81(280): 1903–1927, October 2012. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2012-81-
280/S0025-5718-2012-02585-
3;   http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-
2012-02585-3/S0025-5718-2012-
02585-3.pdf.

Chu:2014:ADS

[CZ14a] Wenchang Chu and Wenlong Zhang. Accelerating Dougall’s _{5}F_{4}-sum and infinite series involving π. Mathematics of Computation, 83 (285):475–512, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-285/
S0025-5718-2013-02701-9;
http://www.ams.org/journals/
mcom/2014-83-285/S0025-5718-
2013-02701-9/S0025-5718-2013-
02701-9.pdf.

Cockburn:2014:PEE

[CZ14b] Bernardo Cockburn and Wujun Zhang. An a posteriori error estimate for the variable-degree Raviart–Thomas method. Mathematics of Computation, 83 (287):1063–1082, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-287/
S0025-5718-2013-02789-5;
http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02789-5/S0025-5718-2013-
02789-5.pdf.

Cao:2016:SLD

[CZ16] Waixiang Cao and Zhimin Zhang. Superconvergence of Local Discontinuous Galerkin methods for one-dimensional linear parabolic equations. Mathematics of Computation, 85(297):63–84, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-297/
S0025-5718-2015-02975-5;

http://www.ams.org/journals/
mcom/2016-85-297/S0025-5718-
2015-02975-5/S0025-5718-2015-
02975-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Cao%
2C%20Waixiang;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=303173.

Dabbaghian:2010:CMR

[DD10] Vahid Dabbaghian and John D.

Dixon. Computing matrix representations. Mathematics of Computation, 79(271):1801– 1810, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-10-02330-6/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
10-02330-6/S0025-5718-10-
02330-6.pdf.

Pietro:2017:HHO

[DD17] Daniele A. Di Pietro and Jérôme Droniou. A Hybrid High-Order method for Leray– Lions elliptic equations on general meshes. Mathematics of Computation, 86(307): 2159–2191, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-307/
S0025-5718-2016-03180-4;
http://www.ams.org/journals/
mcom/2017-86-307/S0025-5718-
2016-03180-4/S0025-5718-2016-
03180-4.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Di%
20Pietro%2C%20Daniele%20A.

                                              ;   http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              655312.

Dokchitser:2019:TCG

[DD19] Tim Dokchitser and Christopher Doris. 3-torsion and conductor of genus 2 curves. Mathematics of Computation, 88(318): 1913–1927, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-318/
                                              S0025-5718-2018-03387-7;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-318/S0025-5718-
                                              2018-03387-7/S0025-5718-2018-
                                              03387-7.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Doris%
                                              2C%20Christopher;   https:
                                              //www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              733080.

deDios:2014:MPD

[dDHZZ14] Blanca Ayuso de Dios, Michael Holst, Yunrong Zhu, and Ludmil Zikatanov. Multilevel preconditioners for discontinuous, Galerkin approximations of elliptic problems, with jump coefficients. Mathematics of Computation, 83(287): 1083–1120, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-287/
                                              S0025-5718-2013-02760-3;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-287/S0025-5718-
                                              2013-02760-3/S0025-5718-2013-
                                              02760-3.pdf.

Darmon:2015:ACH

[DDLR15] Henri Darmon, Michael Daub, Sam Lichtenstein, and Victor Rotger. Algorithms for Chow–Heegner points via iterated integrals. Mathematics of Computation, 84(295): 2505–2547, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-295/
S0025-5718-2015-02927-5;
http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-
2015-02927-5/S0025-5718-2015-
02927-5.pdf.

DiPietro:2010:DFA

[DE10] Daniele A. Di Pietro and Alexandre Ern. Discrete functional analysis tools for Discontinuous Galerkin methods with application to the incompressible Navier–Stokes equations. Mathematics of Computation, 79(271):1303– 1330, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-10-02333-1/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
10-02333-1/S0025-5718-10-
02333-1.pdf.

Dziuk:2013:EES

[DE13] Gerhard Dziuk and Charles M.

Elliott. L^{2}-estimates for the evolving surface finite element method. Mathematics of Computation, 82(281): 1–24, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-281/
                                              S0025-5718-2012-02601-9;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-281/S0025-5718-
                                              2012-02601-9/S0025-5718-2012-
                                              02601-9.pdf.

Debussche:2011:WAS

[Deb11] Arnaud Debussche. Weak approximation of stochastic partial differential equations: the nonlinear case. Mathematics of Computation, 80(273): 89–117, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              273/S0025-5718-2010-02395-
                                              6/;   http://www.ams.org/
                                              journals/mcom/2011-80-273/
                                              S0025-5718-2010-02395-6/S0025-
                                              5718-2010-02395-6.pdf.

Derenthal:2014:FAP

[DEJ14] Ulrich Derenthal, Andreas-Stephan Elsenhans, and Jörg Jahnel. On the factor alpha in Peyre’s constant. Mathematics of Computation, 83 (286):965–977, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-286/
                                              S0025-5718-2013-02772-X;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-286/S0025-5718-
                                              2013-02772-X/S0025-5718-2013-
                                              02772-X.pdf.

Dembele:2014:CAM

[Dem14] Lassina Dembélé. On the computation of algebraic modular forms on compact inner forms of GSp_{4}. Mathematics of Computation, 83(288): 1931–1950, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-288/
S0025-5718-2014-02374-0;
http://www.ams.org/journals/
mcom/2014-83-288/S0025-5718-
2014-02374-0/S0025-5718-2014-
02374-0.pdf.

Deckelnick:2019:HJE

[DEMS19] Klaus Deckelnick, Charles M.

Elliott, Tatsu-Hiko Miura, and Vanessa Styles. Hamilton– Jacobi equations on an evolving surface. Mathematics of Computation, 88(320):2635– 2664, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-320/
S0025-5718-2019-03420-8;
https://www.ams.org/journals/
mcom/2019-88-320/S0025-5718-
2019-03420-8/S0025-5718-2019-
03420-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1219669;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
318167;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=62960;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
653122.

Dieulefait:2014:FTE

[DF14] Luis Dieulefait and Nuno Freitas. The Fermat-type equations x^{5} + y^{5} = 2z^{p} or 3z^{p} solved through ℚ-curves. Mathematics of Computation, 83 (286):917–933, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-286/
                                              S0025-5718-2013-02731-7;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-286/S0025-5718-
                                              2013-02731-7/S0025-5718-2013-
                                              02731-7.pdf.

Delgado:2013:GFR

[DFGSL13] M. Delgado, J. I. Farrán, P. A. García-Sánchez, and D. Llena. On the generalized Feng–Rao numbers of numerical semigroups generated by intervals. Mathematics of Computation, 82(283):1813– 1836, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-283/
                                              S0025-5718-2013-02673-7;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-283/S0025-5718-
                                              2013-02673-7/S0025-5718-2013-
                                              02673-7.pdf.

Detinko:2018:ZDC

[DFH18] A. Detinko, D. L. Flannery, and A. Hulpke. Zariski density and computing in arithmetic groups. Mathematics of Computation, 87(310):967–986, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03236-1;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03236-1/S0025-5718-2017-
03236-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Hulpke%
2C%20A.;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=335525;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
350842.

Dahlke:2012:MPA

[DFR12] Stephan Dahlke, Massimo Fornasier, and Thorsten Raasch. Multilevel preconditioning and adaptive sparse solution of inverse problems. Mathematics of Computation, 81(277):419– 446, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02507-X;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02507-X/S0025-5718-2011-
02507-X.pdf.

Dopico:2016:PML

[DGKS16] Froilán M. Dopico, Javier González, Daniel Kressner, and Valeria Simoncini. Projection methods for large-scale T - Sylvester equations. Mathematics of Computation, 85(301): 2427–2455, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2016-85-

301/S0025-5718-2016-03081-
1;   http://www.ams.org/journals/
mcom/2016-85-301/S0025-5718-
2016-03081-1/S0025-5718-2016-
03081-1.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Gonzalez%
2C%20Javier;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Kressner%
2C%20Daniel;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Simoncini%
2C%20Valeria;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=664010.

Donatelli:2016:SAS

[DGM^{+}16] Marco Donatelli, Carlo Garoni, Carla Manni, Stefano Serra-Capizzano, and Hendrik Speleers. Spectral analysis and spectral symbol of matrices in isogeometric collocation methods. Mathematics of Computation, 85(300):1639– 1680, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-300/
S0025-5718-2015-03027-0;
http://www.ams.org/journals/
mcom/2016-85-300/S0025-5718-
2015-03027-0/S0025-5718-2015-
03027-0.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Serra-
Capizzano%2C%20Stefano;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
1021672;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=119310;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=

751539;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=778374.

deGraaf:2014:CSO

[dGO14] Willem A. de Graaf and Francesco Oriente. Classifying semisimple orbits of θ-groups. Mathematics of Computation, 83(289):2509–2526, 2014. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2014-83-
289/S0025-5718-2014-02812-
3;   http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-
2014-02812-3/S0025-5718-2014-
02812-3.pdf.

Dieulefait:2010:PMG

[DGP10] Luis Dieulefait, Lucio Guerberoff, and Ariel Pacetti. Proving modularity for a given elliptic curve over an imaginary quadratic field. Mathematics of Computation, 79(270):1145– 1170, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02291-1/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02291-1/S0025-5718-09-
02291-1.pdf.

Demlow:2011:LEE

[DGS11] Alan Demlow, Johnny Guzmán, and Alfred H. Schatz. Local energy estimates for the finite element method on sharply varying grids. Mathematics of Computation, 80(273):1–9, January 2011.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02353-
1/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02353-1/S0025-
5718-2010-02353-1.pdf.

Demkowicz:2012:PEO

[DGS12] L. Demkowicz, J. Gopalakrishnan, and J. Schöberl. Polynomial extension operators. Part III. Mathematics of Computation, 81(279):1289– 1326, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2011-02536-6;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2011-02536-6/S0025-5718-2011-
02536-6.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Demkowicz%
2C%20L;   http://www.ams.
org/mathscinet/search/authors.
html?authorName=Gopalakrishnan%
2C%20J;   http://www.ams.
org/mathscinet/search/authors.
html?authorName=Schoberl%
2C%20J..

deHoog:2011:DMF

[dHAL11] F. R. de Hoog, R. S. Anderssen, and M. A. Lukas. Differentiation of matrix functionals using triangular factorization. Mathematics of Computation, 80(275):1585–1600, July 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.

org/journals/mcom/2011-80-
275/S0025-5718-2011-02451-
8/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2011-02451-8/S0025-
5718-2011-02451-8.pdf.

Dubickas:2017:NTN

[DHJ17] Artūras Dubickas, Kevin G.

Hare, and Jonas Jankauskas. No two non-real conjugates of a Pisot number have the same imaginary part. Mathematics of Computation, 86 (304):935–950, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03103-8;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03103-8/S0025-5718-2016-
03103-8.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Dubickas%
2C%20Arturas;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Hare%
2C%20Kevin%20G;   http://
www.ams.org/mathscinet/search/
author.html?mrauthid=825362.

Donat:2011:SII

[DHMG11] R. Donat, I. Higueras, and A. Martínez-Gavara. On stability issues for IMEX schemes applied to 1D scalar hyperbolic equations with stiff reaction terms. Mathematics of Computation, 80(276):2097– 2126, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-

                                              276/S0025-5718-2011-02463-
                                              4/;   http://www.ams.org/
                                              journals/mcom/2011-80-276/
                                              S0025-5718-2011-02463-4/S0025-
                                              5718-2011-02463-4.pdf;   http:
                                              //www.ams.org/mathscinet-
                                              getitem?mr=2813350.

Dai:2017:SUS

[DHYZ17] Yu-Hong Dai, Deren Han, Xiaoming Yuan, and Wenxing Zhang. A sequential updating scheme of the Lagrange multiplier for separable convex programming. Mathematics of Computation, 86(303):315– 343, January 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-303/
                                              S0025-5718-2016-03104-X;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-303/S0025-5718-
                                              2016-03104-X/S0025-5718-2016-
                                              03104-X.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Dai%
                                              2C%20Yu-Hong;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Zhang%
                                              2C%20Wenxing;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=664477;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              729439.

Demaret:2015:OTA

[DI15] Laurent Demaret and Armin Iske. Optimal N-term approximation by linear splines over anisotropic Delaunay triangulations. Mathematics of Computation, 84(293):1241–1264, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2015-84-
293/S0025-5718-2014-02908-
6;   http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02908-6/S0025-5718-2014-
02908-6.pdf.

Diem:2011:DLP

[Die11]  Claus Diem.  On the dis-

crete logarithm problem in class groups of curves. Mathematics of Computation, 80(273): 443–475, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02281-
1/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02281-1/S0025-
5718-2010-02281-1.pdf.

DiPasquale:2018:DMS

[DiP18]  Michael DiPasquale.  Di-

mension of mixed splines on polytopal cells. Mathematics of Computation, 87(310): 905–939, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03224-5;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03224-5/S0025-5718-2017-
03224-5.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=DiPasquale%
2C%20Michael.

Dittmer:2014:SOP

[Dit14] Samuel J. Dittmer. Spoof odd perfect numbers. Mathematics of Computation, 83(289): 2575–2582, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-289/
                                              S0025-5718-2013-02793-7;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-289/S0025-5718-
                                              2013-02793-7/S0025-5718-2013-
                                              02793-7.pdf.

Dutkay:2012:FDF

[DJ12a] Dorin Ervin Dutkay and Palle E. T. Jorgensen. Fourier duality for fractal measures with affine scales. Mathematics of Computation, 81(280):2253– 2273, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-280/
                                              S0025-5718-2012-02580-4;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-280/S0025-5718-
                                              2012-02580-4/S0025-5718-2012-
                                              02580-4.pdf.

Dutkay:2012:SMC

[DJ12b] Dorin Ervin Dutkay and Palle E. T. Jorgensen. Spectral measures and Cuntz algebras. Mathematics of Computation, 81(280):2275–2301, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-280/
                                              S0025-5718-2012-02589-0;

http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-
2012-02589-0/S0025-5718-2012-
02589-0.pdf.

Debrabant:2013:SLS

[DJ13] Kristian Debrabant and Espen R. Jakobsen. Semi-Lagrangian schemes for linear and fully non-linear diffusion equations. Mathematics of Computation, 82(283):1433– 1462, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2012-02632-9;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-
2012-02632-9/S0025-5718-2012-
02632-9.pdf.

Du:2019:DGM

[DJL19] Qiang Du, Lili Ju, and Jianfang Lu. A discontinuous Galerkin method for one-dimensional time-dependent nonlocal diffusion problems. Mathematics of Computation, 88(315): 123–147, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03333-6;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03333-6/S0025-5718-2018-
03333-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1050703;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
191080;   https://www.ams.

org/mathscinet/search/authors.
html?mrauthid=645968.

Drungilas:2018:LNP

[DJŠ18] P. Drungilas, J. Jankauskas, and J. Šiurys. On Littlewood and Newman polynomial multiples of Borwein polynomials. Mathematics of Computation, 87(311): 1523–1541, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-311/
S0025-5718-2017-03258-0;
http://www.ams.org/journals/
mcom/2018-87-311/S0025-5718-
2017-03258-0/S0025-5718-2017-
03258-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Drungilas%
2C%20P.;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Siurys%2C%
20J.;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=825362.

Du:2013:PEA

[DJTZ13] Qiang Du, Lili Ju, Li Tian, and Kun Zhou. A posteriori error analysis of finite element method for linear nonlocal diffusion and peridynamic models. Mathematics of Computation, 82(284): 1889–1922, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-284/
S0025-5718-2013-02708-1;
http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02708-1/S0025-5718-2013-
02708-1.pdf.

Dick:2010:DTP

[DK10] Josef Dick and Peter Kritzer.

Duality theory and propagation rules for generalized digital nets. Mathematics of Computation, 79(270):993– 1017, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02315-1/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02315-1/S0025-5718-09-
02315-1.pdf.

Doyle:2015:CAN

[DK15] John R. Doyle and David Krumm. Computing algebraic numbers of bounded height. Mathematics of Computation, 84(296):2867–2891, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2015-84-
296/S0025-5718-2015-02954-
8;   http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02954-8/S0025-5718-2015-
02954-8.pdf.

Dick:2013:CCS

[DKMW13] Alexander Dick, Othmar Koch, Roswitha März, and Ewa Weinmüller. Convergence of collocation schemes for boundary value problems in nonlinear index 1 DAEs with a singular point. Mathematics of Computation, 82(282):893– 918, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-282/
                                              S0025-5718-2012-02637-8;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-282/S0025-5718-
                                              2012-02637-8/S0025-5718-2012-
                                              02637-8.pdf.

Deckelnick:2015:CFE

[DKS15] Klaus Deckelnick, Jakob Katz, and Friedhelm Schieweck. A C^{1}-finite element method for the Willmore flow of twodimensional graphs. Mathematics of Computation, 84(296): 2617–2643, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-296/
                                              S0025-5718-2015-02973-1;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-296/S0025-5718-
                                              2015-02973-1/S0025-5718-2015-
                                              02973-1.pdf.

Demlow:2013:LPP

[DL13] Alan Demlow and Stig Larsson. Local pointwise a posteriori gradient error bounds for the Stokes equations. Mathematics of Computation, 82(282): 625–649, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-282/
                                              S0025-5718-2012-02647-0;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-282/S0025-5718-
                                              2012-02647-0/S0025-5718-2012-
                                              02647-0.pdf.

DiPietro:2015:ECR

[DL15] Daniele A. Di Pietro and Simon Lemaire. An extension of the Crouzeix–Raviart space to general meshes with application to quasi-incompressible linear elasticity and Stokes flow. Mathematics of Computation, 84(291):1–31, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-291/
S0025-5718-2014-02861-5;
http://www.ams.org/journals/
mcom/2015-84-291/S0025-5718-
2014-02861-5/S0025-5718-2014-
02861-5.pdf.

Dick:2011:ECT

[DLPW11] Josef Dick, Gerhard Larcher, Friedrich Pillichshammer, and Henryk Woźniakowski. Exponential convergence and tractability of multivariate integration for Korobov spaces. Mathematics of Computation, 80(274):905–930, April 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02433-
0/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02433-0/S0025-
5718-2010-02433-0.pdf.

Daniels:2018:TSR

[DLRNS18] Harris B. Daniels, Álvaro Lozano-Robledo, Filip Najman, and Andrew V. Sutherland. Torsion subgroups of rational elliptic curves over the compositum of all cubic fields. Mathematics of Computation, 87 (310):425–458, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03213-0;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03213-0/S0025-5718-2017-
03213-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Lozano-
Robledo%2C%20Alvaro;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
1105200;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=852273;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
886852.

Demlow:2012:BAP

[DLSW12] A. Demlow, D. Leykekhman, A. H. Schatz, and L. B. Wahlbin. Best approximation property in the W_{∞}^{1} norm for finite element methods on graded meshes. Mathematics of Computation, 81(278): 743–764, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02546-9;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02546-9/S0025-5718-2011-
02546-9.pdf.

Dayton:2011:MZN

[DLZ11] Barry H. Dayton, Tien-Yien Li, and Zhonggang Zeng. Multiple zeros of nonlinear systems. Mathematics of Computation, 80(276):2143– 2168, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02462-
2/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02462-2/S0025-
5718-2011-02462-2.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813352.

Demlow:2010:SLP

[DM10] Alan Demlow and Charalambos Makridakis. Sharply local pointwise a posteriori error estimates for parabolic problems. Mathematics of Computation, 79(271): 1233–1262, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-10-02346-X/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
10-02346-X/S0025-5718-10-
02346-X.pdf.

DeBonis:2014:NTC

[DM14] M. C. De Bonis and G. Mastroianni. Numerical treatment of a class of systems of Fredholm integral equations on the real line. Mathematics of Computation, 83 (286):771–788, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-286/
S0025-5718-2013-02727-5;
http://www.ams.org/journals/
mcom/2014-83-286/S0025-5718-

                                              2013-02727-5/S0025-5718-2013-
                                              02727-5.pdf.

DeSilva:2019:LLC

[DMPP19] Dilum De Silva, Michael J. Mossinghoff, Vincent Pigno, and Christopher Pinner. The Lind–Lehmer constant for certain p-groups. Mathematics of Computation, 88(316):949– 972, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-316/
                                              S0025-5718-2018-03350-6;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-316/S0025-5718-
                                              2018-03350-6/S0025-5718-2018-
                                              03350-6.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1052058;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              1057528;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=319822;   https:
                                              //www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              630072.

Derickx:2019:TEC

[DN19] Maarten Derickx and Filip Najman. Torsion of elliptic curves over cyclic cubic fields. Mathematics of Computation, 88(319): 2443–2459, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-319/
                                              S0025-5718-2018-03408-1;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-319/S0025-5718-
                                              2018-03408-1/S0025-5718-2018-
                                              03408-1.pdf;   https://www.

ams.org/mathscinet/search/
authors.html?mrauthid=1040992;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
886852.

Doornik:2015:NEG

[Doo15] Jurgen A. Doornik. Numerical evaluation of the Gauss hypergeometric function by power summations. Mathematics of Computation, 84(294): 1813–1833, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-294/
S0025-5718-2014-02905-0;
http://www.ams.org/journals/
mcom/2015-84-294/S0025-5718-
2014-02905-0/S0025-5718-2014-
02905-0.pdf.

Dopico:2013:SBF

[DOZ13] Froilán M. Dopico, Vadim Olshevsky, and Pavel Zhlobich. Stability of QR-based fast system solvers for a subclass of quasiseparable rank one matrices. Mathematics of Computation, 82(284): 2007–2034, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-284/
S0025-5718-2013-02710-X;
http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02710-X/S0025-5718-2013-
02710-X.pdf.

Dieci:2015:HMT

[DP15] Luca Dieci and Alessandro Pugliese. Hermitian matrices of three parameters: perturbing coalescing eigenvalues and a numerical method. Mathematics of Computation, 84(296): 2763–2790, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-296/
S0025-5718-2015-02977-9;
http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02977-9/S0025-5718-2015-
02977-9.pdf.

Davydov:2014:TSA

[DPR14] O. Davydov, J. Prasiswa, and U. Reif. Two-stage approximation methods with extended B-splines. Mathematics of Computation, 83 (286):809–833, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-286/
S0025-5718-2013-02734-2;
http://www.ams.org/journals/
mcom/2014-83-286/S0025-5718-
2013-02734-2/S0025-5718-2013-
02734-2.pdf.

Dopico:2019:SBE

[DPV19] Froilán M. Dopico, Javier Pérez, and Paul Van Dooren. Structured backward error analysis of linearized structured polynomial eigenvalue problems. Mathematics of Computation, 88(317):1189–1228, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-317/

S0025-5718-2018-03360-9;
https://www.ams.org/journals/
mcom/2019-88-317/S0025-5718-
2018-03360-9/S0025-5718-2018-
03360-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1022232;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
176855;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=664010.

deReyna:2011:HPC

[dR11] J. Arias de Reyna. High precision computation of Riemann’s zeta function by the Riemann-Siegel formula, I. Mathematics of Computation, 80(274): 995–1009, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02426-
3/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02426-3/S0025-
5718-2010-02426-3.pdf.

Droniou:2010:NMF

[Dro10] Jérôme Droniou. A numerical method for fractal conservation laws. Mathematics of Computation, 79(269):95– 124, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-269/
S0025-5718-09-02293-5/;
http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02293-5/S0025-5718-09-
02293-5.pdf.

Duran:2011:NAF

[DRS11] Ricardo G. Durán, Rodolfo Rodríguez, and Frank Sanhueza. Numerical analysis of a finite element method to compute the vibration modes of a Reissner–Mindlin laminated plate. Mathematics of Computation, 80(275): 1239–1264, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              275/S0025-5718-2011-02456-
                                              7/;   http://www.ams.org/
                                              journals/mcom/2011-80-275/
                                              S0025-5718-2011-02456-7/S0025-
                                              5718-2011-02456-7.pdf.

Doliskani:2014:TRH

[DS14] Javad Doliskani and Éric Schost. Taking roots over high extensions of finite fields. Mathematics of Computation, 83 (285):435–446, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-285/
                                              S0025-5718-2013-02715-9;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-285/S0025-5718-
                                              2013-02715-9/S0025-5718-2013-
                                              02715-9.pdf.

Dyn:2017:GAR

[DS17] Nira Dyn and Nir Sharon. A global approach to the refinement of manifold data. Mathematics of Computation, 86(303): 375–395, January 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-
303/S0025-5718-2016-03087-
2;   http://www.ams.org/journals/
mcom/2017-86-303/S0025-5718-
2016-03087-2/S0025-5718-2016-
03087-2.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=61245;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
974347.

Dabrowski:2018:OTS

[DS18] Andrzej Dabrowski and Lucjan Szymaszkiewicz. Orders of Tate–Shafarevich groups for the Neumann–Setzer type elliptic curves. Mathematics of Computation, 87(311):1509– 1522, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-311/
S0025-5718-2017-03248-8;
http://www.ams.org/journals/
mcom/2018-87-311/S0025-5718-
2017-03248-8/S0025-5718-2017-
03248-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=357378;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
754209.

Duan:2011:PFI

[DT11] Huo-Yuan Duan and Roger C. E. Tan. On the Poincaré– Friedrichs inequality for piecewise H^{1} functions in anisotropic discontinuous Galerkin finite element methods. Mathematics of Computation, 80(273): 119–140, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              273/S0025-5718-2010-02296-
                                              3/;   http://www.ams.org/
                                              journals/mcom/2011-80-273/
                                              S0025-5718-2010-02296-3/S0025-
                                              5718-2010-02296-3.pdf.

Duan:2014:FEM

[Dua14] Huoyuan Duan. A finite element method for Reissner– Mindlin plates. Mathematics of Computation, 83(286): 701–733, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-286/
                                              S0025-5718-2013-02767-6;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-286/S0025-5718-
                                              2013-02767-6/S0025-5718-2013-
                                              02767-6.pdf.

Dumke:2017:AZQ

[Dum17] Jan H. Dumke. p-adic zeros of quintic forms. Mathematics of Computation, 86(307): 2469–2478, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-307/
                                              S0025-5718-2017-03182-3;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-307/S0025-5718-
                                              2017-03182-3/S0025-5718-2017-
                                              03182-3.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Dumke%
                                              2C%20Jan%20H..

Dupont:2011:FEM

[Dup11] Régis Dupont. Fast evaluation of modular functions using Newton iterations and the AGM. Mathematics of Computation, 80(275):1823–1847, July 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2011-01880-
6/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2011-01880-6/S0025-
5718-2011-01880-6.pdf.

Dusart:2016:ELV

[Dus16] Pierre Dusart. Estimates of ψ, θ for large values of x without the Riemann hypothesis. Mathematics of Computation, 85(298):875–888, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-03005-1;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-03005-1/S0025-5718-2015-
03005-1.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Dusart%
2C%20Pierre.

deVilliers:2016:LIO

[dVG16] Johan de Villiers and Mpfareleni Rejoyce Gavhi. Local interpolation with optimal polynomial exactness in refinement spaces. Mathematics of Computation, 85 (298):759–782, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-03006-3;

                                              http://www.ams.org/journals/
                                              mcom/2016-85-298/S0025-5718-
                                              2015-03006-3/S0025-5718-2015-
                                              03006-3.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Gavhi%
                                              2C%20Mpfareleni%20Rejoyce;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              228876.

daVeiga:2019:VES

[dVMR19] L. Beirão da Veiga, D. Mora, and G. Rivera. Virtual elements for a shear-deflection formulation of Reissner–Mindlin plates. Mathematics of Computation, 88(315):149–178, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

                                              www.ams.org/journals/mcom/
                                              2019-88-315/S0025-5718-2018-
                                              03331-2;   https://www.ams.
                                              org/journals/mcom/2019-88-
                                              315/S0025-5718-2018-03331-
                                              2/S0025-5718-2018-03331-2.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?mrauthid=1111154;   https:
                                              //www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              696855;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=876029.

DiPietro:2015:ARL

[DVY15] Daniele A. Di Pietro, Martin Vohralík, and Soleiman Yousef. Adaptive regularization, linearization, and discretization and a posteriori error control for the two-phase Stefan problem. Mathematics of Computation, 84(291):153–186, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-291/
S0025-5718-2014-02854-8;
http://www.ams.org/journals/
mcom/2015-84-291/S0025-5718-
2014-02854-8/S0025-5718-2014-
02854-8.pdf.

deWet:2011:SCS

[dW11] Deter de Wet. Subsequence convergence in subdivision. Mathematics of Computation, 80 (274):973–994, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02380-
4/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02380-4/S0025-
5718-2010-02380-4.pdf.

Demanet:2012:FWC

[DY12] Laurent Demanet and Lexing Ying. Fast wave computation via Fourier integral operators. Mathematics of Computation, 81(279):1455– 1486, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2012-02557-9;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2012-02557-9/S0025-5718-2012-
02557-9.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Demanet%
2C%20Laurent;   http://www.
ams.org/mathscinet/search/

                                              authors.html?authorName=Ying%
                                              2C%20Lexing.

Dong:2014:SHH

[DYY14] Bo Dong, Bo Yu, and Yan Yu. A symmetric homotopy and hybrid polynomial system solving method for mixed trigonometric polynomial systems. Mathematics of Computation, 83 (288):1847–1868, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-288/
                                              S0025-5718-2013-02763-9;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-288/S0025-5718-
                                              2013-02763-9/S0025-5718-2013-
                                              02763-9.pdf.

Demanet:2016:ELC

[DZ16] Laurent Demanet and Xiangxiong Zhang. Eventual linear convergence of the Douglas–Rachford iteration for basis pursuit. Mathematics of Computation, 85 (297):209–238, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-297/
                                              S0025-5718-2015-02965-2;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-297/S0025-5718-
                                              2015-02965-2/S0025-5718-2015-
                                              02965-2.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Demanet%
                                              2C%20Laurent;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Zhang%
                                              2C%20Xiangxiong.

Desveaux:2016:WBS

[DZBK16] Vivien Desveaux, Markus Zenk, Christophe Berthon, and Christian Klingenberg. Wellbalanced schemes to capture non-explicit steady states: Ripa model. Mathematics of Computation, 85(300):1571– 1602, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-300/
S0025-5718-2016-03069-0;
http://www.ams.org/journals/
mcom/2016-85-300/S0025-5718-
2016-03069-0/S0025-5718-2016-
03069-0.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Zenk%
2C%20Markus;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=221691;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
654277;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=961115.

Elsey:2018:TDA

[EE18] Matt Elsey and Selim Esedoğlu.

Threshold dynamics for anisotropic surface energies. Mathematics of Computation, 87(312):1721– 1756, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2017-03268-3;
http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03268-3/S0025-5718-2017-
03268-3.pdf;   https://www.
ams.org/mathscinet/search/

                                              authors.html?authorName=Esedoglu%
                                              2C%20Selim;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=869207.

Elliott:2019:SOS

[EFH19] Charles M. Elliott, Hans Fritz, and Graham Hobbs. Second order splitting for a class of fourth order equations. Mathematics of Computation, 88(320):2605– 2634, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-320/
                                              S0025-5718-2019-03425-7;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-320/S0025-5718-
                                              2019-03425-7/S0025-5718-2019-
                                              03425-7.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1106591;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              1176722;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=62960.

Elman:2010:PMF

[EFP10] Howard C. Elman, Darran G. Furnival, and Catherine E. Powell. H(div) preconditioning for a mixed finite element formulation of the diffusion problem with random data. Mathematics of Computation, 79(270): 733–760, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-270/
                                              S0025-5718-09-02274-1/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-270/S0025-5718-

09-02274-1/S0025-5718-09-
02274-1.pdf.

Eymard:2010:CFE

[EGHL10] R. Eymard, T. Gallouët, R. Herbin, and J. C. Latché. A convergent finite element-finite volume scheme for the compressible Stokes problem. Part II: the isentropic case. Mathematics of Computation, 79(270): 649–675, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02310-2/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02310-2/S0025-5718-09-
02310-2.pdf.

Eymard:2012:FVS

[EGHL12] R. Eymard, T. Gallouët, R. Herbin, and A. Linke. Finite volume schemes for the biharmonic problem on general meshes. Mathematics of Computation, 81(280):2019– 2048, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-280/
S0025-5718-2012-02608-1;
http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-
2012-02608-1/S0025-5718-2012-
02608-1.pdf.

Ervin:2018:RSD

[EHR18] V. J. Ervin, N. Heuer, and J. P.

Roop. Regularity of the solution to 1-D fractional order diffusion equations. Mathematics of Computation, 87(313):2273– 2294, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-313/
S0025-5718-2018-03295-1;
http://www.ams.org/journals/
mcom/2018-87-313/S0025-5718-
2018-03295-1/S0025-5718-2018-
03295-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Roop%
2C%20J.%20P.;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=314970;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
64070.

Elsenhans:2012:RPD

[Els12] Andreas-Stephan Elsenhans. Rational points on diagonal quartic surfaces. Mathematics of Computation, 81(277):481– 492, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02500-7;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02500-7/S0025-5718-2011-
02500-7.pdf.

Efremenko:2018:MSP

[ELSW18] Klim Efremenko, J. M. Landsberg, Hal Schenck, and Jerzy Weyman. The method of shifted partial derivatives cannot separate the permanent from the determinant. Mathematics of Computation, 87(312):2037– 2045, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2017-03284-1;
http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03284-1/S0025-5718-2017-
03284-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Efremenko%
2C%20Klim;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Landsberg%
2C%20J.%20M.;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Weyman%
2C%20Jerzy;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=621581.

Errthum:2014:MPS

[Err14] Eric Errthum. Minimal polynomials of singular moduli. Mathematics of Computation, 83 (285):411–420, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-285/
S0025-5718-2013-02709-3;
http://www.ams.org/journals/
mcom/2014-83-285/S0025-5718-
2013-02709-3/S0025-5718-2013-
02709-3.pdf.

Ern:2016:DGM

[ES16] Alexandre Ern and Friedhelm Schieweck. Discontinuous Galerkin method in time combined with a stabilized finite element method in space for linear first-order PDEs. Mathematics of Computation, 85(301):2099– 2129, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-301/
                                              S0025-5718-2016-03073-2;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-301/S0025-5718-
                                              2016-03073-2/S0025-5718-2016-
                                              03073-2.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=155960;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              349433.

Espinosa:2010:EMS

[Esp10] Olivier Espinosa. On the evaluation of Matsubara sums. Mathematics of Computation, 79(271): 1709–1725, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-271/
                                              S0025-5718-09-02307-2/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-271/S0025-5718-
                                              09-02307-2/S0025-5718-09-
                                              02307-2.pdf.

Emmrich:2010:SAR

[ET10] Etienne Emmrich and Mechthild Thalhammer. Stiffly accurate Runge–Kutta methods for nonlinear evolution problems governed by a monotone operator. Mathematics of Computation, 79(270):785– 806, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-270/
                                              S0025-5718-09-02285-6/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-270/S0025-5718-
                                              09-02285-6/S0025-5718-09-
                                              02285-6.pdf.

Eichel:2011:SSS

[ETX11] Hagen Eichel, Lutz Tobiska, and Hehu Xie. Supercloseness and superconvergence of stabilized low-order finite element discretizations of the Stokes Problem. Mathematics of Computation, 80(274): 697–722, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02404-
4/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02404-4/S0025-
5718-2010-02404-4.pdf.

Egan:2016:EMS

[EW16] Judith Egan and Ian M. Wanless. Enumeration of MOLS of small order. Mathematics of Computation, 85(298): 799–824, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-03010-5;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-03010-5/S0025-5718-2015-
03010-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Wanless%
2C%20Ian%20M;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=300990.

Fayed:2014:EIP

[FA14a] Hatem A. Fayed and Amir F.

Atiya. An evaluation of the integral of the product of the error function and the normal probability density with application to the bivariate normal integral. Mathematics of Computation, 83(285):235–250, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-285/
S0025-5718-2013-02720-2;
http://www.ams.org/journals/
mcom/2014-83-285/S0025-5718-
2013-02720-2/S0025-5718-2013-
02720-2.pdf.

Fayed:2014:NSE

[FA14b] Hatem A. Fayed and Amir F. Atiya. A novel series expansion for the multivariate normal probability integrals based on Fourier series. Mathematics of Computation, 83(289): 2385–2402, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-289/
S0025-5718-2014-02844-5;
http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-
2014-02844-5/S0025-5718-2014-
02844-5.pdf.

Fan:2012:MLM

[Fan12] Jinyan Fan. The modified Levenberg–Marquardt method for nonlinear equations with cubic convergence. Mathematics of Computation, 81(277):447– 466, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02496-8;
http://www.ams.org/journals/

mcom/2012-81-277/S0025-5718-
2011-02496-8/S0025-5718-2011-
02496-8.pdf.

Fan:2014:AML

[Fan14] Jinyan Fan. Accelerating the modified Levenberg–Marquardt method for nonlinear equations. Mathematics of Computation, 83(287):1173–1187, 2014. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2014-83-
287/S0025-5718-2013-02752-
4;   http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02752-4/S0025-5718-2013-
02752-4.pdf.

Feng:2019:PAE

[FCW19] Yong Feng, Jingwei Chen, and Wenyuan Wu. The PSLQ algorithm for empirical data. Mathematics of Computation, 88(317):1479– 1501, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-317/
S0025-5718-2018-03356-7;
https://www.ams.org/journals/
mcom/2019-88-317/S0025-5718-
2018-03356-7/S0025-5718-2018-
03356-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Chen%
2C%20Jingwei;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=612618;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
802008.

Feng:2018:CGG

[Fen18] Ruyong Feng. On the computation of the Galois group of linear difference equations. Mathematics of Computation, 87(310):941–965, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-310/
                                              S0025-5718-2017-03232-4;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-310/S0025-5718-
                                              2017-03232-4/S0025-5718-2017-
                                              03232-4.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=756535.

Farashahi:2013:IDH

[FFS^{+}13] Reza R. Farashahi, Pierre-Alain Fouque, Igor E. Shparlinski, Mehdi Tibouchi, and J. Felipe Voloch. Indifferentiable deterministic hashing to elliptic and hyperelliptic curves. Mathematics of Computation, 82(281):491– 512, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-281/
                                              S0025-5718-2012-02606-8;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-281/S0025-5718-
                                              2012-02606-8/S0025-5718-2012-
                                              02606-8.pdf.

Fernandez:2017:FDN

[FG17] Julio Fernández and Josep González. Functions and differentials on the non-split Cartan modular curve of level 11. Mathematics of Computation, 86(303):437–454, January 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-303/
S0025-5718-2016-03109-9;
http://www.ams.org/journals/
mcom/2017-86-303/S0025-5718-
2016-03109-9/S0025-5718-2016-
03109-9.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=319937;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
704910.

Farouki:2012:DRR

[FGMS12] Rida T. Farouki, Carlotta Giannelli, Carla Manni, and Alessandra Sestini. Design of rational rotation–minimizing rigid body motions by Hermite interpolation. Mathematics of Computation, 81(278): 879–903, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02519-6;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02519-6/S0025-5718-2011-
02519-6.pdf.

Fromentin:2016:ETN

[FH16] Jean Fromentin and Florent Hivert. Exploring the tree of numerical semigroups. Mathematics of Computation, 85(301): 2553–2568, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2016-85-
301/S0025-5718-2015-03075-

                                              0;   http://www.ams.org/journals/
                                              mcom/2016-85-301/S0025-5718-
                                              2015-03075-0/S0025-5718-2015-
                                              03075-0.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Fromentin%
                                              2C%20Jean;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=639193.

Fuhrer:2017:CDB

[FHK17] Thomas Führer, Norbert Heuer, and Michael Karkulik. On the coupling of DPG and BEM. Mathematics of Computation, 86(307):2261–2284, 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2017-86-
                                              307/S0025-5718-2016-03170-
                                              1;   http://www.ams.org/journals/
                                              mcom/2017-86-307/S0025-5718-
                                              2016-03170-1/S0025-5718-2016-
                                              03170-1.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Heuer%
                                              2C%20Norbert;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=1017746;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              965821.

Feng:2017:FEM

[FHN17] Xiaobing Feng, Lauren Hennings, and Michael Neilan. Finite element methods for second order linear elliptic partial differential equations in nondivergence form. Mathematics of Computation, 86(307): 2025–2051, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/

journals/mcom/2017-86-307/
S0025-5718-2017-03168-9;
http://www.ams.org/journals/
mcom/2017-86-307/S0025-5718-
2017-03168-9/S0025-5718-2017-
03168-9.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Hennings%
2C%20Lauren;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=351561;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
824091.

Fuhrer:2019:UFK

[FHN19] Thomas Führer, Norbert Heuer, and Antti H. Niemi. An ultraweak formulation of the Kirchhoff–Love plate bending model and DPG approximation. Mathematics of Computation, 88(318):1587– 1619, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-318/
S0025-5718-2018-03381-6;
https://www.ams.org/journals/
mcom/2019-88-318/S0025-5718-
2018-03381-6/S0025-5718-2018-
03381-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Niemi%
2C%20Antti%20H.;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
1017746;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=314970.

Fisher:2017:HDE

[Fis17] Tom Fisher. Higher descents on an elliptic curve with a rational 2-torsion point. Mathematics of Computation, 86 (307):2493–2518, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-307/
S0025-5718-2016-03163-4;
http://www.ams.org/journals/
mcom/2017-86-307/S0025-5718-
2016-03163-4/S0025-5718-2016-
03163-4.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=678544.

Friedman:2011:EBH

[FJK11] J. S. Friedman, J. Jorgenson, and J. Kramer. An effective bound for the Huber constant for cofinite Fuchsian groups. Mathematics of Computation, 80(274):1163–1196, April 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02430-
5/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02430-5/S0025-
5718-2010-02430-5.pdf.

Fort:2015:CWL

[FJK^{+}15] Gersende Fort, Benjamin Jourdain, Estelle Kuhn, Tony Leliévre, and Gabriel Stoltz. Convergence of the Wang– Landau algorithm. Mathematics of Computation, 84(295): 2297–2327, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-295/
S0025-5718-2015-02952-4;
http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-

2015-02952-4/S0025-5718-2015-
02952-4.pdf.

Fan:2016:DGA

[FJS16] Zhitao Fan, Hui Ji, and Zuowei Shen. Dual Gramian analysis: Duality principle and unitary extension principle. Mathematics of Computation, 85 (297):239–270, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-297/
S0025-5718-2015-02987-1;
http://www.ams.org/journals/
mcom/2016-85-297/S0025-5718-
2015-02987-1/S0025-5718-2015-
02987-1.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Fan%
2C%20Zhitao;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Ji%
2C%20Hui;   http://www.ams.
org/mathscinet/search/author.
html?authorName=Shen%2C%20Zuowei.

Faber:2015:NBX

[FK15] Laura Faber and Habiba Kadiri.

New bounds for ψ(x). Mathematics of Computation, 84 (293):1339–1357, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02886-X;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02886-X/S0025-5718-2014-

02886-X.pdf. See corrigendum [FK18a].

Faber:2018:CNB

[FK18a] Laura Faber and Habiba Kadiri. Corrigendum to New bounds for ψ(x). Mathematics of Computation, 87(311):1451– 1455, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-311/
S0025-5718-2018-03340-3;
http://www.ams.org/journals/
mcom/2018-87-311/S0025-5718-
2018-03340-3/S0025-5718-2018-
03340-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1097651;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=

760548. See [FK15].

Fung:2018:SCC

[FK18b] King Cheong Fung and Ben Kane. On sign changes of cusp forms and the halting of an algorithm to construct a supersingular elliptic curve with a given endomorphism ring. Mathematics of Computation, 87(310):501–514, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03206-3;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03206-3/S0025-5718-2017-
03206-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Fung%
2C%20King%20Cheong;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
789505.

Franke:2017:PAM

[FKBJ17] Jens Franke, Thorsten Kleinjung, Jan Büthe, and Alexander Jost. A practical analytic method for calculating π(x). Mathematics of Computation, 86(308):2889–2909, 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-
308/S0025-5718-2017-03038-
6;   http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03038-6/S0025-5718-2017-
03038-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Jost%
2C%20Alexander;   http://
www.ams.org/mathscinet/search/
author.html?mrauthid=1017601;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
237434;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=704259.

Franz:2012:GSD

[FKS12] Sebastian Franz, R. Bruce Kellogg, and Martin Stynes. Galerkin and streamline diffusion finite element methods on a Shishkin mesh for a convectiondiffusion problem with corner singularities. Mathematics of Computation, 81(278): 661–685, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02526-3;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-

                                              2011-02526-3/S0025-5718-2011-
                                              02526-3.pdf.

Friedland:2018:NNH

[FL18] Shmuel Friedland and Lek-Heng Lim. Nuclear norm of higherorder tensors. Mathematics of Computation, 87(311):1255– 1281, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-311/
                                              S0025-5718-2017-03239-7;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-311/S0025-5718-
                                              2017-03239-7/S0025-5718-2017-
                                              03239-7.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=680138;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              69405.

Ford:2014:VEF

[FLM14] Kevin Ford, Florian Luca, and Pieter Moree. Values of the Euler φ-function not divisible by a given odd prime, and the distribution of Euler– Kronecker constants for cyclotomic fields. Mathematics of Computation, 83(287): 1447–1476, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-287/
                                              S0025-5718-2013-02749-4;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-287/S0025-5718-
                                              2013-02749-4/S0025-5718-2013-
                                              02749-4.pdf.

Filaseta:2012:DIP

[FM12] Michael Filaseta and Michael J.

Mossinghoff. The distance to an irreducible polynomial, II. Mathematics of Computation, 81(279):1571– 1585, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2011-02555-X;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2011-02555-X/S0025-5718-2011-
02555-X.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Filaseta%
2C%20Michael;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Mossinghoff%
2C%20Michael%20J..

Franz:2018:UFT

[FM18] Sebastian Franz and Gunar Matthies. A unified framework for time-dependent singularly perturbed problems with discontinuous Galerkin methods in time. Mathematics of Computation, 87(313):2113– 2132, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-313/
S0025-5718-2018-03326-9;
http://www.ams.org/journals/
mcom/2018-87-313/S0025-5718-
2018-03326-9/S0025-5718-2018-
03326-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=641700;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
745061.

Forsgaard:2019:LAA

[FMMdW19] Jens Forsgård, Laura Felicia Matusevich, Nathan Mehlhop, and Timo de Wolff. Lopsided approximation of amoebas. Mathematics of Computation, 88(315):485–500, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

                                              www.ams.org/journals/mcom/
                                              2019-88-315/S0025-5718-2018-
                                              03323-3;   https://www.ams.
                                              org/journals/mcom/2019-88-
                                              315/S0025-5718-2018-03323-
                                              3/S0025-5718-2018-03323-3.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Forsgard%
                                              2C%20Jens;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Mehlhop%
                                              2C%20Nathan;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1019872;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              632562.

Faustmann:2016:EMA

[FMP16] Markus Faustmann, Jens Markus Melenk, and Dirk Praetorius. Existence of H-matrix approximants to the inverses of BEM matrices: The simplelayer operator. Mathematics of Computation, 85(297): 119–152, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-297/
                                              S0025-5718-2015-02990-1;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-297/S0025-5718-

2015-02990-1/S0025-5718-2015-
02990-1.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Faustmann%
2C%20Markus;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Melenk%
2C%20Jens%20Markus;   http:
//www.ams.org/mathscinet/
search/author.html?authorName=
Praetorius%2C%20Dirk.

Feng:2011:DFE

[FN11] Xiaobing Feng and Michael Neilan. Discontinuous finite element methods for a biwave equation modeling d-wave superconductors. Mathematics of Computation, 80(275): 1303–1333, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2010-02436-
6/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2010-02436-6/S0025-
5718-2010-02436-6.pdf.

Fletcher:2012:SMO

[FNO12] S. Adam Fletcher, Pace P.

Nielsen, and Pascal Ochem. Sieve methods for odd perfect numbers. Mathematics of Computation, 81(279):1753– 1776, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2011-02576-7;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2011-02576-7/S0025-5718-2011-
02576-7.pdf;   http://www.

                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Fletcher%
                                              2C%20S.%20Adam;   http://
                                              www.ams.org/mathscinet/search/
                                              authors.html?authorName=Nielsen%
                                              2C%20Pace%20P;   http://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Ochem%
                                              2C%20Pascal.

Fontein:2011:IGF

[Fon11] Felix Fontein. The infrastructure of a global field of arbitrary unit rank. Mathematics of Computation, 80(276): 2325–2357, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              276/S0025-5718-2011-02490-
                                              7/;   http://www.ams.org/
                                              journals/mcom/2011-80-276/
                                              S0025-5718-2011-02490-7/S0025-
                                              5718-2011-02490-7.pdf;   http:
                                              //www.ams.org/mathscinet-
                                              getitem?mr=2813364.

Flammang:2015:AMM

[FR15a] V. Flammang and G. Rhin. On the absolute Mahler measure of polynomials having all zeros in a sector. III. Mathematics of Computation, 84(296): 2927–2938, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-296/
                                              S0025-5718-2015-02959-7;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-296/S0025-5718-
                                              2015-02959-7/S0025-5718-2015-
                                              02959-7.pdf.

Franek:2015:ETD

[FR15b] Peter Franek and Stefan Ratschan.

Effective topological degree computation based on interval arithmetic. Mathematics of Computation, 84(293): 1265–1290, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02877-9;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02877-9/S0025-5718-2014-
02877-9.pdf.

Freitas:2018:SBM

[Fre18] Pedro Freitas. Sharp bounds for the modulus and phase of Hankel functions with applications to Jaeger integrals. Mathematics of Computation, 87 (310):289–308, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03267-1;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03267-1/S0025-5718-2017-
03267-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Freitas%
2C%20Pedro.

Fisher:2012:LSH

[FS12] T. A. Fisher and G. F. Sills.

Local solubility and height bounds for coverings of elliptic curves. Mathematics of Computation, 81(279):1635– 1662, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2012-02587-7;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2012-02587-7/S0025-5718-2012-
02587-7.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Fisher%
2C%20T.%20A;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Sills%
2C%20G.%20F..

Fu:2019:ALE

[FSX19] Pei Fu, Gero Schnücke, and Yinhua Xia. Arbitrary Lagrangian– Eulerian discontinuous Galerkin method for conservation laws on moving simplex meshes. Mathematics of Computation, 88(319): 2221–2255, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-319/
S0025-5718-2019-03417-8;
https://www.ams.org/journals/
mcom/2019-88-319/S0025-5718-
2019-03417-8/S0025-5718-2019-
03417-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Fu%
2C%20Pei;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Schnucke%
2C%20Gero;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=819188.

Fukushima:2011:PFC

[Fuk11] Toshio Fukushima. Precise and fast computation of the general complete elliptic integral of the second kind. Mathematics of Computation, 80(275): 1725–1743, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2011-02455-
5/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2011-02455-5/S0025-
5718-2011-02455-5.pdf.

Fukushima:2012:SES

[Fuk12] Toshio Fukushima. Series expansions of symmetric elliptic integrals. Mathematics of Computation, 81(278):957– 990, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02531-7;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02531-7/S0025-5718-2011-
02531-7.pdf.

Feng:2011:DGM

[FW11] Xiaobing Feng and Haijun Wu. hp-discontinuous Galerkin methods for the Helmholtz equation with large wave number. Mathematics of Computation, 80(276):1997– 2024, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02475-
0/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02475-0/S0025-
5718-2011-02475-0.pdf;   http:

                                              //www.ams.org/mathscinet-
                                              getitem?mr=2813347.

Falk:2014:LBC

[FW14] Richard S. Falk and Ragnar Winther. Local bounded cochain projections. Mathematics of Computation, 83(290): 2631–2656, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-290/
                                              S0025-5718-2014-02827-5;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-290/S0025-5718-
                                              2014-02827-5/S0025-5718-2014-
                                              02827-5.pdf.

Feng:2013:ASL

[FX13] Xiaobing Feng and Yulong Xing. Absolutely stable local discontinuous Galerkin methods for the Helmholtz equation with large wave number. Mathematics of Computation, 82(283): 1269–1296, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-283/
                                              S0025-5718-2012-02652-4;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-283/S0025-5718-
                                              2012-02652-4/S0025-5718-2012-
                                              02652-4.pdf.

Gallistl:2017:SSP

[Gal17] Dietmar Gallistl. Stable splitting of polyharmonic operators by generalized Stokes systems. Mathematics of Computation, 86(308):2555–2577, 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-
308/S0025-5718-2017-03208-
7;   http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03208-7/S0025-5718-2017-
03208-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1020312.

Gallistl:2019:NAP

[Gal19a] Dietmar Gallistl. Numerical approximation of planar oblique derivative problems in nondivergence form. Mathematics of Computation, 88(317):1091– 1119, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-317/
S0025-5718-2018-03371-3;
https://www.ams.org/journals/
mcom/2019-88-317/S0025-5718-
2018-03371-3/S0025-5718-2018-
03371-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1020312.

Gallistl:2019:RRA

[Gal19b] Dietmar Gallistl. Rayleigh– Ritz approximation of the inf– sup constant for the divergence. Mathematics of Computation, 88(315):73–89, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

www.ams.org/journals/mcom/
2019-88-315/S0025-5718-2018-
03327-0;   https://www.ams.
org/journals/mcom/2019-88-
315/S0025-5718-2018-03327-
0/S0025-5718-2018-03327-0.
pdf;   https://www.ams.org/

                                              mathscinet/search/authors.
                                              html?mrauthid=1020312.

Gautschi:2017:MPZ

[Gau17]  Walter Gautschi.  Mono-

tonicity properties of the zeros of Freud and sub-range Freud polynomials: Analytic and empirical results. Mathematics of Computation, 86 (304):855–864, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-304/
                                              S0025-5718-2016-03181-6;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-304/S0025-5718-
                                              2016-03181-6/S0025-5718-2016-
                                              03181-6.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=71975.

Gavrilyuk:2010:BRI

[Gav10a] I. P. Gavrilyuk. Book review: Introduction to interval analysis. Mathematics of Computation, 79(269):615– 616, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-269/
                                              S0025-5718-09-02327-8/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-269/S0025-5718-
                                              09-02327-8/S0025-5718-09-
                                              02327-8.pdf.

Gavrilyuk:2010:BRL

[Gav10b] I. P. Gavrilyuk. Book review: Lagrange multiplier approach to variational problems and applications. Mathematics of Computation, 79(269):617– 618, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-269/
S0025-5718-09-02328-X/;
http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02328-X/S0025-5718-09-
02328-X.pdf.

Gavrilyuk:2011:BRF

[Gav11] I. P. Gavrilyuk. Book review: Fast multipole boundary element method. Mathematics of Computation, 80(275): 1867–1869, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2011-02516-
0/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2011-02516-0/S0025-
5718-2011-02516-0.pdf.

Guo:2019:DAD

[GCL19] Zhen-Chen Guo, Eric King-Wah Chu, and Wen-Wei Lin. Doubling algorithm for the discretized Bethe–Salpeter eigenvalue problem. Mathematics of Computation, 88(319):2325– 2350, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-319/
S0025-5718-2019-03398-7;
https://www.ams.org/journals/
mcom/2019-88-319/S0025-5718-
2019-03398-7/S0025-5718-2019-
03398-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1095300;

                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              232126;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=49125.

Gomez:2014:RZP

[GG14] Domingo Gómez and Jaime Gutierrez. Recovering zeros of polynomials modulo a prime. Mathematics of Computation, 83(290):2953–2965, 2014. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2014-83-
                                              290/S0025-5718-2014-02808-
                                              1;   http://www.ams.org/journals/
                                              mcom/2014-83-290/S0025-5718-
                                              2014-02808-1/S0025-5718-2014-
                                              02808-1.pdf.

Gerencser:2017:LES

[GG17] Máté Gerencsér and István Gyöngy. Localization errors in solving stochastic partial differential equations in the whole space. Mathematics of Computation, 86(307): 2373–2397, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-307/
                                              S0025-5718-2016-03201-9;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-307/S0025-5718-
                                              2016-03201-9/S0025-5718-2016-
                                              03201-9.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Gerencser%
                                              2C%20Mate;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=230651.

Grillmeier:2019:NPC

[GG19] Hubertus Grillmeier and Günther Grün. Nonnegativity preserving convergent schemes for stochastic porous-medium equations. Mathematics of Computation, 88(317):1021– 1059, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-317/
S0025-5718-2018-03372-5;
https://www.ams.org/journals/
mcom/2019-88-317/S0025-5718-
2018-03372-5/S0025-5718-2018-
03372-5.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Grillmeier%
2C%20Hubertus;   https://
www.ams.org/mathscinet/search/
authors.html?authorName=Grun%
2C%20Gunther.

Girault:2011:MFA

[GGG11] V. Girault and F. Guillén-González. Mixed formulation, approximation and decoupling algorithm for a penalized nematic liquid crystals model. Mathematics of Computation, 80(274):781–819, April 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02429-
9/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02429-9/S0025-
5718-2010-02429-9.pdf.

Gottlieb:2015:OES

[GGH15] Sigal Gottlieb, Zachary Grant, and Daniel Higgs. Optimal explicit strong stability preserving Runge–Kutta methods with high linear order and optimal nonlinear order. Mathematics of Computation, 84(296):2743– 2761, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-296/
S0025-5718-2015-02966-4;
http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02966-4/S0025-5718-2015-
02966-4.pdf.

Guillen-Gonzalez:2019:UES

[GGRBRG19] F. Guillén-González, M. A. Rodríguez-Bellido, and D. A. Rueda-Gómez. Unconditionally energy stable fully discrete schemes for a chemorepulsion model. Mathematics of Computation, 88(319): 2069–2099, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-319/
S0025-5718-2019-03418-X;
https://www.ams.org/journals/
mcom/2019-88-319/S0025-5718-
2019-03418-X/S0025-5718-2019-
03418-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Rodriguez-
Bellido%2C%20M.%20A.;   https:
//www.ams.org/mathscinet/
search/authors.html?authorName=
Rueda-Gomez%2C%20D.%20A.;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
326792.

Griebel:2013:CST

[GH13] Michael Griebel and Helmut Harbrecht. On the construction of sparse tensor product spaces. Mathematics of Computation, 82(282):975– 994, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02638-X;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02638-X/S0025-5718-2012-
02638-X.pdf.

Geck:2015:KLC

[GH15] Meinolf Geck and Abbie Halls.

On the Kazhdan–Lusztig cells in type E_{8}. Mathematics of Computation, 84(296):3029– 3049, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-296/
S0025-5718-2015-02963-9;
http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02963-9/S0025-5718-2015-
02963-9.pdf.

Gander:2018:ASM

[GH18] Martin J. Gander and Soheil Hajian. Analysis of Schwarz methods for a hybridizable discontinuous Galerkin discretization: the many-subdomain case. Mathematics of Computation, 87(312):1635–1657, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2017-03293-2;

http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03293-2/S0025-5718-2017-
03293-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Gander%
2C%20Martin%20J.;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
1096749.

Gallouet:2018:CMS

[GHLM18] T. Gallouët, R. Herbin, J.- C. Latché, and D. Maltese. Convergence of the MAC scheme for the compressible stationary Navier– Stokes equations. Mathematics of Computation, 87(311): 1127–1163, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-311/
S0025-5718-2017-03260-9;
http://www.ams.org/journals/
mcom/2018-87-311/S0025-5718-
2017-03260-9/S0025-5718-2017-
03260-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Gallouet%
2C%20T.;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Herbin%2C%
20R.;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1074312;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
715367.

Gatica:2010:DCL

[GHS10] Gabriel N. Gatica, Norbert Heuer, and Francisco-Javier Sayas. A direct coupling of local discontinuous Galerkin and boundary element methods. Mathematics of Computation, 79(271):1369–1394, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2010-79-
271/S0025-5718-10-02309-4/
;   http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
10-02309-4/S0025-5718-10-
02309-4.pdf.

Goetz:2016:ASG

[GI16] Claus R. Goetz and Armin Iske. Approximate solutions of generalized Riemann problems for nonlinear systems of hyperbolic conservation laws. Mathematics of Computation, 85(297):35–62, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-297/
S0025-5718-2015-02970-6;
http://www.ams.org/journals/
mcom/2016-85-297/S0025-5718-
2015-02970-6/S0025-5718-2015-
02970-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Goetz%
2C%20Claus%20R;   http://
www.ams.org/mathscinet/search/
author.html?authorName=Iske%
2C%20Armin.

Gittelson:2013:ASG

[Git13] Claude Jeffrey Gittelson. An adaptive stochastic Galerkin method for random elliptic operators. Mathematics of Computation, 82(283):1515– 1541, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-283/
                                              S0025-5718-2013-02654-3;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-283/S0025-5718-
                                              2013-02654-3/S0025-5718-2013-
                                              02654-3.pdf.

Gologlu:2019:SAR

[GJ19] Faruk Göloğlu and Antoine Joux. A simplified approach to rigorous degree 2 elimination in discrete logarithm algorithms. Mathematics of Computation, 88(319):2485– 2496, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-319/
                                              S0025-5718-2018-03404-4;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-319/S0025-5718-
                                              2018-03404-4/S0025-5718-2018-
                                              03404-4.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Gologlu%
                                              2C%20Faruk;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=316495.

Gonzalez-Jimenez:2018:TRE

[GJLR18] Enrique González-Jiménez and Álvaro Lozano-Robledo. On the torsion of rational elliptic curves over quartic fields. Mathematics of Computation, 87(311): 1457–1478, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-311/
                                              S0025-5718-2017-03235-X;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-311/S0025-5718-
                                              2017-03235-X/S0025-5718-2017-

03235-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Lozano-
Robledo%2C%20Alvaro;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
703386.

Gyongy:2011:AFD

[GK11] István Gyöngy and Nicolai Krylov. Accelerated finite difference schemes for second order degenerate elliptic and parabolic problems in the whole space. Mathematics of Computation, 80(275): 1431–1458, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2011-02478-
6/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2011-02478-6/S0025-
5718-2011-02478-6.pdf.

Gillette:2019:TSF

[GK19] Andrew Gillette and Tyler Kloefkorn. Trimmed serendipity finite element differential forms. Mathematics of Computation, 88(316):583– 606, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03354-3;
https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03354-3/S0025-5718-2018-
03354-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1086521;
 https://www.ams.org/mathscinet/

                                              search/authors.html?mrauthid=
                                              991138.

Gallet:2017:PLF

[GKL^{+}17] Matteo Gallet, Christoph Koutschan, Zijia Li, Georg Regensburger, Josef Schicho, and Nelly Villamizar. Planar linkages following a prescribed motion. Mathematics of Computation, 86(303): 473–506, January 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2017-86-
                                              303/S0025-5718-2016-03120-
                                              8;   http://www.ams.org/journals/
                                              mcom/2017-86-303/S0025-5718-
                                              2016-03120-8/S0025-5718-2016-
                                              03120-8.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Koutschan%
                                              2C%20Christoph;   http://
                                              www.ams.org/mathscinet/search/
                                              author.html?authorName=Regensburger%
                                              2C%20Georg;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Villamizar%
                                              2C%20Nelly;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=1094242;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              332588;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=978116.

Griebel:2013:SEI

[GKS13] Michael Griebel, Frances Y. Kuo, and Ian H. Sloan. The smoothing effect of integration in ℝ^{d} and the ANOVA decomposition. Mathematics of Computation, 82(281):383– 400, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-281/
S0025-5718-2012-02578-6;
http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-
2012-02578-6/S0025-5718-2012-
02578-6.pdf.  See note

[GKS17b].

Griebel:2017:ADN

[GKS17a] Michael Griebel, Frances Y.

Kuo, and Ian H. Sloan. The ANOVA decomposition of a non-smooth function of infinitely many variables can have every term smooth. Mathematics of Computation, 86(306): 1855–1876, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-306/
S0025-5718-2016-03171-3;
http://www.ams.org/journals/
mcom/2017-86-306/S0025-5718-
2016-03171-3/S0025-5718-2016-
03171-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=163675;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
270664;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=703418.

Griebel:2017:NSE

[GKS17b] Michael Griebel, Frances Y.

Kuo, and Ian H. Sloan. Note on “The smoothing effect of integration in ℝ^{d} and the ANOVA decomposition”. Mathematics of Computation, 86(306): 1847–1854, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-306/
                                              S0025-5718-2016-03172-5;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-306/S0025-5718-
                                              2016-03172-5/S0025-5718-2016-
                                              03172-5.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=163675;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              270664;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=703418. See

[GKS13].

Guzman:2012:PEE

[GL12] J. Guzmán and D. Leykekhman. Pointwise error estimates of finite element approximations to the Stokes problem on convex polyhedra. Mathematics of Computation, 81(280):1879– 1902, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-280/
                                              S0025-5718-2012-02603-2;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-280/S0025-5718-
                                              2012-02603-2/S0025-5718-2012-
                                              02603-2.pdf.

Gonzalez:2015:CTC

[GL15] Oscar Gonzalez and Jun Li. A convergence theorem for a class of Nyström methods for weakly singular integral equations on surfaces in ℝ^{3}. Mathematics of Computation, 84 (292):675–714, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/

journals/mcom/2015-84-292/
S0025-5718-2014-02869-X;
http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02869-X/S0025-5718-2014-
02869-X.pdf.

Gallouet:2012:CFV

[GLL12] T. Gallouët, A. Larcher, and J. C. Latché. Convergence of a finite volume scheme for the convection-diffusion equation with L^{1} data. Mathematics of Computation, 81(279): 1429–1454, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2011-02571-8;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2011-02571-8/S0025-5718-2011-
02571-8.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Gallouet%
2C%20T;   http://www.ams.
org/mathscinet/search/authors.
html?authorName=Larcher%2C%
20A;   http://www.ams.org/
mathscinet/search/authors.
html?authorName=Latche%2C%
20J.%20C.

Griewank:2014:AEC

[GLLZ14] Andreas Griewank, Lutz Lehmann, Hernan Leovey, and Marat Zilberman. Automatic evaluations of cross-derivatives. Mathematics of Computation, 83 (285):251–274, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-285/
S0025-5718-2013-02717-2;

                                              http://www.ams.org/journals/
                                              mcom/2014-83-285/S0025-5718-
                                              2013-02717-2/S0025-5718-2013-
                                              02717-2.pdf.

Gauckler:2019:TIQ

[GLM^{+}19a] Ludwig Gauckler, Jianfeng Lu, Jeremy L. Marzuola, Frédéric Rousset, and Katharina Schratz. Trigonometric integrators for quasilinear wave equations. Mathematics of Computation, 88(316):717– 749, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-316/
                                              S0025-5718-2018-03339-7;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-316/S0025-5718-
                                              2018-03339-7/S0025-5718-2018-
                                              03339-7.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Rousset%
                                              2C%20Frederic;   https://
                                              www.ams.org/mathscinet/search/
                                              authors.html?mrauthid=787291;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              822782;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=832307;   https:
                                              //www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              990639.

Guidi:2019:FOS

[GLM19b] Federico Amadio Guidi, Sofia Lindqvist, and Giacomo Micheli. Full orbit sequences in affine spaces via fractional jumps and pseudorandom number generation. Mathematics of Computation, 88(318):2005– 2025, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-318/
S0025-5718-2018-03400-7;
https://www.ams.org/journals/
mcom/2019-88-318/S0025-5718-
2018-03400-7/S0025-5718-2018-
03400-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Amadio%
20Guidi%2C%20Federico;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
1078793;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=1177141.

Gopalakrishnan:2015:SAH

[GLNP15] J. Gopalakrishnan, F. Li, N.- C. Nguyen, and J. Peraire. Spectral approximations by the HDG method. Mathematics of Computation, 84(293): 1037–1059, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02885-8;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02885-8/S0025-5718-2014-
02885-8.pdf.

Ganesh:2011:PQM

[GLS11] M. Ganesh, Q. T. Le Gia, and I. H. Sloan. A pseudospectral quadrature method for Navier–Stokes equations on rotating spheres. Mathematics of Computation, 80(275): 1397–1430, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.

                                              org/journals/mcom/2011-80-
                                              275/S0025-5718-2010-02440-
                                              8/;   http://www.ams.org/
                                              journals/mcom/2011-80-275/
                                              S0025-5718-2010-02440-8/S0025-
                                              5718-2010-02440-8.pdf.

Gunzburger:2019:SCR

[GLW19] Max Gunzburger, Buyang Li, and Jilu Wang. Sharp convergence rates of time discretization for stochastic time-fractional PDEs subject to additive space-time white noise. Mathematics of Computation, 88(318):1715– 1741, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-318/
                                              S0025-5718-2018-03397-X;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-318/S0025-5718-
                                              2018-03397-X/S0025-5718-2018-
                                              03397-X.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1059885;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              78360;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=910552.

Ghorbel:2010:WPN

[GM10] A. Ghorbel and R. Monneau. Well-posedness and numerical analysis of a onedimensional non-local transport equation modelling dislocations dynamics. Mathematics of Computation, 79(271): 1535–1564, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/

journals/mcom/2010-79-271/
S0025-5718-10-02326-4/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
10-02326-4/S0025-5718-10-
02326-4.pdf.

Granger:2013:GMN

[GM13a] Robert Granger and Andrew Moss. Generalised Mersenne numbers revisited. Mathematics of Computation, 82(284): 2389–2420, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-284/
S0025-5718-2013-02704-4;
http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02704-4/S0025-5718-2013-
02704-4.pdf.

Guitart:2013:CFS

[GM13b] Xavier Guitart and Marc Masdeu. Continued fractions in 2-stage Euclidean quadratic fields. Mathematics of Computation, 82(282):1223– 1233, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02620-2;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02620-2/S0025-5718-2012-
02620-2.pdf.

Gaspoz:2014:ACA

[GM14] Fernando D. Gaspoz and Pedro Morin. Approximation classes for adaptive higher order finite element approximation. Mathematics of Computation, 83 (289):2127–2160, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-289/
S0025-5718-2013-02777-9;
http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-
2013-02777-9/S0025-5718-2013-
02777-9.pdf.  See errata

[GM17].

Guitart:2015:EMD

[GM15] Xavier Guitart and Marc Masdeu. Elementary matrix decomposition and the computation of Darmon points with higher conductor. Mathematics of Computation, 84 (292):875–893, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-292/
S0025-5718-2014-02853-6;
http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02853-6/S0025-5718-2014-
02853-6.pdf.

Grenie:2016:ESP

[GM16a] Loïc Grenié and Giuseppe Molteni. Explicit smoothed prime ideals theorems under GRH. Mathematics of Computation, 85(300):1875– 1899, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-300/
S0025-5718-2015-03039-7;
http://www.ams.org/journals/
mcom/2016-85-300/S0025-5718-

2015-03039-7/S0025-5718-2015-
03039-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=357391;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
712882.

Grenie:2016:EVP

[GM16b] Loïc Grenié and Giuseppe Molteni. Explicit versions of the prime ideal theorem for Dedekind zeta functions under GRH. Mathematics of Computation, 85 (298):889–906, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-03031-2;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-03031-2/S0025-5718-2015-
03031-2.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=357391;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
712882.

Grenie:2016:ZDZ

[GM16c] Loïc Grenié and Giuseppe Molteni. Zeros of Dedekind zeta functions under GRH. Mathematics of Computation, 85 (299):1503–1522, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2015-03024-5;
http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2015-03024-5/S0025-5718-2015-
03024-5.pdf;   http://www.

ams.org/mathscinet/search/
author.html?mrauthid=357391;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
712882.

Gaspoz:2017:EAC

[GM17] Fernando D. Gaspoz and Pedro Morin. Errata to “Approximation classes for adaptive higher order finite element approximation”. Mathematics of Computation, 86(305): 1525–1526, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-305/
S0025-5718-2016-03243-3;
http://www.ams.org/journals/
mcom/2017-86-305/S0025-5718-
2016-03243-3/S0025-5718-2016-
03243-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Morin%
2C%20Pedro;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=884857.

See [GM14].

Grenie:2018:EBG

[GM18] Loïc Grenié and Giuseppe Molteni. Explicit bounds for generators of the class group. Mathematics of Computation, 87(313):2483–2511, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-313/
S0025-5718-2017-03281-6;
http://www.ams.org/journals/
mcom/2018-87-313/S0025-5718-
2017-03281-6/S0025-5718-2017-
03281-6.pdf;   https://www.

ams.org/mathscinet/search/
authors.html?mrauthid=357391;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
712882.

Giesselmann:2014:ECD

[GMP14] Jan Giesselmann, Charalambos Makridakis, and Tristan Pryer. Energy consistent discontinuous Galerkin methods for the Navier–Stokes– Korteweg system. Mathematics of Computation, 83(289): 2071–2099, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-289/
S0025-5718-2014-02792-0;
http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-
2014-02792-0/S0025-5718-2014-
02792-0.pdf.

Gil:2018:EMF

[GMRL18] José Ignacio Burgos Gil, Ricardo Menares, and Juan Rivera-Letelier. On the essential minimum of Faltings’ height. Mathematics of Computation, 87(313):2425–2459, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-313/
S0025-5718-2018-03286-0;
http://www.ams.org/journals/
mcom/2018-87-313/S0025-5718-
2018-03286-0/S0025-5718-2018-
03286-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=349969;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=

670564;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=880333.

Guermond:2012:CAC

[GMS12] Jean-Luc Guermond, Peter D. Minev, and Abner J. Salgado. Convergence analysis of a class of massively parallel direction splitting algorithms for the Navier–Stokes equations in simple domains. Mathematics of Computation, 81(280):1951– 1977, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-280/
S0025-5718-2012-02588-9;
http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-
2012-02588-9/S0025-5718-2012-
02588-9.pdf.

Gonzalez-Meneses:2019:PBC

[GMS19] Juan González-Meneses and Marithania Silvero. Polynomial braid combing. Mathematics of Computation, 88(318): 2027–2045, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-318/
S0025-5718-2018-03392-0;
https://www.ams.org/journals/
mcom/2019-88-318/S0025-5718-
2018-03392-0/S0025-5718-2018-
03392-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Gonzalez-
Meneses%2C%20Juan;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
1008060.

Garoni:2017:SAS

[GMSC^{+}17] Carlo Garoni, Carla Manni, Stefano Serra-Capizzano, Debora Sesana, and Hendrik Speleers. Spectral analysis and spectral symbol of matrices in isogeometric Galerkin methods. Mathematics of Computation, 86 (305):1343–1373, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-305/
S0025-5718-2016-03143-9;
http://www.ams.org/journals/
mcom/2017-86-305/S0025-5718-
2016-03143-9/S0025-5718-2016-
03143-9.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1021672;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
119310;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=332436;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
778374;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=855861.

Gallot:2011:EME

[GMZ11] Yves Gallot, Pieter Moree, and Wadim Zudilin. The Erdős– Moser equation 1^{k} + 2^{k} + ⋯ + (m − 1)^{k} = m^{k} revisited using continued fractions. Mathematics of Computation, 80(274): 1221–1237, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02439-
1/;   http://www.ams.org/

                                              journals/mcom/2011-80-274/
                                              S0025-5718-2010-02439-1/S0025-
                                              5718-2010-02439-1.pdf.

Guzman:2014:CDF

[GN14] Johnny Guzmán and Michael Neilan. Conforming and divergence-free Stokes elements on general triangular meshes. Mathematics of Computation, 83(285):15–36, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-285/
                                              S0025-5718-2013-02753-6;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-285/S0025-5718-
                                              2013-02753-6/S0025-5718-2013-
                                              02753-6.pdf.

Gnewuch:2012:IDI

[Gne12] Michael Gnewuch. Infinitedimensional integration on weighted Hilbert spaces. Mathematics of Computation, 81(280): 2175–2205, October 2012. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2012-81-
                                              280/S0025-5718-2012-02583-
                                              X;   http://www.ams.org/journals/
                                              mcom/2012-81-280/S0025-5718-
                                              2012-02583-X/S0025-5718-2012-
                                              02583-X.pdf.

Gracia:2016:NAS

[GO16] J. L. Gracia and E. O’Riordan. Numerical approximation of solution derivatives of singularly perturbed parabolic problems of convection-diffusion type. Mathematics of Computation, 85 (298):581–599, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-02998-6;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-02998-6/S0025-5718-2015-
02998-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Gracia%
2C%20J.%20L;   http://www.
ams.org/mathscinet/search/
author.html?authorName=O'
Riordan%2C%20E..

Guzman:2018:ISS

[GO18] Johnny Guzmán and Maxim Olshanskii. Inf-sup stability of geometrically unfitted Stokes finite elements. Mathematics of Computation, 87(313):2091– 2112, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-313/
S0025-5718-2017-03288-9;
http://www.ams.org/journals/
mcom/2018-87-313/S0025-5718-
2017-03288-9/S0025-5718-2017-
03288-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=343398;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
775211.

Grau:2011:TPT

[GOM11] José María Grau and Antonio M. Oller-Marcén. An Õ(log^{2}(N)) time primality test for generalized Cullen numbers. Mathematics of Computation, 80(276):2315– 2323, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              276/S0025-5718-2011-02489-
                                              0/;   http://www.ams.org/
                                              journals/mcom/2011-80-276/
                                              S0025-5718-2011-02489-0/S0025-
                                              5718-2011-02489-0.pdf;   http:
                                              //www.ams.org/mathscinet-
                                              getitem?mr=2813363.

Grau:2015:PTN

[GOMS15] José María Grau, Antonio M. Oller-Marcén, and Daniel Sadornil. A primality test for Kp^{n} + 1 numbers. Mathematics of Computation, 84(291):505–512, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

                                              www.ams.org/journals/mcom/
                                              2015-84-291/S0025-5718-2014-
                                              02849-4;   http://www.ams.
                                              org/journals/mcom/2015-84-
                                              291/S0025-5718-2014-02849-
                                              4/S0025-5718-2014-02849-4.
                                              pdf.

Gong:2013:EEF

[Gon13] Wei Gong. Error estimates for finite element approximations of parabolic equations with measure data. Mathematics of Computation, 82(281): 69–98, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-281/
                                              S0025-5718-2012-02630-5;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-281/S0025-5718-
                                              2012-02630-5/S0025-5718-2012-
                                              02630-5.pdf.

Gould:2016:DTS

[GOP16] N. Gould, C. Ortner, and D. Packwood. A dimertype saddle search algorithm with preconditioning and linesearch. Mathematics of Computation, 85(302):2939–2966, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-302/
S0025-5718-2016-03096-3;
http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-
2016-03096-3/S0025-5718-2016-
03096-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Packwood%
2C%20D;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=75720;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
803698.

Gatica:2011:AFM

[GOS11] Gabriel N. Gatica, Ricardo Oyarzúa, and Francisco-Javier Sayas. Analysis of fullymixed finite element methods for the Stokes–Darcy coupled problem. Mathematics of Computation, 80(276):1911– 1948, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02466-
X/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02466-X/S0025-
5718-2011-02466-X.pdf;   http:

                                              //www.ams.org/mathscinet-
                                              getitem?mr=2813344.

Glaubitz:2018:AMF

[GÖS18] Jan Glaubitz, Philipp Öffner, and Thomas Sonar. Application of modal filtering to a spectral difference method. Mathematics of Computation, 87 (310):175–207, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-310/
                                              S0025-5718-2017-03257-9;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-310/S0025-5718-
                                              2017-03257-9/S0025-5718-2017-
                                              03257-9.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Glaubitz%
                                              2C%20Jan;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?authorName=Offner%2C%
                                              20Philipp;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=324309.

Gudi:2014:PEC

[GP14] Thirupathi Gudi and Kamana Porwal. A posteriori error control of discontinuous Galerkin methods for elliptic obstacle problems. Mathematics of Computation, 83 (286):579–602, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-286/
                                              S0025-5718-2013-02728-7;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-286/S0025-5718-
                                              2013-02728-7/S0025-5718-2013-
                                              02728-7.pdf.

Gil:2017:HSB

[GP17] Jose Ignacio Burgos Gil and Ariel Pacetti. Hecke and Sturm bounds for Hilbert modular forms over real quadratic fields. Mathematics of Computation, 86(306):1949–1978, 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-
306/S0025-5718-2016-03187-
7;   http://www.ams.org/journals/
mcom/2017-86-306/S0025-5718-
2016-03187-7/S0025-5718-2016-
03187-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=349969;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
759256.

Gomez-Perez:2014:AEA

[GPOS14] Domingo Gómez-Pérez, Alina Ostafe, and Igor Shparlinski. Algebraic entropy, automorphisms and sparsity of algebraic dynamical systems and pseudorandom number generators. Mathematics of Computation, 83(287):1535–1550, 2014. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2014-83-
287/S0025-5718-2013-02780-
9;   http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02780-9/S0025-5718-2013-
02780-9.pdf.

Galbraith:2013:CDL

[GPR13] Steven D. Galbraith, John M.

Pollard, and Raminder S. Ruprai. Computing discrete logarithms in an interval. Mathematics of Computation, 82(282): 1181–1195, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02641-X;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02641-X/S0025-5718-2012-
02641-X.pdf.

Gopalakrishnan:2014:APD

[GQ14] J. Gopalakrishnan and W. Qiu. An analysis of the practical DPG method. Mathematics of Computation, 83 (286):537–552, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-286/
S0025-5718-2013-02721-4;
http://www.ams.org/journals/
mcom/2014-83-286/S0025-5718-
2013-02721-4/S0025-5718-2013-
02721-4.pdf.

Graham:2011:BRH

[Gra11] Sid Graham. Book review: A higher-dimensional sieve method with procedures for computing sieve functions. Mathematics of Computation, 80(273): 623–624, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02421-
4/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02421-4/S0025-
5718-2010-02421-4.pdf.

Grande:2019:RGR

[Gra19a] Jörg Grande. Red–green refinement of simplicial meshes in d dimensions. Mathematics of Computation, 88(316):751– 782, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03383-X;
https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03383-X/S0025-5718-2018-
03383-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Grande%
2C%20Jorg.

Gras:2019:HCD

[Gra19b]  Georges Gras.  Heuristics

and conjectures in the direction of a p-adic Brauer– Siegel Theorem. Mathematics of Computation, 88(318):1929– 1965, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-318/
S0025-5718-2018-03395-6;
https://www.ams.org/journals/
mcom/2019-88-318/S0025-5718-
2018-03395-6/S0025-5718-2018-
03395-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Gras%
2C%20Georges.

Gatica:2016:MFE

[GRBT16] Gabriel N. Gatica, Ricardo Ruiz-Baier, and Giordano Tierra. A mixed finite element method for Darcy’s equations with pressure dependent porosity. Mathematics of Computation, 85 (297):1–33, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-297/
S0025-5718-2015-02980-9;
http://www.ams.org/journals/
mcom/2016-85-297/S0025-5718-
2015-02980-9/S0025-5718-2015-
02980-9.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Gatica%
2C%20Gabriel%20N;   http:
//www.ams.org/mathscinet/
search/author.html?authorName=
Ruiz-Baier%2C%20Ricardo;
http://www.ams.org/mathscinet/
search/author.html?authorName=
Tierra%2C%20Giordano.

Greaves:2015:SSH

[Gre15] Gary Greaves. Small-span Hermitian matrices over quadratic integer rings. Mathematics of Computation, 84(291): 409–424, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-291/
S0025-5718-2014-02836-6;
http://www.ams.org/journals/
mcom/2015-84-291/S0025-5718-
2014-02836-6/S0025-5718-2014-
02836-6.pdf.

Greither:2004:BSC

[GRT04] Cornelius Greither, Xavier-François Roblot, and Brett A. Tangedal. The Brumer–Stark conjecture in some families of extensions of specified degree. Mathematics of Computation, 73(245):297–315, January 2004. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http:/

/www.ams.org/mcom/2004-73-
245/S0025-5718-03-01565-5/
home.html;   http://www.ams.
org/mcom/2004-73-245/S0025-
5718-03-01565-5/S0025-5718-
03-01565-5.dvi;   http://
www.ams.org/mcom/2004-73-
245/S0025-5718-03-01565-5/
S0025-5718-03-01565-5.pdf;
http://www.ams.org/mcom/
2004-73-245/S0025-5718-03-
01565-5/S0025-5718-03-01565-
5.ps;   http://www.ams.org/
mcom/2004-73-245/S0025-5718-
03-01565-5/S0025-5718-03-

01565-5.tex. See corrigendum [GRT15].

Greither:2015:CBS

[GRT15] Cornelius Greither, Xavier-François Roblot, and Brett A. Tangedal. Corrigendum to “The Brumer–Stark conjecture in some families of extensions of specified degree”. Mathematics of Computation, 84 (292):955–957, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-292/
S0025-5718-2014-02851-2;
http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02851-2/S0025-5718-2014-

02851-2.pdf. See [GRT04].

Garcia:2010:AGE

[GS10] A. López García and E. B.

Saff. Asymptotics of greedy energy points. Mathematics of Computation, 79(272):2287– 2316, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02358-6/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02358-6/S0025-5718-10-
02358-6.pdf.

Guzman:2019:SVF

[GS19] Johnny Guzmán and L. Ridgway Scott. The Scott– Vogelius finite elements revisited. Mathematics of Computation, 88(316):515–529, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03346-4;
https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03346-4/S0025-5718-2018-
03346-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=157720;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
775211.

Gutierrez-Santacreu:2012:UTE

[GSRM12] J. V. Gutiérrez-Santacreu and M. A. Rojas-Medar. Uniformin-time error estimates for spectral Galerkin approximations of a mass diffusion model. Mathematics of Computation, 81(277):191– 218, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/

journals/mcom/2012-81-277/
S0025-5718-2011-02491-9;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02491-9/S0025-5718-2011-
02491-9.pdf.

Guzman:2016:AFE

[GSS16] Johnny Guzmán, Manuel A.

Sánchez, and Marcus Sarkis. On the accuracy of finite element approximations to a class of interface problems. Mathematics of Computation, 85(301):2071– 2098, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-301/
S0025-5718-2015-03051-8;
http://www.ams.org/journals/
mcom/2016-85-301/S0025-5718-
2015-03051-8/S0025-5718-2015-
03051-8.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=358674;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
775211;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=895873.

Graham:2017:DDP

[GSV17] I. G. Graham, E. A. Spence, and E. Vainikko. Domain decomposition preconditioning for high-frequency Helmholtz problems with absorption. Mathematics of Computation, 86 (307):2089–2127, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-307/
S0025-5718-2017-03190-2;
http://www.ams.org/journals/

mcom/2017-86-307/S0025-5718-
2017-03190-2/S0025-5718-2017-
03190-2.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Graham%
2C%20I.%20G.;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Spence%
2C%20E.%20A.;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=176540.

Gobet:2016:LRM

[GT16] Emmanuel Gobet and Plamen Turkedjiev. Linear regression MDP scheme for discrete backward stochastic differential equations under general conditions. Mathematics of Computation, 85(299): 1359–1391, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2015-03013-0;
http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2015-03013-0/S0025-5718-2015-
03013-0.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Turkedjiev%
2C%20Plamen;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=639623.

Gudi:2010:NEA

[Gud10] Thirupathi Gudi. A new error analysis for discontinuous finite element methods for linear elliptic problems. Mathematics of Computation, 79(272):2169– 2189, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/

journals/mcom/2010-79-272/
S0025-5718-10-02360-4/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02360-4/S0025-5718-10-
02360-4.pdf.

Guillevic:2019:FID

[Gui19] Aurore Guillevic. Faster individual discrete logarithms in finite fields of composite extension degree. Mathematics of Computation, 88(317):1273– 1301, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-317/
S0025-5718-2018-03376-2;
https://www.ams.org/journals/
mcom/2019-88-317/S0025-5718-
2018-03376-2/S0025-5718-2018-
03376-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=963265.

Greenberg:2011:CSH

[GV11] Matthew Greenberg and John Voight. Computing systems of Hecke eigenvalues associated to Hilbert modular forms. Mathematics of Computation, 80(274): 1071–1092, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02423-
8/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02423-8/S0025-
5718-2010-02423-8.pdf.

Georgoulis:2015:ADG

[GV15] Emmanuil H. Georgoulis and Juha M. Virtanen. Adaptive discontinuous Galerkin approximations to fourth order parabolic problems. Mathematics of Computation, 84(295): 2163–2190, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-295/
S0025-5718-2015-02936-6;
http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-
2015-02936-6/S0025-5718-2015-
02936-6.pdf.

Guarnieri:2017:SBY

[GV17] L. Guarnieri and L. Vendramin. Skew braces and the Yang– Baxter equation. Mathematics of Computation, 86(307): 2519–2534, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-307/
S0025-5718-2016-03161-0;
http://www.ams.org/journals/
mcom/2017-86-307/S0025-5718-
2016-03161-0/S0025-5718-2016-
03161-0.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Guarnieri%
2C%20L.;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=829575.

Gao:2016:NFC

[GVW16] Shuhong Gao, Frank Volny IV, and Mingsheng Wang. A new framework for computing Gröbner bases. Mathematics of Computation, 85 (297):449–465, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/

journals/mcom/2016-85-297/
S0025-5718-2015-02969-X;
http://www.ams.org/journals/
mcom/2016-85-297/S0025-5718-
2015-02969-X/S0025-5718-2015-
02969-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Wang%
2C%20Mingsheng;   http://
www.ams.org/mathscinet/search/
author.html?mrauthid=291308;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
978150.

Gander:2017:OSM

[GX17] Martin J. Gander and Yingxiang Xu. Optimized Schwarz methods with nonoverlapping circular domain decomposition. Mathematics of Computation, 86(304):637–660, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03127-0;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03127-0/S0025-5718-2016-
03127-0.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Gander%
2C%20Martin%20J;   http://
www.ams.org/mathscinet/search/
author.html?mrauthid=730883.

Greaves:2019:ELD

[GY19] Gary R. W. Greaves and Pavlo Yatsyna. On equiangular lines in 17 dimensions and the characteristic polynomial of a Seidel matrix. Mathematics of Computation, 88(320):3041– 3061, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-320/
                                              S0025-5718-2019-03433-6;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-320/S0025-5718-
                                              2019-03433-6/S0025-5718-2019-
                                              03433-6.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1047455;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              986306.

Gao:2013:ANS

[GZ13] Hao Gao and Hongkai Zhao. Analysis of a numerical solver for radiative transport equation. Mathematics of Computation, 82(281):153–172, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2013-82-
                                              281/S0025-5718-2012-02605-
                                              6;   http://www.ams.org/journals/
                                              mcom/2013-82-281/S0025-5718-
                                              2012-02605-6/S0025-5718-2012-
                                              02605-6.pdf.

Guo:2017:HRF

[GZZ17] Hailong Guo, Zhimin Zhang, and Ren Zhao. Hessian recovery for finite element methods. Mathematics of Computation, 86(306):1671–1692, 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2017-86-
                                              306/S0025-5718-2016-03186-
                                              5;   http://www.ams.org/journals/
                                              mcom/2017-86-306/S0025-5718-
                                              2016-03186-5/S0025-5718-2016-
                                              03186-5.pdf;   http://www.
                                              ams.org/mathscinet/search/

author.html?mrauthid=1022936;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
1142013;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=303173.

Hegland:2011:DII

[HA11] Markus Hegland and Robert S.

Anderssen. Dilational interpolatory inequalities. Mathematics of Computation, 80(274): 1019–1036, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02431-
7/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02431-7/S0025-
5718-2010-02431-7.pdf.

Hairer:2010:BRN

[Hai10]  Ernst Hairer.  Book re-

view: Numerical methods for evolutionary differential equations. Mathematics of Computation, 79(269):613–614, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-269/
S0025-5718-09-02321-7/;
http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02321-7/S0025-5718-09-
02321-7.pdf.

Hakberg:2013:DKT

[Hak13] Bengt Hakberg. A discrete KPP-theory for Fisher’s equation. Mathematics of Computation, 82(282):781–802, ????

2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/
2013-82-282/S0025-5718-2012-
02642-1;   http://www.ams.
org/journals/mcom/2013-82-
282/S0025-5718-2012-02642-
1/S0025-5718-2012-02642-1.
pdf.

Han:2010:SBM

[Han10] Bin Han. The structure of balanced multivariate biorthogonal multiwavelets and dual multiframelets. Mathematics of Computation, 79(270):917– 951, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02320-5/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02320-5/S0025-5718-09-
02320-5.pdf.

Han:2015:ACS

[Han15]  Bin Han.  Algorithm for

constructing symmetric dual framelet filter banks. Mathematics of Computation, 84 (292):767–801, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-292/
S0025-5718-2014-02856-1;
http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02856-1/S0025-5718-2014-
02856-1.pdf.

Harvey:2010:MAC

[Har10] David Harvey. A multimodular algorithm for computing Bernoulli numbers. Mathematics of Computation, 79(272): 2361–2370, October 2010. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2010-79-
272/S0025-5718-2010-02367-
1/;   http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-2010-02367-1/S0025-
5718-2010-02367-1.pdf.

Harvey:2011:FAS

[Har11] David Harvey. Faster algorithms for the square root and reciprocal of power series. Mathematics of Computation, 80 (273):387–394, January 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02392-
0/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02392-0/S0025-
5718-2010-02392-0.pdf.

Harvey:2014:SAC

[Har14] David Harvey. A subquadratic algorithm for computing the nth Bernoulli number. Mathematics of Computation, 83 (289):2471–2477, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-289/
S0025-5718-2014-02832-9;
http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-

                                              2014-02832-9/S0025-5718-2014-
                                              02832-9.pdf.

Hashemi:2012:ECC

[Has12] Amir Hashemi. Efficient computation of Castelnuovo–Mumford regularity. Mathematics of Computation, 81(278):1163– 1177, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-278/
                                              S0025-5718-2011-02515-9;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-278/S0025-5718-
                                              2011-02515-9/S0025-5718-2011-
                                              02515-9.pdf.

Hu:2011:ATG

[HC11] Xiaozhe Hu and Xiaoliang Cheng. Acceleration of a two-grid method for eigenvalue problems. Mathematics of Computation, 80(275): 1287–1301, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              275/S0025-5718-2011-02458-
                                              0/;   http://www.ams.org/
                                              journals/mcom/2011-80-275/
                                              S0025-5718-2011-02458-0/S0025-
                                              5718-2011-02458-0.pdf. See

corrigendum [HC15].

Hu:2015:CAT

[HC15] Xiaozhe Hu and Xiaoliang Cheng. Corrigendum to: “Acceleration of a two-grid method for eigenvalue problems”. Mathematics of Computation, 84 (296):2701–2704, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2015-84-
296/S0025-5718-2015-02967-
6;   http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02967-6/S0025-5718-2015-
02967-6.pdf. See [HC11].

He:2018:QMC

[He18]  Zhijian He.  Quasi-Monte

Carlo for discontinuous integrands with singularities along the boundary of the unit cube. Mathematics of Computation, 87(314):2857– 2870, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-314/
S0025-5718-2018-03324-5;
http://www.ams.org/journals/
mcom/2018-87-314/S0025-5718-
2018-03324-5/S0025-5718-2018-
03324-5.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1056493.

Heijne:2012:MRE

[Hei12]  Bas Heijne.  The maxi-

mal rank of elliptic Delsarte surfaces. Mathematics of Computation, 81(278):1111– 1130, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02529-9;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02529-9/S0025-5718-2011-
02529-9.pdf.

He:2018:SDF

[HFC18] Ruijian He, Xinlong Feng, and Zhangxin Chen. H^{1}- superconvergence of a difference finite element method based on the P_{1}–P_{1}-conforming element on non-uniform meshes for the 3D Poisson equation. Mathematics of Computation, 87(312):1659–1688, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2017-03266-X;
http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03266-X/S0025-5718-2017-
03266-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1190404;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
246747;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=670358.

Han:2013:TFP

[HH13] Houde Han and Zhongyi Huang. Tailored finite point method based on exponential bases for convection-diffusion-reaction equation. Mathematics of Computation, 82(281):213–226, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-281/
S0025-5718-2012-02616-0;
http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-
2012-02616-0/S0025-5718-2012-
02616-0.pdf.

Hart:2017:IPT

[HHO17] William Hart, David Harvey, and Wilson Ong. Irregular primes to two billion. Mathematics of Computation, 86 (308):3031–3049, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-308/
S0025-5718-2017-03211-7;
http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03211-7/S0025-5718-2017-
03211-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1001021;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
734771;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=777249.

Hajdu:2019:FWA

[HHT19] A. Hajdu, L. Hajdu, and R. Tijdeman. Finding well approximating lattices for a finite set of points. Mathematics of Computation, 88(315): 369–387, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03320-8;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03320-8/S0025-5718-2018-
03320-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=172600;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
339279;   https://www.ams.

                                              org/mathscinet/search/authors.
                                              html?mrauthid=652741.

Hiary:2011:ACM

[Hia11] Ghaith A. Hiary. An amortizedcomplexity method to compute the Riemann zeta function. Mathematics of Computation, 80(275):1785–1796, July 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              275/S0025-5718-2011-02452-
                                              X/;   http://www.ams.org/
                                              journals/mcom/2011-80-275/
                                              S0025-5718-2011-02452-X/S0025-
                                              5718-2011-02452-X.pdf.

Hiary:2016:ARS

[Hia16a]  Ghaith A. Hiary.  An al-

ternative to Riemann–Siegel type formulas. Mathematics of Computation, 85(298): 1017–1032, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-298/
                                              S0025-5718-2015-03019-1;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-298/S0025-5718-
                                              2015-03019-1/S0025-5718-2015-
                                              03019-1.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=930454.

Hiary:2016:DAI

[Hia16b] Ghaith A. Hiary. A deterministic algorithm for integer factorization. Mathematics of Computation, 85(300): 2065–2069, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-300/
S0025-5718-2015-03037-3;
http://www.ams.org/journals/
mcom/2016-85-300/S0025-5718-
2015-03037-3/S0025-5718-2015-
03037-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=930454.

Hirn:2013:FEA

[Hir13] Adrian Hirn. Finite element approximation of singular powerlaw systems. Mathematics of Computation, 82(283):1247– 1268, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2013-02668-3;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-
2013-02668-3/S0025-5718-2013-
02668-3.pdf.

Hittmeir:2017:DFS

[Hit17] Markus Hittmeir. Deterministic factorization of sums and differences of powers. Mathematics of Computation, 86 (308):2947–2954, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-308/
S0025-5718-2017-03197-5;
http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03197-5/S0025-5718-2017-
03197-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Hittmeir%
2C%20Markus.

Hittmeir:2018:BGM

[Hit18] Markus Hittmeir. A babystep– giantstep method for faster deterministic integer factorization. Mathematics of Computation, 87(314):2915– 2935, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-314/
                                              S0025-5718-2018-03313-0;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-314/S0025-5718-
                                              2018-03313-0/S0025-5718-2018-
                                              03313-0.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Hittmeir%
                                              2C%20Markus.

Hiptmair:2015:EZD

[HJHM15] Ralf Hiptmair, Carlos Jerez-Hanckes, and Shipeng Mao. Extension by zero in discrete trace spaces: Inverse estimates. Mathematics of Computation, 84(296):2589– 2615, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-296/
                                              S0025-5718-2015-02955-X;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-296/S0025-5718-
                                              2015-02955-X/S0025-5718-2015-
                                              02955-X.pdf.

Han:2018:SCQ

[HJSZ18] Bin Han, Qingtang Jiang, Zuowei Shen, and Xiaosheng Zhuang. Symmetric canonical quincunx tight framelets with high vanishing moments and smoothness. Mathematics of Computation, 87(310): 347–379, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03205-1;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03205-1/S0025-5718-2017-
03205-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Han%
2C%20Bin;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Jiang%2C%
20Qingtang;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Zhuang%
2C%20Xiaosheng;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=292105.

Hutzenthaler:2018:EIP

[HJW18] Martin Hutzenthaler, Arnulf Jentzen, and Xiaojie Wang. Exponential integrability properties of numerical approximation processes for nonlinear stochastic differential equations. Mathematics of Computation, 87(311):1353– 1413, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-311/
S0025-5718-2017-03146-X;
http://www.ams.org/journals/
mcom/2018-87-311/S0025-5718-
2017-03146-X/S0025-5718-2017-
03146-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=809631;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=

                                              824543;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=898911.

Haynes:2017:DAD

[HK17] Ronald D. Haynes and Felix Kwok. Discrete analysis of domain decomposition approaches for mesh generation via the equidistribution principle. Mathematics of Computation, 86(303):233– 273, January 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-303/
                                              S0025-5718-2016-03095-1;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-303/S0025-5718-
                                              2016-03095-1/S0025-5718-2016-
                                              03095-1.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Haynes%
                                              2C%20Ronald%20D;   http://
                                              www.ams.org/mathscinet/search/
                                              author.html?mrauthid=825760.

Hadjimichael:2018:SSP

[HK18a] Yiannis Hadjimichael and David I. Ketcheson. Strong-stability-preserving additive linear multistep methods. Mathematics of Computation, 87(313):2295– 2320, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-313/
                                              S0025-5718-2018-03296-3;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-313/S0025-5718-
                                              2018-03296-3/S0025-5718-2018-
                                              03296-3.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Ketcheson%

2C%20David%20I.;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
1029413.

Huang:2018:MNM

[HK18b] Weizhang Huang and Lennard Kamenski. On the mesh nonsingularity of the moving mesh PDE method. Mathematics of Computation, 87(312):1887– 1911, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2017-03271-3;
http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03271-3/S0025-5718-2017-
03271-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Kamenski%
2C%20Lennard;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=326320.

Holden:2013:OST

[HKK13] Helge Holden, Kenneth H.

Karlsen, and Trygve Karper. Operator splitting for twodimensional incompressible fluid equations. Mathematics of Computation, 82(282):719–748, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/
2013-82-282/S0025-5718-2012-
02626-3;   http://www.ams.
org/journals/mcom/2013-82-
282/S0025-5718-2012-02626-
3/S0025-5718-2012-02626-3.
pdf.

Hausen:2016:CCR

[HKL16] Jürgen Hausen, Simon Keicher, and Antonio Laface. Computing Cox rings. Mathematics of Computation, 85 (297):467–502, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-297/
                                              S0025-5718-2015-02989-5;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-297/S0025-5718-
                                              2015-02989-5/S0025-5718-2015-
                                              02989-5.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Hausen%
                                              2C%20Jurgen;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Keicher%
                                              2C%20Simon;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Laface%
                                              2C%20Antonio.

Hulpke:2011:NLS

[HKÖ11] Alexander Hulpke, Petteri Kaski, and Patric R. J. Östergård. The number of Latin squares of order 11. Mathematics of Computation, 80(274): 1197–1219, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              274/S0025-5718-2010-02420-
                                              2/;   http://www.ams.org/
                                              journals/mcom/2011-80-274/
                                              S0025-5718-2010-02420-2/S0025-
                                              5718-2010-02420-2.pdf.

Holden:2011:OSK

[HKRT11] Helge Holden, Kenneth H. Karlsen, Nils Henrik Risebro, and Terence Tao. Operator splitting for the KdV equation. Mathematics of Computation, 80(274):821–846, April 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02402-
0/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02402-0/S0025-
5718-2010-02402-0.pdf.

Hausen:2017:CAM

[HKW17] Jürgen Hausen, Simon Keicher, and Rüdiger Wolf. Computing automorphisms of Mori dream spaces. Mathematics of Computation, 86(308): 2955–2974, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-308/
S0025-5718-2017-03185-9;
http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03185-9/S0025-5718-2017-
03185-9.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Wolf%
2C%20Rudiger;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1001701;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
361664.

Hubert:2016:CIF

[HL16] Evelyne Hubert and George Labahn. Computation of invariants of finite abelian groups. Mathematics of Computation, 85(302):3029–3050, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-302/
                                              S0025-5718-2016-03076-8;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-302/S0025-5718-
                                              2016-03076-8/S0025-5718-2016-
                                              03076-8.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=108900;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              626362.

Huang:2017:EER

[HL17] Hui Huang and Jian-Guo Liu. Error estimate of a random particle blob method for the Keller–Segel equation. Mathematics of Computation, 86 (308):2719–2744, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-308/
                                              S0025-5718-2017-03174-4;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-308/S0025-5718-
                                              2017-03174-4/S0025-5718-2017-
                                              03174-4.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Huang%
                                              2C%20Hui;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=233036.

Hungria:2016:RNA

[HLJ16] Allan Hungria, Jean-Philippe Lessard, and J. D. Mireles James. Rigorous numerics for analytic solutions of differential equations: the radii polynomial approach. Mathematics of Computation, 85(299): 1427–1459, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2015-03046-4;
http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2015-03046-4/S0025-5718-2015-
03046-4.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Hungria%
2C%20Allan;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Lessard%
2C%20Jean-Philippe;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
912491.

Hu:2016:LSB

[HLL16] Lili Hu, Yao Li, and Yingjie Liu.

A limiting strategy for the back and forth error compensation and correction method for solving advection equations. Mathematics of Computation, 85 (299):1263–1280, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2016-03026-4;
http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2016-03026-4/S0025-5718-2016-
03026-4.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Hu%
2C%20Lili;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=349901;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
928517.

Holden:2013:OSP

[HLR13] Helge Holden, Christian Lubich, and Nils Henrik Risebro. Operator splitting for partial differential equations with Burgers nonlinearity. Mathematics of Computation, 82(281):173– 185, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-281/
                                              S0025-5718-2012-02624-X;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-281/S0025-5718-
                                              2012-02624-X/S0025-5718-2012-
                                              02624-X.pdf.

He:2015:IPO

[HLZ15] Simai He, Zhening Li, and Shuzhong Zhang. Inhomogeneous polynomial optimization over a convex set: an approximation approach. Mathematics of Computation, 84 (292):715–741, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-292/
                                              S0025-5718-2014-02875-5;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-292/S0025-5718-
                                              2014-02875-5/S0025-5718-2014-
                                              02875-5.pdf.

Harada:2014:EET

[HM14] Masaaki Harada and Tsuyoshi Miezaki. On the existence of extremal Type II ℤ_{2k}-codes. Mathematics of Computation, 83(287):1427–1446, 2014. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.

org/journals/mcom/2014-83-
287/S0025-5718-2013-02750-
0;   http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02750-0/S0025-5718-2013-
02750-0.pdf.

Hiptmair:2013:EAT

[HMP13] Ralf Hiptmair, Andrea Moiola, and Ilaria Perugia. Error analysis of Trefftz-discontinuous Galerkin methods for the timeharmonic Maxwell equations. Mathematics of Computation, 82(281):247–268, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2013-82-
281/S0025-5718-2012-02627-
5;   http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-
2012-02627-5/S0025-5718-2012-
02627-5.pdf.

Harase:2011:FLR

[HMS11] Shin Harase, Makoto Matsumoto, and Mutsuo Saito. Fast lattice reduction for F_{2}- linear pseudorandom number generators. Mathematics of Computation, 80(273): 395–407, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02391-
9/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02391-9/S0025-
5718-2010-02391-9.pdf.

Hovhannisyan:2014:AMD

[HMS14] Nune Hovhannisyan, Siegfried Müller, and Roland Schäfer. Adaptive multiresolution discontinuous Galerkin schemes for conservation laws. Mathematics of Computation, 83 (285):113–151, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-285/
S0025-5718-2013-02732-9;
http://www.ams.org/journals/
mcom/2014-83-285/S0025-5718-
2013-02732-9/S0025-5718-2013-
02732-9.pdf.

Hanke:2019:LSC

[HMT19] Michael Hanke, Roswitha März, and Caren Tischendorf. Leastsquares collocation for higherindex linear differential-algebraic equations: Estimating the instability threshold. Mathematics of Computation, 88(318): 1647–1683, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-318/
S0025-5718-2018-03393-2;
https://www.ams.org/journals/
mcom/2019-88-318/S0025-5718-
2018-03393-2/S0025-5718-2018-
03393-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Hanke%
2C%20Michael;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Marz%
2C%20Roswitha;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=346469.

Hangelbroek:2018:ITC

[HNRW18] T. Hangelbroek, F. J. Narcowich, C. Rieger, and J. D.

Ward. An inverse theorem for compact Lipschitz regions in ℝ^{d} using localized kernel bases. Mathematics of Computation, 87(312):1949– 1989, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2017-03256-7;
http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03256-7/S0025-5718-2017-
03256-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Hangelbroek%
2C%20T.;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Rieger%2C%
20C.;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=129435;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
180590.

Hinrichs:2014:CDN

[HNUW14] A. Hinrichs, E. Novak, M. Ullrich, and H. Woz.niakowski. The curse of dimensionality for numerical integration of smooth functions. Mathematics of Computation, 83(290): 2853–2863, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02855-X;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02855-X/S0025-5718-2014-
02855-X.pdf.

Hiary:2012:ZFC

[HO12] Ghaith A. Hiary and Andrew M. Odlyzko. The zeta function on the critical line: Numerical evidence for moments and random matrix theory models. Mathematics of Computation, 81(279): 1723–1752, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-279/
                                              S0025-5718-2011-02573-1;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-279/S0025-5718-
                                              2011-02573-1/S0025-5718-2011-
                                              02573-1.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Hiary%
                                              2C%20Ghaith%20A;   http://
                                              www.ams.org/mathscinet/search/
                                              authors.html?authorName=Odlyzko%
                                              2C%20Andrew%20M..

Haanpaa:2014:CHC

[HÖ14] Harri Haanpää and Patric R. J. Östergård. Counting Hamiltonian cycles in bipartite graphs. Mathematics of Computation, 83(286):979–995, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-286/
                                              S0025-5718-2013-02741-X;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-286/S0025-5718-
                                              2013-02741-X/S0025-5718-2013-
                                              02741-X.pdf.

Hoang:2016:NDI

[Hoa16] N. S. Hoang. On node distributions for interpolation and spectral methods. Mathematics of Computation, 85 (298):667–692, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-03018-X;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-03018-X/S0025-5718-2015-
03018-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=796419.

Haehnle:2010:ANW

[HP10] Jonas Haehnle and Andreas Prohl. Approximation of nonlinear wave equations with nonstandard anisotropic growth conditions. Mathematics of Computation, 79(269):189–208, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-269/
S0025-5718-09-02231-5/;
http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02231-5/S0025-5718-09-
02231-5.pdf.

Huhtanen:2012:NSL

[HP12] Marko Huhtanen and Allan Perämäki. Numerical solution of the ℝ-linear Beltrami equation. Mathematics of Computation, 81(277):387– 397, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02541-X;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-

                                              2011-02541-X/S0025-5718-2011-
                                              02541-X.pdf.

Hare:2013:SCG

[HP13] Kevin G. Hare and Maysum Panju. Some comments on Garsia numbers. Mathematics of Computation, 82(282):1197– 1221, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-282/
                                              S0025-5718-2012-02636-6;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-282/S0025-5718-
                                              2012-02636-6/S0025-5718-2012-
                                              02636-6.pdf.

Hejda:2016:SPC

[HP16] Tomáš Hejda and Edita Pelantová. Spectral properties of cubic complex Pisot units. Mathematics of Computation, 85 (297):401–421, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-297/
                                              S0025-5718-2015-02983-4;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-297/S0025-5718-
                                              2015-02983-4/S0025-5718-2015-
                                              02983-4.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Hejda%
                                              2C%20Tomas;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Pelantova%
                                              2C%20Edita.

Harbrecht:2017:QMC

[HPS17] Helmut Harbrecht, Michael Peters, and Markus Siebenmorgen. On the quasi-Monte Carlo method with Halton points for elliptic PDEs with log-normal diffusion. Mathematics of Computation, 86 (304):771–797, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03107-5;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03107-5/S0025-5718-2016-
03107-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Harbrecht%
2C%20Helmut;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Siebenmorgen%
2C%20Markus;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=975028.

Hoang:2010:DSM

[HR10] N. S. Hoang and A. G.

Ramm. Dynamical systems method for solving nonlinear equations with monotone operators. Mathematics of Computation, 79(269):239– 258, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-269/
S0025-5718-09-02260-1/;
http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02260-1/S0025-5718-09-
02260-1.pdf.

Hundsdorfer:2011:BSS

[HS11] W. Hundsdorfer and M. N.

Spijker. Boundedness and strong stability of Runge– Kutta methods. Mathematics of Computation, 80(274): 863–886, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              274/S0025-5718-2010-02422-
                                              6/;   http://www.ams.org/
                                              journals/mcom/2011-80-274/
                                              S0025-5718-2010-02422-6/S0025-
                                              5718-2010-02422-6.pdf.

Hajdu:2012:DCJ

[HS12] L. Hajdu and N. Saradha. Disproof of a conjecture of Jacobsthal. Mathematics of Computation, 81(280):2461– 2471, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-280/
                                              S0025-5718-2012-02581-6;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-280/S0025-5718-
                                              2012-02581-6/S0025-5718-2012-
                                              02581-6.pdf.

Hansen:2013:CIE

[HS13] Eskil Hansen and Tony Stillfjord. Convergence of the implicitexplicit Euler scheme applied to perturbed dissipative evolution equations. Mathematics of Computation, 82(284): 1975–1985, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-284/
                                              S0025-5718-2013-02702-0;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-284/S0025-5718-
                                              2013-02702-0/S0025-5718-2013-
                                              02702-0.pdf.

Heuer:2015:ANS

[HS15] Norbert Heuer and Francisco-Javier Sayas. Analysis of a non-symmetric coupling of Interior Penalty DG and BEM. Mathematics of Computation, 84(292):581–598, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-292/
S0025-5718-2014-02918-9;
http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02918-9/S0025-5718-2014-
02918-9.pdf.

Hochbruck:2019:UDG

[HS19a] Marlis Hochbruck and Andreas Sturm. Upwind discontinuous Galerkin space discretization and locally implicit time integration for linear Maxwell’s equations. Mathematics of Computation, 88(317):1121– 1153, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-317/
S0025-5718-2018-03365-8;
https://www.ams.org/journals/
mcom/2019-88-317/S0025-5718-
2018-03365-8/S0025-5718-2018-
03365-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1183900;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
332872.

Hoste:2019:EPR

[HS19b] Jim Hoste and Patrick D.

Shanahan. An enumeration process for racks. Mathematics of Computation, 88(317):1427– 1448, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-317/
S0025-5718-2018-03374-9;
https://www.ams.org/journals/
mcom/2019-88-317/S0025-5718-
2018-03374-9/S0025-5718-2018-
03374-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=537475;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
88610.

Hu:2010:NDD

[HSW10] Qiya Hu, Shi Shu, and Junxian Wang. Nonoverlapping domain decomposition methods with a simple coarse space for elliptic problems. Mathematics of Computation, 79(272):2059– 2078, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02361-6/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02361-6/S0025-5718-10-
02361-6.pdf.

Hauenstein:2011:RHS

[HSW11] Jonathan D. Hauenstein, Andrew J. Sommese, and Charles W. Wampler. Regeneration homotopies for solving systems of polynomials. Mathematics of Computation, 80(273): 345–377, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.

org/journals/mcom/2011-80-
273/S0025-5718-2010-02399-
3/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02399-3/S0025-
5718-2010-02399-3.pdf.

Hurst:2018:CMF

[Hur18] Greg Hurst. Computations of the Mertens function and improved bounds on the Mertens conjecture. Mathematics of Computation, 87(310):1013– 1028, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/
2018-87-310/S0025-5718-2017-
03275-0;   http://www.ams.
org/journals/mcom/2018-87-
310/S0025-5718-2017-03275-
0/S0025-5718-2017-03275-0.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Hurst%2C%
20Greg.

Hutz:2015:DAR

[Hut15] Benjamin Hutz. Determination of all rational preperiodic points for morphisms of PN. Mathematics of Computation, 84(291):289–308, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-291/
S0025-5718-2014-02850-0;
http://www.ams.org/journals/
mcom/2015-84-291/S0025-5718-
2014-02850-0/S0025-5718-2014-
02850-0.pdf.

Harvey:2019:FIM

[HvdH19] David Harvey and Joris van der Hoeven. Faster integer multiplication using plain vanilla FFT primes. Mathematics of Computation, 88(315):501– 514, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03328-2;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03328-2/S0025-5718-2018-
03328-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=621578;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
734771.

Houston:2018:AND

[HW18] Paul Houston and Thomas P. Wihler. An hp-adaptive Newton-discontinuous-Galerkin finite element approach for semilinear elliptic boundary value problems. Mathematics of Computation, 87(314):2641– 2674, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-314/
S0025-5718-2018-03308-7;
http://www.ams.org/journals/
mcom/2018-87-314/S0025-5718-
2018-03308-7/S0025-5718-2018-
03308-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=635107;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
662940.

Huang:2014:NDI

[HWZ14] Xiaowei Huang, Chuansheng Wu, and Jun Zhou. Numerical differentiation by integration. Mathematics of Computation, 83(286):789–807, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-286/
S0025-5718-2013-02722-6;
http://www.ams.org/journals/
mcom/2014-83-286/S0025-5718-
2013-02722-6/S0025-5718-2013-
02722-6.pdf.

Hu:2019:SPF

[HX19] Kaibo Hu and Jinchao Xu.

Structure-preserving finite element methods for stationary MHD models. Mathematics of Computation, 88(316):553– 581, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03341-5;
https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03341-5/S0025-5718-2018-
03341-5.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1161425;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
228866.

Hu:2015:FAE

[HY15] Jingwei Hu and Lexing Ying.

A fast algorithm for the energy space boson Boltzmann collision operator. Mathematics of Computation, 84 (291):271–288, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/

                                              journals/mcom/2015-84-291/
                                              S0025-5718-2014-02824-X;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-291/S0025-5718-
                                              2014-02824-X/S0025-5718-2014-
                                              02824-X.pdf.

Han:2014:ALB

[HYZ14] Deren Han, Xiaoming Yuan, and Wenxing Zhang. An augmented Lagrangian based parallel splitting method for separable convex minimization with applications to image processing. Mathematics of Computation, 83(289): 2263–2291, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-289/
                                              S0025-5718-2014-02829-9;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-289/S0025-5718-
                                              2014-02829-9/S0025-5718-2014-
                                              02829-9.pdf.

Han:2013:AME

[HZ13] Bin Han and Xiaosheng Zhuang. Algorithms for matrix extension and orthogonal wavelet filter banks over algebraic number fields. Mathematics of Computation, 82(281):459– 490, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-281/
                                              S0025-5718-2012-02618-4;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-281/S0025-5718-
                                              2012-02618-4/S0025-5718-2012-
                                              02618-4.pdf.

Hajdu:2014:DUG

[HZ14] Lajos Hajdu and Volker Ziegler.

Distinct unit generated totally complex quartic fields. Mathematics of Computation, 83 (287):1495–1512, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-287/
S0025-5718-2013-02751-2;
http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02751-2/S0025-5718-2013-
02751-2.pdf.

Hu:2015:MCF

[HZ15] Jun Hu and Shangyou Zhang.

The minimal conforming H^{k} finite element spaces on R^{n} rectangular grids. Mathematics of Computation, 84 (292):563–579, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-292/
S0025-5718-2014-02871-8;
http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02871-8/S0025-5718-2014-
02871-8.pdf.

He:2017:SFE

[HZ17] Wenming He and Zhimin Zhang. 2k superconvergence of Q_{k} finite elements by anisotropic mesh approximation in weighted Sobolev spaces. Mathematics of Computation, 86(306):1693–1718, 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-

306/S0025-5718-2016-03159-
2;   http://www.ams.org/journals/
mcom/2017-86-306/S0025-5718-
2016-03159-2/S0025-5718-2016-
03159-2.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=303173;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
671119.

Ionica:2013:PV

[IJ13] Sorina Ionica and Antoine Joux. Pairing the volcano. Mathematics of Computation, 82(281): 581–603, January 2013. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2013-82-
281/S0025-5718-2012-02622-
6;   http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-
2012-02622-6/S0025-5718-2012-
02622-6.pdf.

Ichim:2017:HCS

[IKMF17] Bogdan Ichim, Lukas Katthän, and Julio José Moyano-Fernández. How to compute the Stanley depth of a module. Mathematics of Computation, 86(303):455– 472, January 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-303/
S0025-5718-2016-03106-3;
http://www.ams.org/journals/
mcom/2017-86-303/S0025-5718-
2016-03106-3/S0025-5718-2016-
03106-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Ichim%
2C%20Bogdan;   http://www.
ams.org/mathscinet/search/

author.html?authorName=Katthan%
2C%20Lukas;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Moyano-
Fernandez%2C%20Julio%20Jose.

Ivanyos:2012:TGA

[IKRS12] Gábor Ivanyos, Marek Karpinski, Lajos Rónyai, and Nitin Saxena. Trading GRH for algebra: Algorithms for factoring polynomials and related structures. Mathematics of Computation, 81(277):493– 531, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02505-6;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02505-6/S0025-5718-2011-
02505-6.pdf.

Iserles:2011:AEQ

[IL11] Arieh Iserles and David Levin.

Asymptotic expansion and quadrature of composite highly oscillatory integrals. Mathematics of Computation, 80(273): 279–296, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02386-
5/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02386-5/S0025-
5718-2010-02386-5.pdf.

Ingram:2013:NLE

[Ing13] Ross Ingram. A new linearly extrapolated Crank–Nicolson time-stepping scheme for the Navier–Stokes equations. Mathematics of Computation, 82 (284):1953–1973, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-284/
S0025-5718-2013-02678-6;
http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02678-6/S0025-5718-2013-
02678-6.pdf.

Ignat:2015:LTA

[IPZ15] Liviu I. Ignat, Alejandro Pozo, and Enrique Zuazua. Largetime asymptotics, vanishing viscosity and numerics for 1-D scalar conservation laws. Mathematics of Computation, 84 (294):1633–1662, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-294/
S0025-5718-2014-02915-3;
http://www.ams.org/journals/
mcom/2015-84-294/S0025-5718-
2014-02915-3/S0025-5718-2014-
02915-3.pdf.

Ishizaka:2017:SEM

[Ish17] Kanya Ishizaka. A solution to the energy minimization problem constrained by a density function. Mathematics of Computation, 86(303):275– 314, January 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-303/
S0025-5718-2016-03136-1;
http://www.ams.org/journals/

mcom/2017-86-303/S0025-5718-
2016-03136-1/S0025-5718-2016-
03136-1.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=767775.

Iwata:2012:TIH

[ITT12] Satoru Iwata, Mizuyo Takamatsu, and Caren Tischendorf. Tractability index of hybrid equations for circuit simulation. Mathematics of Computation, 81(278): 923–939, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02558-5;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02558-5/S0025-5718-2011-
02558-5.pdf.

Iliescu:2013:VMP

[IW13] Traian Iliescu and Zhu Wang.

Variational multiscale proper orthogonal decomposition: Convectiondominated convection-diffusion-reaction equations. Mathematics of Computation, 82(283): 1357–1378, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2013-02683-X;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-
2013-02683-X/S0025-5718-2013-
02683-X.pdf.

Jager:2012:MRA

[Jag12] H. Jager. A metrical result on the approximation by continued fractions. Mathematics of Computation, 81(280):2377– 2382, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-280/
S0025-5718-2012-02621-4;
http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-
2012-02621-4/S0025-5718-2012-
02621-4.pdf.

Johansson:2019:CSC

[JB19] Fredrik Johansson and Iaroslav V. Blagouchine. Computing Stieltjes constants using complex integration. Mathematics of Computation, 88(318):1829– 1850, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-318/
S0025-5718-2018-03401-9;
https://www.ams.org/journals/
mcom/2019-88-318/S0025-5718-
2018-03401-9/S0025-5718-2018-
03401-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=906772;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
999321.

Jiang:2014:SPF

[JD14] Yupeng Jiang and Yingpu Deng. Strong pseudoprimes to the first eight prime bases. Mathematics of Computation, 83(290): 2915–2924, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02830-5;
http://www.ams.org/journals/

mcom/2014-83-290/S0025-5718-
2014-02830-5/S0025-5718-2014-
02830-5.pdf.

Jespers:2014:WUI

[JdRV14] Eric Jespers, Ángel del Río, and Inneke Van Gelder. Writing units of integral group rings of finite abelian groups as a product of Bass units. Mathematics of Computation, 83 (285):461–473, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-285/
S0025-5718-2013-02718-4;
http://www.ams.org/journals/
mcom/2014-83-285/S0025-5718-
2013-02718-4/S0025-5718-2013-
02718-4.pdf.

Jeon:2016:FEC

[Jeo16] Daeyeol Jeon. Families of elliptic curves over cyclic cubic number fields with prescribed torsion. Mathematics of Computation, 85(299): 1485–1502, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2015-03012-9;
http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2015-03012-9/S0025-5718-2015-
03012-9.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=658790.

Jin:2013:DDM

[JgLW13] Shi Jin, Jian guo Liu, and Li Wang. A domain decomposition method for semilinear hyperbolic systems with twoscale relaxations. Mathematics of Computation, 82(282): 749–779, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02643-3;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02643-3/S0025-5718-2012-
02643-3.pdf.

Jin:2010:CFR

[Jin10]  Qinian Jin.  On a class

of frozen regularized Gauss– Newton methods for nonlinear inverse problems. Mathematics of Computation, 79(272):2191– 2211, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02359-8/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02359-8/S0025-5718-10-
02359-8.pdf.

Jin:2013:FCR

[Jin13] Qinian Jin. Further convergence results on the general iteratively regularized Gauss– Newton methods under the discrepancy principle. Mathematics of Computation, 82(283): 1647–1665, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2012-02665-2;
http://www.ams.org/journals/

mcom/2013-82-283/S0025-5718-
2012-02665-2/S0025-5718-2012-
02665-2.pdf.

Jespers:2015:PAT

[JJK^{+}15] E. Jespers, S. O. Juriaans, A. Kiefer, A. de A.e Silva, and A. C. Souza Filho. From the Poincaré Theorem to generators of the unit group of integral group rings of finite groups. Mathematics of Computation, 84(293): 1489–1520, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02865-2;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02865-2/S0025-5718-2014-
02865-2.pdf.

Jaquette:2017:APD

[JK17] Jonathan Jaquette and Miroslav Kramár. On ϵ approximations of persistence diagrams. Mathematics of Computation, 86 (306):1887–1912, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-306/
S0025-5718-2016-03137-3;
http://www.ams.org/journals/
mcom/2017-86-306/S0025-5718-
2016-03137-3/S0025-5718-2016-
03137-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Jaquette%
2C%20Jonathan;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=747856.

Jespers:2016:PGA

[JKdR16] E. Jespers, A. Kiefer, and Á. del Río. Presentations of groups acting discontinuously on direct products of hyperbolic spaces. Mathematics of Computation, 85(301):2515– 2552, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-301/
S0025-5718-2015-03071-3;
http://www.ams.org/journals/
mcom/2016-85-301/S0025-5718-
2015-03071-3/S0025-5718-2015-
03071-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Kiefer%
2C%20A;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=288713;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
94560.

Jaklic:2010:IPC

[JKK^{+}10] Gasper Jaklic, Jernej Kozak, Marjeta Krajnc, Vito Vitrih, and Emil Zagar. On interpolation by planar cubic G^{2} Pythagorean-hodograph spline curves. Mathematics of Computation, 79(269):305– 326, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-269/
S0025-5718-09-02298-4/;
http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02298-4/S0025-5718-09-
02298-4.pdf.

Jeon:2011:FECa

[JKL11a] Daeyeol Jeon, Chang Heon Kim, and Yoonjin Lee. Families of elliptic curves over cubic number fields with prescribed torsion subgroups. Mathematics of Computation, 80(273):579– 591, January 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-10-02369-0/;
http://www.ams.org/journals/
mcom/2011-80-273/S0025-5718-
10-02369-0/S0025-5718-10-
02369-0.pdf.

Jeon:2011:FECb

[JKL11b] Daeyeol Jeon, Chang Heon Kim, and Yoonjin Lee. Families of elliptic curves over quartic number fields with prescribed torsion subgroups. Mathematics of Computation, 80(276): 2395–2410, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02493-
2/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02493-2/S0025-
5718-2011-02493-2.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813367.

Jeannerod:2017:EBC

[JKLM17] Claude-Pierre Jeannerod, Peter Kornerup, Nicolas Louvet, and Jean-Michel Muller. Error bounds on complex floating-point multiplication with an FMA. Mathematics of Computation, 86(304): 881–898, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03123-3;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03123-3/S0025-5718-2016-
03123-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Kornerup%
2C%20Peter;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Muller%
2C%20Jean-Michel;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
644190;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=893389.

Jin:2012:NIR

[JL12] Bangti Jin and Xiliang Lu. Numerical identification of a Robin coefficient in parabolic problems. Mathematics of Computation, 81(279):1369– 1398, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2012-02559-2;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2012-02559-2/S0025-5718-2012-
02559-2.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Jin%
2C%20Bangti;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Lu%
2C%20Xiliang.

Jiang:2013:FAB

[JLH13] Shidong Jiang, Zhi Liang, and Jingfang Huang. A fast algorithm for Brownian dynamics simulation with hydrodynamic interactions. Mathematics of Computation, 82(283): 1631–1645, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2013-02672-5;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-
2013-02672-5/S0025-5718-2013-
02672-5.pdf.

Jeannerod:2013:FAK

[JLM13] Claude-Pierre Jeannerod, Nicolas Louvet, and Jean-Michel Muller. Further analysis of Kahan’s algorithm for the accurate computation of 2 × 2 determinants. Mathematics of Computation, 82(284): 2245–2264, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-284/
S0025-5718-2013-02679-8;
http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02679-8/S0025-5718-2013-
02679-8.pdf.

Jin:2015:VFP

[JLPR15] Bangti Jin, Raytcho Lazarov, Joseph Pasciak, and William Rundell. Variational formulation of problems involving fractional order differential operators. Mathematics of Computation, 84(296):2665– 2700, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-296/
S0025-5718-2015-02960-3;
http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02960-3/S0025-5718-2015-
02960-3.pdf.

Ju:2018:ESE

[JLQZ18] Lili Ju, Xiao Li, Zhonghua Qiao, and Hui Zhang. Energy stability and error estimates of exponential time differencing schemes for the epitaxial growth model without slope selection. Mathematics of Computation, 87(312):1859– 1885, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2017-03262-2;
http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03262-2/S0025-5718-2017-
03262-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Li%
2C%20Xiao;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Zhang%
2C%20Hui;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=645968;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
711384.

Jin:2017:NPF

[JLTZ17] Bangti Jin, Raytcho Lazarov, Vidar Thomée, and Zhi Zhou.

On nonnegativity preservation in finite element methods for subdiffusion equations. Mathematics of Computation, 86 (307):2239–2260, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-307/
S0025-5718-2016-03167-1;
http://www.ams.org/journals/
mcom/2017-86-307/S0025-5718-
2016-03167-1/S0025-5718-2016-
03167-1.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Jin%
2C%20Bangti;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1011798;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
111240;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=172250.

Jin:2019:STD

[JLZ19] Bangti Jin, Buyang Li, and Zhi Zhou. Subdiffusion with a timedependent coefficient: Analysis and numerical solution. Mathematics of Computation, 88(319): 2157–2186, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-319/
S0025-5718-2019-03413-0;
https://www.ams.org/journals/
mcom/2019-88-319/S0025-5718-
2019-03413-0/S0025-5718-2019-
03413-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1011798;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
741824;   https://www.ams.

                                              org/mathscinet/search/authors.
                                              html?mrauthid=910552.

Johansson:2015:FAR

[Joh15] Fredrik Johansson. A fast algorithm for reversion of power series. Mathematics of Computation, 84(291): 475–484, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-291/
                                              S0025-5718-2014-02857-3;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-291/S0025-5718-
                                              2014-02857-3/S0025-5718-2014-
                                              02857-3.pdf.

Jorgensen:2012:ESF

[Jor12] Palle E. T. Jorgensen. Ergodic scales in fractal measures. Mathematics of Computation, 81(278):941–955, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2012-81-
                                              278/S0025-5718-2011-02517-
                                              2;   http://www.ams.org/journals/
                                              mcom/2012-81-278/S0025-5718-
                                              2011-02517-2/S0025-5718-2011-
                                              02517-2.pdf.

Jagels:2013:SMR

[JR13] Carl Jagels and Lothar Reichel. The structure of matrices in rational Gauss quadrature. Mathematics of Computation, 82 (284):2035–2060, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-284/
                                              S0025-5718-2013-02695-6;

http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02695-6/S0025-5718-2013-
02695-6.pdf.

Jeannerod:2018:REF

[JR18] Claude-Pierre Jeannerod and Siegfried M. Rump. On relative errors of floating-point operations: Optimal bounds and applications. Mathematics of Computation, 87 (310):803–819, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03234-8;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03234-8/S0025-5718-2017-
03234-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Rump%
2C%20Siegfried%20M.;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
644190.

Jorgenson:2014:DEM

[JST14] Jay Jorgenson, Lejla Smajlović, and Holger Then. On the distribution of eigenvalues of Maass forms on certain moonshine groups. Mathematics of Computation, 83 (290):3039–3070, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02823-8;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02823-8/S0025-5718-2014-
02823-8.pdf.

Jarso:2019:QNR

[JT19] Tamiru Jarso and Tim Trudgian. Quadratic non-residues that are not primitive roots. Mathematics of Computation, 88(317):1251–1260, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

                                              www.ams.org/journals/mcom/
                                              2019-88-317/S0025-5718-2018-
                                              03378-6;   https://www.ams.
                                              org/journals/mcom/2019-88-
                                              317/S0025-5718-2018-03378-
                                              6/S0025-5718-2018-03378-6.
                                              pdf;   https://www.ams.org/
                                              mathscinet/search/authors.
                                              html?authorName=Jarso%2C%
                                              20Tamiru;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=909247.

Ji:2014:SEE

[JvSRV14] Liangyue Ji, Paulien van Slingerland, Jennifer K. Ryan, and Kees Vuik. Superconvergent error estimates for position-dependent smoothnessincreasing accuracy-conserving (SIAC) post-processing of discontinuous Galerkin solutions. Mathematics of Computation, 83(289):2239–2262, 2014. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2014-83-
                                              289/S0025-5718-2014-02835-
                                              4;   http://www.ams.org/journals/
                                              mcom/2014-83-289/S0025-5718-
                                              2014-02835-4/S0025-5718-2014-
                                              02835-4.pdf.

Jones:2012:NFS

[JW12] John W. Jones and Rachel Wallington. Number fields with solvable Galois groups and small Galois root discriminants. Mathematics of Computation, 81(277):555–567, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02511-1;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02511-1/S0025-5718-2011-
02511-1.pdf.

Jiang:2014:FCM

[JX14] Ying Jiang and Yuesheng Xu.

Fast computation of the multidimensional discrete Fourier transform and discrete backward Fourier transform on sparse grids. Mathematics of Computation, 83(289): 2347–2384, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-289/
S0025-5718-2014-02785-3;
http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-
2014-02785-3/S0025-5718-2014-
02785-3.pdf.

Ji:2012:AED

[JXR12] Liangyue Ji, Yan Xu, and Jennifer K. Ryan. Accuracyenhancement of discontinuous Galerkin solutions for convection-diffusion equations in multiple-dimensions. Mathematics of Computation, 81(280): 1929–1950, October 2012. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2012-81-
                                              280/S0025-5718-2012-02586-
                                              5;   http://www.ams.org/journals/
                                              mcom/2012-81-280/S0025-5718-
                                              2012-02586-5/S0025-5718-2012-
                                              02586-5.pdf.

Jia:2011:RBW

[JZ11] Rong-Qing Jia and Wei Zhao. Riesz bases of wavelets and applications to numerical solutions of elliptic equations. Mathematics of Computation, 80(275): 1525–1556, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              275/S0025-5718-2011-02448-
                                              8/;   http://www.ams.org/
                                              journals/mcom/2011-80-275/
                                              S0025-5718-2011-02448-8/S0025-
                                              5718-2011-02448-8.pdf.

Karabina:2013:SCS

[Kar13] Koray Karabina. Squaring in cyclotomic subgroups. Mathematics of Computation, 82(281): 555–579, January 2013. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2013-82-
                                              281/S0025-5718-2012-02625-
                                              1;   http://www.ams.org/journals/
                                              mcom/2013-82-281/S0025-5718-
                                              2012-02625-1/S0025-5718-2012-
                                              02625-1.pdf.

Korobov:2016:APS

[KB16] V. I. Korobov and A. N. Bugaevskaya. Almost power sum systems. Mathematics of Computation, 85(298): 717–736, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-02994-9;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-02994-9/S0025-5718-2015-
02994-9.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Bugaevskaya%
2C%20A.%20N;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=215695.

Knessl:2011:AFS

[KC11a] Charles Knessl and Mark W.

Coffey. An asymptotic form for the Stieltjes constants γ_{k}(a) and for a sum S_{γ}(n) appearing under the Li criterion. Mathematics of Computation, 80(276): 2197–2217, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02497-
X/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02497-X/S0025-
5718-2011-02497-X.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813355.

Knessl:2011:EAF

[KC11b] Charles Knessl and Mark W.

Coffey. An effective asymptotic formula for the Stieltjes constants. Mathematics of Computation, 80(273):379– 386, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              273/S0025-5718-2010-02390-
                                              7/;   http://www.ams.org/
                                              journals/mcom/2011-80-273/
                                              S0025-5718-2010-02390-7/S0025-
                                              5718-2010-02390-7.pdf.

Kim:2015:CPD

[KC15] Minkyu Kim and Jung Hee Cheon. Computing prime divisors in an interval. Mathematics of Computation, 84 (291):339–354, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-291/
                                              S0025-5718-2014-02840-8;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-291/S0025-5718-
                                              2014-02840-8/S0025-5718-2014-
                                              02840-8.pdf.

Kim:2014:AGA

[KCL14] Minkyu Kim, Jung Hee Cheon, and In-Sok Lee. Analysis on a generalized algorithm for the strong discrete logarithm problem with auxiliary inputs. Mathematics of Computation, 83 (288):1993–2004, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-288/
                                              S0025-5718-2014-02813-5;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-288/S0025-5718-
                                              2014-02813-5/S0025-5718-2014-
                                              02813-5.pdf.

Kemper:2010:BRC

[Kem10] Gregor Kemper. Book review: A computational introduction to number theory and algebra. Mathematics of Computation, 79(270):1231– 1232, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02357-6/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02357-6/S0025-5718-09-
02357-6.pdf.

Kreuzer:2018:CAD

[KG18] Christian Kreuzer and Emmanuil H. Georgoulis. Convergence of adaptive discontinuous Galerkin methods. Mathematics of Computation, 87(314): 2611–2640, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-314/
S0025-5718-2018-03318-X;
http://www.ams.org/journals/
mcom/2018-87-314/S0025-5718-
2018-03318-X/S0025-5718-2018-
03318-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Georgoulis%
2C%20Emmanuil%20H.;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
833122.

Kohatsu-Higa:2014:OSS

[KHOLT14] Arturo Kohatsu-Higa, Salvador Ortiz-Latorre, and Peter Tankov. Optimal simulation schemes for Lévy driven stochastic differential equations. Mathematics of Computation, 83(289):2293–2324, 2014. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2014-83-
                                              289/S0025-5718-2013-02786-
                                              X;   http://www.ams.org/journals/
                                              mcom/2014-83-289/S0025-5718-
                                              2013-02786-X/S0025-5718-2013-
                                              02786-X.pdf.

Kamenski:2014:CFE

[KHX14] Lennard Kamenski, Weizhang Huang, and Hongguo Xu. Conditioning of finite element equations with arbitrary anisotropic meshes. Mathematics of Computation, 83(289): 2187–2211, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-289/
                                              S0025-5718-2014-02822-6;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-289/S0025-5718-
                                              2014-02822-6/S0025-5718-2014-
                                              02822-6.pdf.

Kim:2017:MAC

[Kim17]  Dohyeong Kim.  A modu-

lar approach to cubic Thue– Mahler equations. Mathematics of Computation, 86(305): 1435–1471, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-305/
                                              S0025-5718-2016-03139-7;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-305/S0025-5718-
                                              2016-03139-7/S0025-5718-2016-
                                              03139-7.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=970842.

Kirat:2010:DLT

[Kir10] Ibrahim Kirat. Disk-like tiles and self-affine curves with noncollinear digits. Mathematics of Computation, 79(270):1019– 1045, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02301-1/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02301-1/S0025-5718-09-
02301-1.pdf.

Kirschmer:2012:NFD

[Kir12] Markus Kirschmer. A normal form for definite quadratic forms over 𝔽_{q}[t]. Mathematics of Computation, 81(279): 1619–1634, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2011-02570-6;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2011-02570-6/S0025-5718-2011-
02570-6.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Kirschmer%
2C%20Markus.

Kim:2010:GFF

[KK10] Chang Heon Kim and Ja Kyung Koo. Generators of function fields of the modular curves X_{1}(5) and X_{1}(6). Mathematics of Computation, 79(270): 1047–1066, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/

                                              journals/mcom/2010-79-270/
                                              S0025-5718-09-02303-5/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-270/S0025-5718-
                                              09-02303-5/S0025-5718-09-
                                              02303-5.pdf.

Karlsen:2011:CMM

[KK11] Kenneth H. Karlsen and Trygve K. Karper. Convergence of a mixed method for a semi-stationary compressible Stokes system. Mathematics of Computation, 80(275):1459–1498, July 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              275/S0025-5718-2010-02446-
                                              9/;   http://www.ams.org/
                                              journals/mcom/2011-80-275/
                                              S0025-5718-2010-02446-9/S0025-
                                              5718-2010-02446-9.pdf.

Kalla:2014:CTT

[KK14] C. Kalla and C. Klein. Computation of the topological type of a real Riemann surface. Mathematics of Computation, 83 (288):1823–1846, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-288/
                                              S0025-5718-2014-02817-2;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-288/S0025-5718-
                                              2014-02817-2/S0025-5718-2014-
                                              02817-2.pdf.

Kim:2010:FTU

[KKP10] Byeong Moon Kim, Ji Young Kim, and Poo-Sung Park. The fifteen theorem for universal Hermitian lattices over imaginary quadratic fields. Mathematics of Computation, 79(270): 1123–1144, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02287-X/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02287-X/S0025-5718-09-
02287-X.pdf.

Koumandos:2010:SCM

[KL10] Stamatis Koumandos and Martin Lamprecht. Some completely monotonic functions of positive order. Mathematics of Computation, 79(271):1697– 1707, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-09-02313-8/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
09-02313-8/S0025-5718-09-
02313-8.pdf.

Koumandos:2013:CMR

[KL13] Stamatis Koumandos and Martin Lamprecht. Complete monotonicity and related properties of some special functions. Mathematics of Computation, 82(282):1097–1120, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2013-82-
282/S0025-5718-2012-02629-
9;   http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-

                                              2012-02629-9/S0025-5718-2012-
                                              02629-9.pdf.

Klein:2013:EFH

[Kle13]  Georges Klein.  An exten-

sion of the Floater–Hormann family of barycentric rational interpolants. Mathematics of Computation, 82(284): 2273–2292, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-284/
                                              S0025-5718-2013-02688-9;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-284/S0025-5718-
                                              2013-02688-9/S0025-5718-2013-
                                              02688-9.pdf.

Kleinjung:2016:QS

[Kle16] Thorsten Kleinjung. Quadratic sieving. Mathematics of Computation, 85(300):1861– 1873, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-300/
                                              S0025-5718-2015-03058-0;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-300/S0025-5718-
                                              2015-03058-0/S0025-5718-2015-
                                              03058-0.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=704259.

Karakashian:2015:PEE

[KM15] Ohannes Karakashian and Charalambos Makridakis. A posteriori error estimates for discontinuous Galerkin methods for the generalized Korteweg– de Vries equation. Mathematics of Computation, 84(293):

1145–1167, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02878-0;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02878-0/S0025-5718-2014-
02878-0.pdf.

Krejic:2016:IRA

[KM16] Nataša Krejić and J. M.

Martínez. Inexact Restoration approach for minimization with inexact evaluation of the objective function. Mathematics of Computation, 85(300): 1775–1791, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-300/
S0025-5718-2015-03025-7;
http://www.ams.org/journals/
mcom/2016-85-300/S0025-5718-
2015-03025-7/S0025-5718-2015-
03025-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Martinez%
2C%20J.%20M;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=307392.

Kim:2019:CNL

[KM19] Kwang-Seob Kim and John C.

Miller. Class numbers of large degree nonabelian number fields. Mathematics of Computation, 88(316):973–981, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03335-X;

                                              https://www.ams.org/journals/
                                              mcom/2019-88-316/S0025-5718-
                                              2018-03335-X/S0025-5718-2018-
                                              03335-X.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1074298;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              1079146.

Kojima:2017:SDI

[KMF17] Hiroki Kojima, Takayasu Matsuo, and Daisuke Furihata. Some discrete inequalities for central-difference type operators. Mathematics of Computation, 86(306):1719–1739, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2017-86-
                                              306/S0025-5718-2016-03154-
                                              3;   http://www.ams.org/journals/
                                              mcom/2017-86-306/S0025-5718-
                                              2016-03154-3/S0025-5718-2016-
                                              03154-3.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Kojima%
                                              2C%20Hiroki;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=601502;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              664782.

Koch:2010:CMI

[KMPW10] Othmar Koch, Roswitha März, Dirk Praetorius, and Ewa Weinmüller. Collocation methods for index 1 DAEs with a singularity of the first kind. Mathematics of Computation, 79(269): 281–304, January 2010. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.

org/journals/mcom/2010-79-
269/S0025-5718-09-02267-4/
;   http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02267-4/S0025-5718-09-
02267-4.pdf.

Keller:2012:CMS

[KMSwaAbMDS12] Wolfgang Keller, Jacques Martinet, Achill Schürmann, and with an Appendix by Mathieu Dutour Sikirić. On classifying Minkowskian sublattices. Mathematics of Computation, 81(278):1063– 1092, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02528-7;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02528-7/S0025-5718-2011-
02528-7.pdf.

Karabas:2012:AMH

[KN12] Ján Karabáš and Roman Nedela. Archimedean maps of higher genera. Mathematics of Computation, 81(277):569– 583, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02502-0;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02502-0/S0025-5718-2011-
02502-0.pdf.

Kaltenbacher:2012:CAR

[KO12] Barbara Kaltenbacher and Jonas Offtermatt. A convergence analysis of regularization by discretization in preimage space. Mathematics of Computation, 81(280):2049– 2069, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-280/
S0025-5718-2012-02596-8;
http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-
2012-02596-8/S0025-5718-2012-
02596-8.pdf.

Kokkala:2018:CNS

[KÖ18] Janne I. Kokkala and Patric R. J. Östergård. The chromatic number of the square of the 8-cube. Mathematics of Computation, 87(313):2551– 2561, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-313/
S0025-5718-2017-03291-9;
http://www.ams.org/journals/
mcom/2018-87-313/S0025-5718-
2017-03291-9/S0025-5718-2017-
03291-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Kokkala%
2C%20Janne%20I.;   https:
//www.ams.org/mathscinet/
search/authors.html?authorName=
Ostergard%2C%20Patric%20R.
%20J..

Kontogeorgis:2014:CCI

[Kon14] Aristides Kontogeorgis. Constructing class invariants. Mathematics of Computation, 83 (287):1477–1488, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-287/
S0025-5718-2013-02769-X;
http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02769-X/S0025-5718-2013-
02769-X.pdf.

Konig:2017:CHS

[Kön17] Joachim König. Computation of Hurwitz spaces and new explicit polynomials for almost simple Galois groups. Mathematics of Computation, 86 (305):1473–1498, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-305/
S0025-5718-2016-03116-6;
http://www.ams.org/journals/
mcom/2017-86-305/S0025-5718-
2016-03116-6/S0025-5718-2016-
03116-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Konig%
2C%20Joachim.

Kopteva:2014:LFE

[Kop14] Natalia Kopteva. Linear finite elements may be only first-order pointwise accurate on anisotropic triangulations. Mathematics of Computation, 83(289):2061–2070, 2014. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2014-83-
289/S0025-5718-2014-02820-
2;   http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-
2014-02820-2/S0025-5718-2014-
02820-2.pdf.

Kaski:2015:EST

[KÖP15] Petteri Kaski, Patric R. J. Östergård, and Alexandru Popa. Enumeration of Steiner triple systems with subsystems. Mathematics of Computation, 84(296):3051–3067, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2015-84-
                                              296/S0025-5718-2015-02945-
                                              7;   http://www.ams.org/journals/
                                              mcom/2015-84-296/S0025-5718-
                                              2015-02945-7/S0025-5718-2015-
                                              02945-7.pdf.

Kopteva:2019:EAL

[Kop19a] Natalia Kopteva. Error analysis of the L1 method on graded and uniform meshes for a fractionalderivative problem in two and three dimensions. Mathematics of Computation, 88(319): 2135–2155, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-319/
                                              S0025-5718-2019-03410-5;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-319/S0025-5718-
                                              2019-03410-5/S0025-5718-2019-
                                              03410-5.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=610720.

Kopteva:2019:LCR

[Kop19b] Natalia Kopteva. Logarithm cannot be removed in maximum norm error estimates for linear finite elements in 3D. Mathematics of Computation, 88(318): 1527–1532, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-318/
S0025-5718-2018-03384-1;
https://www.ams.org/journals/
mcom/2019-88-318/S0025-5718-
2018-03384-1/S0025-5718-2018-
03384-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=610720.

Kreuzer:2011:SBB

[KP11] Martin Kreuzer and Henk Poulisse. Subideal border bases. Mathematics of Computation, 80(274):1135–1154, April 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02432-
9/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02432-9/S0025-
5718-2010-02432-9.pdf.

Kopteva:2012:SOO

[KP12] Natalia Kopteva and Maria Pickett. A second-order overlapping Schwarz method for a 2D singularly perturbed semilinear reactiondiffusion problem. Mathematics of Computation, 81(277):81– 105, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02521-4;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02521-4/S0025-5718-2011-
02521-4.pdf.

Kovacs:2014:SOC

[KP14] Mihály Kovács and Jacques Printems. Strong order of convergence of a fully discrete approximation of a linear stochastic Volterra type evolution equation. Mathematics of Computation, 83(289): 2325–2346, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-289/
                                              S0025-5718-2014-02803-2;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-289/S0025-5718-
                                              2014-02803-2/S0025-5718-2014-
                                              02803-2.pdf.

Katsoulakis:2014:CGS

[KPRBT14] Markos A. Katsoulakis, Petr Plecháč, Luc Rey-Bellet, and Dimitrios K. Tsagkarogiannis. Coarse-graining schemes for stochastic lattice systems with short and longrange interactions. Mathematics of Computation, 83(288): 1757–1793, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-288/
                                              S0025-5718-2014-02806-8;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-288/S0025-5718-
                                              2014-02806-8/S0025-5718-2014-
                                              02806-8.pdf.

King:2018:UKD

[KPSY18] Oliver D. King, Cris Poor, Jerry Shurman, and David S. Yuen. Using Katsurada’s determination of the Eisenstein series to compute Siegel eigenforms.

Mathematics of Computation, 87(310):879–892, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03218-X;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03218-X/S0025-5718-2017-
03218-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=270719;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
291737;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=364614;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
685320.

Kritzer:2014:MII

[KPW14] Peter Kritzer, Friedrich Pillichshammer, and Henryk Woźniakowski. Multivariate integration of infinitely many times differentiable functions in weighted Korobov spaces. Mathematics of Computation, 83(287): 1189–1206, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-287/
S0025-5718-2013-02739-1;
http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02739-1/S0025-5718-2013-
02739-1.pdf.

Kornhuber:2018:ACV

[KPY18] Ralf Kornhuber, Daniel Peterseim, and Harry Yserentant. An analysis of a class of variational multiscale methods based on subspace decomposition. Mathematics of Computation, 87(314):2765– 2774, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-314/
                                              S0025-5718-2018-03302-6;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-314/S0025-5718-
                                              2018-03302-6/S0025-5718-2018-
                                              03302-6.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Kornhuber%
                                              2C%20Ralf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=185935;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              848711.

Kreuzer:2012:AAU

[Kre12] Christian Kreuzer. Analysis of an adaptive Uzawa finite element method for the nonlinear Stokes problem. Mathematics of Computation, 81(277): 21–55, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-277/
                                              S0025-5718-2011-02524-X;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-277/S0025-5718-
                                              2011-02524-X/S0025-5718-2011-
                                              02524-X.pdf.

Karlsen:2014:EED

[KRS14] K. H. Karlsen, N. H. Risebro, and E. B. Storrøsten. L^{1} error estimates for difference approximations of degenerate convection-diffusion equations. Mathematics of Computation, 83(290):2717–2762, 2014. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2014-83-
290/S0025-5718-2014-02818-
4;   http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02818-4/S0025-5718-2014-
02818-4.pdf.

Krumm:2016:CPB

[Kru16]  David Krumm.  Comput-

ing points of bounded height in projective space over a number field. Mathematics of Computation, 85(297): 423–447, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-297/
S0025-5718-2015-02984-6;
http://www.ams.org/journals/
mcom/2016-85-297/S0025-5718-
2015-02984-6/S0025-5718-2015-
02984-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Krumm%
2C%20David.

Krylov:2013:IES

[Kry13] N. V. Krylov. Interior estimates for second-order differences of solutions of finitedifference elliptic Bellman’s equations. Mathematics of Computation, 82(283):1463– 1487, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2013-02684-1;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-

                                              2013-02684-1/S0025-5718-2013-
                                              02684-1.pdf.

Koornwinder:2015:IBO

[KS15] Tom H. Koornwinder and Stefan A. Sauter. The intersection of bivariate orthogonal polynomials on triangle patches. Mathematics of Computation, 84 (294):1795–1812, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-294/
                                              S0025-5718-2014-02910-4;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-294/S0025-5718-
                                              2014-02910-4/S0025-5718-2014-
                                              02910-4.pdf.

Kamalinejad:2017:CDE

[KS17] Ali Kamalinejad and Mehrdad Shahshahani. On computations with dessins d’enfants. Mathematics of Computation, 86(303): 419–436, January 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2017-86-
                                              303/S0025-5718-2016-03110-
                                              5;   http://www.ams.org/journals/
                                              mcom/2017-86-303/S0025-5718-
                                              2016-03110-5/S0025-5718-2016-
                                              03110-5.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Kamalinejad%
                                              2C%20Ali;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=159485.

Ko:2019:FEA

[KS19] Seungchan Ko and Endre Süli. Finite element approximation of steady flows of generalized Newtonian fluids with concentration-dependent powerlaw index. Mathematics of Computation, 88(317):1061– 1090, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-317/
S0025-5718-2018-03379-8;
https://www.ams.org/journals/
mcom/2019-88-317/S0025-5718-
2018-03379-8/S0025-5718-2018-
03379-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Ko%
2C%20Seungchan;   https://
www.ams.org/mathscinet/search/
authors.html?authorName=Suli%
2C%20Endre.

Kuo:2017:MQM

[KSS^{+}17] Frances Y. Kuo, Robert Scheichl, Christoph Schwab, Ian H. Sloan, and Elisabeth Ullmann. Multilevel quasi-Monte Carlo methods for lognormal diffusion problems. Mathematics of Computation, 86(308): 2827–2860, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-308/
S0025-5718-2017-03207-5;
http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03207-5/S0025-5718-2017-
03207-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Scheichl%
2C%20Robert;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Ullmann%
2C%20Elisabeth;   http://
www.ams.org/mathscinet/search/
author.html?mrauthid=163675;

                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              305221;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=703418.

Kestler:2016:EST

[KSU16] Sebastian Kestler, Kristina Steih, and Karsten Urban. An efficient space-time adaptive wavelet Galerkin method for time-periodic parabolic partial differential equations. Mathematics of Computation, 85 (299):1309–1333, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-299/
                                              S0025-5718-2015-03009-9;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-299/S0025-5718-
                                              2015-03009-9/S0025-5718-2015-
                                              03009-9.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Kestler%
                                              2C%20Sebastian;   http://
                                              www.ams.org/mathscinet/search/
                                              author.html?authorName=Steih%
                                              2C%20Kristina;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Urban%
                                              2C%20Karsten.

Kuo:2017:MIA

[KSW17] Frances Y. Kuo, Ian H. Sloan, and Henryk Woźniakowski. Multivariate integration for analytic functions with Gaussian kernels. Mathematics of Computation, 86(304): 829–853, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-304/

S0025-5718-2016-03144-0;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03144-0/S0025-5718-2016-
03144-0.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Wozniakowski%
2C%20Henryk;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=163675;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
703418.

Klagsbrun:2019:ECR

[KSW19] Zev Klagsbrun, Travis Sherman, and James Weigandt. The Elkies curve has rank 28 subject only to GRH. Mathematics of Computation, 88(316):837– 846, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03348-8;
https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03348-8/S0025-5718-2018-
03348-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Sherman%
2C%20Travis;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1016010;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
1175569.

Kuo:2010:DMF

[KSWW10] F. Y. Kuo, I. H. Sloan, G. W.

Wasilkowski, and H. Wozniakowski. On decompositions of multivariate functions. Mathematics of Computation, 79(270): 953–966, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02319-9/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02319-9/S0025-5718-09-
02319-9.pdf.

Klingenberg:2017:ALE

[KSX17] Christian Klingenberg, Gero Schnücke, and Yinhua Xia. Arbitrary Lagrangian–Eulerian discontinuous Galerkin method for conservation laws: Analysis and application in one dimension. Mathematics of Computation, 86(305):1203–1232, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-
305/S0025-5718-2016-03126-
9;   http://www.ams.org/journals/
mcom/2017-86-305/S0025-5718-
2016-03126-9/S0025-5718-2016-
03126-9.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Schnucke%
2C%20Gero;   http://www.ams.
org/mathscinet/search/author.
html?authorName=Xia%2C%20Yinhua;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
221691.

Kouznetsov:2010:PFR

[KT10] Dmitrii Kouznetsov and Henryk Trappmann. Portrait of the four regular super-exponentials to base \sqrt{2}. Mathematics of Computation, 79(271):1727– 1756, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-10-02342-2/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
10-02342-2/S0025-5718-10-
02342-2.pdf.

Kublik:2018:EAI

[KT18] Catherine Kublik and Richard Tsai. An extrapolative approach to integration over hypersurfaces in the level set framework. Mathematics of Computation, 87(313):2365– 2392, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-313/
S0025-5718-2018-03282-3;
http://www.ams.org/journals/
mcom/2018-87-313/S0025-5718-
2018-03282-3/S0025-5718-2018-
03282-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=731088;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
916969.

Kemper:2018:TTM

[KTA18] Gregor Kemper, Ngo Viet Trung, and Nguyen Thi Van Anh. Toward a theory of monomial preorders. Mathematics of Computation, 87(313):2513– 2537, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-313/
S0025-5718-2017-03289-0;
http://www.ams.org/journals/
mcom/2018-87-313/S0025-5718-
2017-03289-0/S0025-5718-2017-

                                              03289-0.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Van%
                                              20Anh%2C%20Nguyen%20Thi;
                                              https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              207806;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=608681.

Korevaar:2010:APP

[KtR10] Jaap Korevaar and Herman te Riele. Average prime-pair counting formula. Mathematics of Computation, 79(270): 1209–1229, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-270/
                                              S0025-5718-09-02312-6/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-270/S0025-5718-
                                              09-02312-6/S0025-5718-09-
                                              02312-6.pdf.

Kucuksakalli:2011:CNR

[Kuc11] Omer Kucuksakalli. Class numbers of ray class fields of imaginary quadratic fields. Mathematics of Computation, 80(274): 1099–1122, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              274/S0025-5718-2010-02413-
                                              5/;   http://www.ams.org/
                                              journals/mcom/2011-80-274/
                                              S0025-5718-2010-02413-5/S0025-
                                              5718-2010-02413-5.pdf.

Kuszmaul:2018:FAF

[Kus18] William Kuszmaul. Fast algorithms for finding pattern avoiders and counting pattern occurrences in permutations. Mathematics of Computation, 87(310):987–1011, 2018. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2018-87-
310/S0025-5718-2017-03216-
6;   http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03216-6/S0025-5718-2017-
03216-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=997955.

Kuznetsov:2015:CTT

[Kuz15] A. Kuznetsov. Computing the truncated theta function via Mordell integral. Mathematics of Computation, 84(296):2911– 2926, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-296/
S0025-5718-2015-02953-6;
http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02953-6/S0025-5718-2015-
02953-6.pdf.

Kvernadze:2010:ADF

[Kve10] George Kvernadze. Approximation of the discontinuities of a function by its classical orthogonal polynomial Fourier coefficients. Mathematics of Computation, 79(272):2265– 2285, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02366-5/;
http://www.ams.org/journals/

                                              mcom/2010-79-272/S0025-5718-
                                              10-02366-5/S0025-5718-10-
                                              02366-5.pdf.

Kruse:2019:RFD

[KW19] Raphael Kruse and Yue Wu. A randomized and fully discrete Galerkin finite element method for semilinear stochastic evolution equations. Mathematics of Computation, 88(320):2793– 2825, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-320/
                                              S0025-5718-2019-03421-X;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-320/S0025-5718-
                                              2019-03421-X/S0025-5718-2019-
                                              03421-X.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1161278;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              904215.

Kornhuber:2011:ERH

[KZ11] Ralf Kornhuber and Qingsong Zou. Efficient and reliable hierarchical error estimates for the discretization error of elliptic obstacle problems. Mathematics of Computation, 80(273): 69–88, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              273/S0025-5718-2010-02394-
                                              4/;   http://www.ams.org/
                                              journals/mcom/2011-80-273/
                                              S0025-5718-2010-02394-4/S0025-
                                              5718-2010-02394-4.pdf.

Krenn:2018:NMW

[KZ18] Daniel Krenn and Volker Ziegler. Non-minimality of the width-w non-adjacent form in conjunction with trace one τ-adic digit expansions and Koblitz curves in characteristic two. Mathematics of Computation, 87(310): 821–854, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03227-0;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03227-0/S0025-5718-2017-
03227-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Krenn%
2C%20Daniel;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=744740.

LaScala:2015:GBG

[La 15] Roberto La Scala. Gröbner bases and gradings for partial difference ideals. Mathematics of Computation, 84 (292):959–985, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-292/
S0025-5718-2014-02859-7;
http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02859-7/S0025-5718-2014-
02859-7.pdf.

Loy:2014:CDS

[LA14] R. J. Loy and R. S. Anderssen.

On the construction of Dirichlet series approximations for completely monotone functions. Mathematics of Computation, 83(286):835–846, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-286/
S0025-5718-2013-02725-1;
http://www.ams.org/journals/
mcom/2014-83-286/S0025-5718-
2013-02725-1/S0025-5718-2013-
02725-1.pdf.

Labrande:2018:CJT

[Lab18]  Hugo Labrande.  Comput-

ing Jacobi’s theta in quasilinear time. Mathematics of Computation, 87(311):1479– 1508, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-311/
S0025-5718-2017-03245-2;
http://www.ams.org/journals/
mcom/2018-87-311/S0025-5718-
2017-03245-2/S0025-5718-2017-
03245-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=989660.

Lairez:2016:CPR

[Lai16] Pierre Lairez. Computing periods of rational integrals. Mathematics of Computation, 85(300): 1719–1752, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-300/
S0025-5718-2015-03054-3;
http://www.ams.org/journals/
mcom/2016-85-300/S0025-5718-
2015-03054-3/S0025-5718-2015-
03054-3.pdf;   http://www.
ams.org/mathscinet/search/

author.html?authorName=Lairez%
2C%20Pierre.

Lanphier:2011:VSC

[Lan11]  Dominic Lanphier.  Val-

ues of symmetric cube Lfunctions and Fourier coefficients of Siegel Eisenstein series of degree-3. Mathematics of Computation, 80(273):409– 428, January 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-10-02350-1/;
http://www.ams.org/journals/
mcom/2011-80-273/S0025-5718-
10-02350-1/S0025-5718-10-
02350-1.pdf.

Lu:2017:ASH

[LCQ17] Peipei Lu, Huangxin Chen, and Weifeng Qiu. An absolutely stable hp-HDG method for the time-harmonic Maxwell equations with high wave number. Mathematics of Computation, 86(306):1553–1577, 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-
306/S0025-5718-2016-03150-
6;   http://www.ams.org/journals/
mcom/2017-86-306/S0025-5718-
2016-03150-6/S0025-5718-2016-
03150-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1030345;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
845089;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=893977.

Leppala:2016:ELB

[Lep16] Kalle Leppälä. Explicit lower bounds for linear forms. Mathematics of Computation, 85(302): 2995–3008, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2016-85-
                                              302/S0025-5718-2016-03078-
                                              1;   http://www.ams.org/journals/
                                              mcom/2016-85-302/S0025-5718-
                                              2016-03078-1/S0025-5718-2016-
                                              03078-1.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Leppala%
                                              2C%20Kalle.

Leroux:2012:CTP

[Ler12]  Louis Leroux.  Computing

the torsion points of a variety defined by lacunary polynomials. Mathematics of Computation, 81(279):1587– 1607, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-279/
                                              S0025-5718-2011-02548-2;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-279/S0025-5718-
                                              2011-02548-2/S0025-5718-2011-
                                              02548-2.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Leroux%
                                              2C%20Louis.

Lezowski:2014:CEM

[Lez14] Pierre Lezowski. Computation of the Euclidean minimum of algebraic number fields. Mathematics of Computation, 83 (287):1397–1426, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-287/
S0025-5718-2013-02746-9;
http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02746-9/S0025-5718-2013-
02746-9.pdf.

Lopez-Fernandez:2015:FSC

[LFS15] M. Lopez-Fernandez and S. Sauter.

Fast and stable contour integration for high order divided differences via elliptic functions. Mathematics of Computation, 84(293): 1291–1315, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02890-1;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02890-1/S0025-5718-2014-
02890-1.pdf.

Lui:2015:CIA

[LGY15] Lok Ming Lui, Xianfeng Gu, and Shing-Tung Yau. Convergence of an iterative algorithm for Teichmüller maps via harmonic energy optimization. Mathematics of Computation, 84(296):2823– 2842, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-296/
S0025-5718-2015-02962-7;
http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02962-7/S0025-5718-2015-
02962-7.pdf.

Li:2015:WPS

[LHY15] Jichun Li, Yunqing Huang, and Wei Yang. Well-posedness study and finite element simulation of time-domain cylindrical and elliptical cloaks. Mathematics of Computation, 84 (292):543–562, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-292/
                                              S0025-5718-2014-02911-6;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-292/S0025-5718-
                                              2014-02911-6/S0025-5718-2014-
                                              02911-6.pdf.

Li:2010:SRC

[Li10] Ren-Cang Li. Sharpness in rates of convergence for the symmetric Lanczos method. Mathematics of Computation, 79(269): 419–435, January 2010. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2010-79-
                                              269/S0025-5718-09-02258-3/
                                              ;   http://www.ams.org/journals/
                                              mcom/2010-79-269/S0025-5718-
                                              09-02258-3/S0025-5718-09-
                                              02258-3.pdf.

Li:2015:RMA

[Li15a]  Hengguang Li.  Regular-

ity and multigrid analysis for Laplace-type axisymmetric equations. Mathematics of Computation, 84(293): 1113–1144, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-293/

S0025-5718-2014-02879-2;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02879-2/S0025-5718-2014-
02879-2.pdf.

Li:2015:CIB

[Li15b] Zhilin Li. On convergence of the immersed boundary method for elliptic interface problems. Mathematics of Computation, 84(293): 1169–1188, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02932-3;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02932-3/S0025-5718-2014-
02932-3.pdf.

Li:2017:SRP

[Li17] Hengguang Li. The W_{p}^{1} stability of the Ritz projection on graded meshes. Mathematics of Computation, 86(303): 49–74, January 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-303/
S0025-5718-2016-03101-4;
http://www.ams.org/journals/
mcom/2017-86-303/S0025-5718-
2016-03101-4/S0025-5718-2016-
03101-4.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=835341.

Li:2018:AFE

[Li18] Hengguang Li. An anisotropic finite element method on polyhedral domains: interpolation error analysis. Mathematics of Computation, 87(312):1567– 1600, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2017-03290-7;
http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03290-7/S0025-5718-2017-
03290-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=835341.

Li:2019:AMR

[Li19] Buyang Li. Analyticity, maximal regularity and maximumnorm stability of semi-discrete finite element solutions of parabolic equations in nonconvex polyhedra. Mathematics of Computation, 88(315): 1–44, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03316-6;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03316-6/S0025-5718-2018-
03316-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=910552.

Licht:2019:SPM

[Lic19a] Martin W. Licht. Smoothed projections and mixed boundary conditions. Mathematics of Computation, 88(316):607– 635, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/

S0025-5718-2018-03330-0;
https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03330-0/S0025-5718-2018-
03330-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1225084.

Licht:2019:SPW

[Lic19b] Martin W. Licht. Smoothed projections over weakly Lipschitz domains. Mathematics of Computation, 88(315): 179–210, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03329-4;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03329-4/S0025-5718-2018-
03329-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1225084.

Lin:2017:LSS

[Lin17]  Lin Lin.  Localized spec-

trum slicing. Mathematics of Computation, 86(307): 2345–2371, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-307/
S0025-5718-2017-03166-5;
http://www.ams.org/journals/
mcom/2017-86-307/S0025-5718-
2017-03166-5/S0025-5718-2017-
03166-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=884840.

Liu:2015:OEE

[Liu15] Hailiang Liu. Optimal error estimates of the direct discontinuous Galerkin method for convection-diffusion equations. Mathematics of Computation, 84(295):2263–2295, 2015. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2015-84-
295/S0025-5718-2015-02923-
8;   http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-
2015-02923-8/S0025-5718-2015-
02923-8.pdf.

Li:2012:FAN

[LK12] Ren-Cang Li and William Kahan. A family of Anadromic numerical methods for matrix Riccati differential equations. Mathematics of Computation, 81(277):233–265, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2012-81-
277/S0025-5718-2011-02498-
1;   http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02498-1/S0025-5718-2011-
02498-1.pdf.

Lai:2015:PEL

[LL15] Ming-Jun Lai and Yang Liu. The probabilistic estimates on the largest and smallest q-singular values of random matrices. Mathematics of Computation, 84(294): 1775–1794, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-294/
S0025-5718-2014-02895-0;
http://www.ams.org/journals/

mcom/2015-84-294/S0025-5718-
2014-02895-0/S0025-5718-2014-
02895-0.pdf.

Larsson:2017:CDG

[LL17] Karl Larsson and Mats G. Larson. A continuous/discontinuous Galerkin method and a priori error estimates for the biharmonic problem on surfaces. Mathematics of Computation, 86 (308):2613–2649, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-308/
S0025-5718-2017-03179-3;
http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03179-3/S0025-5718-2017-
03179-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Larsson%
2C%20Karl;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=648688.

Lebowitz-Lockard:2016:GRF

[LLP16] Noah Lebowitz-Lockard and Carl Pomerance. Generating random factored Gaussian integers, easily. Mathematics of Computation, 85 (297):503–516, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-297/
S0025-5718-2015-03000-2;
http://www.ams.org/journals/
mcom/2016-85-297/S0025-5718-
2015-03000-2/S0025-5718-2015-
03000-2.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Lebowitz-
Lockard%2C%20Noah;   http:

                                              //www.ams.org/mathscinet/
                                              search/author.html?authorName=
                                              Pomerance%2C%20Carl.

Lamzouri:2015:CBL

[LLS15] Youness Lamzouri, Xiannan Li, and Kannan Soundararajan. Conditional bounds for the least quadratic non-residue and related problems. Mathematics of Computation, 84 (295):2391–2412, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-295/
                                              S0025-5718-2015-02925-1;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-295/S0025-5718-
                                              2015-02925-1/S0025-5718-2015-

02925-1.pdf. See corrigendum [LLS17a].

Lamzouri:2017:CLL

[LLS17a] Youness Lamzouri, Xiannan Li, and Kannan Soundararajan. Corrigendum to “Conditional bounds for the least quadratic non-residue and related problems”. Mathematics of Computation, 86(307): 2551–2554, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-307/
                                              S0025-5718-2017-03261-0;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-307/S0025-5718-
                                              2017-03261-0/S0025-5718-2017-
                                              03261-0.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=319775;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              804642;   http://www.ams.

org/mathscinet/search/author.
html?mrauthid=867056. See

[LLS15].

Li:2017:HSK

[LLS17b] Qin Li, Jianfeng Lu, and Weiran Sun. Half-space kinetic equations with general boundary conditions. Mathematics of Computation, 86 (305):1269–1301, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-305/
S0025-5718-2016-03155-5;
http://www.ams.org/journals/
mcom/2017-86-305/S0025-5718-
2016-03155-5/S0025-5718-2016-
03155-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Sun%
2C%20Weiran;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1016753;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
822782.

Lin:2011:FDS

[LLX11] Yumin Lin, Xianjuan Li, and Chuanju Xu. Finite difference/spectral approximations for the fractional cable equation. Mathematics of Computation, 80(275):1369–1396, July 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2010-02438-
X/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2010-02438-X/S0025-
5718-2010-02438-X.pdf.

Lu:2013:HPS

[LM13] Shuai Lu and Peter Mathé. Heuristic parameter selection based on functional minimization: Optimality and model function approach. Mathematics of Computation, 82(283): 1609–1630, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-283/
                                              S0025-5718-2013-02674-9;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-283/S0025-5718-
                                              2013-02674-9/S0025-5718-2013-
                                              02674-9.pdf.

Lubich:2015:VDW

[LM15] Christian Lubich and Dhia Mansour. Variational discretization of wave equations on evolving surfaces. Mathematics of Computation, 84 (292):513–542, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-292/
                                              S0025-5718-2014-02882-2;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-292/S0025-5718-
                                              2014-02882-2/S0025-5718-2014-
                                              02882-2.pdf.

Lezowski:2017:EAQ

[LM17] Pierre Lezowski and Kevin J. McGown. The Euclidean algorithm in quintic and septic cyclic fields. Mathematics of Computation, 86(307): 2535–2549, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-307/
S0025-5718-2017-03169-0;
http://www.ams.org/journals/
mcom/2017-86-307/S0025-5718-
2017-03169-0/S0025-5718-2017-
03169-0.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=768800;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
988126.

Lopez:2018:TTE

[LM18] L. Lopez and S. Maset.

Time-transformations for the event location in discontinuous ODEs. Mathematics of Computation, 87(313):2321– 2341, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-313/
S0025-5718-2017-03305-6;
http://www.ams.org/journals/
mcom/2018-87-313/S0025-5718-
2017-03305-6/S0025-5718-2017-
03305-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=198645;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
658579.

Lepe:2019:AVP

[LMMR19] Felipe Lepe, Salim Meddahi, David Mora, and Rodolfo Rodríguez. Acoustic vibration problem for dissipative fluids. Mathematics of Computation, 88(315): 45–71, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/

journals/mcom/2019-88-315/
S0025-5718-2018-03336-1;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03336-1/S0025-5718-2018-
03336-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Rodriguez%
2C%20Rodolfo;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1070451;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
331506;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=876029.

Linke:2018:QOP

[LMNN18] A. Linke, C. Merdon, M. Neilan, and F. Neumann. Quasioptimality of a pressure-robust nonconforming finite element method for the Stokes Problem. Mathematics of Computation, 87(312):1543–1566, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2018-03344-0;
http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2018-03344-0/S0025-5718-2018-
03344-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Neumann%
2C%20F.;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=824091;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
853921;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=901357.

Lakkis:2015:CDE

[LMP15] Omar Lakkis, Charalambos Makridakis, and Tristan Pryer. A comparison of duality and energy a posteriori estimates for L_{∞}(0, T; L_{2}(Ω)) in parabolic problems. Mathematics of Computation, 84(294):1537–1569, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/
2015-84-294/S0025-5718-2014-
02912-8;   http://www.ams.
org/journals/mcom/2015-84-
294/S0025-5718-2014-02912-
8/S0025-5718-2014-02912-8.
pdf.

Lanzara:2011:FCH

[LMS11] Flavia Lanzara, Vladimir Maz’ya, and Gunther Schmidt. On the fast computation of high dimensional volume potentials. Mathematics of Computation, 80 (274):887–904, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02425-
1/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02425-1/S0025-
5718-2010-02425-1.pdf.

Ledoux:2010:GSS

[LMT10] Veerle Ledoux, Simon J. A.

Malham, and Vera Thümmler. Grassmannian spectral shooting. Mathematics of Computation, 79(271):1585–1619, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

                                              www.ams.org/journals/mcom/
                                              2010-79-271/S0025-5718-10-
                                              02323-9/;   http://www.ams.
                                              org/journals/mcom/2010-79-
                                              271/S0025-5718-10-02323-9/
                                              S0025-5718-10-02323-9.pdf.

Lenczner:2012:DRS

[LMY12] Michel Lenczner, Gérard Montseny, and Youssef Yakoubi. Diffusive realizations for solutions of some operator equations: The onedimensional case. Mathematics of Computation, 81(277):319– 344, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-277/
                                              S0025-5718-2011-02485-3;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-277/S0025-5718-
                                              2011-02485-3/S0025-5718-2011-
                                              02485-3.pdf.

Luca:2011:LPF

[LN11] Florian Luca and Filip Najman. On the largest prime factor of x^{2} − 1. Mathematics of Computation, 80(273): 429–435, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              273/S0025-5718-2010-02381-
                                              6/;   http://www.ams.org/
                                              journals/mcom/2011-80-273/
                                              S0025-5718-2010-02381-6/S0025-
                                              5718-2010-02381-6.pdf. See

errata [LN14].

Luca:2014:ELP

[LN14] Florian Luca and Filip Najman. Errata to “On the largest prime factor of x^{2} − 1”. Mathematics of Computation, 83(285):337, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-285/
S0025-5718-2013-02696-8;
http://www.ams.org/journals/
mcom/2014-83-285/S0025-5718-
2013-02696-8/S0025-5718-2013-
02696-8.pdf. See [LN11].

Lehoucq:2018:MGM

[LNRW18] R. B. Lehoucq, F. J. Narcowich, S. T. Rowe, and J. D. Ward. A meshless Galerkin method for non-local diffusion using localized kernel bases. Mathematics of Computation, 87(313):2233– 2258, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-313/
S0025-5718-2018-03294-X;
http://www.ams.org/journals/
mcom/2018-87-313/S0025-5718-
2018-03294-X/S0025-5718-2018-
03294-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=129435;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
180590;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=611433;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
975133.

Lombardo:2019:CGE

[Lom19]  Davide Lombardo.  Com-

puting the geometric endomorphism ring of a genus-2 Jacobian. Mathematics of Computation, 88(316):889– 929, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03358-0;
https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03358-0/S0025-5718-2018-
03358-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1143804.

Louboutin:2011:UBR

[Lou11] Stéphane R. Louboutin. Upper bounds for residues of Dedekind zeta functions and class numbers of cubic and quartic number fields. Mathematics of Computation, 80(275): 1813–1822, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2011-02457-
9/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2011-02457-9/S0025-
5718-2011-02457-9.pdf.

Ling:2012:CIP

[LOX12] San Ling, Enver Ozdemir, and Chaoping Xing. Constructing irreducible polynomials over finite fields. Mathematics of Computation, 81(279): 1663–1668, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2011-02567-6;
http://www.ams.org/journals/

mcom/2012-81-279/S0025-5718-
2011-02567-6/S0025-5718-2011-
02567-6.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Ling%
2C%20San;   http://www.ams.
org/mathscinet/search/authors.
html?authorName=Ozdemir%2C%
20Enver;   http://www.ams.
org/mathscinet/search/authors.
html?authorName=Xing%2C%20Chaoping.

Ling:2014:RBE

[LOX14] San Ling, Enver Ozdemir, and Chaoping Xing. A relation between embedding degrees and class numbers of binary quadratic forms. Mathematics of Computation, 83(290): 3001–3004, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02831-7;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02831-7/S0025-5718-2014-
02831-7.pdf.

Leykekhman:2017:PDH

[LP17a] D. Leykekhman and M. Pruitt.

On the positivity of discrete harmonic functions and the discrete Harnack inequality for piecewise linear finite elements. Mathematics of Computation, 86 (305):1127–1145, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-305/
S0025-5718-2016-03117-8;
http://www.ams.org/journals/
mcom/2017-86-305/S0025-5718-

                                              2016-03117-8/S0025-5718-2016-
                                              03117-8.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Leykekhman%
                                              2C%20D;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?authorName=Pruitt%2C%
                                              20M..

Leykin:2017:DRH

[LP17b] Anton Leykin and Daniel Plaumann. Determinantal representations of hyperbolic curves via polynomial homotopy continuation. Mathematics of Computation, 86(308): 2877–2888, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-308/
                                              S0025-5718-2017-03194-X;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-308/S0025-5718-
                                              2017-03194-X/S0025-5718-2017-
                                              03194-X.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=687160;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              894950.

Lopez:2017:AFE

[LP17c] José L. López and Pedro J. Pagola. Analytic formulas for the evaluation of the Pearcey integral. Mathematics of Computation, 86(307): 2399–2407, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-307/
                                              S0025-5718-2016-03164-6;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-307/S0025-5718-

2016-03164-6/S0025-5718-2016-
03164-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Lopez%
2C%20Jose%20L.;   http://
www.ams.org/mathscinet/search/
author.html?mrauthid=806866.

Lee:2018:CFO

[LP18a] Yoonjin Lee and Yoon Kyung Park. A continued fraction of order twelve as a modular function. Mathematics of Computation, 87(312):2011– 2036, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2017-03259-2;
http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03259-2/S0025-5718-2017-
03259-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=689346;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
836403.

Lichtman:2018:IEB

[LP18b] Jared Duker Lichtman and Carl Pomerance. Improved error bounds for the Fermat primality test on random inputs. Mathematics of Computation, 87(314): 2871–2890, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-314/
S0025-5718-2018-03314-2;
http://www.ams.org/journals/
mcom/2018-87-314/S0025-5718-
2018-03314-2/S0025-5718-2018-
03314-2.pdf;   https://www.

                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Lichtman%
                                              2C%20Jared%20Duker;   https:
                                              //www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              140915.

Lipman:2013:CWD

[LPD13] Y. Lipman, J. Puente, and I. Daubechies. Conformal Wasserstein distance: II. computational aspects and extensions. Mathematics of Computation, 82(281):331–381, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-281/
                                              S0025-5718-2012-02569-5;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-281/S0025-5718-
                                              2012-02569-5/S0025-5718-2012-
                                              02569-5.pdf.

Lagomasino:2010:CCR

[LPRY10] Guillermo López Lagomasino, Domingo Pestana, José M. Rodríguez, and Dmitry Yakubovich. Computation of conformal representations of compact Riemann surfaces. Mathematics of Computation, 79(269):365– 381, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-269/
                                              S0025-5718-09-02265-0/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-269/S0025-5718-
                                              09-02265-0/S0025-5718-09-
                                              02265-0.pdf.

Li:2011:LSA

[LPSH11] Chi-Kwong Li, Yiu-Tung Poon, and Thomas Schulte-Herbrüggen. Least-squares approximation by elements from matrix orbits achieved by gradient flows on compact Lie groups. Mathematics of Computation, 80(275): 1601–1621, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2010-02450-
0/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2010-02450-0/S0025-
5718-2010-02450-0.pdf.

Lange:2019:SEP

[LR19a] Marko Lange and Siegfried M.

Rump. Sharp estimates for perturbation errors in summations. Mathematics of Computation, 88(315):349– 368, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03355-5;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03355-5/S0025-5718-2018-
03355-5.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1082372;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
151815.

Lissy:2019:OFA

[LR19b] Pierre Lissy and Ionel Rovenţa.

Optimal filtration for the approximation of boundary controls for the one-dimensional wave equation using a finitedifference method. Mathematics of Computation, 88(315): 273–291, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03345-2;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03345-2/S0025-5718-2018-
03345-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Roventa%
2C%20Ionel;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=987205.

Lelievre:2012:LDC

[LRS12] Tony Lelièvre, Mathias Rousset, and Gabriel Stoltz. Langevin dynamics with constraints and computation of free energy differences. Mathematics of Computation, 81(280):2071– 2125, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-280/
S0025-5718-2012-02594-4;
http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-
2012-02594-4/S0025-5718-2012-
02594-4.pdf.

Lercier:2016:EGO

[LRS16] Reynald Lercier, Christophe Ritzenthaler, and Jeroen Sijsling. Explicit Galois obstruction and descent for hyperelliptic curves with tamely cyclic reduced automorphism group. Mathematics of Computation, 85(300):2011–2045, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2016-85-
300/S0025-5718-2015-03032-
4;   http://www.ams.org/journals/
mcom/2016-85-300/S0025-5718-
2015-03032-4/S0025-5718-2015-
03032-4.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Lercier%
2C%20Reynald;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=702917;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
974789.

Liu:2013:EEG

[LRT13] Hailiang Liu, Olof Runborg, and Nicolay M. Tanushev. Error estimates for Gaussian beam superpositions. Mathematics of Computation, 82(282):919– 952, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02656-1;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02656-1/S0025-5718-2012-
02656-1.pdf.

Li:2010:OEE

[LS10a] Huiyuan Li and Jie Shen. Optimal error estimates in Jacobiweighted Sobolev spaces for polynomial approximations on the triangle. Mathematics of Computation, 79(271):1621– 1646, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-09-02308-4/;

                                              http://www.ams.org/journals/
                                              mcom/2010-79-271/S0025-5718-
                                              09-02308-4/S0025-5718-09-
                                              02308-4.pdf.

Lopez:2010:TPT

[LS10b] José L. López and Ester Pérez Sinusía. Two-point Taylor expansions and one-dimensional boundary value problems. Mathematics of Computation, 79 (272):2103–2115, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2010-79-
                                              272/S0025-5718-10-02370-7/
                                              ;   http://www.ams.org/journals/
                                              mcom/2010-79-272/S0025-5718-
                                              10-02370-7/S0025-5718-10-
                                              02370-7.pdf.

Lubinsky:2010:PIQ

[LS10c] D. S. Lubinsky and A. Sidi. Positive interpolatory quadrature rules generated by some biorthogonal polynomials. Mathematics of Computation, 79 (270):845–855, April 2010. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2010-79-
                                              270/S0025-5718-09-02299-6/
                                              ;   http://www.ams.org/journals/
                                              mcom/2010-79-270/S0025-5718-
                                              09-02299-6/S0025-5718-09-
                                              02299-6.pdf.

Li:2017:MAF

[LS17] Buyang Li and Weiwei Sun. Maximal L^{p} analysis of finite element solutions for parabolic equations with nonsmooth coefficients in convex polyhedra. Mathematics of Computation, 86(305): 1071–1102, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-305/
S0025-5718-2016-03133-6;
http://www.ams.org/journals/
mcom/2017-86-305/S0025-5718-
2016-03133-6/S0025-5718-2016-
03133-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Sun%
2C%20Weiwei;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=910552.

Latche:2018:CSS

[LS18] J. C. Latché and K. Saleh. A convergent staggered scheme for the variable density incompressible Navier–Stokes equations. Mathematics of Computation, 87(310):581–632, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03241-5;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03241-5/S0025-5718-2017-
03241-5.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=715367;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
960776.

Li:2019:ESC

[LSR19] Xiaoli Li, Jie Shen, and Hongxing Rui. Energy stability and convergence of SAV block-centered finite difference method for gradient flows. Mathematics of Computation, 88(319):2047– 2068, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-319/
                                              S0025-5718-2019-03428-2;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-319/S0025-5718-
                                              2019-03428-2/S0025-5718-2019-
                                              03428-2.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1152951;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              257933;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=268523.

Ling:2014:MDN

[LSW14] San Ling, Igor Shparlinski, and Huaxiong Wang. On the multidimensional distribution of the Naor–Reingold pseudorandom function. Mathematics of Computation, 83(289): 2429–2434, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-289/
                                              S0025-5718-2014-02794-4;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-289/S0025-5718-
                                              2014-02794-4/S0025-5718-2014-
                                              02794-4.pdf.

Li:2012:MPF

[LSXZ12] Yonghai Li, Shi Shu, Yuesheng Xu, and Qingsong Zou. Multilevel preconditioning for the finite volume method. Mathematics of Computation, 81(279): 1399–1428, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2012-02582-8;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2012-02582-8/S0025-5718-2012-
02582-8.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Li%
2C%20Yonghai;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Shu%
2C%20Shi;   http://www.ams.
org/mathscinet/search/authors.
html?authorName=Xu%2C%20Yuesheng;
http://www.ams.org/mathscinet/
search/authors.html?authorName=
Zou%2C%20Qingsong.

Loeffler:2012:CLC

[LW12] David Loeffler and Jared Weinstein. On the computation of local components of a newform. Mathematics of Computation, 81(278):1179– 1200, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02530-5;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02530-5/S0025-5718-2011-
02530-5.pdf.  See erratum

[LW15].

Loeffler:2015:ECL

[LW15] David Loeffler and Jared Weinstein. Erratum: “On the computation of local components of a newform”. Mathematics of Computation, 84 (291):355–356, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-291/
S0025-5718-2014-02867-6;
http://www.ams.org/journals/
mcom/2015-84-291/S0025-5718-
2014-02867-6/S0025-5718-2014-
02867-6.pdf. See [LW12].

Liu:2016:ARK

[LW16] Ji Liu and Stephen J. Wright. An accelerated randomized Kaczmarz algorithm. Mathematics of Computation, 85 (297):153–178, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-297/
S0025-5718-2015-02971-8;
http://www.ams.org/journals/
mcom/2016-85-297/S0025-5718-
2015-02971-8/S0025-5718-2015-
02971-8.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Liu%
2C%20Ji;   http://www.ams.
org/mathscinet/search/author.
html?authorName=Wright%2C%
20Stephen%20J..

Lee:2018:LCA

[LW18a] Jeonghun J. Lee and Ragnar Winther. Local coderivatives and approximation of Hodge Laplace problems. Mathematics of Computation, 87(314): 2709–2735, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-314/
S0025-5718-2018-03315-4;
http://www.ams.org/journals/

mcom/2018-87-314/S0025-5718-
2018-03315-4/S0025-5718-2018-
03315-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1067639;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
183665.

Liu:2018:EEA

[LW18b] Hailiang Liu and Hairui Wen.

Error estimates for the AEDG method to one-dimensional linear convection–diffusion equations. Mathematics of Computation, 87(310):123–148, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2018-87-
310/S0025-5718-2017-03226-
9;   http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03226-9/S0025-5718-2017-
03226-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Liu%
2C%20Hailiang;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=753749.

Liu:2013:NRR

[LWCI13] Xiaoji Liu, Shuxia Wu, and Dragana S. Cvetković-Ilić. New results on reverse order law for 1, 2, 3- and 1, 2, 4-inverses of bounded operators. Mathematics of Computation, 82(283): 1597–1607, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2013-02660-9;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-

                                              2013-02660-9/S0025-5718-2013-
                                              02660-9.pdf.

Liu:2019:OEE

[LWL19] Wenjie Liu, Li-Lian Wang, and Huiyuan Li. Optimal error estimates for Chebyshev approximations of functions with limited regularity in fractional Sobolev-type spaces. Mathematics of Computation, 88(320):2857–2895, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-320/
                                              S0025-5718-2019-03456-7;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-320/S0025-5718-
                                              2019-03456-7/S0025-5718-2019-
                                              03456-7.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1077719;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              681795;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=708582.

Liu:2018:PPA

[LWZ18] Jian-Guo Liu, Li Wang, and Zhennan Zhou. Positivitypreserving and asymptotic preserving method for 2D Keller– Segal equations. Mathematics of Computation, 87(311): 1165–1189, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-311/
                                              S0025-5718-2017-03250-6;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-311/S0025-5718-
                                              2017-03250-6/S0025-5718-2017-

03250-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1067205;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
233036;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=239750.

Lin:2015:MLC

[LX15] Qun Lin and Hehu Xie. A multi-level correction scheme for eigenvalue problems. Mathematics of Computation, 84 (291):71–88, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-291/
S0025-5718-2014-02825-1;
http://www.ams.org/journals/
mcom/2015-84-291/S0025-5718-
2014-02825-1/S0025-5718-2014-
02825-1.pdf.

Loureiro:2019:VTC

[LX19] Ana F. Loureiro and Kuan Xu. Volterra-type convolution of classical polynomials. Mathematics of Computation, 88(319): 2351–2381, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-319/
S0025-5718-2019-03427-0;
https://www.ams.org/journals/
mcom/2019-88-319/S0025-5718-
2019-03427-0/S0025-5718-2019-
03427-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1010592;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
793430.

Lin:2014:LBD

[LXX14] Qun Lin, Hehu Xie, and Jinchao Xu. Lower bounds of the discretization error for piecewise polynomials. Mathematics of Computation, 83 (285):1–13, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-285/
                                              S0025-5718-2013-02724-X;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-285/S0025-5718-
                                              2013-02724-X/S0025-5718-2013-
                                              02724-X.pdf.

Liu:2017:RPB

[LY17] Jian-Guo Liu and Rong Yang. A random particle blob method for the Keller–Segel equation and convergence analysis. Mathematics of Computation, 86 (304):725–745, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-304/
                                              S0025-5718-2016-03118-X;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-304/S0025-5718-
                                              2016-03118-X/S0025-5718-2016-
                                              03118-X.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=1060676;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              233036.

Liu:2013:EBU

[LZ13] Youming Liu and Junjian Zhao. An extension of Bittner and Urban’s theorem. Mathematics of Computation, 82(281):401– 411, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-281/
S0025-5718-2012-02592-0;
http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-
2012-02592-0/S0025-5718-2012-
02592-0.pdf.

Lescarret:2015:NAS

[LZ15] Vincent Lescarret and Enrique Zuazua. Numerical approximation schemes for multidimensional wave equations in asymmetric spaces. Mathematics of Computation, 84 (291):119–152, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-291/
S0025-5718-2014-02887-1;
http://www.ams.org/journals/
mcom/2015-84-291/S0025-5718-
2014-02887-1/S0025-5718-2014-
02887-1.pdf.

Lecaros:2016:CSC

[LZ16] Rodrigo Lecaros and Enrique Zuazua. Control of 2D scalar conservation laws in the presence of shocks. Mathematics of Computation, 85(299): 1183–1224, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2015-03015-4;
http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2015-03015-4/S0025-5718-2015-
03015-4.pdf;   http://www.
ams.org/mathscinet/search/

                                              author.html?authorName=Lecaros%
                                              2C%20Rodrigo;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=187655.

Li:2017:MNA

[LZ17] Buyang Li and Zhimin Zhang. Mathematical and numerical analysis of the time-dependent Ginzburg–Landau equations in nonconvex polygons based on Hodge decomposition. Mathematics of Computation, 86 (306):1579–1608, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-306/
                                              S0025-5718-2016-03177-4;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-306/S0025-5718-
                                              2016-03177-4/S0025-5718-2016-
                                              03177-4.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=303173;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              910552.

Lu:2018:FGA

[LZ18] Jianfeng Lu and Zhennan Zhou. Frozen Gaussian approximation with surface hopping for mixed quantum-classical dynamics: a mathematical justification of fewest switches surface hopping algorithms. Mathematics of Computation, 87(313):2189– 2232, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-313/
                                              S0025-5718-2017-03310-X;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-313/S0025-5718-

2017-03310-X/S0025-5718-2017-
03310-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1067205;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
822782.

Mahe:2014:PPT

[Mah14] Valéry Mahé. Prime power terms in elliptic divisibility sequences. Mathematics of Computation, 83(288): 1951–1991, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-288/
S0025-5718-2013-02790-1;
http://www.ams.org/journals/
mcom/2014-83-288/S0025-5718-
2013-02790-1/S0025-5718-2013-
02790-1.pdf.

Mascot:2018:CMG

[Mas18] Nicolas Mascot. Certification of modular Galois representations. Mathematics of Computation, 87(310):381–423, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03215-4;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03215-4/S0025-5718-2017-
03215-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1040021.

Mathe:2019:BIP

[Mat19] Peter Mathé. Bayesian inverse problems with non-commuting operators. Mathematics of Computation, 88(320):2897– 2912, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-320/
S0025-5718-2019-03439-7;
https://www.ams.org/journals/
mcom/2019-88-320/S0025-5718-
2019-03439-7/S0025-5718-2019-
03439-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Mathe%
2C%20Peter.

Mirebeau:2012:GBG

[MC12] Jean-Marie Mirebeau and Albert Cohen. Greedy bisection generates optimally adapted triangulations. Mathematics of Computation, 81(278):811– 837, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02459-2;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02459-2/S0025-5718-2011-
02459-2.pdf.

Martinhao:2013:SCC

[MC13] Anderson N. Martinhão and Emerson L. Monte Carmelo. Short covering codes arising from matchings in weighted graphs. Mathematics of Computation, 82(281):605–616, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-281/
S0025-5718-2012-02613-5;

http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-
2012-02613-5/S0025-5718-2012-
02613-5.pdf.

Morais:2015:QZS

[MC15] J. Morais and I. Caçao.

Quaternion Zernike spherical polynomials. Mathematics of Computation, 84(293): 1317–1337, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02888-3;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02888-3/S0025-5718-2014-
02888-3.pdf.

McKee:2011:CTP

[McK11] James McKee. Computing totally positive algebraic integers of small trace. Mathematics of Computation, 80(274): 1041–1052, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02424-
X/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02424-X/S0025-
5718-2010-02424-X.pdf.

McNew:2015:SIW

[McN15]  Nathan McNew.  On sets

of integers which contain no three terms in geometric progression. Mathematics of Computation, 84(296):2893– 2910, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-296/
S0025-5718-2015-02979-2;
http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02979-2/S0025-5718-2015-
02979-2.pdf.

Makarov:2013:FMS

[MDK13] V. L. Makarov, D. V. Dragunov, and Ya. V. Klimenko. The FD-method for solving Sturm–Liouville problems with special singular differential operator. Mathematics of Computation, 82(282):953– 973, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02634-2;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02634-2/S0025-5718-2012-
02634-2.pdf.

Melman:2016:PTC

[Mel16]  A. Melman.  On Pellet’s

Theorem for a class of lacunary polynomials. Mathematics of Computation, 85 (298):707–716, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-03011-7;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-03011-7/S0025-5718-2015-
03011-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=293268.

Melman:2018:EBM

[Mel18] A. Melman. Eigenvalue bounds for matrix polynomials in generalized bases. Mathematics of Computation, 87(312):1935– 1948, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2017-03252-X;
http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03252-X/S0025-5718-2017-
03252-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=293268.

Mestrovic:2014:SPS

[Meš14] Romeo Meštrović. A search for primes p such that the Euler number E_{p−3} is divisible by p. Mathematics of Computation, 83(290): 2967–2976, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02814-7;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02814-7/S0025-5718-2014-
02814-7.pdf.

Mezhericher:2011:EWF

[Mez11] Borislav Mezhericher. Evaluating Whittaker functions and Maass forms for SL(3, ℤ). Mathematics of Computation, 80 (276):2299–2313, October 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-

276/S0025-5718-2011-02499-
3/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02499-3/S0025-
5718-2011-02499-3.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813362.

Muntingh:2011:DDI

[MF11] Georg Muntingh and Michael Floater. Divided differences of implicit functions. Mathematics of Computation, 80(276): 2185–2195, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02486-
5/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02486-5/S0025-
5718-2011-02486-5.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813354.

Martinez-Finkelshtein:2018:EZS

[MFRV18] A. Martínez-Finkelshtein, A. Sri Ranga, and D. O. Veronese. Extreme zeros in a sequence of para-orthogonal polynomials and bounds for the support of the measure. Mathematics of Computation, 87 (310):261–288, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03210-5;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03210-5/S0025-5718-2017-
03210-5.pdf;   https://www.
ams.org/mathscinet/search/

authors.html?mrauthid=238837;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
248069;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=928590.

Miller:2015:RCF

[Mil15] John C. Miller. Real cyclotomic fields of prime conductor and their class numbers. Mathematics of Computation, 84 (295):2459–2469, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-295/
S0025-5718-2015-02924-X;
http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-
2015-02924-X/S0025-5718-2015-
02924-X.pdf.

Mosunov:2016:UCG

[MJ16] A. S. Mosunov and M. J.

Jacobson, Jr. Unconditional class group tabulation of imaginary quadratic fields to |Δ| < 2^{40}. Mathematics of Computation, 85(300): 1983–2009, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-300/
S0025-5718-2015-03050-6;
http://www.ams.org/journals/
mcom/2016-85-300/S0025-5718-
2015-03050-6/S0025-5718-2015-
03050-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Mosunov%
2C%20A.%20S;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=606532.

Mekchay:2011:ALB

[MMN11] Khamron Mekchay, Pedro Morin, and Ricardo H. Nochetto. AFEM for the Laplace– Beltrami operator on graphs: Design and conditional contraction property. Mathematics of Computation, 80(274): 625–648, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              274/S0025-5718-2010-02435-
                                              4/;   http://www.ams.org/
                                              journals/mcom/2011-80-274/
                                              S0025-5718-2010-02435-4/S0025-
                                              5718-2010-02435-4.pdf.

McLachlan:2017:MVS

[MMV17] Robert McLachlan, Klas Modin, and Olivier Verdier. A minimalvariable symplectic integrator on spheres. Mathematics of Computation, 86(307): 2325–2344, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-307/
                                              S0025-5718-2016-03153-1;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-307/S0025-5718-
                                              2016-03153-1/S0025-5718-2016-
                                              03153-1.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Modin%
                                              2C%20Klas;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=321838;   http:
                                              //www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              873226.

Mastroianni:2010:LTP

[MN10] G. Mastroianni and I. Notarangelo. A Lagrange-type projector on the real line. Mathematics of Computation, 79(269): 327–352, January 2010. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2010-79-
269/S0025-5718-09-02278-9/
;   http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02278-9/S0025-5718-09-
02278-9.pdf.

Moore:2014:ASS

[MN14] Dennis Moore and Uwe Nagel.

Algorithms for strongly stable ideals. Mathematics of Computation, 83(289):2527–2552, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/
2014-83-289/S0025-5718-2014-
02784-1;   http://www.ams.
org/journals/mcom/2014-83-
289/S0025-5718-2014-02784-
1/S0025-5718-2014-02784-1.
pdf.

Molin:2019:CPM

[MN19a] Pascal Molin and Christian Neurohr. Computing period matrices and the Abel– Jacobi map of superelliptic curves. Mathematics of Computation, 88(316):847– 888, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03351-8;

https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03351-8/S0025-5718-2018-
03351-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Neurohr%
2C%20Christian;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=884381.

Moret:2019:KSM

[MN19b] Igor Moret and Paolo Novati. Krylov subspace methods for functions of fractional differential operators. Mathematics of Computation, 88(315): 293–312, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03332-4;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03332-4/S0025-5718-2018-
03332-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Moret%
2C%20Igor;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=679699.

Mhaskar:2010:BEA

[MNPW10] H. N. Mhaskar, F. J. Narcowich, J. Prestin, and J. D. Ward. L^{p} Bernstein estimates and approximation by spherical basis functions. Mathematics of Computation, 79(271): 1647–1679, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-09-02322-9/;
http://www.ams.org/journals/

mcom/2010-79-271/S0025-5718-
09-02322-9/S0025-5718-09-
02322-9.pdf.

Myasnikov:2015:KPG

[MNU15] Alexei Myasnikov, Andrey Nikolaev, and Alexander Ushakov. Knapsack problems in groups. Mathematics of Computation, 84(292):987–1016, 2015. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2015-84-
292/S0025-5718-2014-02880-
9;   http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02880-9/S0025-5718-2014-
02880-9.pdf.

Montgomery:2010:PBN

[MNW10] Peter L. Montgomery, Sangil Nahm, and Samuel S. Wagstaff, Jr. The period of the Bell numbers modulo a prime. Mathematics of Computation, 79(271): 1793–1800, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-10-02340-9/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
10-02340-9/S0025-5718-10-
02340-9.pdf.

Melquiond:2013:NAM

[MNZ13] Guillaume Melquiond, W. Georg Nowak, and Paul Zimmermann. Numerical approximation of the Masser–Gramain constant to four decimal digits: δ = 1.819 . . .. Mathematics of Computation, 82(282):

1235–1246, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02635-4;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02635-4/S0025-5718-2012-
02635-4.pdf.

Molchanov:2013:CHP

[MO13] Vladimir Molchanov and Marcel Oliver. Convergence of the Hamiltonian particlemesh method for barotropic fluid flow. Mathematics of Computation, 82(282):861– 891, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02648-2;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02648-2/S0025-5718-2012-
02648-2.pdf.

Monien:2010:GQS

[Mon10a] H. Monien. Gaussian quadrature for sums: a rapidly convergent summation scheme. Mathematics of Computation, 79(270): 857–869, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02289-3/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02289-3/S0025-5718-09-
02289-3.pdf.

Monteillet:2010:CAS

[Mon10b] Aurélien Monteillet. Convergence of approximation schemes for nonlocal front propagation equations. Mathematics of Computation, 79(269):125– 146, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-269/
S0025-5718-09-02270-4/;
http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02270-4/S0025-5718-09-
02270-4.pdf.

Mortici:2011:SBL

[Mor11] Cristinel Mortici. Sharp bounds of the Landau constants. Mathematics of Computation, 80(274): 1011–1018, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02428-
7/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02428-7/S0025-
5718-2010-02428-7.pdf.

Morra:2013:ACR

[Mor13]  Anna Morra.  An algo-

rithm to compute relative cubic fields. Mathematics of Computation, 82(284): 2343–2361, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-284/
S0025-5718-2013-02686-5;
http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-

                                              2013-02686-5/S0025-5718-2013-
                                              02686-5.pdf.

Morrow:2019:CIG

[Mor19] Jackson S. Morrow. Composite images of Galois for elliptic curves over Q and entanglement fields. Mathematics of Computation, 88(319): 2389–2421, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-319/
                                              S0025-5718-2019-03426-9;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-319/S0025-5718-
                                              2019-03426-9/S0025-5718-2019-
                                              03426-9.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1099817.

Magnus:1966:FTS

[MOS66] Wilhelm Magnus, Fritz Oberhettinger, and Raj Pal Soni. Formulas and theorems for the special functions of mathematical physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1966. viii + 508 pp. LCCN QA1 G88 v. 52, 1966. See errata [Coh12].

Mourrain:2014:DSS

[Mou14] Bernard Mourrain. On the dimension of spline spaces on planar T-meshes. Mathematics of Computation, 83 (286):847–871, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-286/
                                              S0025-5718-2013-02738-X;

http://www.ams.org/journals/
mcom/2014-83-286/S0025-5718-
2013-02738-X/S0025-5718-2013-
02738-X.pdf.

Mehmetoglu:2012:MPC

[MP12] Orhan Mehmetoğlu and Bojan Popov. Maximum principle and convergence of central schemes based on slope limiters. Mathematics of Computation, 81(277):219–231, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02514-7;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02514-7/S0025-5718-2011-
02514-7.pdf.

Maalqvist:2014:LEM

[MP14] Axel Målqvist and Daniel Peterseim. Localization of elliptic multiscale problems. Mathematics of Computation, 83 (290):2583–2603, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02868-8;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02868-8/S0025-5718-2014-
02868-8.pdf.

Marin:2017:BFC

[MP17] Ivan Marin and Götz Pfeiffer. The BMR freeness conjecture for the 2-reflection groups. Mathematics of Computation, 86(306):2005–2023, 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-
306/S0025-5718-2016-03142-
7;   http://www.ams.org/journals/
mcom/2017-86-306/S0025-5718-
2016-03142-7/S0025-5718-2016-
03142-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Pfeiffer%
2C%20Gotz;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=664485.

Malnic:2019:SLS

[MP19] Aleksander Malnič and Rok Požar. On split liftings with sectional complements. Mathematics of Computation, 88 (316):983–1005, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

www.ams.org/journals/mcom/
2019-88-316/S0025-5718-2018-
03352-X;   https://www.ams.
org/journals/mcom/2019-88-
316/S0025-5718-2018-03352-
X/S0025-5718-2018-03352-X.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?authorName=Malnic%2C%
20Aleksander;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Pozar%
2C%20Rok.

Morini:2018:AND

[MPT18] Benedetta Morini, Margherita Porcelli, and Philippe L. Toint. Approximate norm descent methods for constrained nonlinear systems. Mathematics of Computation, 87(311):1327– 1351, July 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-311/
S0025-5718-2017-03251-8;
http://www.ams.org/journals/
mcom/2018-87-311/S0025-5718-
2017-03251-8/S0025-5718-2017-
03251-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Toint%
2C%20Philippe%20L.;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
608586;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=867592.

Moore:2010:PEE

[MR10] Peter K. Moore and Marina Rangelova. A posteriori error estimation for hp-adaptivity for fourth-order equations. Mathematics of Computation, 79(270): 677–705, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02290-X/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02290-X/S0025-5718-09-
02290-X.pdf.

Mazza:2019:SAS

[MRSC19] Mariarosa Mazza, Ahmed Ratnani, and Stefano Serra-Capizzano. Spectral analysis and spectral symbol for the 2D curl–curl (stabilized) operator with applications to the related iterative solutions. Mathematics of Computation, 88(317):1155– 1188, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-317/
                                              S0025-5718-2018-03366-X;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-317/S0025-5718-
                                              2018-03366-X/S0025-5718-2018-
                                              03366-X.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Ratnani%
                                              2C%20Ahmed;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1079112;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              332436.

Matthews:2010:MCS

[MRW10] Keith Matthews, John Robertson, and Jim White. Midpoint criteria for solving Pell’s equation using the nearest square continued fraction. Mathematics of Computation, 79(269):485– 499, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-269/
                                              S0025-5718-09-02286-8/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-269/S0025-5718-
                                              09-02286-8/S0025-5718-09-
                                              02286-8.pdf.

Melenk:2010:CAF

[MS10] J. M. Melenk and S. Sauter. Convergence analysis for finite element discretizations of the Helmholtz equation with Dirichlet-to-Neumann boundary conditions. Mathematics of Computation, 79(272):1871– 1914, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02362-8/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02362-8/S0025-5718-10-
02362-8.pdf.

Mishra:2012:STM

[MS12] S. Mishra and Ch. Schwab.

Sparse tensor multi-level Monte Carlo finite volume methods for hyperbolic conservation laws with random initial data. Mathematics of Computation, 81(280):1979– 2018, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-280/
S0025-5718-2012-02574-9;
http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-
2012-02574-9/S0025-5718-2012-
02574-9.pdf.

Miller:2013:EID

[MS13] Robert L. Miller and Michael Stoll. Explicit isogeny descent on elliptic curves. Mathematics of Computation, 82(281):513– 529, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-281/
S0025-5718-2012-02619-6;
http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-
2012-02619-6/S0025-5718-2012-
02619-6.pdf.

Milovanovic:2014:KEM

[MS14] Gradimir V. Milovanović and Miodrag M. Spalević. Kronrod extensions with multiple nodes of quadrature formulas for Fourier coefficients. Mathematics of Computation, 83 (287):1207–1231, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-287/
S0025-5718-2013-02761-5;
http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02761-5/S0025-5718-2013-
02761-5.pdf.

Moody:2016:AVF

[MS16] Dustin Moody and Daniel Shumow. Analogues of Vélu’s formulas for isogenies on alternate models of elliptic curves. Mathematics of Computation, 85(300): 1929–1951, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-300/
S0025-5718-2015-03036-1;
http://www.ams.org/journals/
mcom/2016-85-300/S0025-5718-
2015-03036-1/S0025-5718-2015-
03036-1.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1072502;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
870964.

Matsumoto:2014:CFM

[MSM14] Makoto Matsumoto, Mutsuo Saito, and Kyle Matoba. A computable figure of merit for quasi-Monte Carlo point sets. Mathematics of Computation, 83 (287):1233–1250, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-287/
S0025-5718-2013-02774-3;
http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02774-3/S0025-5718-2013-
02774-3.pdf.

Meddahi:2011:NCB

[MSS11] Salim Meddahi, Francisco-Javier Sayas, and Virginia Selgás. Nonsymmetric coupling of BEM and mixed FEM on polyhedral interfaces. Mathematics of Computation, 80(273): 43–68, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02401-
9/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02401-9/S0025-
5718-2010-02401-9.pdf.

McGown:2019:NEH

[MSV19] Kevin J. McGown, Jonathan W.

Sands, and Daniel Vallières. Numerical evidence for higher order Stark-type conjectures. Mathematics of Computation, 88(315):389–420, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

www.ams.org/journals/mcom/
2019-88-315/S0025-5718-2018-
03337-3;   https://www.ams.
org/journals/mcom/2019-88-
315/S0025-5718-2018-03337-
3/S0025-5718-2018-03337-3.
pdf;   https://www.ams.org/
mathscinet/search/authors.

                                              html?authorName=Vallieres%
                                              2C%20Daniel;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=154195;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              768800.

Meng:2016:OEE

[MSW16] Xiong Meng, Chi-Wang Shu, and Boying Wu. Optimal error estimates for discontinuous Galerkin methods based on upwind-biased fluxes for linear hyperbolic equations. Mathematics of Computation, 85 (299):1225–1261, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-299/
                                              S0025-5718-2015-03022-1;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-299/S0025-5718-
                                              2015-03022-1/S0025-5718-2015-
                                              03022-1.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Wu%
                                              2C%20Boying;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=242268;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              998988.

Murru:2019:AMC

[MT19] Nadir Murru and Lea Terracini. On p-adic multidimensional continued fractions. Mathematics of Computation, 88(320):2913–2934, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-320/

S0025-5718-2019-03450-6;
https://www.ams.org/journals/
mcom/2019-88-320/S0025-5718-
2019-03450-6/S0025-5718-2019-
03450-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=261537;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
905269.

Muller:2010:ENE

[Mül10] Siguna Müller. On the existence and non-existence of elliptic pseudoprimes. Mathematics of Computation, 79(270): 1171–1190, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02275-3/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02275-3/S0025-5718-09-
02275-3.pdf.

Muller:2014:CCH

[Mül14] Jan Steffen Müller. Computing canonical heights using arithmetic intersection theory. Mathematics of Computation, 83 (285):311–336, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-285/
S0025-5718-2013-02719-6;
http://www.ams.org/journals/
mcom/2014-83-285/S0025-5718-
2013-02719-6/S0025-5718-2013-
02719-6.pdf.

Mustapha:2013:SDG

[Mus13] Kassem Mustapha. A superconvergent discontinuous Galerkin method for Volterra integrodifferential equations, smooth and non-smooth kernels. Mathematics of Computation, 82 (284):1987–2005, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-284/
S0025-5718-2013-02689-0;
http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02689-0/S0025-5718-2013-
02689-0.pdf.

Mustapha:2018:FTF

[Mus18]  Kassem Mustapha.  FEM

for time-fractional diffusion equations, novel optimal error analyses. Mathematics of Computation, 87(313):2259– 2272, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-313/
S0025-5718-2018-03304-X;
http://www.ams.org/journals/
mcom/2018-87-313/S0025-5718-
2018-03304-X/S0025-5718-2018-
03304-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=727133.

Ma:2018:CHO

[MX18] Yunyun Ma and Yuesheng Xu. Computing highly oscillatory integrals. Mathematics of Computation, 87 (310):309–345, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03214-2;
http://www.ams.org/journals/

mcom/2018-87-310/S0025-5718-
2017-03214-2/S0025-5718-2017-
03214-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Xu%
2C%20Yuesheng;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=975223.

Martins:2012:VAM

[MYS12] M. M. Martins, W. Yousif, and J. L. Santos. A variant of the AOR method for augmented systems. Mathematics of Computation, 81(277):399– 417, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02483-X;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02483-X/S0025-5718-2011-
02483-X.pdf.

Mu:2010:DSN

[MZ10] Mo Mu and Xiaohong Zhu.

Decoupled schemes for a nonstationary mixed Stokes–Darcy model. Mathematics of Computation, 79(270):707–731, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/
2010-79-270/S0025-5718-09-
02302-3/;   http://www.ams.
org/journals/mcom/2010-79-
270/S0025-5718-09-02302-3/
S0025-5718-09-02302-3.pdf.

Moeller:2016:FSR

[MZ16] Michael Moeller and Xiaoqun Zhang. Fast sparse reconstruction: Greedy inverse scale space flows. Mathematics of Computation, 85 (297):179–208, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-297/
S0025-5718-2015-03004-X;
http://www.ams.org/journals/
mcom/2016-85-297/S0025-5718-
2015-03004-X/S0025-5718-2015-
03004-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Moeller%
2C%20Michael;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Zhang%
2C%20Xiaoqun.

Nakatsukasa:2011:GTG

[Nak11] Yuji Nakatsukasa. Gerschgorin’s theorem for generalized eigenvalue problems in the Euclidean metric. Mathematics of Computation, 80(276): 2127–2142, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02482-
8/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02482-8/S0025-
5718-2011-02482-8.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813351.

Nart:2014:LCD

[Nar14]  Enric Nart.  Local com-

putation of differents and discriminants. Mathematics of Computation, 83(287): 1513–1534, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-287/
S0025-5718-2013-02754-8;
http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02754-8/S0025-5718-2013-
02754-8.pdf.

Nazareth:2010:BRI

[Naz10] J. L. Nazareth. Book review:

Introduction to derivative-free optimization. Mathematics of Computation, 79(271):1867– 1869, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-10-02379-3/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
10-02379-3/S0025-5718-10-
02379-3.pdf.

Nemes:2019:AEI

[ND19] Gergő Nemes and Adri B. Olde Daalhuis. Asymptotic expansions for the incomplete gamma function in the transition regions. Mathematics of Computation, 88(318):1805– 1827, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-318/
S0025-5718-2018-03391-9;
https://www.ams.org/journals/
mcom/2019-88-318/S0025-5718-
2018-03391-9/S0025-5718-2018-
03391-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Nemes%
2C%20Gergo;   https://www.

                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=293428.

Neilan:2015:DCS

[Nei15] Michael Neilan. Discrete and conforming smooth de Rham complexes in three dimensions. Mathematics of Computation, 84(295):2059–2081, 2015. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2015-84-
                                              295/S0025-5718-2015-02958-
                                              5;   http://www.ams.org/journals/
                                              mcom/2015-84-295/S0025-5718-
                                              2015-02958-5/S0025-5718-2015-
                                              02958-5.pdf.

Nelson:2015:EMF

[Nel15]  Paul D. Nelson.  Eval-

uating modular forms on Shimura curves. Mathematics of Computation, 84(295): 2471–2503, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-295/
                                              S0025-5718-2015-02943-3;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-295/S0025-5718-
                                              2015-02943-3/S0025-5718-2015-
                                              02943-3.pdf.

Nguyen:2016:CNG

[Ngu16] Hoi H. Nguyen. On a condition number of general random polynomial systems. Mathematics of Computation, 85 (298):737–757, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-298/
                                              S0025-5718-2015-02993-7;

http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-02993-7/S0025-5718-2015-
02993-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Nguyen%
2C%20Hoi%20H..

Narayan:2011:GWR

[NH11] Akil C. Narayan and Jan S.

Hesthaven. A generalization of the Wiener rational basis functions on infinite intervals: Part I — derivation and properties. Mathematics of Computation, 80(275):1557–1583, July 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2010-02437-
8/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2010-02437-8/S0025-
5718-2010-02437-8.pdf.

Nielsen:2015:OPN

[Nie15] Pace P. Nielsen. Odd perfect numbers, Diophantine equations, and upper bounds. Mathematics of Computation, 84 (295):2549–2567, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-295/
S0025-5718-2015-02941-X;
http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-
2015-02941-X/S0025-5718-2015-
02941-X.pdf.

Narayan:2017:CFW

[NJZ17] Akil Narayan, John D. Jakeman, and Tao Zhou. A Christoffel function weighted least squares algorithm for collocation approximations. Mathematics of Computation, 86 (306):1913–1947, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-306/
S0025-5718-2016-03192-0;
http://www.ams.org/journals/
mcom/2017-86-306/S0025-5718-
2016-03192-0/S0025-5718-2016-
03192-0.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=862962;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
908982;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=932761.

Niemimaki:2017:SPM

[NKK17] Ossi Niemimäki, Stefan Kurz, and Lauri Kettunen. Structurepreserving mesh coupling based on the Buffa–Christiansen complex. Mathematics of Computation, 86(304):507–524, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-
304/S0025-5718-2016-03121-
X;   http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03121-X/S0025-5718-2016-
03121-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Niemimaki%
2C%20Ossi;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=630615;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
768977.

Nordstrom:2019:ESB

[NL19] Jan Nordström and Cristina La Cognata. Energy stable boundary conditions for the nonlinear incompressible Navier– Stokes equations. Mathematics of Computation, 88(316):665– 690, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03375-0;
https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03375-0/S0025-5718-2018-
03375-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=La%
20Cognata%2C%20Cristina;
https://www.ams.org/mathscinet/
search/authors.html?authorName=
Nordstrom%2C%20Jan.

Neumann:2012:ECD

[NM12] V. G. Lopez Neumann and Constantin Manoil. Explicit computations on the desingularized Kummer surface. Mathematics of Computation, 81(278): 1149–1161, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02547-0;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02547-0/S0025-5718-2011-
02547-0.pdf.

Nicaise:2017:MDF

[NM17] S. Nicaise and I. Merabet.

A mixed discontinuous finite element method for folded Naghdi’s shell in Cartesian coordinates. Mathematics of Computation, 86(303):1– 47, January 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-303/
S0025-5718-2016-03094-X;
http://www.ams.org/journals/
mcom/2017-86-303/S0025-5718-
2016-03094-X/S0025-5718-2016-
03094-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Nicaise%
2C%20S;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=1011564.

Nakatsukasa:2016:SCP

[NN16] Yuji Nakatsukasa and Vanni Noferini. On the stability of computing polynomial roots via confederate linearizations. Mathematics of Computation, 85(301):2391–2425, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-301/
S0025-5718-2015-03049-X;
http://www.ams.org/journals/
mcom/2016-85-301/S0025-5718-
2015-03049-X/S0025-5718-2015-
03049-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Nakatsukasa%
2C%20Yuji;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=936379.

Nochetto:2019:TSM

[NNZ19] R. H. Nochetto, D. Ntogkas, and W. Zhang. Two-scale method for the Monge–Ampère equation: Convergence to the viscosity solution. Mathematics of Computation, 88(316):637– 664, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03353-1;
https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03353-1/S0025-5718-2018-
03353-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Ntogkas%
2C%20D.;   https://www.ams.
org/mathscinet/search/authors.
html?authorName=Zhang%2C%
20W.;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=131850.

Notaris:2015:ENG

[Not15] Sotirios E. Notaris. The error norm of Gauss–Radau quadrature formulae for Bernstein– Szegő weight functions. Mathematics of Computation, 84(296): 2843–2865, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-296/
S0025-5718-2015-02944-5;
http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02944-5/S0025-5718-2015-
02944-5.pdf.

Naughton:2012:CTM

[NP12] L. Naughton and G. Pfeiffer.

Computing the table of marks of a cyclic extension. Mathematics of Computation, 81(280):2419– 2438, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-280/
                                              S0025-5718-2012-02600-7;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-280/S0025-5718-
                                              2012-02600-7/S0025-5718-2012-
                                              02600-7.pdf.

Nebe:2014:EEU

[NP14] Gabriele Nebe and Richard Parker. On extremal even unimodular 72-dimensional lattices. Mathematics of Computation, 83(287):1489–1494, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2014-83-
                                              287/S0025-5718-2013-02744-
                                              5;   http://www.ams.org/journals/
                                              mcom/2014-83-287/S0025-5718-
                                              2013-02744-5/S0025-5718-2013-
                                              02744-5.pdf.

Noferini:2017:CRC

[NP17] Vanni Noferini and Javier Pérez. Chebyshev rootfinding via computing eigenvalues of colleague matrices: when is it stable? Mathematics of Computation, 86(306): 1741–1767, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-306/
                                              S0025-5718-2016-03149-X;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-306/S0025-5718-
                                              2016-03149-X/S0025-5718-2016-
                                              03149-X.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=1022232;
                                               http://www.ams.org/mathscinet/

search/author.html?mrauthid=
936379.

Nguyen:2019:RSA

[NP19] Hanh My Nguyen and Carl Pomerance. The reciprocal sum of the amicable numbers. Mathematics of Computation, 88(317):1503–1526, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-317/
S0025-5718-2018-03362-2;
https://www.ams.org/journals/
mcom/2019-88-317/S0025-5718-
2018-03362-2/S0025-5718-2018-
03362-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Nguyen%
2C%20Hanh%20My;   https://
www.ams.org/mathscinet/search/
authors.html?mrauthid=140915.

Niemeyer:2012:SPF

[NPPY12] Alice C. Niemeyer, Tomasz Popiel, Cheryl E. Praeger, and Şükrü Yalçınkaya. On semiregular permutations of a finite set. Mathematics of Computation, 81(277):605– 622, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02506-8;
http://www.ams.org/journals/
mcom/2012-81-277/S0025-5718-
2011-02506-8/S0025-5718-2011-
02506-8.pdf.

Navas:2012:AEA

[NRV12] Luis M. Navas, Francisco J.

Ruiz, and Juan L. Varona.

Asymptotic estimates for Apostol– Bernoulli and Apostol–Euler polynomials. Mathematics of Computation, 81(279):1707– 1722, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2012-02568-3;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2012-02568-3/S0025-5718-2012-
02568-3.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Navas%
2C%20Luis%20M;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Ruiz%
2C%20Francisco%20J;   http:
//www.ams.org/mathscinet/
search/authors.html?authorName=
Varona%2C%20Juan%20L..

Navas:2015:SFR

[NRV15] Luis M. Navas, Francisco J. Ruiz, and Juan L. Varona. Some functional relations derived from the Lindelöf–Wirtinger expansion of the Lerch transcendent function. Mathematics of Computation, 84 (292):803–813, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-292/
S0025-5718-2014-02864-0;
http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02864-0/S0025-5718-2014-
02864-0.pdf.

Narcowich:2017:NGM

[NRW17] Francis J. Narcowich, Stephen T. Rowe, and Joseph D. Ward.

A novel Galerkin method for solving PDES on the sphere using highly localized kernel bases. Mathematics of Computation, 86(303):197– 231, January 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-303/
S0025-5718-2016-03097-5;
http://www.ams.org/journals/
mcom/2017-86-303/S0025-5718-
2016-03097-5/S0025-5718-2016-
03097-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=129435;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
180590;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=975133.

Nielsen:2013:EIN

[NS13] Johan Sejr Brinch Nielsen and Jakob Grue Simonsen. An experimental investigation of the normality of irrational algebraic numbers. Mathematics of Computation, 82(283):1837– 1858, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2013-02675-0;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-
2013-02675-0/S0025-5718-2013-
02675-0.pdf.

Nogneng:2018:ESS

[NS18] Dorian Nogneng and Éric Schost. On the evaluation of some sparse polynomials. Mathematics of Computation, 87 (310):893–904, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-310/
                                              S0025-5718-2017-03231-2;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-310/S0025-5718-
                                              2017-03231-2/S0025-5718-2017-
                                              03231-2.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1122361;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              672551.

Ngo:2016:SRC

[NT16] Hoang-Long Ngo and Dai Taguchi. Strong rate of convergence for the Euler–Maruyama approximation of stochastic differential equations with irregular coefficients. Mathematics of Computation, 85(300): 1793–1819, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-300/
                                              S0025-5718-2015-03042-7;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-300/S0025-5718-
                                              2015-03042-7/S0025-5718-2015-
                                              03042-7.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Taguchi%
                                              2C%20Dai;   http://www.ams.
                                              org/mathscinet/search/author.
                                              html?mrauthid=804677.

Nugent:2017:ADT

[NV17] Steve Nugent and John Voight. On the arithmetic dimension of triangle groups. Mathematics of Computation, 86(306): 1979–2004, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-306/
S0025-5718-2016-03147-6;
http://www.ams.org/journals/
mcom/2017-86-306/S0025-5718-
2016-03147-6/S0025-5718-2016-
03147-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Nugent%
2C%20Steve;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Voight%
2C%20John.

OliveiraeSilva:2014:EVE

[OHP14] Tomás Oliveira e Silva, Siegfried Herzog, and Silvio Pardi. Empirical verification of the even Goldbach conjecture and computation of prime gaps up to 4 · 10^{18}. Mathematics of Computation, 83(288): 2033–2060, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-288/
S0025-5718-2013-02787-1;
http://www.ams.org/journals/
mcom/2014-83-288/S0025-5718-
2013-02787-1/S0025-5718-2013-
02787-1.pdf.

Olshanskii:2012:MAT

[Ols12] Maxim A. Olshanskii. Multigrid analysis for the time dependent Stokes problem. Mathematics of Computation, 81(277): 57–79, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-277/
S0025-5718-2011-02494-4;

                                              http://www.ams.org/journals/
                                              mcom/2012-81-277/S0025-5718-
                                              2011-02494-4/S0025-5718-2011-
                                              02494-4.pdf.

Olver:2011:CHT

[Olv11]  Sheehan Olver.  Comput-

ing the Hilbert transform and its inverse. Mathematics of Computation, 80(275): 1745–1767, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              275/S0025-5718-2011-02418-
                                              X/;   http://www.ams.org/
                                              journals/mcom/2011-80-275/
                                              S0025-5718-2011-02418-X/S0025-
                                              5718-2011-02418-X.pdf.

Okayama:2015:TAS

[OMS15] Tomoaki Okayama, Takayasu Matsuo, and Masaaki Sugihara. Theoretical analysis of Sinc– Nyström methods for Volterra integral equations. Mathematics of Computation, 84(293): 1189–1215, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-293/
                                              S0025-5718-2014-02929-3;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-293/S0025-5718-
                                              2014-02929-3/S0025-5718-2014-
                                              02929-3.pdf.

Ochem:2012:OPN

[OR12] Pascal Ochem and Michaël Rao. Odd perfect numbers are greater than 10^{1500}. Mathematics of Computation, 81(279): 1869–1877, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2012-02563-4;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2012-02563-4/S0025-5718-2012-
02563-4.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Ochem%
2C%20Pascal;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Rao%
2C%20Michael.

Ochem:2014:NPF

[OR14] Pascal Ochem and Michał Rao.

On the number of prime factors of an odd perfect number. Mathematics of Computation, 83 (289):2435–2439, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-289/
S0025-5718-2013-02776-7;
http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-
2013-02776-7/S0025-5718-2013-
02776-7.pdf.

Ortner:2011:PPA

[Ort11] Christoph Ortner. A priori and a posteriori analysis of the quasinonlocal quasicontinuum method in 1D. Mathematics of Computation, 80(275): 1265–1285, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2010-02453-
6/;   http://www.ams.org/
journals/mcom/2011-80-275/

                                              S0025-5718-2010-02453-6/S0025-
                                              5718-2010-02453-6.pdf.

Ostafe:2010:DGS

[OS10] Alina Ostafe and Igor E. Shparlinski. On the degree growth in some polynomial dynamical systems and nonlinear pseudorandom number generators. Mathematics of Computation, 79(269):501– 511, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-269/
                                              S0025-5718-09-02271-6/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-269/S0025-5718-
                                              09-02271-6/S0025-5718-09-
                                              02271-6.pdf.

Ortner:2013:AEB

[OS13] C. Ortner and A. V. Shapeev. Analysis of an energy-based atomistic/continuum approximation of a vacancy in the 2D triangular lattice. Mathematics of Computation, 82 (284):2191–2236, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-284/
                                              S0025-5718-2013-02687-7;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-284/S0025-5718-
                                              2013-02687-7/S0025-5718-2013-
                                              02687-7.pdf.

Olshanskii:2016:NBU

[OS16a] Maxim A. Olshanskii and Danil Safin. A narrow-band unfitted finite element method for elliptic PDEs posed on surfaces. Mathematics of Computation, 85(300):1549– 1570, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-300/
S0025-5718-2015-03030-0;
http://www.ams.org/journals/
mcom/2016-85-300/S0025-5718-
2015-03030-0/S0025-5718-2015-
03030-0.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Safin%
2C%20Danil;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=343398.

OSullivan:2016:ZD

[O’S16b] Cormac O’Sullivan. Zeros of the dilogarithm. Mathematics of Computation, 85(302):2967– 2993, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-302/
S0025-5718-2016-03065-3;
http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-
2016-03065-3/S0025-5718-2016-
03065-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=658848.

Ozman:2019:QPM

[OS19] Ekin Ozman and Samir Siksek. Quadratic points on modular curves. Mathematics of Computation, 88(319): 2461–2484, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-319/
S0025-5718-2018-03407-X;

                                              https://www.ams.org/journals/
                                              mcom/2019-88-319/S0025-5718-
                                              2018-03407-X/S0025-5718-2018-
                                              03407-X.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Siksek%
                                              2C%20Samir;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=955558.

Oh:2018:BAD

[OWZD18] Duk-Soon Oh, Olof B. Widlund, Stefano Zampini, and Clark R. Dohrmann. BDDC Algorithms with deluxe scaling and adaptive selection of primal constraints for Raviart– Thomas vector fields. Mathematics of Computation, 87 (310):659–692, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-310/
                                              S0025-5718-2017-03254-3;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-310/S0025-5718-
                                              2017-03254-3/S0025-5718-2017-
                                              03254-3.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Zampini%
                                              2C%20Stefano;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1011821;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              182600;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=625826.

Page:2015:CAK

[Pag15] Aurel Page. Computing arithmetic Kleinian groups. Mathematics of Computation, 84 (295):2361–2390, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-295/
S0025-5718-2015-02939-1;
http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-
2015-02939-1/S0025-5718-2015-
02939-1.pdf.

Pan:2017:FAC

[Pan17]  Victor Y. Pan.  Fast ap-

proximate computations with Cauchy matrices and polynomials. Mathematics of Computation, 86(308):2799–2826, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-
308/S0025-5718-2017-03204-
X;   http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03204-X/S0025-5718-2017-
03204-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=135585.

Pappalardi:2015:DRG

[Pap15] Francesco Pappalardi. Divisibility of reduction in groups of rational numbers. Mathematics of Computation, 84 (291):385–407, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-291/
S0025-5718-2014-02872-X;
http://www.ams.org/journals/
mcom/2015-84-291/S0025-5718-
2014-02872-X/S0025-5718-2014-
02872-X.pdf.

rez-DAaz:2015:CSR

[PDSV15] S. Pérez-Díaz, J. R. Sendra, and C. Villarino. Computing the singularities of rational surfaces. Mathematics of Computation, 84(294): 1991–2021, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-294/
S0025-5718-2014-02907-4;
http://www.ams.org/journals/
mcom/2015-84-294/S0025-5718-
2014-02907-4/S0025-5718-2014-
02907-4.pdf.

Peherstorfer:2011:PTQ

[Peh11]  Franz Peherstorfer.  Posi-

tive trigonometric quadrature formulas and quadrature on the unit circle. Mathematics of Computation, 80(275): 1685–1701, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2011-02414-
2/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2011-02414-2/S0025-
5718-2011-02414-2.pdf.

Peterseim:2014:CFE

[Pet14] Daniel Peterseim. Composite finite elements for elliptic interface problems. Mathematics of Computation, 83(290): 2657–2674, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02815-9;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-

2014-02815-9/S0025-5718-2014-
02815-9.pdf.

Peterseim:2017:EPE

[Pet17] Daniel Peterseim. Eliminating the pollution effect in Helmholtz problems by local subscale correction. Mathematics of Computation, 86(305): 1005–1036, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-305/
S0025-5718-2016-03156-7;
http://www.ams.org/journals/
mcom/2017-86-305/S0025-5718-
2016-03156-7/S0025-5718-2016-
03156-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=848711.

Pan:2014:GFB

[PKLC14] Shaohua Pan, Sangho Kum, Yongdo Lim, and Jein-Shan Chen. On the generalized Fischer–Burmeister merit function for the second-order cone complementarity problem. Mathematics of Computation, 83(287):1143–1171, 2014. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2014-83-
287/S0025-5718-2013-02742-
1;   http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02742-1/S0025-5718-2013-
02742-1.pdf.

Platt:2015:CXE

[Pla15] David J. Platt. Computing π(x) analytically. Mathematics of Computation, 84(293):

1521–1535, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02884-6;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02884-6/S0025-5718-2014-
02884-6.pdf.

Platt:2016:NCC

[Pla16] David J. Platt. Numerical computations concerning the GRH. Mathematics of Computation, 85(302):3009–3027, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-302/
S0025-5718-2016-03077-X;
http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-
2016-03077-X/S0025-5718-2016-
03077-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1045993.

Platt:2017:ISN

[Pla17] David J. Platt. Isolating some non-trivial zeros of zeta. Mathematics of Computation, 86 (307):2449–2467, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-307/
S0025-5718-2017-03198-7;
http://www.ams.org/journals/
mcom/2017-86-307/S0025-5718-
2017-03198-7/S0025-5718-2017-
03198-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1045993.

Pizarro-Madariaga:2011:LBA

[PM11] Amalia Pizarro-Madariaga. Lower bounds for the Artin conductor. Mathematics of Computation, 80(273):539–561, January 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02403-
2/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02403-2/S0025-
5718-2010-02403-2.pdf.

Plato:2018:ORL

[PMH18] Robert Plato, Peter Mathé, and Bernd Hofmann. Optimal rates for Lavrentiev regularization with adjoint source conditions. Mathematics of Computation, 87(310):785–801, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03237-3;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03237-3/S0025-5718-2017-
03237-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Hofmann%
2C%20Bernd;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Mathe%
2C%20Peter;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=140260.

Papageorgiou:2013:FAA

[PPTZ13] A. Papageorgiou, I. Petras, J. F.

Traub, and C. Zhang. A fast algorithm for approximating the ground state energy on a quantum computer. Mathematics of Computation, 82 (284):2293–2304, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-284/
S0025-5718-2013-02714-7;
http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02714-7/S0025-5718-2013-
02714-7.pdf.

Pieltant:2015:NUA

[PR15] Julia Pieltant and Hugues Randriam. New uniform and asymptotic upper bounds on the tensor rank of multiplication in extensions of finite fields. Mathematics of Computation, 84 (294):2023–2045, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-294/
S0025-5718-2015-02921-4;
http://www.ams.org/journals/
mcom/2015-84-294/S0025-5718-
2015-02921-4/S0025-5718-2015-
02921-4.pdf.

Protasov:2017:EBP

[Pro17] Vladimir Yu. Protasov. The Euler binary partition function and subdivision schemes. Mathematics of Computation, 86 (305):1499–1524, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-305/
S0025-5718-2016-03128-2;
http://www.ams.org/journals/
mcom/2017-86-305/S0025-5718-
2016-03128-2/S0025-5718-2016-

03128-2.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Protasov%
2C%20Vladimir%20Yu..

Pacetti:2014:CIC

[PS14] Ariel Pacetti and Nicolás Sirolli.

Computing ideal classes representatives in quaternion algebras. Mathematics of Computation, 83(289):2479–2507, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2014-83-
289/S0025-5718-2014-02796-
8;   http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-
2014-02796-8/S0025-5718-2014-
02796-8.pdf.

Pinto:2016:GCG

[PS16] Martin Campos Pinto and Eric Sonnendrücker. Gausscompatible Galerkin schemes for time-dependent Maxwell equations. Mathematics of Computation, 85(302):2651–2685, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-302/
S0025-5718-2016-03079-3;
http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-
2016-03079-3/S0025-5718-2016-
03079-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=360263;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
758627.

Pollack:2019:DPT

[PS19] Paul Pollack and Peter Schorn.

Dirichlet’s proof of the threesquare theorem: an algorithmic perspective. Mathematics of Computation, 88(316):1007– 1019, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03349-X;
https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03349-X/S0025-5718-2018-
03349-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=249112;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
830585.

Platt:2016:FSC

[PT16] D. J. Platt and T. S. Trudgian. On the first sign change of θ(x) − x. Mathematics of Computation, 85(299): 1539–1547, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2015-03021-X;
http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2015-03021-X/S0025-5718-2015-
03021-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Trudgian%
2C%20T.%20S;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1045993.

Pitoun:2015:CTR

[PV15] Frédéric Pitoun and Firmin Varescon. Computing the torsion of the p-ramified module of a number field. Mathematics of Computation, 84 (291):371–383, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-291/
S0025-5718-2014-02838-X;
http://www.ams.org/journals/
mcom/2015-84-291/S0025-5718-
2014-02838-X/S0025-5718-2014-
02838-X.pdf.

Paredes:2017:RMH

[PVV17] Diego Paredes, Frédéric Valentin, and Henrique M. Versieux. On the robustness of multiscale hybrid-mixed methods. Mathematics of Computation, 86 (304):525–548, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03108-7;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03108-7/S0025-5718-2016-
03108-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Paredes%
2C%20Diego;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Valentin%
2C%20Frederic;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Versieux%
2C%20Henrique%20M..

Pomerance:2014:VTE

[PY14] Carl Pomerance and Hee-Sung Yang. Variant of a theorem of Erdős on the sum-of-proper-divisors function. Mathematics of Computation, 83(288): 1903–1913, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-288/
                                              S0025-5718-2013-02775-5;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-288/S0025-5718-
                                              2013-02775-5/S0025-5718-2013-
                                              02775-5.pdf.

Poor:2015:PCF

[PY15] Cris Poor and David S. Yuen. Paramodular cusp forms. Mathematics of Computation, 84 (293):1401–1438, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-293/
                                              S0025-5718-2014-02870-6;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-293/S0025-5718-
                                              2014-02870-6/S0025-5718-2014-
                                              02870-6.pdf.

Pan:2011:IF

[PZ11] Hao Pan and Wei Zhang. On the integers of the form p^{2} + b^{2} + 2^{n} and b_{1}^{2} + b_{2}^{2} + 2^{n^{2}}. Mathematics of Computation, 80(275): 1849–1864, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              275/S0025-5718-2011-02445-
                                              2/;   http://www.ams.org/
                                              journals/mcom/2011-80-275/
                                              S0025-5718-2011-02445-2/S0025-
                                              5718-2011-02445-2.pdf.

Porretta:2017:NHK

[PZ17] Alessio Porretta and Enrique Zuazua. Numerical hypocoercivity for the Kolmogorov equation. Mathematics of Computation, 86(303):97–119, January 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-303/
S0025-5718-2016-03157-9;
http://www.ams.org/journals/
mcom/2017-86-303/S0025-5718-
2016-03157-9/S0025-5718-2016-
03157-9.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=187655;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
631455.

Qiu:2016:CSS

[QS16] Tianyu Qiu and Francisco-Javier Sayas. The Costabel– Stephan system of boundary integral equations in the time domain. Mathematics of Computation, 85(301):2341–2364, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-301/
S0025-5718-2015-03053-1;
http://www.ams.org/journals/
mcom/2016-85-301/S0025-5718-
2015-03053-1/S0025-5718-2015-
03053-1.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Qiu%
2C%20Tianyu;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=621885.

Qiu:2018:HML

[QSS18] Weifeng Qiu, Jiguang Shen, and Ke Shi. An HDG method for linear elasticity with strong symmetric stresses. Mathematics of Computation, 87 (310):69–93, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-310/
                                              S0025-5718-2017-03249-X;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-310/S0025-5718-
                                              2017-03249-X/S0025-5718-2017-
                                              03249-X.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Shen%
                                              2C%20Jiguang;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=845089;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              904733.

Qureshi:2019:BML

[Qur19] Muhammad Imran Qureshi. Biregular models of log Del Pezzo surfaces with rigid singularities. Mathematics of Computation, 88(319):2497– 2521, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-319/
                                              S0025-5718-2019-03432-4;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-319/S0025-5718-
                                              2019-03432-4/S0025-5718-2019-
                                              03432-4.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=948483.

Wang:2017:VLJ

[qWlGjY17] Zhong qing Wang, Yu ling Guo, and Li jun Yi. An hp-version Legendre–Jacobi spectral collocation method for Volterra integro-differential equations with smooth and weakly singular kernels. Mathematics of Computation, 86(307): 2285–2324, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-307/
S0025-5718-2017-03183-5;
http://www.ams.org/journals/
mcom/2017-86-307/S0025-5718-
2017-03183-5/S0025-5718-2017-
03183-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Guo%
2C%20Yu-ling;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=678011;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
883098.

Qiao:2015:SCS

[QzSZ15] Zhonghua Qiao, Zhi zhong Sun, and Zhengru Zhang. Stability and convergence of secondorder schemes for the nonlinear epitaxial growth model without slope selection. Mathematics of Computation, 84 (292):653–674, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-292/
S0025-5718-2014-02874-3;
http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02874-3/S0025-5718-2014-
02874-3.pdf.

Rahman:2014:AET

[Rah14] Sharif Rahman. Approximation errors in truncated dimensional decompositions. Mathematics of Computation, 83 (290):2799–2819, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-290/
                                              S0025-5718-2014-02883-4;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-290/S0025-5718-
                                              2014-02883-4/S0025-5718-2014-
                                              02883-4.pdf.

Ramare:2015:EES

[Ram15] Olivier Ramaré. Explicit estimates on several summatory functions involving the Moebius function. Mathematics of Computation, 84(293): 1359–1387, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-293/
                                              S0025-5718-2014-02914-1;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-293/S0025-5718-
                                              2014-02914-1/S0025-5718-2014-

02914-1.pdf. See corrigendum [Ram19].

Ramare:2016:EDE

[Ram16] O. Ramaré. An explicit density estimate for Dirichlet L-series. Mathematics of Computation, 85(297):325–356, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-297/
                                              S0025-5718-2015-02991-3;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-297/S0025-5718-
                                              2015-02991-3/S0025-5718-2015-
                                              02991-3.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=360330.

Ramare:2019:CEE

[Ram19] Olivier Ramaré. Corrigendum to Explicit estimates on several summatory functions involving the Moebius function. Mathematics of Computation, 88(319): 2383–2388, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-319/
S0025-5718-2019-03449-X;
https://www.ams.org/journals/
mcom/2019-88-319/S0025-5718-
2019-03449-X/S0025-5718-2019-
03449-X.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Ramare%
2C%20Olivier. See [Ram15].

Raum:2016:CGJ

[Rau16]  Martin Raum.  Computing

genus 1 Jacobi forms. Mathematics of Computation, 85 (298):931–960, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-02992-5;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-02992-5/S0025-5718-2015-
02992-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Raum%
2C%20Martin.

Rand:2014:QSF

[RGB14] Alexander Rand, Andrew Gillette, and Chandrajit Bajaj. Quadratic serendipity finite elements on polygons using generalized barycentric coordinates. Mathematics of Computation, 83 (290):2691–2716, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02807-X;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02807-X/S0025-5718-2014-
02807-X.pdf.

Rebenich:2018:APN

[RGNS18] Niko Rebenich, T. Aaron Gulliver, Stephen Neville, and Ulrich Speidel. An analog of the prime number theorem for finite fields via truncated polylogarithm expansions. Mathematics of Computation, 87 (310):855–877, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03247-6;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03247-6/S0025-5718-2017-
03247-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Neville%
2C%20Stephen;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1079851;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
294190;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=821380.

Rito:2010:EDP

[Rit10] Carlos Rito. On equations of double planes with p_{g} = q = 1. Mathematics of Computation, 79(270):1091– 1108, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02283-2/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02283-2/S0025-5718-09-
02283-2.pdf.

Rivin:2016:HPR

[Riv16] Igor Rivin. How to pick a random integer matrix? (and other questions). Mathematics of Computation, 85(298): 783–797, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-02986-X;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-02986-X/S0025-5718-2015-
02986-X.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Rivin%
2C%20Igor.

Rozenhart:2012:TCF

[RJS12] Pieter Rozenhart, Michael Jacobson Jr., and Renate Scheidler. Tabulation of cubic function fields via polynomial binary cubic forms. Mathematics of Computation, 81(280):2335– 2359, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-280/
S0025-5718-2012-02591-9;
http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-

                                              2012-02591-9/S0025-5718-2012-
                                              02591-9.pdf.

Roblot:2015:CAF

[Rob15] Xavier-François Roblot. Computing p-adic L-functions of totally real number fields. Mathematics of Computation, 84 (292):831–874, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-292/
                                              S0025-5718-2014-02889-5;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-292/S0025-5718-
                                              2014-02889-5/S0025-5718-2014-
                                              02889-5.pdf.

Robles-Perez:2012:FPN

[RPR12] Aureliano M. Robles-Pérez and José Carlos Rosales. The Frobenius problem for numerical semigroups with embedding dimension equal to three. Mathematics of Computation, 81(279): 1609–1617, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-279/
                                              S0025-5718-2011-02561-5;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-279/S0025-5718-
                                              2011-02561-5/S0025-5718-2011-
                                              02561-5.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Robles-
                                              Perez%2C%20Aureliano%20M;
                                               http://www.ams.org/mathscinet/
                                              search/authors.html?authorName=
                                              Rosales%2C%20Jose%20Carlos.

Romero:2013:HGS

[RR13] Ana Romero and Julio Rubio. Homotopy groups of suspended classifying spaces: an experimental approach. Mathematics of Computation, 82 (284):2237–2244, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-284/
S0025-5718-2013-02680-4;
http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02680-4/S0025-5718-2013-
02680-4.pdf.

Rao:2016:ALA

[RR16] Michaël Rao and Matthieu Rosenfeld. Avoidability of long k-abelian repetitions. Mathematics of Computation, 85(302): 3051–3060, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2016-85-
302/S0025-5718-2016-03085-
9;   http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-
2016-03085-9/S0025-5718-2016-
03085-9.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Rosenfeld%
2C%20Matthieu;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=714149.

Rubin:2010:CCE

[RS10] K. Rubin and A. Silverberg. Choosing the correct elliptic curve in the CM method. Mathematics of Computation, 79(269):545– 561, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/

                                              journals/mcom/2010-79-269/
                                              S0025-5718-09-02266-2/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-269/S0025-5718-
                                              09-02266-2/S0025-5718-09-
                                              02266-2.pdf.

Rubinstein-Salzedo:2014:PCC

[RS14] Simon Rubinstein-Salzedo. Period computations for covers of elliptic curves. Mathematics of Computation, 83(289): 2455–2470, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-289/
                                              S0025-5718-2014-02797-X;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-289/S0025-5718-
                                              2014-02797-X/S0025-5718-2014-
                                              02797-X.pdf.

Reddy:2016:CNA

[RS16] G. Murali Mohan Reddy and Rajen K. Sinha. On the Crank–Nicolson anisotropic a posteriori error analysis for parabolic integro-differential equations. Mathematics of Computation, 85(301):2365– 2390, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2016-85-301/
                                              S0025-5718-2015-03067-1;
                                              http://www.ams.org/journals/
                                              mcom/2016-85-301/S0025-5718-
                                              2015-03067-1/S0025-5718-2015-
                                              03067-1.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Reddy%
                                              2C%20G.%20Murali%20Mohan;
                                               http://www.ams.org/mathscinet/

search/author.html?authorName=
Sinha%2C%20Rajen%20K..

Rauhut:2017:CSP

[RS17a] Holger Rauhut and Christoph Schwab. Compressive sensing Petrov–Galerkin approximation of high-dimensional parametric operator equations. Mathematics of Computation, 86 (304):661–700, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03113-0;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03113-0/S0025-5718-2016-
03113-0.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=305221;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
720385.

Rutka:2017:EBF

[RS17b] Przemysław Rutka and Ryszard Smarzewski. Explicit barycentric formulae for osculatory interpolation at roots of classical orthogonal polynomials. Mathematics of Computation, 86 (307):2409–2427, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-307/
S0025-5718-2017-03184-7;
http://www.ams.org/journals/
mcom/2017-86-307/S0025-5718-
2017-03184-7/S0025-5718-2017-
03184-7.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=163855;
 http://www.ams.org/mathscinet/

                                              search/author.html?mrauthid=
                                              890344.

Rutka:2019:DIB

[RS19] Przemysław Rutka and Ryszard Smarzewski. Difference inequalities and barycentric identities for classical discrete iterated weights. Mathematics of Computation, 88(318):1791– 1804, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-318/
                                              S0025-5718-2018-03396-8;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-318/S0025-5718-
                                              2018-03396-8/S0025-5718-2018-
                                              03396-8.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=163855;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              890344.

Ryan:2012:NCC

[RSS12] Nathan C. Ryan, Nils-Peter Skoruppa, and Fredrik Strömberg. Numerical computation of a certain Dirichlet series attached to Siegel modular forms of degree two. Mathematics of Computation, 81(280):2361– 2376, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-280/
                                              S0025-5718-2012-02584-1;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-280/S0025-5718-
                                              2012-02584-1/S0025-5718-2012-
                                              02584-1.pdf.

Recio:2010:GCA

[RSTV10] T. Recio, J. R. Sendra, L. F.

Tabera, and C. Villarino. Generalizing circles over algebraic extensions. Mathematics of Computation, 79(270):1067– 1089, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02284-4/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02284-4/S0025-5718-09-
02284-4.pdf.

Ryan:2016:FCV

[RTwaabRS16] Nathan C. Ryan, Gonzalo Tornaría, and with an appendix by Ralf Schmidt. Formulas for central values of twisted spin L-functions attached to paramodular forms. Mathematics of Computation, 85 (298):907–929, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-02988-3;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-02988-3/S0025-5718-2015-
02988-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Ryan%
2C%20Nathan%20C;   http://
www.ams.org/mathscinet/search/
author.html?authorName=Schmidt%
2C%20Ralf;   http://www.ams.
org/mathscinet/search/author.
html?authorName=Tornaria%
2C%20Gonzalo.

Rodriguez:2014:NAS

[RV14] Rodolfo Rodríguez and Pablo Venegas. Numerical approximation of the spectrum of the curl operator. Mathematics of Computation, 83 (286):553–577, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-286/
                                              S0025-5718-2013-02745-7;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-286/S0025-5718-
                                              2013-02745-7/S0025-5718-2013-
                                              02745-7.pdf.

Rebholz:2019:ENI

[RVX19] Leo G. Rebholz, Alex Viguerie, and Mengying Xiao. Efficient nonlinear iteration schemes based on algebraic splitting for the incompressible Navier– Stokes equations. Mathematics of Computation, 88(318): 1533–1557, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-318/
                                              S0025-5718-2019-03411-7;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-318/S0025-5718-
                                              2019-03411-7/S0025-5718-2019-
                                              03411-7.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=1120573;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              1255895;   https://www.ams.
                                              org/mathscinet/search/authors.
                                              html?mrauthid=787731.

Ramming:2018:KBD

[RW18] Tobias Ramming and Holger Wendland. A kernelbased discretisation method for first order partial differential equations. Mathematics of Computation, 87(312):1757– 1781, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2017-03265-8;
http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03265-8/S0025-5718-2017-
03265-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Ramming%
2C%20Tobias;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=602098.

Rubinstein:2015:CMP

[RY15] Michael O. Rubinstein and Shuntaro Yamagishi. Computing the moment polynomials of the zeta function. Mathematics of Computation, 84 (291):425–454, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-291/
S0025-5718-2014-02845-7;
http://www.ams.org/journals/
mcom/2015-84-291/S0025-5718-
2014-02845-7/S0025-5718-2014-
02845-7.pdf.

Sadek:2014:ECW

[Sad14] Mohammad Sadek. On elliptic curves whose conductor is a product of two prime powers. Mathematics of Computation, 83(285):447–460, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-285/
                                              S0025-5718-2013-02726-3;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-285/S0025-5718-
                                              2013-02726-3/S0025-5718-2013-
                                              02726-3.pdf.

Sanchez:2019:ECD

[Sán19] Omar León Sánchez. Estimates for the coefficients of differential dimension polynomials. Mathematics of Computation, 88(320):2959– 2985, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-320/
                                              S0025-5718-2019-03429-4;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-320/S0025-5718-
                                              2019-03429-4/S0025-5718-2019-
                                              03429-4.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Sanchez%
                                              2C%20Omar%20Leon.

Schirokauer:2010:NFS

[Sch10] Oliver Schirokauer. The number field sieve for integers of low weight. Mathematics of Computation, 79(269):583– 602, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-269/
                                              S0025-5718-09-02198-X/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-269/S0025-5718-
                                              09-02198-X/S0025-5718-09-
                                              02198-X.pdf.

Scheerer:2017:CAN

[Sch17] Adrian-Maria Scheerer. Computable absolutely normal numbers and discrepancies. Mathematics of Computation, 86 (308):2911–2926, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-308/
S0025-5718-2017-03189-6;
http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03189-6/S0025-5718-2017-
03189-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Scheerer%
2C%20Adrian-Maria.

Saouter:2010:SRW

[SD10] Yannick Saouter and Patrick Demichel. A sharp region where π(x) − li(x) is positive. Mathematics of Computation, 79(272):2395–2405, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02351-3/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02351-3/S0025-5718-10-
02351-3.pdf.

Stoll:2011:ICZ

[SD11] Douglas A. Stoll and Patrick Demichel. The impact of ζ(s) complex zeros on π(x) for x < 10^{10^{13}}. Mathematics of Computation, 80(276): 2381–2394, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02477-
4/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02477-4/S0025-
5718-2011-02477-4.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813366.

Sdika:2019:DBS

[Sdi19]  Michaël Sdika.  Diffeomor-

phic B-spline vector fields with a tractable set of inequalities. Mathematics of Computation, 88(320):2827– 2856, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-320/
S0025-5718-2019-03419-1;
https://www.ams.org/journals/
mcom/2019-88-320/S0025-5718-
2019-03419-1/S0025-5718-2019-
03419-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Sdika%
2C%20Michael.

Segura:2017:SNM

[Seg17] Javier Segura. The Schwarzian– Newton method for solving nonlinear equations, with applications. Mathematics of Computation, 86(304):865–879, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-
304/S0025-5718-2016-03119-
1;   http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03119-1/S0025-5718-2016-
03119-1.pdf;   http://www.

ams.org/mathscinet/search/
author.html?mrauthid=627158.

Seri:2017:SPA

[Ser17] Raffaello Seri. Statistical properties of b-adic diaphonies. Mathematics of Computation, 86(304):799–828, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03148-8;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03148-8/S0025-5718-2016-
03148-8.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=710036.

Sertoz:2019:CPH

[Ser19]  Emre Can Sertöz.  Com-

puting periods of hypersurfaces. Mathematics of Computation, 88(320):2987–3022, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-320/
S0025-5718-2019-03430-0;
https://www.ams.org/journals/
mcom/2019-88-320/S0025-5718-
2019-03430-0/S0025-5718-2019-
03430-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Sertoz%
2C%20Emre%20Can.

Sfakianakis:2013:AMR

[Sfa13] Nikolaos Sfakianakis. Adaptive mesh reconstruction for hyperbolic conservation laws with total variation bound. Mathematics of Computation, 82(281):

129–151, January 2013. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2013-82-
281/S0025-5718-2012-02615-
9;   http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-
2012-02615-9/S0025-5718-2012-
02615-9.pdf.

Saouter:2011:ILB

[SGD11] Yannick Saouter, Xavier Gourdon, and Patrick Demichel. An improved lower bound for the de Bruijn–Newman constant. Mathematics of Computation, 80(276):2281– 2287, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02472-
5/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02472-5/S0025-
5718-2011-02472-5.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813360.

Sugiura:2011:PIP

[SH11] Hiroshi Sugiura and Takemitsu Hasegawa. A polynomial interpolation process at quasi-Chebyshev nodes with the FFT. Mathematics of Computation, 80(276):2169–2184, October 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02484-
1/;   http://www.ams.org/

journals/mcom/2011-80-276/
S0025-5718-2011-02484-1/S0025-
5718-2011-02484-1.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813353.

Shardlow:2019:WOS

[Sha19]  Tony Shardlow.  A Walk

Outside Spheres for the fractional Laplacian: Fields and first eigenvalue. Mathematics of Computation, 88(320):2767– 2792, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-320/
S0025-5718-2019-03422-1;
https://www.ams.org/journals/
mcom/2019-88-320/S0025-5718-
2019-03422-1/S0025-5718-2019-
03422-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=356141.

Shiu:1986:CST

[Shi86] Peter Shiu. Counting sums of two squares: the Meissel– Lehmer method. Mathematics of Computation, 47(175):351– 360, July 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [Shi19].

Shibuta:2013:ICA

[Shi13]  Takafumi Shibuta.  Irre-

ducibility criterion for algebroid curves. Mathematics of Computation, 82(281):531– 554, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-281/
S0025-5718-2012-02607-X;

                                              http://www.ams.org/journals/
                                              mcom/2013-82-281/S0025-5718-
                                              2012-02607-X/S0025-5718-2012-
                                              02607-X.pdf.

Shiu:2019:CCS

[Shi19]  Peter Shiu.  Corrigendum

to “Counting sums of two squares: the Meissel–Lehmer method”. Mathematics of Computation, 88(320):2935– 2938, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-320/
                                              S0025-5718-2019-03477-4;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-320/S0025-5718-
                                              2019-03477-4/S0025-5718-2019-
                                              03477-4.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=160965.

Shparlinski:2011:ADP

[Shp11] Igor E. Shparlinski. On the average distribution of pseudorandom numbers generated by nonlinear permutations. Mathematics of Computation, 80(274): 1053–1061, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              274/S0025-5718-2010-02408-
                                              1/;   http://www.ams.org/
                                              journals/mcom/2011-80-274/
                                              S0025-5718-2010-02408-1/S0025-
                                              5718-2010-02408-1.pdf.

Sidi:2010:AAG

[Sid10] Avram Sidi. Asymptotic analysis of a generalized Richardson extrapolation process on linear sequences. Mathematics of Computation, 79(271): 1681–1695, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-09-02318-7/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
09-02318-7/S0025-5718-09-
02318-7.pdf.

Sidi:2011:AEL

[Sid11] Avram Sidi. Asymptotic expansions of Legendre series coefficients for functions with interior and endpoint singularities. Mathematics of Computation, 80(275):1663–1684, July 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2010-02454-
8/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2010-02454-8/S0025-
5718-2010-02454-8.pdf.

Sidi:2012:EME

[Sid12] Avram Sidi. Euler–Maclaurin expansions for integrals with arbitrary algebraic endpoint singularities. Mathematics of Computation, 81(280):2159– 2173, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-280/
S0025-5718-2012-02597-X;
http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-

                                              2012-02597-X/S0025-5718-2012-
                                              02597-X.pdf.

Sijsling:2012:ACB

[Sij12] Jeroen Sijsling. Arithmetic (1; e)-curves and Belĭ maps. Mathematics of Computation, 81(279):1823–1855, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2012-81-
                                              279/S0025-5718-2012-02560-
                                              9;   http://www.ams.org/journals/
                                              mcom/2012-81-279/S0025-5718-
                                              2012-02560-9/S0025-5718-2012-
                                              02560-9.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Sijsling%
                                              2C%20Jeroen.

Shan:2017:TSE

[SL17] Weikun Shan and Huiyuan Li. The triangular spectral element method for Stokes eigenvalues. Mathematics of Computation, 86(308):2579–2611, 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2017-86-
                                              308/S0025-5718-2017-03173-
                                              2;   http://www.ams.org/journals/
                                              mcom/2017-86-308/S0025-5718-
                                              2017-03173-2/S0025-5718-2017-
                                              03173-2.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Shan%
                                              2C%20Weikun;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=708582.

Soga:2015:SVA

[Sog15] Kohei Soga. Stochastic and variational approach to the Lax–Friedrichs scheme. Mathematics of Computation, 84 (292):629–651, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-292/
S0025-5718-2014-02863-9;
http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02863-9/S0025-5718-2014-
02863-9.pdf.

Soga:2016:MSV

[Sog16] Kohei Soga. More on stochastic and variational approach to the Lax–Friedrichs scheme. Mathematics of Computation, 85(301): 2161–2193, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2016-85-
301/S0025-5718-2016-03061-
6;   http://www.ams.org/journals/
mcom/2016-85-301/S0025-5718-
2016-03061-6/S0025-5718-2016-
03061-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=909684.

Sorevik:2016:GLD

[Sør16] Tor Sørevik. Good low degree rank-1 lattice rules of high dimension. Mathematics of Computation, 85(300): 1821–1835, July 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-300/
S0025-5718-2016-03047-1;
http://www.ams.org/journals/
mcom/2016-85-300/S0025-5718-
2016-03047-1/S0025-5718-2016-
03047-1.pdf;   http://www.

                                              ams.org/mathscinet/search/
                                              author.html?authorName=Sorevik%
                                              2C%20Tor.

Spalevic:2013:EBG

[Spa13] Miodrag M. Spalević. Error bounds of Gaussian quadrature formulae for one class of Bernstein–Szegő weights. Mathematics of Computation, 82 (282):1037–1056, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2013-82-
                                              282/S0025-5718-2012-02667-
                                              6;   http://www.ams.org/journals/
                                              mcom/2013-82-282/S0025-5718-
                                              2012-02667-6/S0025-5718-2012-
                                              02667-6.pdf.

Spalevic:2017:GAG

[Spa17]  Miodrag M. Spalević.  On

generalized averaged Gaussian formulas. II. Mathematics of Computation, 86(306): 1877–1885, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-306/
                                              S0025-5718-2016-03225-1;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-306/S0025-5718-
                                              2016-03225-1/S0025-5718-2016-
                                              03225-1.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=600543.

Spijker:2017:SBN

[Spi17]  M. N. Spijker.  Stability

and boundedness in the numerical solution of initial value problems. Mathematics of Computation, 86(308): 2777–2798, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-308/
S0025-5718-2017-03191-4;
http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03191-4/S0025-5718-2017-
03191-4.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=165560.

Sawyer:2011:PPE

[SR11] Jorge F. Sawyer and Clifford A. Reiter. Perfect parallelepipeds exist. Mathematics of Computation, 80(274): 1037–1040, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02400-
7/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02400-7/S0025-
5718-2010-02400-7.pdf.

Schicho:2014:TWC

[SS14] Josef Schicho and David Sevilla.

Tschirnhaus–Weierstrass curves. Mathematics of Computation, 83(290):3005–3015, 2014. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2014-83-
290/S0025-5718-2014-02801-
9;   http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02801-9/S0025-5718-2014-
02801-9.pdf.

Sauer:2015:LAS

[SS15] Martin Sauer and Wilhelm Stannat. Lattice approximation for stochastic reaction diffusion equations with one-sided Lipschitz condition. Mathematics of Computation, 84 (292):743–766, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-292/
S0025-5718-2014-02873-1;
http://www.ams.org/journals/
mcom/2015-84-292/S0025-5718-
2014-02873-1/S0025-5718-2014-
02873-1.pdf.

Sauer:2016:AAS

[SS16] Martin Sauer and Wilhelm Stannat. Analysis and approximation of stochastic nerve axon equations. Mathematics of Computation, 85(301):2457– 2481, September 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-301/
S0025-5718-2016-03068-9;
http://www.ams.org/journals/
mcom/2016-85-301/S0025-5718-
2016-03068-9/S0025-5718-2016-
03068-9.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Stannat%
2C%20Wilhelm;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=1016033.

Schenck:2014:SST

[SSV14] Hal Schenck, Alexandra Seceleanu, and Javid Validashti. Syzygies and singularities of tensor product surfaces of bidegree (2, 1). Mathematics of Computation, 83(287): 1337–1372, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-287/
S0025-5718-2013-02764-0;
http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02764-0/S0025-5718-2013-
02764-0.pdf.

Sendra:2017:CRR

[SSV17] J. Rafael Sendra, David Sevilla, and Carlos Villarino. Covering rational ruled surfaces. Mathematics of Computation, 86 (308):2861–2875, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-308/
S0025-5718-2017-03193-8;
http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03193-8/S0025-5718-2017-
03193-8.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=260673;
 http://www.ams.org/mathscinet/
search/author.html?mrauthid=
683262;   http://www.ams.
org/mathscinet/search/author.
html?mrauthid=700228.

Schotzau:2016:DSO

[SSW16] Dominik Schötzau, Christoph Schwab, and Thomas P. Wihler. hp-dGFEM for secondorder mixed elliptic problems in polyhedra. Mathematics of Computation, 85(299): 1051–1083, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-299/
S0025-5718-2015-03062-2;
http://www.ams.org/journals/

                                              mcom/2016-85-299/S0025-5718-
                                              2015-03062-2/S0025-5718-2015-
                                              03062-2.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Schotzau%
                                              2C%20Dominik;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Wihler%
                                              2C%20Thomas%20P;   http://
                                              www.ams.org/mathscinet/search/
                                              author.html?mrauthid=305221.

Sands:2018:CAC

[ST18a] Jonathan W. Sands and Brett A. Tangedal. Computing annihilators of class groups from derivatives of L-functions. Mathematics of Computation, 87(314): 2937–2953, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-314/
                                              S0025-5718-2018-03297-5;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-314/S0025-5718-
                                              2018-03297-5/S0025-5718-2018-
                                              03297-5.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=154195;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              612497.

Stuart:2018:PCG

[ST18b] Andrew M. Stuart and Aretha L. Teckentrup. Posterior consistency for Gaussian process approximations of Bayesian posterior distributions. Mathematics of Computation, 87 (310):721–753, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-310/

S0025-5718-2017-03244-0;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03244-0/S0025-5718-2017-
03244-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Stuart%
2C%20Andrew%20M.;   https:
//www.ams.org/mathscinet/
search/authors.html?authorName=
Teckentrup%2C%20Aretha%20L.

.

Saouter:2015:SSR

[STD15] Yannick Saouter, Timothy Trudgian, and Patrick Demichel. A still sharper region where π(x) − li(x) is positive. Mathematics of Computation, 84 (295):2433–2446, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-295/
S0025-5718-2015-02930-5;
http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-
2015-02930-5/S0025-5718-2015-
02930-5.pdf.

Stevenson:2011:DFW

[Ste11] Rob Stevenson. Divergence-free wavelet bases on the hypercube: Free-slip boundary conditions, and applications for solving the instationary Stokes equations. Mathematics of Computation, 80(275):1499–1523, July 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2011-02471-
3/;   http://www.ams.org/
journals/mcom/2011-80-275/

                                              S0025-5718-2011-02471-3/S0025-
                                              5718-2011-02471-3.pdf.

Strauss:2013:SOS

[Str13]  Michael Strauss.  The sec-

ond order spectrum and optimal convergence. Mathematics of Computation, 82(284): 2305–2325, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-284/
                                              S0025-5718-2013-02693-2;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-284/S0025-5718-
                                              2013-02693-2/S0025-5718-2013-
                                              02693-2.pdf.

Saouter:2014:IRM

[StR14a] Yannick Saouter and Herman te Riele. Improved results on the Mertens conjecture. Mathematics of Computation, 83 (285):421–433, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-285/
                                              S0025-5718-2013-02716-0;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-285/S0025-5718-
                                              2013-02716-0/S0025-5718-2013-
                                              02716-0.pdf.

Streng:2014:CIC

[Str14b]  Marco Streng.  Computing

Igusa class polynomials. Mathematics of Computation, 83 (285):275–309, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-285/
                                              S0025-5718-2013-02712-3;

http://www.ams.org/journals/
mcom/2014-83-285/S0025-5718-
2013-02712-3/S0025-5718-2013-
02712-3.pdf.

Stynes:2014:SAI

[Sty14] Martin Stynes. Sharp anisotropic interpolation error estimates for rectangular Raviart–Thomas elements. Mathematics of Computation, 83(290):2675–2689, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/
2014-83-290/S0025-5718-2014-
02826-3;   http://www.ams.
org/journals/mcom/2014-83-
290/S0025-5718-2014-02826-
3/S0025-5718-2014-02826-3.
pdf.

Sutherland:2011:CHC

[Sut11a] Andrew V. Sutherland. Computing Hilbert class polynomials with the Chinese remainder theorem. Mathematics of Computation, 80(273): 501–538, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02373-
7/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02373-7/S0025-
5718-2010-02373-7.pdf.

Sutherland:2011:SCD

[Sut11b] Andrew V. Sutherland. Structure computation and discrete logarithms in finite abelian p-groups. Mathematics of Computation, 80(273):477–500, January 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-10-02356-2/;
http://www.ams.org/journals/
mcom/2011-80-273/S0025-5718-
10-02356-2/S0025-5718-10-
02356-2.pdf.

Sutherland:2012:CEC

[Sut12] Andrew V. Sutherland. Constructing elliptic curves over finite fields with prescribed torsion. Mathematics of Computation, 81(278):1131– 1147, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02538-X;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02538-X/S0025-5718-2011-
02538-X.pdf.

Simis:2012:CCM

[SV12] Aron Simis and Rafael H. Villarreal. Combinatorics of Cremona monomial maps. Mathematics of Computation, 81(279): 1857–1867, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2011-02556-1;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2011-02556-1/S0025-5718-2011-
02556-1.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Simis%

2C%20Aron;   http://www.ams.
org/mathscinet/search/authors.
html?authorName=Villarreal%
2C%20Rafael%20H..

Sadiq:2014:FDW

[SV14] Burhan Sadiq and Divakar Viswanath. Finite difference weights, spectral differentiation, and superconvergence. Mathematics of Computation, 83 (289):2403–2427, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-289/
S0025-5718-2014-02798-1;
http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-
2014-02798-1/S0025-5718-2014-
02798-1.pdf.

Sokolowsky:2014:IFP

[SVVR14] Benjamin D. Sokolowsky, Amy G.

VanHooft, Rachel M. Volkert, and Clifford A. Reiter. An infinite family of perfect parallelepipeds. Mathematics of Computation, 83(289): 2441–2454, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-289/
S0025-5718-2013-02791-3;
http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-
2013-02791-3/S0025-5718-2013-
02791-3.pdf.

Stamm:2010:ODG

[SW10] Benjamin Stamm and Thomas P.

Wihler. hp-optimal discontinuous Galerkin methods for linear elliptic problems. Mathematics of Computation, 79(272):2117– 2133, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02335-5/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02335-5/S0025-5718-10-
02335-5.pdf.

Schrader:2011:HOA

[SW11] Daniela Schräder and Holger Wendland. A highorder, analytically divergencefree discretization method for Darcy’s problem. Mathematics of Computation, 80(273): 263–277, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
273/S0025-5718-2010-02388-
9/;   http://www.ams.org/
journals/mcom/2011-80-273/
S0025-5718-2010-02388-9/S0025-
5718-2010-02388-9.pdf.

Stein:2013:AAE

[SW13] William Stein and Christian Wuthrich. Algorithms for the arithmetic of elliptic curves using Iwasawa theory. Mathematics of Computation, 82(283): 1757–1792, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2012-02649-4;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-

2012-02649-4/S0025-5718-2012-
02649-4.pdf.

Sorenson:2017:SPT

[SW17] Jonathan Sorenson and Jonathan Webster. Strong pseudoprimes to twelve prime bases. Mathematics of Computation, 86 (304):985–1003, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-304/
S0025-5718-2016-03134-8;
http://www.ams.org/journals/
mcom/2017-86-304/S0025-5718-
2016-03134-8/S0025-5718-2016-
03134-8.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=334195;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
903429.

Sze:2011:TSR

[SwaAbLCW11] Tsz-Wo Sze and with an Appendix by Lawrence C. Washington. On taking square roots without quadratic nonresidues over finite fields. Mathematics of Computation, 80(275): 1797–1811, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2011-02419-
1/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2011-02419-1/S0025-
5718-2011-02419-1.pdf.

Shen:2019:FSJ

[SWX19] Jie Shen, Yingwei Wang, and Jianlin Xia. Fast structured Jacobi–Jacobi transforms.

Mathematics of Computation, 88(318):1743–1772, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

www.ams.org/journals/mcom/
2019-88-318/S0025-5718-2018-
03377-4;   https://www.ams.
org/journals/mcom/2019-88-
318/S0025-5718-2018-03377-
4/S0025-5718-2018-03377-4.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=257933;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
638529;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=858648.

Sawa:2014:PCR

[SX14] Masanori Sawa and Yuan Xu. On positive cubature rules on the simplex and isometric embeddings. Mathematics of Computation, 83(287): 1251–1277, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-287/
S0025-5718-2013-02762-7;
http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02762-7/S0025-5718-2013-
02762-7.pdf.

Szyld:2016:PEL

[SX16] Daniel B. Szyld and Fei Xue. Preconditioned eigensolvers for large-scale nonlinear Hermitian eigenproblems with variational characterizations. I. Extreme eigenvalues. Mathematics of Computation, 85(302):2887– 2918, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-302/
S0025-5718-2016-03083-5;
http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-
2016-03083-5/S0025-5718-2016-
03083-5.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=244424;
http://www.ams.org/mathscinet/
search/author.html?mrauthid=
880232.

Shen:2016:RDB

[SY16] Yun-Qiu Shen and Tjalling J.

Ypma. Rank deficiencies and bifurcation into affine subspaces for separable parameterized equations. Mathematics of Computation, 85 (297):271–293, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-297/
S0025-5718-2015-02968-8;
http://www.ams.org/journals/
mcom/2016-85-297/S0025-5718-
2015-02968-8/S0025-5718-2015-
02968-8.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Ypma%
2C%20Tjalling%20J;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
191125.

Szpruch:2018:VVV

[SZ18] Lukasz Szpruch and Xilíng Zhang. V -integrability, asymptotic stability and comparison property of explicit numerical schemes for non-linear SDEs. Mathematics of Computation, 87(310):755–783, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03219-1;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03219-1/S0025-5718-2017-
03219-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Zhang%
2C%20Xiling;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=895132.

Szabo:2011:VCR

[Sza11] Sándor Szabó. Verifying a conjecture of L. Rédei for p = 13. Mathematics of Computation, 80(274):1155–1162, April 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02417-
2/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02417-2/S0025-
5718-2010-02417-2.pdf.

Szmytkowski:2013:EBT

[Szm13] Radosław Szmytkowski. Erratum to Formulas and Theorems for the Special Functions of Mathematical Physics by W. Magnus, F. Oberhettinger, R. P. Soni. Mathematics of Computation, 82(283):1709– 1710, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2013-02671-3;
http://www.ams.org/journals/

mcom/2013-82-283/S0025-5718-
2013-02671-3/S0025-5718-2013-
02671-3.pdf.

Srivastava:2011:ADZ

[SZW11] H. M. Srivastava, Jian-Rong Zhou, and Zhi-Gang Wang. Asymptotic distributions of the zeros of certain classes of hypergeometric functions and polynomials. Mathematics of Computation, 80(275): 1769–1784, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
275/S0025-5718-2011-02409-
9/;   http://www.ams.org/
journals/mcom/2011-80-275/
S0025-5718-2011-02409-9/S0025-
5718-2011-02409-9.pdf.

Tabera:2017:PVH

[Tab17]  Luis Felipe Tabera.  A

parametric version of the Hilbert–Hurwitz theorem using hypercircles. Mathematics of Computation, 86(308): 3001–3018, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-308/
S0025-5718-2017-03202-6;
http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03202-6/S0025-5718-2017-
03202-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Tabera%
2C%20Luis%20Felipe.

Tao:2014:EON

[Tao14] Terence Tao. Every odd number greater than 1 is the sum of at most five primes. Mathematics of Computation, 83 (286):997–1038, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-286/
S0025-5718-2013-02733-0;
http://www.ams.org/journals/
mcom/2014-83-286/S0025-5718-
2013-02733-0/S0025-5718-2013-
02733-0.pdf.

Tao:2012:DMF

[TCH12] Terence Tao, Ernest Croot III, and Harald Helfgott. Deterministic methods to find primes. Mathematics of Computation, 81(278):1233– 1246, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02542-1;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02542-1/S0025-5718-2011-
02542-1.pdf.

Theobald:2015:AAC

[TdW15] Thorsten Theobald and Timo de Wolff. Approximating amoebas and coamoebas by sums of squares. Mathematics of Computation, 84 (291):455–473, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-291/
S0025-5718-2014-02828-7;
http://www.ams.org/journals/
mcom/2015-84-291/S0025-5718-

2014-02828-7/S0025-5718-2014-
02828-7.pdf.

Teng:2010:EBB

[Ten10] Zhen-Huan Teng. Error bound between monotone difference schemes and their modified equations. Mathematics of Computation, 79(271):1473– 1491, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-271/
S0025-5718-09-02306-0/;
http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
09-02306-0/S0025-5718-09-
02306-0.pdf.

Thongjunthug:2010:CLB

[Tho10] Thotsaphon Thongjunthug. Computing a lower bound for the canonical height on elliptic curves over number fields. Mathematics of Computation, 79(272):2431–2449, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02352-5/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02352-5/S0025-5718-10-
02352-5.pdf.

Trappmann:2012:CTR

[TK12] Henryk Trappmann and Dmitrii Kouznetsov. Computation of the two regular superexponentials to base exp(1/e). Mathematics of Computation, 81(280):2207–2227, October 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/
2012-81-280/S0025-5718-2012-
02590-7;   http://www.ams.
org/journals/mcom/2012-81-
280/S0025-5718-2012-02590-
7/S0025-5718-2012-02590-7.
pdf.

Trevino:2012:LIP

[Tre12] Enrique Treviño. The least inert prime in a real quadratic field. Mathematics of Computation, 81(279):1777–1797, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2012-81-
279/S0025-5718-2012-02579-
8;   http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2012-02579-8/S0025-5718-2012-
02579-8.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Trevino%
2C%20Enrique.

Trudgian:2011:ITM

[Tru11] Timothy Trudgian. Improvements to Turing’s method. Mathematics of Computation, 80(276):2259–2279, October 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journals/mcom/
2011-80-276/S0025-5718-2011-
02470-1/;   http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02470-
1/S0025-5718-2011-02470-1.
pdf;   http://www.ams.org/
mathscinet-getitem?mr=2813359.

Trudgian:2012:IUB

[Tru12] Timothy Trudgian. An improved upper bound for the argument of the Riemann zeta-function on the critical line. Mathematics of Computation, 81(278):1053– 1061, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02537-8;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02537-8/S0025-5718-2011-
02537-8.pdf.

Trudgian:2015:IUB

[Tru15] T. S. Trudgian. An improved upper bound for the error in the zero-counting formulae for Dirichlet L-functions and Dedekind zeta-functions. Mathematics of Computation, 84 (293):1439–1450, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02898-6;
http://www.ams.org/journals/
mcom/2015-84-293/S0025-5718-
2014-02898-6/S0025-5718-2014-
02898-6.pdf.

Tabata:2018:ECL

[TU18] Masahisa Tabata and Shinya Uchiumi. An exactly computable Lagrange–Galerkin scheme for the Navier–Stokes equations and its error estimates. Mathematics of Computation, 87(310):39–67, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03222-1;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03222-1/S0025-5718-2017-
03222-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1139936;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
240674.

Tuitman:2016:CPC

[Tui16] Jan Tuitman. Counting points on curves using a map to P^{1}. Mathematics of Computation, 85(298):961–981, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-02996-2;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-02996-2/S0025-5718-2015-
02996-2.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Tuitman%
2C%20Jan.

Tuitman:2019:CZF

[Tui19]  Jan Tuitman.  Computing

zeta functions of generic projective hypersurfaces in larger characteristic. Mathematics of Computation, 88(315):439– 451, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03325-7;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-

2018-03325-7/S0025-5718-2018-
03325-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=941045.

Tao:2010:SAC

[TV10] Terence Tao and Van Vu.

Smooth analysis of the condition number and the least singular value. Mathematics of Computation, 79(272):2333– 2352, October 2010. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2010-79-
272/S0025-5718-2010-02396-
8/;   http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-2010-02396-8/S0025-
5718-2010-02396-8.pdf.

Townsend:2018:FPT

[TWO18] Alex Townsend, Marcus Webb, and Sheehan Olver. Fast polynomial transforms based on Toeplitz and Hankel matrices. Mathematics of Computation, 87(312):1913–1934, October 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-312/
S0025-5718-2017-03277-4;
http://www.ams.org/journals/
mcom/2018-87-312/S0025-5718-
2017-03277-4/S0025-5718-2017-
03277-4.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Townsend%
2C%20Alex;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1004196;
 https://www.ams.org/mathscinet/

                                              search/authors.html?mrauthid=
                                              783322.

Tao:2017:AUM

[TY17] Min Tao and Xiaoming Yuan. Accelerated Uzawa methods for convex optimization. Mathematics of Computation, 86 (306):1821–1845, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-306/
                                              S0025-5718-2016-03145-2;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-306/S0025-5718-
                                              2016-03145-2/S0025-5718-2016-
                                              03145-2.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?mrauthid=729439;
                                               http://www.ams.org/mathscinet/
                                              search/author.html?mrauthid=
                                              987506.

Tian:2019:ADM

[TY19] Wenyi Tian and Xiaoming Yuan. An alternating direction method of multipliers with a worst-case O(1/n^{2}) convergence rate. Mathematics of Computation, 88(318):1685– 1713, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-318/
                                              S0025-5718-2018-03388-9;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-318/S0025-5718-
                                              2018-03388-9/S0025-5718-2018-
                                              03388-9.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=729439;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              999800.

Tian:2015:CSO

[TZD15] WenYi Tian, Han Zhou, and Weihua Deng. A class of second order difference approximations for solving space fractional diffusion equations. Mathematics of Computation, 84(294): 1703–1727, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-294/
S0025-5718-2015-02917-2;
http://www.ams.org/journals/
mcom/2015-84-294/S0025-5718-
2015-02917-2/S0025-5718-2015-
02917-2.pdf.

Ulas:2014:SDS

[Ula14] Maciej Ulas. On some Diophantine systems involving symmetric polynomials. Mathematics of Computation, 83(288): 1915–1930, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-288/
S0025-5718-2013-02778-0;
http://www.ams.org/journals/
mcom/2014-83-288/S0025-5718-
2013-02778-0/S0025-5718-2013-
02778-0.pdf.

Urroz:2012:NIC

[ULS12] Jorge Jiménez Urroz, Florian Luca, and Igor E. Shparlinski. On the number of isogeny classes of pairing-friendly elliptic curves and statistics of MNT curves. Mathematics of Computation, 81(278):1093– 1110, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2012-81-278/
                                              S0025-5718-2011-02543-3;
                                              http://www.ams.org/journals/
                                              mcom/2012-81-278/S0025-5718-
                                              2011-02543-3/S0025-5718-2011-
                                              02543-3.pdf.

Urban:2014:IEB

[UP14] Karsten Urban and Anthony T. Patera. An improved error bound for reduced basis approximation of linear parabolic problems. Mathematics of Computation, 83(288): 1599–1615, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-288/
                                              S0025-5718-2013-02782-2;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-288/S0025-5718-
                                              2013-02782-2/S0025-5718-2013-
                                              02782-2.pdf.

Vabishchevich:2012:NCA

[Vab12] Petr N. Vabishchevich. On a new class of additive (splitting) operator-difference schemes. Mathematics of Computation, 81(277):267–276, January 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2012-81-
                                              277/S0025-5718-2011-02492-
                                              0;   http://www.ams.org/journals/
                                              mcom/2012-81-277/S0025-5718-
                                              2011-02492-0/S0025-5718-2011-
                                              02492-0.pdf.

Vabishchevich:2014:ASS

[Vab14] Petr N. Vabishchevich. Additive schemes (splitting schemes) for some systems of evolutionary equations. Mathematics of Computation, 83(290): 2787–2797, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02799-3;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02799-3/S0025-5718-2014-
02799-3.pdf.

vandenBerg:2010:GSS

[vdBLM10] Jan Bouwe van den Berg, Jean-Philippe Lessard, and Konstantin Mischaikow. Global smooth solution curves using rigorous branch following. Mathematics of Computation, 79(271):1565–1584, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2010-79-
271/S0025-5718-10-02325-2/
;   http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-
10-02325-2/S0025-5718-10-
02325-2.pdf.

vanDiejen:2019:ECR

[vDE19] J. F. van Diejen and E. Emsiz.

Exact cubature rules for symmetric functions. Mathematics of Computation, 88(317):1229– 1249, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-317/
S0025-5718-2018-03380-4;
https://www.ams.org/journals/
mcom/2019-88-317/S0025-5718-

                                              2018-03380-4/S0025-5718-2018-
                                              03380-4.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=306808;
                                               https://www.ams.org/mathscinet/
                                              search/authors.html?mrauthid=
                                              781405.

Verdier:2014:ROP

[Ver14] Olivier Verdier. Reductions of operator pencils. Mathematics of Computation, 83 (285):189–214, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-285/
                                              S0025-5718-2013-02740-8;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-285/S0025-5718-
                                              2013-02740-8/S0025-5718-2013-
                                              02740-8.pdf.

Velichka:2014:CDL

[VJS14] M. D. Velichka, M. J. Jacobson Jr., and A. Stein. Computing discrete logarithms in the Jacobian of high-genus hyperelliptic curves over even characteristic finite fields. Mathematics of Computation, 83 (286):935–963, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-286/
                                              S0025-5718-2013-02748-2;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-286/S0025-5718-
                                              2013-02748-2/S0025-5718-2013-
                                              02748-2.pdf.

Vohralik:2010:UPF

[Voh10] Martin Vohralík. Unified primal formulation-based a priori and a posteriori error analysis of mixed finite element methods. Mathematics of Computation, 79(272): 2001–2032, October 2010. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2010-79-
272/S0025-5718-2010-02375-
0/;   http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-2010-02375-0/S0025-
5718-2010-02375-0.pdf.

Vojtechovsky:2019:ERQ

[VY19] Petr Vojtěchovský and Seung Yeop Yang. Enumeration of racks and quandles up to isomorphism. Mathematics of Computation, 88(319): 2523–2540, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-319/
S0025-5718-2019-03409-9;
https://www.ams.org/journals/
mcom/2019-88-319/S0025-5718-
2019-03409-9/S0025-5718-2019-
03409-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1111587;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
650320.

vonzurGathen:2019:IE

[vzG19] Joachim von zur Gathen. Iteration entropy. Mathematics of Computation, 88(318):1991– 2003, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-318/
S0025-5718-2018-03382-8;

                                              https://www.ams.org/journals/
                                              mcom/2019-88-318/S0025-5718-
                                              2018-03382-8/S0025-5718-2018-
                                              03382-8.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=71800.

Wang:2010:ASA

[Wan10a] Li-Lian Wang. Analysis of spectral approximations using prolate spheroidal wave functions. Mathematics of Computation, 79(270):807–827, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2010-79-
                                              270/S0025-5718-09-02268-6/
                                              ;   http://www.ams.org/journals/
                                              mcom/2010-79-270/S0025-5718-
                                              09-02268-6/S0025-5718-09-
                                              02268-6.pdf.

Wang:2010:ASS

[Wan10b]  Xiaoming Wang.  Approxi-

mation of stationary statistical properties of dissipative dynamical systems: Time discretization. Mathematics of Computation, 79(269):259– 280, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2010-79-269/
                                              S0025-5718-09-02256-X/;
                                              http://www.ams.org/journals/
                                              mcom/2010-79-269/S0025-5718-
                                              09-02256-X/S0025-5718-09-
                                              02256-X.pdf.

Wang:2019:NCR

[Wan19]  Yanqiu Wang.  A non-

conforming Crouzeix–Raviart type finite element on polygonal meshes. Mathematics of Computation, 88(315):237– 271, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03334-8;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03334-8/S0025-5718-2018-
03334-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=670715.

Wasilkowski:2014:ACT

[Was14] G. W. Wasilkowski. Average case tractability of approximating ∞-variate functions. Mathematics of Computation, 83 (287):1319–1336, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-287/
S0025-5718-2013-02759-7;
http://www.ams.org/journals/
mcom/2014-83-287/S0025-5718-
2013-02759-7/S0025-5718-2013-
02759-7.pdf.

Weingartner:2019:CFS

[Wei19]  Andreas Weingartner.  On

the constant factor in several related asymptotic estimates. Mathematics of Computation, 88(318):1883– 1902, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-318/
S0025-5718-2018-03402-0;
https://www.ams.org/journals/
mcom/2019-88-318/S0025-5718-
2018-03402-0/S0025-5718-2018-
03402-0.pdf;   https://www.

                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=678374.

Wells:2017:CCH

[Wel17] Elliot Wells. Computing canonical heights on the projective line with no factorization. Mathematics of Computation, 86 (308):3019–3029, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2017-86-308/
                                              S0025-5718-2017-03200-2;
                                              http://www.ams.org/journals/
                                              mcom/2017-86-308/S0025-5718-
                                              2017-03200-2/S0025-5718-2017-
                                              03200-2.pdf;   http://www.
                                              ams.org/mathscinet/search/
                                              author.html?authorName=Wells%
                                              2C%20Elliot.

Wendland:2013:HOA

[Wen13] Holger Wendland. A highorder approximation method for semilinear parabolic equations on spheres. Mathematics of Computation, 82(281):227– 245, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2013-82-281/
                                              S0025-5718-2012-02623-8;
                                              http://www.ams.org/journals/
                                              mcom/2013-82-281/S0025-5718-
                                              2012-02623-8/S0025-5718-2012-
                                              02623-8.pdf.

Wang:2019:CFV

[WHL19] Xiang Wang, Weizhang Huang, and Yonghai Li. Conditioning of the finite volume element method for diffusion problems with general simplicial meshes. Mathematics of Computation, 88(320):2665– 2696, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-320/
S0025-5718-2019-03423-3;
https://www.ams.org/journals/
mcom/2019-88-320/S0025-5718-
2019-03423-3/S0025-5718-2019-
03423-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=326320;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
363086;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=936154.

Wang:2014:EBW

[WHV14] Haiyong Wang, Daan Huybrechs, and Stefan Vandewalle. Explicit barycentric weights for polynomial interpolation in the roots or extrema of classical orthogonal polynomials. Mathematics of Computation, 83 (290):2893–2914, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-290/
S0025-5718-2014-02821-4;
http://www.ams.org/journals/
mcom/2014-83-290/S0025-5718-
2014-02821-4/S0025-5718-2014-
02821-4.pdf.

Williams:2018:ESH

[Wil18]  D. M. Williams.  An en-

tropy stable, hybridizable discontinuous Galerkin method for the compressible Navier– Stokes equations. Mathematics of Computation, 87 (310):95–121, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03199-9;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03199-9/S0025-5718-2017-
03199-9.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1044919.

Watanabe:2013:PEI

[WKN13] Yoshitaka Watanabe, Takehiko Kinoshita, and Mitsuhiro T. Nakao. A posteriori estimates of inverse operators for boundary value problems in linear elliptic partial differential equations. Mathematics of Computation, 82(283):1543– 1557, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2013-02676-2;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-
2013-02676-2/S0025-5718-2013-
02676-2.pdf.

Wang:2017:MCG

[WLZ17] Tongke Wang, Zhifang Liu, and Zhiyue Zhang. The modified composite Gauss type rules for singular integrals using Puiseux expansions. Mathematics of Computation, 86(303):345– 373, January 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-303/
S0025-5718-2016-03105-1;
http://www.ams.org/journals/

mcom/2017-86-303/S0025-5718-
2016-03105-1/S0025-5718-2016-
03105-1.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Liu%
2C%20Zhifang;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Wang%
2C%20Tongke;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=645950.

Wang:2019:LDG

[WLZS19] Haijin Wang, Yunxian Liu, Qiang Zhang, and Chi-Wang Shu. Local discontinuous Galerkin methods with implicitexplicit time-marching for timedependent incompressible fluid flow. Mathematics of Computation, 88(315):91–121, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

www.ams.org/journals/mcom/
2019-88-315/S0025-5718-2018-
03312-9;   https://www.ams.
org/journals/mcom/2019-88-
315/S0025-5718-2018-03312-
9/S0025-5718-2018-03312-9.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=1022956;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
242268;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=637183;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
648640.

Wang:2017:PMS

[WMxY17] Xiao Wang, Shiqian Ma, and Ya xiang Yuan. Penalty methods with stochastic approximation for stochastic nonlinear programming. Mathematics of Computation, 86(306): 1793–1820, 2017. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2017-86-306/
S0025-5718-2016-03178-6;
http://www.ams.org/journals/
mcom/2017-86-306/S0025-5718-
2016-03178-6/S0025-5718-2016-
03178-6.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Wang%
2C%20Xiao;   http://www.ams.
org/mathscinet/search/author.
html?authorName=Yuan%2C%20Ya-
xiang;   http://www.ams.org/
mathscinet/search/author.
html?mrauthid=826033.

Wang:2018:CFD

[WNK18] Siyang Wang, Anna Nissen, and Gunilla Kreiss. Convergence of finite difference methods for the wave equation in two space dimensions. Mathematics of Computation, 87(314): 2737–2763, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-314/
S0025-5718-2018-03319-1;
http://www.ams.org/journals/
mcom/2018-87-314/S0025-5718-
2018-03319-1/S0025-5718-2018-
03319-1.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1062535;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
238145;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=908946.

Wang:2010:NEB

[WO10] Heping Wang and Sofiya Ostrovska. The norm estimates for the q-Bernstein operator in the case q > 1. Mathematics of Computation, 79(269):353– 363, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-269/
S0025-5718-09-02273-X/;
http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02273-X/S0025-5718-09-
02273-X.pdf.

Wu:2010:SPN

[Wu10a] Qiang Wu. The smallest Perron numbers. Mathematics of Computation, 79(272):2387– 2394, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02345-8/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02345-8/S0025-5718-10-
02345-8.pdf.

Wu:2010:MSS

[Wu10b] Yingquan Wu. More on solving systems of power equations. Mathematics of Computation, 79(272):2317–2332, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02363-X/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-

                                              10-02363-X/S0025-5718-10-
                                              02363-X.pdf.

Wuthrich:2018:NMS

[Wut18] Christian Wuthrich. Numerical modular symbols for elliptic curves. Mathematics of Computation, 87(313):2393– 2423, January 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-313/
                                              S0025-5718-2017-03274-9;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-313/S0025-5718-
                                              2017-03274-9/S0025-5718-2017-
                                              03274-9.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=681572.

Wasilkowski:2010:ED

[WW10] Grzegorz W. Wasilkowski and Henryk Wozniakowski. On the exponent of discrepancies. Mathematics of Computation, 79(270):983–992, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2010-79-
                                              270/S0025-5718-09-02314-X/
                                              ;   http://www.ams.org/journals/
                                              mcom/2010-79-270/S0025-5718-
                                              09-02314-X/S0025-5718-09-
                                              02314-X.pdf.

Wang:2018:PDW

[WW18] Chunmei Wang and Junping Wang. A primal-dual weak Galerkin finite element method for second order elliptic equations in non-divergence form. Mathematics of Computation, 87(310):515–545, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03220-8;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03220-8/S0025-5718-2017-
03220-8.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1018207;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
216677.

Wang:2012:CRL

[WX12] Haiyong Wang and Shuhuang Xiang. On the convergence rates of Legendre approximation. Mathematics of Computation, 81(278):861– 877, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02549-4;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02549-4/S0025-5718-2011-
02549-4.pdf.

Wang:2013:MFE

[WX13] Ming Wang and Jinchao Xu.

Minimal finite element spaces for 2m-th-order partial differential equations in R^{n}. Mathematics of Computation, 82(281): 25–43, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-281/
S0025-5718-2012-02611-1;
http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-

2012-02611-1/S0025-5718-2012-
02611-1.pdf.

Wang:2019:TVB

[WX19a] Sulin Wang and Zhengfu Xu. Total variation bounded flux limiters for high order finite difference schemes solving one-dimensional scalar conservation laws. Mathematics of Computation, 88(316):691– 716, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03364-6;
https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03364-6/S0025-5718-2018-
03364-6.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Wang%
2C%20Sulin;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=705141.

Wang:2019:CTL

[WX19b] Wei Wang and Xuejun Xu. On the convergence of a twolevel preconditioned Jacobi– Davidson method for eigenvalue problems. Mathematics of Computation, 88(319): 2295–2324, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-319/
S0025-5718-2018-03403-2;
https://www.ams.org/journals/
mcom/2019-88-319/S0025-5718-
2018-03403-2/S0025-5718-2018-
03403-2.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?authorName=Wang%

2C%20Wei;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=365400.

Wu:2019:NFE

[WX19c] Shuonan Wu and Jinchao Xu. Nonconforming finite element spaces for 2m th order partial differential equations on ℝ^{n} simplicial grids when m = n + 1. Mathematics of Computation, 88(316):531– 551, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-316/
S0025-5718-2018-03361-0;
https://www.ams.org/journals/
mcom/2019-88-316/S0025-5718-
2018-03361-0/S0025-5718-2018-
03361-0.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1021082;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
228866.

Wang:2014:WGM

[WY14] Junping Wang and Xiu Ye. A weak Galerkin mixed finite element method for second order elliptic problems. Mathematics of Computation, 83 (289):2101–2126, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-289/
S0025-5718-2014-02852-4;
http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-
2014-02852-4/S0025-5718-2014-
02852-4.pdf.

Wang:2019:WGF

[WZZZ19] Junping Wang, Qilong Zhai, Ran Zhang, and Shangyou Zhang. A weak Galerkin finite element scheme for the Cahn– Hilliard equation. Mathematics of Computation, 88(315): 211–235, July 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/2019-88-315/
S0025-5718-2018-03369-5;
https://www.ams.org/journals/
mcom/2019-88-315/S0025-5718-
2018-03369-5/S0025-5718-2018-
03369-5.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=1114983;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
216677;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=261174;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
719427.

Xia:2018:CDS

[Xia18] Binzhou Xia. Cyclotomic difference sets in finite fields. Mathematics of Computation, 87(313): 2461–2482, January 2018. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2018-87-
313/S0025-5718-2018-03311-
7;   http://www.ams.org/journals/
mcom/2018-87-313/S0025-5718-
2018-03311-7/S0025-5718-2018-
03311-7.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=905727.

Xu:2014:UCC

[Xu14a] Da Xu. Uniform l^{1} convergence in the Crank–Nicolson method of a linear integrodifferential equation for viscoelastic rods and plates. Mathematics of Computation, 83 (286):735–769, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-286/
S0025-5718-2013-02756-1;
http://www.ams.org/journals/
mcom/2014-83-286/S0025-5718-
2013-02756-1/S0025-5718-2013-
02756-1.pdf.

Xu:2014:PMP

[Xu14b]  Zhengfu Xu.  Parametrized

maximum principle preserving flux limiters for high order schemes solving hyperbolic conservation laws: one-dimensional scalar problem. Mathematics of Computation, 83(289): 2213–2238, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-289/
S0025-5718-2013-02788-3;
http://www.ams.org/journals/
mcom/2014-83-289/S0025-5718-
2013-02788-3/S0025-5718-2013-
02788-3.pdf.

Xiang:2010:FIH

[XW10] Shuhuang Xiang and Haiyong Wang. Fast integration of highly oscillatory integrals with exotic oscillators. Mathematics of Computation, 79(270): 829–844, April 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-270/
S0025-5718-09-02279-0/;
http://www.ams.org/journals/
mcom/2010-79-270/S0025-5718-
09-02279-0/S0025-5718-09-
02279-0.pdf.

Xie:2013:ECG

[XWZ13] Ziqing Xie, Li-Lian Wang, and Xiaodan Zhao. On exponential convergence of Gegenbauer interpolation and spectral differentiation. Mathematics of Computation, 82(282): 1017–1036, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02645-7;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02645-7/S0025-5718-2012-
02645-7.pdf.

Xue:2013:CEE

[XY13] Jungong Xue and Qiang Ye. Computing exponentials of essentially non-negative matrices entrywise to high relative accuracy. Mathematics of Computation, 82(283):1577– 1596, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2013-02677-4;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-
2013-02677-4/S0025-5718-2013-
02677-4.pdf.

Xie:2010:USA

[XZ10] Ziqing Xie and Zhimin Zhang.

Uniform superconvergence analysis of the discontinuous Galerkin method for a singularly perturbed problem in 1-D. Mathematics of Computation, 79(269): 35–45, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-269/
S0025-5718-09-02297-2/;
http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02297-2/S0025-5718-09-
02297-2.pdf.

Xu:2015:CAF

[XZ15] Yifeng Xu and Jun Zou. Convergence of an adaptive finite element method for distributed flux reconstruction. Mathematics of Computation, 84(296): 2645–2663, ???? 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-296/
S0025-5718-2015-02961-5;
http://www.ams.org/journals/
mcom/2015-84-296/S0025-5718-
2015-02961-5/S0025-5718-2015-
02961-5.pdf.

Yang:2017:NIP

[Yan17] Danping Yang. Non-iterative parallel Schwarz algorithms based on overlapping domain decomposition for parabolic partial differential equations. Mathematics of Computation, 86(308):2687–2718, 2017. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2017-86-
308/S0025-5718-2017-03102-
1;   http://www.ams.org/journals/
mcom/2017-86-308/S0025-5718-
2017-03102-1/S0025-5718-2017-
03102-1.pdf;   http://www.
ams.org/mathscinet/search/
author.html?mrauthid=238349.

Yao:2013:MMF

[Yao13]  Xudong Yao.  A mini-

max method for finding saddle critical points of upper semi-differentiable locally Lipschitz continuous functional in Hilbert space and its convergence. Mathematics of Computation, 82(284): 2087–2136, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-284/
S0025-5718-2013-02669-5;
http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02669-5/S0025-5718-2013-
02669-5.pdf.

Yanfen:2012:ESC

[YDk12] Cui Yanfen and Mao De-kang. Error self-canceling of a difference scheme maintaining two conservation laws for linear advection equation. Mathematics of Computation, 81(278): 715–741, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02523-8;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-

2011-02523-8/S0025-5718-2011-
02523-8.pdf.

Ye:2018:AIC

[Ye18] Qiang Ye. Accurate inverses for computing eigenvalues of extremely ill-conditioned matrices and differential operators. Mathematics of Computation, 87(310):237–259, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-310/
S0025-5718-2017-03223-3;
http://www.ams.org/journals/
mcom/2018-87-310/S0025-5718-
2017-03223-3/S0025-5718-2017-
03223-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=237891.

Yengui:2011:SFM

[Yen11]  Ihsen Yengui.  Stably free

modules over R[X] of rank > dim R are free. Mathematics of Computation, 80(274): 1093–1098, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
274/S0025-5718-2010-02427-
5/;   http://www.ams.org/
journals/mcom/2011-80-274/
S0025-5718-2010-02427-5/S0025-
5718-2010-02427-5.pdf.

Yarman:2015:GPA

[YF15] Can Evren Yarman and Garret M. Flagg. Generalization of Padé approximation from rational functions to arbitrary analytic functions — theory. Mathematics of Computation, 84 (294):1835–1860, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-294/
S0025-5718-2015-02928-7;
http://www.ams.org/journals/
mcom/2015-84-294/S0025-5718-
2015-02928-7/S0025-5718-2015-
02928-7.pdf.

Yang:2013:DSB

[YMO13] Yi Yang, Michael Möller, and Stanley Osher. A dual split Bregman method for fast ℓ^{1} minimization. Mathematics of Computation, 82(284): 2061–2085, 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-284/
S0025-5718-2013-02700-7;
http://www.ams.org/journals/
mcom/2013-82-284/S0025-5718-
2013-02700-7/S0025-5718-2013-
02700-7.pdf.

Yadav:2013:SDG

[YPP13] Sangita Yadav, Amiya K. Pani, and Eun-Jae Park. Superconvergent discontinuous Galerkin methods for nonlinear elliptic equations. Mathematics of Computation, 82(283):1297– 1335, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-283/
S0025-5718-2013-02662-2;
http://www.ams.org/journals/
mcom/2013-82-283/S0025-5718-
2013-02662-2/S0025-5718-2013-
02662-2.pdf.

Yang:2013:LAL

[YY13] Junfeng Yang and Xiaoming Yuan. Linearized augmented Lagrangian and alternating direction methods for nuclear norm minimization. Mathematics of Computation, 82(281): 301–329, January 2013. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2013-82-
281/S0025-5718-2012-02598-
1;   http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-
2012-02598-1/S0025-5718-2012-
02598-1.pdf.

Zhang:2013:TPI

[ZC13] Yong Zhang and Tianxin Cai.

n-tuples of positive integers with the same sum and the same product. Mathematics of Computation, 82(281):617– 623, January 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-281/
S0025-5718-2012-02609-3;
http://www.ams.org/journals/
mcom/2013-82-281/S0025-5718-
2012-02609-3/S0025-5718-2012-
02609-3.pdf.

Zhou:2018:SDA

[ZC18] Yang Zhou and Xiaojun Chen.

Spherical t_{ϵ}-designs for approximations on the sphere. Mathematics of Computation, 87(314): 2831–2855, April 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2018-87-314/

S0025-5718-2018-03306-3;
http://www.ams.org/journals/
mcom/2018-87-314/S0025-5718-
2018-03306-3/S0025-5718-2018-
03306-3.pdf;   https://www.
ams.org/mathscinet/search/
authors.html?mrauthid=196364;
 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
272850.

Zhong:2012:COA

[ZCS^{+}12] Liuqiang Zhong, Long Chen, Shi Shu, Gabriel Wittum, and Jinchao Xu. Convergence and optimality of adaptive edge finite element methods for time-harmonic Maxwell equations. Mathematics of Computation, 81(278): 623–642, April 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-278/
S0025-5718-2011-02544-5;
http://www.ams.org/journals/
mcom/2012-81-278/S0025-5718-
2011-02544-5/S0025-5718-2011-
02544-5.pdf.

Zhu:2014:MDS

[ZD14] Liyong Zhu and Qiang Du. Mesh dependent stability and condition number estimates for finite element approximations of parabolic problems. Mathematics of Computation, 83 (285):37–64, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-285/
S0025-5718-2013-02703-2;
http://www.ams.org/journals/
mcom/2014-83-285/S0025-5718-

2013-02703-2/S0025-5718-2013-
02703-2.pdf.

Zhang:2013:MWF

[ZDL13] Yong Zhang, Bin Dong, and Zhaosong Lu. L_{0} Minimization for wavelet frame based image restoration. Mathematics of Computation, 82(282):995– 1015, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2013-82-282/
S0025-5718-2012-02631-7;
http://www.ams.org/journals/
mcom/2013-82-282/S0025-5718-
2012-02631-7/S0025-5718-2012-
02631-7.pdf.

Zhang:2014:EBT

[ZGFD14] Peng Zhang, Yuxiang Gao, Janet Fierson, and Yuefan Deng. Eigenanalysis-based task mapping on parallel computers with cellular networks. Mathematics of Computation, 83 (288):1727–1756, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-288/
S0025-5718-2013-02770-6;
http://www.ams.org/journals/
mcom/2014-83-288/S0025-5718-
2013-02770-6/S0025-5718-2013-
02770-6.pdf.

Zhang:2011:DFF

[Zha11a] Shangyou Zhang. Divergencefree finite elements on tetrahedral grids for k ≥ 6. Mathematics of Computation, 80 (274):669–695, April 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              274/S0025-5718-2010-02412-
                                              3/;   http://www.ams.org/
                                              journals/mcom/2011-80-274/
                                              S0025-5718-2010-02412-3/S0025-
                                              5718-2010-02412-3.pdf.

Zhang:2011:CCN

[Zha11b]  Zhenxiang Zhang.  Count-

ing Carmichael numbers with small seeds. Mathematics of Computation, 80(273):437– 442, January 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2011-80-
                                              273/S0025-5718-2010-02382-
                                              8/;   http://www.ams.org/
                                              journals/mcom/2011-80-273/
                                              S0025-5718-2010-02382-8/S0025-
                                              5718-2010-02382-8.pdf.

Zhang:2015:ECC

[Zha15] Zhenxiang Zhang. Estimating the counts of Carmichael and Williams numbers with small multiple seeds. Mathematics of Computation, 84 (291):309–337, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2015-84-291/
                                              S0025-5718-2014-02837-8;
                                              http://www.ams.org/journals/
                                              mcom/2015-84-291/S0025-5718-
                                              2014-02837-8/S0025-5718-2014-
                                              02837-8.pdf.

Zheng:2011:NFE

[ZHX11] Bin Zheng, Qiya Hu, and Jinchao Xu. A nonconforming finite element method for fourth order curl equations in ℝ^{3}. Mathematics of Computation, 80(276):1871– 1886, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2011-80-
276/S0025-5718-2011-02480-
4/;   http://www.ams.org/
journals/mcom/2011-80-276/
S0025-5718-2011-02480-4/S0025-
5718-2011-02480-4.pdf;   http:
//www.ams.org/mathscinet-
getitem?mr=2813342.

Ziegler:2010:FTE

[Zie10a] Volker Ziegler. On a family of Thue equations of degree 16. Mathematics of Computation, 79(272):2407– 2429, October 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-272/
S0025-5718-10-02354-9/;
http://www.ams.org/journals/
mcom/2010-79-272/S0025-5718-
10-02354-9/S0025-5718-10-
02354-9.pdf.

Zieve:2010:TGT

[Zie10b] Michael E. Zieve. p^{k}-torsion of genus two curves over 𝔽_{p^{m}}. Mathematics of Computation, 79(271):1833–1838, July 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2010-79-
271/S0025-5718-10-02305-7/
;   http://www.ams.org/journals/
mcom/2010-79-271/S0025-5718-

10-02305-7/S0025-5718-10-
02305-7.pdf.

Zhou:2019:CRE

[ZN19] Ming Zhou and Klaus Neymeyr. Cluster robust estimates for block gradient-type eigensolvers. Mathematics of Computation, 88(320):2737–2765, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

www.ams.org/journals/mcom/
2019-88-320/S0025-5718-2019-
03446-4;   https://www.ams.
org/journals/mcom/2019-88-
320/S0025-5718-2019-03446-
4/S0025-5718-2019-03446-4.
pdf;   https://www.ams.org/
mathscinet/search/authors.
html?mrauthid=672470;   https:
//www.ams.org/mathscinet/
search/authors.html?mrauthid=
943384.

Zhong-qing:2016:SCM

[ZqCt16] Wang Zhong-qing and Sheng Chang-tao. An hp-spectral collocation method for nonlinear Volterra integral equations with vanishing variable delays. Mathematics of Computation, 85 (298):635–666, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2016-85-298/
S0025-5718-2015-03023-3;
http://www.ams.org/journals/
mcom/2016-85-298/S0025-5718-
2015-03023-3/S0025-5718-2015-
03023-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Chang-
tao%2C%20Sheng;   http://
www.ams.org/mathscinet/search/

author.html?authorName=Zhong-
qing%2C%20Wang.

Zhang:2012:CSO

[ZQY12] Xinzhen Zhang, Liqun Qi, and Yinyu Ye. The cubic spherical optimization problems. Mathematics of Computation, 81(279): 1513–1525, July 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2012-81-279/
S0025-5718-2012-02577-4;
http://www.ams.org/journals/
mcom/2012-81-279/S0025-5718-
2012-02577-4/S0025-5718-2012-
02577-4.pdf;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Zhang%
2C%20Xinzhen;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Qi%
2C%20Liqun;   http://www.
ams.org/mathscinet/search/
authors.html?authorName=Ye%
2C%20Yinyu.

Zralek:2010:DVP

[Zra10a] Bartosz Zrałek. A deterministic version of Pollard’s p − 1 algorithm. Mathematics of Computation, 79(269):513– 533, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2010-79-269/
S0025-5718-09-02262-5/;
http://www.ams.org/journals/
mcom/2010-79-269/S0025-5718-
09-02262-5/S0025-5718-09-
02262-5.pdf.

Zralek:2010:UPS

[Zra10b] Bartosz Zrałek. Using partial smoothness of p−1 for factoring polynomials modulo p. Mathematics of Computation, 79(272): 2353–2359, October 2010. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.
                                              org/journals/mcom/2010-79-
                                              272/S0025-5718-2010-02377-
                                              4/;   http://www.ams.org/
                                              journals/mcom/2010-79-272/
                                              S0025-5718-2010-02377-4/S0025-
                                              5718-2010-02377-4.pdf.

Zralek:2019:ERA

[Źra19] Bartosz Źrałek. An extension of a result about divisors in a residue class and its application to reducing integer factorization to computing Euler’s totient. Mathematics of Computation, 88(317):1261– 1272, January 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL https://www.ams.org/
                                              journals/mcom/2019-88-317/
                                              S0025-5718-2018-03368-3;
                                              https://www.ams.org/journals/
                                              mcom/2019-88-317/S0025-5718-
                                              2018-03368-3/S0025-5718-2018-
                                              03368-3.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Zralek%
                                              2C%20Bartosz.

Zhang:2016:TKT

[ZX16] Long Zhang and Kejian Xu. The tame kernel of ℚ(ζ_{5}) is trivial. Mathematics of Computation, 85(299):1523–1538, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/journals/mcom/2016-85-
299/S0025-5718-2015-03003-
8;   http://www.ams.org/journals/
mcom/2016-85-299/S0025-5718-
2015-03003-8/S0025-5718-2015-
03003-8.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Xu%
2C%20Kejian;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Zhang%
2C%20Long.

Zhang:2014:EAS

[ZY14] Lei-Hong Zhang and Wei Hong Yang. An efficient algorithm for second-order cone linear complementarity problems. Mathematics of Computation, 83 (288):1701–1726, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2014-83-288/
S0025-5718-2013-02795-0;
http://www.ams.org/journals/
mcom/2014-83-288/S0025-5718-
2013-02795-0/S0025-5718-2013-
02795-0.pdf.

Zeng:2015:CCM

[ZY15] Jinxiang Zeng and Linsheng Yin. On the computation of coefficients of modular forms: the reduction modulo p approach. Mathematics of Computation, 84(293): 1469–1488, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-293/
S0025-5718-2014-02892-5;
http://www.ams.org/journals/

                                              mcom/2015-84-293/S0025-5718-
                                              2014-02892-5/S0025-5718-2014-
                                              02892-5.pdf.

Zhu:2014:UCL

[ZZ14] Huiqing Zhu and Zhimin Zhang. Uniform convergence of the LDG method for a singularly perturbed problem with the exponential boundary layer. Mathematics of Computation, 83 (286):635–663, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2014-83-286/
                                              S0025-5718-2013-02736-6;
                                              http://www.ams.org/journals/
                                              mcom/2014-83-286/S0025-5718-
                                              2013-02736-6/S0025-5718-2013-
                                              02736-6.pdf.

Zhang:2018:NSL

[ZZJZ18] Qian Zhang, Jiwei Zhang, Shidong Jiang, and Zhimin Zhang. Numerical solution to a linearized time fractional KdV equation on unbounded domains. Mathematics of Computation, 87 (310):693–719, 2018. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/
                                              journals/mcom/2018-87-310/
                                              S0025-5718-2017-03229-4;
                                              http://www.ams.org/journals/
                                              mcom/2018-87-310/S0025-5718-
                                              2017-03229-4/S0025-5718-2017-
                                              03229-4.pdf;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?authorName=Zhang%
                                              2C%20Qian;   https://www.
                                              ams.org/mathscinet/search/
                                              authors.html?mrauthid=303173;

 https://www.ams.org/mathscinet/
search/authors.html?mrauthid=
713357;   https://www.ams.
org/mathscinet/search/authors.
html?mrauthid=843093.

Zhang:2015:PAI

[ZZWZ15] Yi Zhang, Liwei Zhang, Jia Wu, and Jianzhong Zhang. A perturbation approach for an inverse quadratic programming problem over second-order cones. Mathematics of Computation, 84(291):209–236, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
journals/mcom/2015-84-291/
S0025-5718-2014-02848-2;
http://www.ams.org/journals/
mcom/2015-84-291/S0025-5718-
2014-02848-2/S0025-5718-2014-
02848-2.pdf.